9 1

A Complete Bibliography of Publications in the VLDB Journal: Very Large Data Bases

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

02 October 2025 Version 1.111

Title word cross-reference

(α, β) [908]. (p, q) [1083]. 10^{(10^{6})} [455]. 3 [496]. ^{+} [1053, 1098, 854, 1238, 23, 971]. ^{2} [1067]. ^{33} [708]. ^{Π} [91]. ^{q} [1196]. c [940]. ℓ [771]. K [685, 1174, 980, 1200, 1058, 429, 553, 581, 909, 638, 906, 391, 612, 798, 937, 1069, 377, 714, 770, 499, 995, 446, 254, 733, 1068, 304, 961, 488, 500, 939, 1061, 545, 789, 747, 1163, 1034, 1187, 489, 728, 694, 1188, 661, 342, 763, 901, 873, 757, 441, 717, 483, 774]. kk [651, 689]. l [744]. l_{p} [779]. n [164, 409]. pq [570]. QR [1136]. s [964]. t [512, 964]. tt [651]. Orpheus [885]. D⃗AG [799].

* [92, 96, 94, 78].

-anonymity [377, 304, 661]. -approximate [940]. -biclique [1083]. -closeness [512]. -closest [689]. -core [908, 728]. -cores [1200]. -gram [409]. -grams [570]. -hop [638]. -most [733]. -Nearest [164, 770, 499, 1174]. -NN [1069, 391]. -overlap [771]. -Path [714]. -reach [980]. -shortest [906]. -TM [708]. -tree [91]. -truss [961].

1999 [192].

2-hop-labeling [1089]. 2005 [309]. 2007 [427]. 2008 [470]. 2010 [557]. 2011 [590]. 2012 [635]. 2013 [697]. 2014 [709]. 2015 [750, 749]. 2016 [812]. 2017 [883]. 2018 [887, 894]. 2019 [915, 930]. 2020 [1028, 1024]. 2021 [1084]. 2022 [1226].

3D [1178]. 3X [475].

‘98 [144].

ABC [1010]. absolute [854]. ABSTAT [1011]. ABSTAT-HD [1011]. Abstract [122]. Abstraction [132]. Accelerated [1045, 741]. Accelerating [725, 998, 1102, 1175, 830, 799, 1177]. acceleration [869]. Access [109, 52, 85, 130, 115, 147, 56, 1219, 208, 305, 809, 587, 370, 617, 955, 389, 368, 233, 169, 989, 335]. accessing [200]. account [1035]. Accounting [188]. Accuracy [473, 1206, 910, 1105]. accurate [231, 1158, 928, 856, 1033, 394]. achieving [304]. ACM [220]. ACME [666]. acquaintance [780]. across [1035, 805, 183]. Active [180, 108, 921, 74, 698, 1044, 1133, 1135, 1164, 387]. ad [105, 393]. Adapative [1032]. Adaptable [62, 439]. adapting [591, 1012]. Adaptive [824, 255, 1236, 562, 641, 725, 923, 351, 341, 1022, 1043, 648, 328, 716, 1173, 408, 746]. adaptivity [1059, 415]. Adding [813]. address [1079]. Addressable [334]. Aditi [36]. adjacency [1176]. adjustment [927]. ADS [746]. Advanced [118]. Adversarial [430, 1031, 1044, 1106, 1238]. Adversarial-knowledge [430]. advisor [699]. aesthetic [952]. against [1014]. agent [835]. agent-based [835]. aggregate [723, 265, 171, 1194]. aggregates [1085, 236]. aggregating [679]. aggregation [564, 712, 970, 541, 1077, 1104]. aggregations [276, 494]. aggregators [250]. agile [1019]. agriculture [242]. ahead [561, 536]. AI [1070]. aided [1032]. alert [1101]. Alfa [1135]. Algebra [83, 55, 161, 818, 339]. Algebraic [80]. Algorithm [77, 139, 279, 745, 530, 423, 1107, 1205, 484, 189, 574]. Algorithms [1220, 310, 318, 245, 149, 975, 230, 1008, 429, 891, 637, 1036, 410, 674, 923, 732, 1032, 789, 870, 303, 379, 1093, 832, 728, 454, 23, 362, 972, 1152, 793, 1153, 1237, 1117, 774, 977]. alignment [1135, 1017, 1023]. all-purpose [335]. Allen [1097, 947]. Allocation [137, 439, 1005]. almost [1169, 773, 825]. almost-latch-free [1169]. Along [70, 69]. AlphaSort [66]. always [1198]. AMIE [702]. Amit [252]. among [449]. analyses [973, 1004, 310, 318, 729, 1013]. Analysis [160, 62, 139, 484, 93, 1236, 230, 558, 510, 530, 392, 637, 696, 283, 578, 411, 1071, 1072, 836, 1137, 1093, 191, 711, 958, 676, 844, 793, 1124, 1076, 1179]. Analytic [52]. analytical [722, 890, 643, 1204, 917, 1155, 1214]. analytics [669, 962, 520, 943, 850, 1132, 575, 1043, 836, 929, 715, 991, 938]. analyzing [1117]. Anatomy [215]. Anchored [986]. anchors [1232]. ANDES [589]. Anna [932]. annotated [616]. annotation [727]. anomalies [956, 634]. anomalous [389]. anomaly [969, 1054, 1206, 1130, 1072, 1125]. Anonymity [374, 534, 377, 304, 661]. anonymized [505]. Anonymizing [476, 513]. Answer [786, 690, 735]. Answering [297, 193, 747, 980, 657, 554, 245, 446, 660, 189]. answers [233, 343, 1194]. any [773, 825]. Anytime [429, 626, 1099, 1222]. AP [707]. AP-Tree [707]. Apache [960, 813, 1124]. applicability [929]. Application [1041, 405, 680, 158, 693]. Application-driven [1041]. Applications [400, 149, 2, 392, 612, 179, 714, 316, 324, 245, 789, 870, 547, 1134, 78, 620, 741, 435, 1145, 1237]. Applying [97, 1148]. Approach [45, 108, 148, 136, 90, 86, 152, 1225, 496, 421, 992, 558, 1096, 242, 795, 371, 768, 521, 1235, 735, 886, 845, 1144, 630, 927, 407, 788, 842, 384, 790, 1075, 408]. Approaches [132, 962, 877, 580, 1074, 1102, 369, 196, 1019]. Approximate [190, 188, 412, 646, 584, 133, 570, 859, 678, 751, 1081, 410, 396, 1146, 723, 1209, 940, 886, 1230, 352, 1188, 863, 1063, 1033]. Approximation [1042, 626, 923, 199, 352, 203, 668, 700]. approximations [1182]. arbitrarily [437]. ArchIS [407]. architectural [958]. Architecture [907, 228, 184, 169, 1165, 330]. Architectures [75, 745, 1216, 369, 546, 634]. archives [295]. archiving [518]. Area [76]. Argument [772]. array [344]. arrays [202]. art [946]. Articles [70]. ary [1173]. Assertions [3]. Assessment [449]. Assignment [74, 691, 477, 1100]. Assisted [1141]. Association [1228, 206, 290]. Asymptotically [89]. Attribute [512, 366, 1185, 234, 91, 844]. attributed [784, 1174, 1088]. attributes [274, 235, 789]. Auditing [599]. augmentation [1174]. Augmented [1076, 797]. Aurora [228, 264]. Authenticated [438]. authentication [478]. authority [1057]. authorization [201, 999]. Auto [1221]. Auto-tables [1221]. autocompletion [902, 765]. AutoCTS [1165]. AutoG [765]. automata [1095, 527]. automated [1126, 942]. Automatic [925, 257, 1213, 269, 712, 196, 634, 971]. automatically [617]. Automating [594, 271]. automation [593]. AutoML [1051, 1134]. Autonomic [415]. autonomous [461]. Autoscaling [932]. availability [592, 655]. averaging [960]. avoidance [934]. Avoiding [718, 731]. aware [535, 496, 1057, 334, 568, 1035, 840, 611, 911, 795, 779, 1130, 1192, 665, 928, 677, 1040, 1043, 353, 1156, 1196, 1202, 963, 707, 684, 952, 1100, 834]. axioms [811].

B [417, 89, 282, 23]. B-link [282]. B-Tree [89]. B-tries [417]. backchase [649]. backed [361]. BAD [921]. balanced [282, 597]. Balancing [116, 265, 988]. Bandwidth [359, 897]. Bandwidth-constrained [359]. Base [49, 88, 84, 215, 754]. Based [148, 55, 11, 130, 29, 90, 149, 123, 152, 8, 92, 586, 417, 1057, 800, 166, 291, 161, 370, 819, 1096, 957, 234, 286, 678, 696, 247, 501, 935, 82, 569, 973, 764, 1081, 542, 911, 240, 725, 580, 487, 204, 400, 243, 1130, 418, 75, 246, 651, 693, 919, 411, 485, 1235, 361, 880, 662, 660, 1135, 1184, 1183, 383, 303, 835, 1137, 947, 902, 422, 945, 203, 272, 1181, 630, 661, 182, 811, 273, 397, 1142, 407, 358, 525, 713, 1091, 1170, 1094, 842, 793, 1157, 759, 573, 698, 920, 280, 802, 1207, 330, 843, 600, 483, 971, 1063, 1065, 790]. based [1100, 289, 408, 629, 799, 652, 700]. Bases [62, 54, 748, 702, 808, 1060]. basket [205]. batch [1098, 758]. batched [948]. BatchHL [1098]. Bayes [339, 381]. Bayesian [222, 435]. Bdual [356]. Bdual-Tree [356]. BEA [259]. Behavior [93, 1018]. Behavioral [60]. behaviour [160]. benchmark [468, 1162, 1023, 815]. benchmarking [852, 972, 1128]. BerlinMOD [468]. BERT [1137, 1139]. BERT-based [1137]. Best [64, 253, 144, 516, 915, 72, 557, 697, 883, 470, 812, 427, 309, 317, 709, 590, 750, 749, 635, 1024, 894, 873]. Better [948]. Between [90, 256, 524, 873]. Beyond [1231, 455, 1120]. bias [962]. bibsonomy [507]. bichromatic [548]. biclique [1034, 1234, 1083]. bicliques [1123]. bidirectional [978, 811]. big [669, 984, 914, 879, 943, 824, 792, 823, 730, 988, 1065, 878, 921, 710, 662]. billion [951, 732, 963]. billion-edge [963]. billion-scale [951, 732]. binary [1120, 364, 792]. Binding [113]. BioGITOM [1239]. bioinformatics [286, 287]. Biomedical [1239]. bipartite [476, 1172, 908, 1175, 1234, 985, 1045, 1152, 1083, 1123, 1078]. biplex [1175, 1234]. Bit [137, 724]. bit-vectors [724]. bitemporal [197]. bitmap [563]. bitruss [985, 1078]. Bitwise [722]. blending [645]. Block [137]. blockchain [1166, 1127]. blockchains [931]. blocking [954]. blocks [916]. Bloom [1059]. bolt [885]. bolt-on [885]. bone [921]. bookmark [507, 503]. Boolean [849]. bottleneck [169]. bottom [1099, 1222]. bottom-up [1099, 1222]. bound [219]. bounded [347, 623, 1081, 300, 399, 657, 668]. bounded-size [347, 623]. bounds [298, 653]. Boyce [1197]. bricolage [752]. bridging [873]. broadcast [452]. bucket [388]. Bucketization [512]. buddies [533]. budget [939]. Buffer [52, 8, 293]. Buffering [146]. Buffers [68]. BugDoc [1038, 1056]. Building [833, 88, 916]. built [1227, 509]. built-in [1227, 509]. Business [224]. Business-to-business [224]. Butterfly [1078, 1045, 1154, 1157]. ByShard [1092]. Byzantine [1092].

C5 [1198]. Cache [313, 321, 66, 947, 497, 329, 739, 1201, 546]. Cache-conscious [313, 321]. Cache-efficient [947]. Cache-Sensitive [66]. Caching [64, 75, 123, 65, 157, 361, 1168, 546]. Calculus [29]. Calibrating [676]. capabilities [509, 1082]. cardinalities [263]. Cardinality [820, 1109, 809, 235, 630]. Cascade [840]. Cascade-aware [840]. case [960, 226, 970, 653, 771]. Categorical [148, 464]. categorization [1032]. categorized [953]. causal [1022]. CDBTune [971]. cell [731, 412, 432]. CellJoin [432]. centric [570, 696, 547, 715, 1144, 1070, 896, 794]. century [810]. certain [560, 734]. CG [1067]. Chain [123, 1031]. Chairmen [144, 143]. challenges [877, 888, 1070]. change [1094]. changes [464, 258]. Characterization [52, 737]. Checking [162, 552]. checkpoint [1129]. checks [934]. Chronological [27]. citizens [275]. CLARO [579]. class [275, 232, 1079, 718]. classes [718]. Classification [124, 132, 231, 1069, 487, 1079, 1160, 796, 381, 1164]. clause [1119]. Cleaning [900, 949, 1238, 1094]. Client [118, 114, 75, 103]. Client-Server [75, 103]. Client/Server [118, 114]. Clip [131]. clique [905, 717]. cliques [841, 1002]. cloned [1198]. closed [987]. closeness [512, 188]. closest [689, 1033]. closure [586]. cloud [664, 778, 1216, 665, 727, 715, 827, 932, 1161, 971, 999]. ClouDiA [699]. clouds [1128, 675, 699]. clues [679]. cluster [941, 781]. clustered [674]. Clustering [148, 679, 152, 833, 722, 800, 464, 327, 1004, 371, 336, 621, 1174, 1122, 1047, 842]. clusters [443]. co [722]. co-clustering [722]. Coalition [1100]. Coalition-based [1100]. coarse [541]. Coconut [860]. Codd [1197]. code [821, 1195]. cognizant [191]. coherent [851]. Cohort [929]. cold [856]. collaboration [404]. Collaborative [277, 637, 1067, 826]. Collection [125, 422, 1070]. collections [240, 212, 831]. collective [1235, 899]. Color [120]. Color-Spatial [120]. Column [15, 16, 378]. combination [357, 606]. combinations [789]. combines [907]. Combining [172]. commerce [179, 1031, 1020, 182]. Commercial [101, 100, 192]. Commercialize [35]. communication [628]. communities [782, 924]. community [1053, 784, 914, 879, 961, 1020, 866]. Compact [846, 1110, 1037]. CompactLTJ [1241]. Comparable [601]. Comparative [139, 858, 1230, 1124]. Comparing [881]. Comparison [946]. comparisons [729]. competition [831]. competitiveness [1103]. compilation [1025, 967]. compiled [895]. complements [194]. Complete [83, 378, 233, 811]. completion [1099, 1222]. Complex [71, 982, 878, 221, 450, 24, 1076]. Complexity [56, 540]. compliance [349]. components [178]. Composing [241, 287, 242]. composite [411]. composition [354]. comprehensive [1023]. compress [724]. Compressed [818, 1176, 893, 333, 419, 672]. compressing [903]. Compression [331, 92, 680, 1158, 918, 981, 938]. Compressive [672]. computation [586, 881, 817, 905, 951, 979, 310, 318, 1068, 633, 908, 473, 412, 236, 1161, 907]. computational [848]. compute [896]. computers [415]. Computing [233, 602, 530, 613, 733, 1042, 383, 277, 827, 941, 682, 997]. Concept [272]. Concept-based [272]. Concepts [118]. conceptual [783, 803, 607, 1214]. Concurrency [9, 282, 104, 99, 21, 520, 91, 1198, 23, 773, 825, 1117]. Conditional [688, 632]. conductance [529]. confidence [939, 290]. Configurable [110, 111]. Configuration [41, 30]. conflict [1210]. conflicts [406]. Conformity [677]. Conformity-aware [677]. conjunctive [726, 596, 889, 827]. Connection [113]. connectivity [1096]. connectivity-based [1096]. conscious [313, 321, 329]. Considering [20]. Consistency [81, 587, 13, 54, 634, 640, 492, 655, 258]. Consistent [14, 312, 320, 13, 1232, 1129, 546]. consistently [492]. Consolidation [7, 91]. Constant [137, 970]. Constrained [472, 327, 359, 1208, 1134, 835, 964, 1188, 1118]. constraint [401, 327, 1062, 602, 761]. Constraints [40, 61, 117, 13, 820, 39, 817, 531, 980, 1090, 776, 949, 922, 622, 780]. constructing [284]. construction [410, 204, 854, 754]. Consumption [8]. containing [814]. Containment [420, 807, 968, 950]. contended [1029]. Content [45, 130, 504, 1066, 334]. Content-Addressable [334]. content-and-structure [1066]. Content-Based [130]. Context [166, 911, 90, 422, 684, 834]. Context-aware [911, 684, 834]. Context-Based [90, 166, 422]. Continuous [553, 937, 526, 129, 995, 137, 478, 673, 477, 496, 294, 992, 581, 463, 263, 721, 385, 718, 543, 670, 574]. Continuous-Media [129]. continuous-time [992]. contraction [1037]. Control [9, 99, 21, 587, 370, 327, 935, 1198, 282, 895, 368, 939, 23, 773, 825, 1117]. convergence [859]. conversations [413]. Cooperative [2]. coordination [934]. copy [1057]. copy-based [1057]. CORAL [34]. core [921, 908, 1208, 870, 728, 851]. coreness [986]. cores [1026, 1200]. coroutines [845]. corpus [400]. corpus-based [400]. Correction [1054, 914, 1055, 1166, 1056, 1222, 1104, 968, 1105]. correctness [78]. Correlated [656, 1165]. correlating [559]. correlation [743, 824, 528, 1065, 1115]. correlations [688, 457, 1115]. corrupted [1014]. Cost [803, 232, 1202, 766, 819, 1147, 532, 948, 1216, 955, 1192, 919, 768, 251, 222, 1091, 642, 408]. Cost-aware [1202, 1192]. cost-based [919, 1091, 408]. Cost-driven [232]. Cost-effective [803, 819, 642]. cost-efficient [1147]. cost-oriented [768]. cost/performance [948]. costly [397]. Costs [77, 105, 795]. counter [895]. Counting [125, 1218, 1081, 928, 704, 399, 1045, 1154, 1157, 1083, 1075, 1078]. counts [1189]. covering [922]. covers [1203, 714]. CPU [646, 781]. CQL [294]. cracking [711]. crawling [593]. Creating [460, 726]. creation [563, 892]. criteria [78]. critical [1172]. Cross [1031, 919, 922]. Cross-chain [1031]. cross-platform [919]. cross-range [922]. crowd [615, 1032]. crowd-aided [1032]. crowd-sourced [615]. crowds [621]. crowdsensing [1192]. Crowdsourced [939, 819, 920]. crowdsourcing [619, 786, 772, 691, 875, 1151, 1100, 1005]. cryptanalysis [1150]. cube [375, 482, 371]. cuckoo [893]. curious [533]. current [783, 877]. curves [356]. custom [821]. customer [733]. customizable [1195]. customizing [179]. cut [1156]. Cyclic [125].

D [496]. DaMoN [915, 1024, 894].

dampening [1086]. dashboard [506]. Data [209, 44, 703, 1147, 50, 118, 431, 107, 148, 1021, 55, 4, 20, 137, 1030, 150, 976, 1055, 47, 1204, 94, 147, 13, 53, 116, 30, 122, 152, 79, 28, 1070, 228, 428, 695, 664, 933, 230, 925, 1185, 669, 838, 208, 429, 778, 570, 1057, 262, 962, 268, 984, 969, 1054, 1052, 618, 615, 291, 558, 161, 315, 323, 912, 520, 298, 229, 430, 463, 464, 656, 775, 909, 957, 654, 285, 173, 476, 471, 223, 327, 619, 518, 696, 306, 451, 1069, 560, 302, 753, 377, 593, 563, 283, 650, 900, 260, 540, 1229, 537, 270, 460]. data [580, 810, 565, 373, 515, 1009, 813, 219, 351, 243, 1210, 712, 462, 1158, 860, 443, 919, 665, 256, 194, 621, 1060, 364, 500, 936, 1169, 1212, 445, 926, 824, 1014, 1106, 1163, 729, 874, 200, 688, 534, 385, 1183, 375, 205, 475, 1050, 466, 292, 836, 393, 686, 974, 1238, 511, 871, 1006, 872, 892, 1141, 865, 931, 620, 792, 990, 203, 197, 454, 1144, 187, 265, 412, 1150, 630, 491, 601, 661, 676, 796, 513, 342, 855, 823, 1193, 603, 579, 1195, 655, 844, 326, 672, 1091, 1173, 1094, 631, 461, 435, 827, 901, 1124, 573, 1145, 681, 1076, 698, 920]. data [910, 1105, 1207, 486, 248, 384, 941, 896, 1065, 452, 408, 746, 822, 878, 921, 710]. data-adaptive [1173]. data-centric [570, 696, 1144, 1070]. Data-induced [1030]. data-intensive [664]. data-oriented [306]. Database [40, 61, 52, 33, 6, 146, 43, 114, 11, 212, 90, 75, 35, 139, 95, 80, 32, 93, 132, 76, 56, 17, 36, 380, 22, 37, 48, 39, 370, 562, 234, 338, 594, 948, 869, 753, 751, 1015, 1186, 42, 316, 324, 1227, 1133, 400, 214, 332, 138, 256, 174, 758, 752, 599, 361, 1129, 169, 94, 592, 172, 1136, 1159, 106, 186, 31, 78, 458, 701, 711, 1150, 741, 1139, 192, 407, 1145, 971, 1073]. database-backed [361]. database-driven [256]. Databases [51, 59, 1241, 141, 155, 350, 41, 1228, 124, 108, 129, 4, 7, 130, 38, 154, 126, 86, 128, 153, 53, 125, 122, 152, 74, 151, 858, 209, 275, 496, 833, 268, 1131, 481, 846, 163, 96, 301, 285, 344, 337, 965, 840, 247, 627, 468, 609, 82, 853, 626, 232, 267, 526, 453, 460, 446, 885, 254, 495, 396, 712, 389, 1132, 368, 805, 521, 583, 743, 514, 545, 667, 660, 307, 426, 393, 1039, 372, 738, 457, 343, 589, 1108, 584, 904, 461, 604, 308, 438, 658, 765, 330, 600, 683, 797, 571, 577, 554]. datalog [798, 610, 966, 760]. Dataset [876, 1103]. datasets [847, 250, 347, 493, 383, 624]. dataspace [973]. dataspace-based [973]. DAWN [373]. DB [885, 1139]. DB&IR [340]. DB-BERT [1139]. DBMS [428, 480]. DBMSs [1204, 372, 1161, 642]. DBOS [1211]. DBSP [1213]. DBToaster [639]. DCT [373]. Deadlock [139]. Dealing [56]. deals [1031]. debugging [1021, 1038, 1056]. decentralized [1196]. decidability [706]. decision [962, 167, 222]. Declarative [109, 159, 966]. DECLARE [35]. Decomposition [1233, 1096, 1051, 870, 1136, 728, 985, 762, 1237, 1078]. decompression [918]. Decorating [727]. decoupling [1121]. dedicated [415]. Deductive [33, 35, 34, 32, 36, 37, 31]. Deep [1044, 1022, 209, 617, 593, 1074, 1122, 713, 1070, 1080, 1145, 971]. DeepDive [754]. definitions [555]. Delay [353]. Delivering [134]. delta [639]. Demarcation [40, 39]. denial [1090]. denormalized [268]. Dense [636]. densest [1102, 1153]. Density [1237]. Dependencies [27, 596, 555, 1146, 976, 1055, 1010, 1230, 978, 601, 811, 975, 461, 1203]. Dependency [300, 632, 719, 524, 802]. dependent [1061]. Deploying [363]. deployment [699]. Depth [433, 852, 327, 1169, 1179]. derived [551]. descent [1145]. description [358]. Design [33, 555, 62, 2, 53, 84, 783, 472, 529, 803, 405, 1106, 1141, 990]. Designing [178]. designs [609]. desired [1006]. desktop [405, 501]. destination [588, 681]. Detecting [389, 1050, 464, 688]. Detection [139, 279, 1008, 969, 1054, 1206, 721, 764, 953, 1130, 336, 719, 1020, 1183, 490, 956, 1125, 348, 634, 644, 1115]. detectors [1072]. deterministic [298, 527]. devices [877]. DHT [544]. diagram [604, 604]. Dictionary [92, 678, 918]. Dictionary-based [92, 678]. DIFF [933]. different [173, 1115]. Differential [10, 656, 1086, 1015, 704, 661]. Differentially [1194, 625, 796]. differentiated [718]. digest [426]. digital [201, 422]. dimension [420]. Dimensional [154, 47, 46, 279, 722, 463, 168, 614, 854, 1191, 203, 328, 1188, 176, 548, 441, 248, 1033]. dimensionality [328]. dimensions [430, 551]. DIRECT [1129]. Directed [1233, 1102]. Direction [542]. Direction-based [542]. directions [1204, 363]. directly [938]. disaggregated [1170]. disclosure [307]. discord [1112]. discoveries [732]. Discovering [1172, 1146, 1006]. discovery [925, 374, 271, 913, 1210, 1003, 743, 528, 772, 1047, 1090, 620, 978, 221, 811, 290, 603, 975, 1153, 866]. Discrete [44]. discriminant [510, 231]. Disjoint [769]. Disk [116, 417, 489]. disk-based [417]. disk-resident [489]. Disseminating [408]. Dissemination [333, 544]. Dissociation [751]. Distance [149, 981, 881, 588, 530, 391, 996, 1098, 613, 1107, 437, 862, 303, 1089, 1149, 902, 868, 603, 397, 657, 761, 622, 573]. Distance-Based [149, 303, 902]. Distances [164]. Distilling [808]. Distributed [40, 671, 551, 139, 991, 956, 76, 1233, 767, 1225, 39, 1052, 859, 451, 935, 764, 1015, 960, 566, 304, 903, 238, 1201, 686, 849, 720, 439, 978, 630, 325, 646, 988, 802, 687, 781, 861, 452, 907]. Distribution [28, 1147, 392, 1021]. distributional [1125]. Dittrich [345]. divergence [895]. Diverse [744, 1068]. Diversified [912, 717, 1034, 851]. Diversifying [1046]. diversity [654, 685, 662, 1151]. DNA [212]. doctor [1242]. Document [109, 123, 17, 291, 301, 240, 955, 490, 18, 831]. Documents [50, 162, 447, 218, 217, 616, 187]. domain [1057, 1079, 1146, 672]. domain-aware [1057]. dominating [549, 441, 483]. down [186]. DP [1205]. Dragoon [943]. driven [627, 1041, 232, 712, 256, 1166, 1127, 940, 1143, 1151]. drives [740]. DTDs [448]. dual [530, 281, 398, 356]. dual-tree [530]. DumpyOS [1173]. duplicated [460]. during [406, 1014]. Dynamic [531, 164, 293, 28, 658, 423, 841, 937, 996, 1098, 753, 889, 1049, 639, 368, 364, 1064, 1138, 292, 379, 251, 170, 620, 1082, 1101, 328, 922, 830, 837, 767, 1237, 408]. Dynamical [148, 791]. dynamically [597]. Dynamics [95, 106]. DynQ [1082].

e-commerce [179, 1020, 182]. e-services [178, 183, 180]. Early [35, 167, 940]. early-termination [940]. Earth [573]. ECC [762]. Eccentricities [857]. edge [636, 1096, 963]. edit [881, 1107, 902, 761, 622]. edit-distance [761]. editing [560]. Editorial [207, 155, 1028, 1226, 213, 1084, 930, 266, 887, 165, 237, 156, 225, 177, 402, 867, 185, 302, 260, 121]. Editors [346]. Effective [784, 1087, 811, 151, 713, 279, 1206, 819, 803, 954, 1209, 871, 335, 573, 1153, 1088, 843, 642, 1164]. effectively [690]. effectiveness [246]. effects [462]. efficiency [1000, 731, 929, 832, 508, 706, 265, 605]. Efficient [1241, 218, 905, 951, 638, 217, 423, 337, 798, 306, 814, 1004, 954, 689, 1220, 567, 499, 137, 923, 1133, 854, 539, 719, 142, 364, 536, 743, 1061, 445, 908, 1208, 1201, 490, 205, 86, 379, 489, 416, 964, 902, 198, 945, 978, 434, 1150, 1188, 1234, 399, 362, 1152, 842, 498, 573, 899, 1153, 807, 687, 896, 866, 861, 622, 790, 1225, 571, 629, 1115, 279, 496, 455, 497, 745, 606, 992, 1206, 821, 817, 1199, 227, 1147, 344, 637, 1171, 784, 943, 569, 316, 324, 1227, 373, 446, 1068, 961, 578, 485, 633, 174, 1166, 1127, 986, 940, 1178, 785, 816, 426, 172, 1168, 484]. efficient [1089, 947, 871, 544, 931, 589, 342, 855, 1193, 335, 844, 904, 985, 585, 384, 762, 682, 971, 1217, 774]. Efficiently [581, 10, 38, 187, 591, 773, 1218, 1112, 250, 770, 707, 825]. Efforts [35, 786]. EGO [614]. EI-LSH [940]. Elastic [546, 730]. elasticity [655]. elicitation [517]. eliminate [492]. Eliminating [628]. Elite [730]. embedded [1082, 975]. Embedding [693, 1201, 944, 1101, 1088, 1207, 942, 997, 554]. Embedding-based [693]. embeddings [1179]. emerging [850, 313, 321]. empty [735]. empty-answer [735]. enabled [183, 1117]. Enabling [1171, 355, 486, 492, 727, 224]. encoding [1124]. encrypted [1150, 827]. encryption [1101]. end [1051]. end-to-end [1051]. endurable [411]. Energy [208, 1229, 265]. Energy-performance [208]. engagement [866]. engine [219, 1040, 475, 917, 1082, 652]. Engineering [497, 962, 659]. engines [993, 608]. Enhanced [122, 1192]. Enhancing [1057, 388, 661, 777]. ensembling [620]. Ensuring [654]. enterprise [1060]. entities [953]. entity [421, 819, 678, 1044, 621, 1087, 1162, 1137, 1093, 465, 624, 631, 1179, 1017, 797, 1023]. entropy [312, 320]. EntropyDB [886]. Enumeration [964, 905, 768, 1083]. Environment [1, 8, 1092, 277, 251, 408]. Environments [64, 65, 566, 544]. epiC [710]. epoch [1169]. Equivalence [444, 726]. equivalences [1049]. Equivalent [26]. era [878]. Eris [1112]. eRiskCom [1020]. Erratum [750, 106, 825]. error [298, 373, 854, 902, 399, 668, 452]. error-bounded [668]. error-prone [452]. Establishment [449]. estimate [473]. estimates [1036]. Estimating [1189, 68, 396]. Estimation [28, 809, 551, 373, 168, 312, 320, 433, 630, 591, 1142, 830, 1109]. estimators [274, 418]. ETL [783]. eTuner [314, 322]. Euclidean [553, 437, 862, 303]. Evaluating [1176, 853, 250, 872, 269]. Evaluation [33, 496, 367, 1111, 337, 73, 1113, 459, 946, 484, 415, 792, 728, 711, 1144, 607, 589, 868, 171, 829, 498, 598]. Event [882, 114, 528, 644, 982, 523, 878, 1183, 450]. events [882]. eventual [640]. everything [597]. evidence [357]. Evolution [141, 70, 594, 1116, 1015, 69]. Evolutionary [251]. evolving [721, 347]. Exact [723, 1008, 429, 588, 437, 489, 837, 657, 108]. ExactSim [972]. example [555, 1212]. examples [1221]. executing [287]. Execution [97, 20, 294, 821, 1199, 552, 608, 1085, 967, 411, 816, 385, 999]. executions [1056]. Exemplar [742]. experimental [82, 1191, 729, 711, 873, 977]. experiments [1013]. expert [406]. explainable [882, 1108]. Explaining [884]. explanation [933, 1038, 1056]. explanations [1016, 1194]. Explicit [53]. exploitation [824, 1065]. Exploiting [167, 1079, 1209, 738, 725, 457]. exploration [406, 1004, 1133, 174, 1043, 1006]. Exploratory [618, 1004, 836, 904]. Exploring [256, 831, 627]. exponential [1036, 332]. expression [263]. expressions [218, 227, 706]. expressive [637]. expressiveness [456]. Extended [112, 65, 885, 947, 888, 1214]. Extending [610, 372, 479]. Extensible [108, 53, 293, 710, 138]. extension [739, 326]. extensions [1120]. External [731, 66, 968, 950, 331]. extracting [666]. extraction [678, 593]. extreme [663, 435].

F [1132]. F-IVM [1132]. facet [1137].

faceted [1039]. facilities [801]. Factor [20]. factorization [849]. Failure [1, 411]. Failure-Prone [1]. Fairness [994, 1100]. family [1059, 335]. Fast [1008, 231, 532, 996, 702, 756, 1049, 135, 120, 849, 898, 974, 1012, 856, 851, 893, 614, 1107, 967, 918, 801, 963, 672, 910, 1105, 331, 1033]. Fast-adapting [1012]. Faster [918, 398]. fault [859]. favorite [651]. Feature [112, 124, 1079, 1121]. feature-oriented [1121]. features [480, 172, 952]. Federated [141, 4, 7, 56, 1160, 1166, 1127, 1012]. federation [1193]. FedST [1160]. feedback [406, 158, 1130, 462]. FENECIA [411].

FERRARI [904]. Fetches [68]. Fibonacci [59]. FICOM [1164]. file [332, 386, 907, 371]. Files [119]. filling [356]. filter [1059]. filtering [218, 1162, 544, 856]. filterings [606]. Filters [97, 893]. Finding [766, 906, 806, 651, 782, 924, 694, 261, 471, 770, 1163, 988, 794]. Fine [1047, 931]. Fine-grained [1047, 931]. Finite [68, 527]. Firm [146]. first [275, 913, 810]. first-class [275]. first-sight [913]. fixes [560]. Flash [739, 877, 731, 474]. flexibility [1140]. Flexible [81, 344, 114, 1158, 305, 723]. FlexpushdownDB [1161]. Fliggy [1080]. flow [895, 1109]. flows [538]. Flying [967]. focus [355]. forecasting [982, 674, 1165]. forest [1062]. Foreword [502, 143, 144]. Form [361, 787, 595, 1197]. Form-based [361]. Formal [966, 210, 195]. formation [826]. formats [805, 1204]. forms [617]. formulations [793]. forward [827]. forwarding [334]. FOSS [1242]. found [815]. foundations [294]. Four [1218, 235]. Four-Node [1218]. Fourier [303]. Fourier-related [303]. FPGAs [869, 550]. fractional [992]. fractured [226]. fragmentation [1015]. fragmented [140]. framework [230, 692, 618, 512, 819, 464, 1035, 405, 637, 518, 678, 973, 269, 316, 324, 373, 1180, 1177, 1068, 304, 485, 828, 1174, 480, 1067, 835, 439, 963, 1214, 844, 1142, 425, 904, 1091, 585, 988, 987, 920, 765, 761, 479, 1164, 1117, 1033]. frameworks [764]. Free [334, 1036, 1169, 355]. frequent [423, 401, 471, 596, 313, 321, 583, 666, 442, 987, 384]. frequently [639]. fresh [639]. Freshness [153, 256]. friendly [1008]. Fu [1193]. full [1026, 409, 536, 1196, 544, 1145]. full-graph [1196]. full-text [409, 536, 544]. fully [996]. function [726]. Functional [145, 632, 596, 555, 1230, 975, 1203]. Functional-Join [145]. functions [705, 650, 361]. further [729]. fusion [1057, 1183]. future [877, 1204]. fuzzy [429, 536, 582].

G [988]. G-thinker [988]. gain [290]. game [534, 920]. game-based [920]. gap [1036]. gaps [693]. Garbage [125]. GARF [1238]. gbase [578]. GCN [1018]. General [889, 13, 1013, 692, 223, 689, 1103, 1091, 988, 987, 479]. general-purpose [987]. generalized [606, 498]. Generating [821, 13, 1195, 822, 667]. Generation [124, 535, 703, 814, 836, 1143]. generative [1106, 1238]. Generator [115]. generators [1013]. Generic [465, 421, 786, 1068, 1163]. Genericity [380]. genomic [584]. Geo [780, 909, 957, 935, 533, 1188, 705]. geo-distributed [935]. Geo-social [780, 533, 705]. geo-tagged [1188]. geo-textual [909, 957]. geometric [386]. geometry [848]. GeoSparkViz [941]. geospatial [941]. gigabytes [864]. GIST [758]. Given [164]. Givens [1136]. glass [815]. Global [195, 21, 671, 513]. globally [671]. Glue [33]. Glue-Nail [33]. GMAP [79]. GNN [1053]. GNNs [1164]. Go [801]. good [1128, 1191]. Gossip [935, 630]. Gossip-based [935, 630]. GPS [1073]. GPU [745, 1180, 998, 741, 646, 1157, 781]. GPU-accelerated [741]. GPU-based [1157]. GQL [1228]. gradient [903, 1145]. grading [703]. GRADOOP [991]. GRAIL [572]. grained [1047, 931]. gram [409]. grammars [366]. grams [570]. granularities [173, 541]. Graph [1241, 1228, 1239, 776, 683, 1095, 1008, 1110, 881, 984, 529, 1096, 1053, 532, 476, 841, 965, 840, 1041, 721, 764, 913, 1126, 1227, 1180, 175, 928, 583, 1209, 1121, 1135, 1099, 1222, 1067, 989, 1190, 835, 1156, 1196, 715, 991, 888, 1101, 927, 1108, 904, 842, 793, 1013, 988, 765, 843, 600, 762, 942, 1023, 622, 554, 652]. graph-based [652]. graph-theoretic [175]. graph-theoretical [927]. graphic [670]. Graphical [96, 758, 591]. Graphs [1233, 1218, 993, 838, 839, 905, 951, 1037, 879, 770, 1180, 1177, 961, 578, 908, 1071, 1175, 720, 980, 1011, 888, 963, 1197, 716, 1234, 420, 830, 985, 1045, 1152, 1001, 657, 1083, 1123, 572, 1200, 682, 861, 1237, 1078, 605, 774, 851, 914]. GRELA [1209]. Grid [289, 680]. Grid-partition [289]. ground [972]. group [814, 836, 826, 508, 780]. group-by [814]. grouping [1158]. groupings [476]. groupjoin [814, 1085]. groups [694, 362, 829]. Growing [616]. gStore [652]. guaranteed [1000]. guarantees [1069, 1050, 657]. guarding [1014]. Guessing [435]. Guest [346, 165, 237, 155, 156, 225, 177, 402, 185, 302, 213, 121, 930, 260]. Guided [394].

Haar [854]. Hadoop [648]. HaLoop [558]. Handling [61]. hard [724]. hard-to-compress [724]. Hardware [736, 737, 561, 1102, 816, 1167]. Harvesting [519]. Hash [97, 280, 1111, 725, 1059, 945]. Hash-based [280, 725, 945]. hashing [779]. HAVING [458]. hB [91]. HD [1011]. HE-Tree [464]. heavily [1134]. heavy [688]. HERMES [1060]. Heterogeneity [141, 136]. Heterogeneous [6, 3, 1225, 481, 821, 96, 285, 179, 1079, 1192, 178, 1047, 183, 601, 781]. HeteroStamp [1192]. Heuristic [102]. heuristics [881]. HFUL [1035]. hide [845]. Hierarchical [124, 371, 307, 99, 1096, 1097, 753, 390, 336, 1154, 1217]. Hierarchically [419]. Hierarchies [123, 2, 209, 482]. Hierarchy [110, 111, 1089, 897]. High [47, 84, 655, 67, 605, 279, 817, 520, 935, 973, 1229, 592, 897, 203, 328, 1188, 176, 972, 966, 248, 1033]. high-bandwidth [897]. High-Dimensional [47, 279, 203, 328, 1188, 176, 248, 1033]. high-energy [1229]. High-Level [84]. high-order [817]. high-performance [935]. high-precision [972]. high-variety [973]. higher [639]. higher-order [639]. Highly [1052, 753, 566, 1158, 1195, 574]. Hilogx [1130]. HINT [1097]. Hippocratic [307]. histogram [625]. Histograms [358, 388, 530]. Historical [69, 70, 333, 801, 910, 1105, 1200]. History [147]. Hit [52, 983]. hitters [688]. HMAP} [173]. HMI [1217]. hoc [393, 105]. Holism [469]. holistic [250, 764, 735, 1075]. Hop [964, 638, 1062, 1089]. Hop-constrained [964]. hot [261]. HPCache [1168]. Hu [1193]. Hu-Fu [1193]. Hub [700]. huge [1201]. human [249, 540, 1130]. Humming [693]. Hybrid [724, 621, 1035, 943, 616, 1155, 289]. Hyper [1149]. Hyper-distance [1149]. hypercores [1064]. HyperFile [50]. Hypergraph [1120]. Hypergraphs [113, 1064, 1208, 1149]. hyperparameter [1165]. HyperStorM [109]. Hypervideo [130]. hypothesis [836].

I/O [948, 961, 940, 762, 682, 774].

I/O-efficient [774]. ICS [1053]. ICS-GNN [1053]. identification [636, 234, 505, 1234]. identify [962]. Identifying [1123]. IDList [629]. II [7]. Image [45, 120, 172, 424]. Images [132]. imbalance [1079]. Impact [101, 737, 100]. Implantation [141]. Implementation [33, 118, 6, 84, 24, 195, 138, 828]. Implementing [354, 166]. Implication [117]. implications [380]. implicit [1146]. imprecise [315, 323]. Improve [153, 1080]. improvements [729, 1013]. Improving [97]. In-database [1186, 758, 1145]. in-depth [852, 1179]. In-Memory [959, 869, 974, 945, 958, 781, 1033]. in-network [538]. In-order [970]. inapproximability [1042]. incentive [1166, 1127]. incentive-driven [1166, 1127]. including [1076]. inclusion [596]. incomplete [455, 1202, 865, 491, 461]. incorporated [411]. increasing [731]. Incremental [841, 6, 1090, 754, 728, 927, 631, 236, 1213, 1126, 940, 379, 1091]. Independence [79, 810, 351, 771].

Independent [3, 559, 403, 194, 829, 788]. indeterminacy [173]. Index [41, 529, 1240, 47, 104, 110, 86, 68, 227, 532, 1097, 91, 235, 409, 1138, 111, 845, 1181, 325, 1173, 842, 572, 1207, 600, 289, 452, 629, 746]. index-based [842]. indexes [1111, 563, 204, 860, 1191, 1170, 753]. Indexing [775, 877, 164, 295, 281, 86, 538, 197, 288, 176, 1219, 1232, 447, 451, 501, 641, 212, 740, 437, 1208, 1043, 172, 974, 648, 395, 203, 328, 922, 397, 1066, 438, 356]. Indices [135, 17, 388, 18]. individual [1144]. induced [1030]. inequality [756]. inference [847, 748, 552, 751, 487, 758, 1022, 1014, 1217]. influence [992, 923, 677, 1231]. influential [782, 1007]. InfoQuilt [221]. Information [49, 120, 1019, 17, 455, 160, 338, 405, 333, 580, 1030, 246, 416, 1202, 439, 18, 273]. information-theoretic [580]. infrastructure [664, 1113, 730]. Ingress [1126]. insecure [527]. Instance [234, 440, 653]. Instance-based [234]. Instance-level [653]. instant [1129]. Instantiating [38]. integrability [1210]. integrable [1210]. Integrated [123]. Integrating [291, 179, 126, 132, 160, 617, 368, 178]. Integration [3, 406, 1147, 234, 338, 405, 619, 431, 257, 1210, 462, 534, 792, 698]. Integrator [257]. Integrity [61, 85, 117, 162, 192]. intelligence [506]. intensive [664, 443, 691, 1181]. inter [206]. inter-transactional [206]. Interaction [535, 96, 528, 716]. Interaction-aware [535]. interactions [1192, 224, 777]. Interactive [142, 162, 934, 1053, 1133, 645, 872, 424]. Interactively [1006]. interest [944]. interesting [556, 688]. interface [933, 1108]. interfaces [858, 257, 1214]. Interleaving [845]. intermediate [731]. Internal [950, 968]. Internet [181, 466, 630, 182]. Internet-based [182]. Internet-scale [466, 630]. interoperability [1150]. interpretable [1238]. intersection [1182, 670, 574]. intersections [1177]. Interval [959, 769, 1097, 1182, 947]. intervals [173]. InterViso [56]. Intrinsic [136]. Introduction [338, 43, 127, 128, 32, 42, 31]. intrusion [336]. inverse [514]. inverses [586]. Inverted [17, 18]. investments [895]. involving [175]. IoTDB [1124]. IR [344]. Islands [737]. isolation [671, 721]. Isomorphism [1239]. Issue [57, 736, 1226, 43, 1084, 128, 32, 17, 516, 915, 838, 22, 37, 48, 63, 557, 1131, 697, 883, 984, 615, 470, 402, 812, 338, 302, 283, 427, 260, 537, 309, 317, 453, 1009, 709, 590, 750, 749, 1167, 930, 635, 1024, 894]. Issues [1, 118, 9, 623, 224, 369, 706, 363]. items [471, 1163]. itemset [392, 384]. Iterative [1038, 1056, 558]. itinerary [1080]. IVM [1132].

Java [380]. Join [145, 97, 267, 77, 20, 135, 1205, 102, 814, 569, 432, 105, 254, 614, 311, 319, 816, 1142, 837, 807, 968, 950, 574, 815]. join&sort [1205]. Joins [959, 135, 897, 891, 570, 800, 775, 979, 569, 1182, 360, 436, 889, 1107, 756, 998, 816, 1187, 1136, 277, 947, 845, 804, 670]. Joint [624, 1165]. Journal [70, 1091, 220]. journey [882]. JSON [847, 864]. judgment [1210]. judgments [939]. jungle [919]. Juxtaposed [352].

keeps [1198]. kernel [510]. key

[617, 1184, 781]. key-value [1184, 781]. keys [734, 1203]. Keyword [595, 1058, 690, 692, 912, 957, 937, 492, 509, 349, 511, 434, 1108, 763, 899, 698, 1207, 797, 629, 774]. keyword-based [957]. keywords [806, 1235, 361, 343]. Klaus [345]. kNN [484, 365, 414]. Know [733]. Knowledge [88, 46, 84, 54, 993, 430, 748, 249, 702, 808, 1079, 462, 1060, 524, 1099, 1222, 989, 1011, 691, 754, 221, 942, 1023, 1217]. Knowledge-Base [84]. knowledge-intensive [691]. KRISYS [118].

Label [1118, 980]. Label-constrained [1118]. labeling [981, 1062, 1089, 788, 920, 280]. labelling [996, 1098]. lakes [1210]. Language [59, 61, 884, 84, 858, 294, 586, 161, 173, 965, 593, 1212, 1143, 1195, 1108, 1217, 340]. language-based [586]. Languages [71, 1224, 12, 80, 32, 1213, 31]. Large [124, 619, 451, 152, 1233, 1218, 933, 211, 852, 890, 969, 1054, 558, 905, 748, 404, 840, 643, 818, 1098, 1004, 250, 721, 270, 1172, 1180, 446, 685, 645, 212, 396, 578, 1002, 575, 1212, 1064, 1099, 1222, 383, 474, 172, 1175, 980, 386, 1011, 602, 511, 715, 888, 1195, 985, 670, 1001, 759, 1083, 572, 1200, 683, 1237, 571, 605, 554]. Large-Scale [1218, 619, 451, 933, 211, 852, 558, 404, 643, 818, 1098, 1004, 270, 1172, 575, 1064, 1099, 1222, 715, 985]. lasting [743]. latch [1169]. latency [1025, 1050, 845]. latent [416]. later [1211]. lattice [600]. lattice-based [600]. law [392]. layer [851, 907]. layered [405]. layouts [209]. Lazy [153]. leakage [308]. Leapfrog [1241]. Learned [1236, 1075, 1191, 1138, 461, 1242]. Learning [239, 158, 1215, 1063, 962, 1131, 748, 1148, 818, 1021, 960, 1044, 1133, 903, 1128, 1113, 1074, 1209, 1122, 1166, 1127, 524, 1014, 1048, 1135, 1099, 1222, 989, 832, 1142, 713, 966, 1070, 308, 1145, 698, 1207, 971, 1023, 1164]. Learning-based [1063, 1142, 713, 971]. learning-dependency [524]. learning-to-rank [1148]. length [358]. Lero [1148]. Level [84, 24, 1072, 897, 653]. Leveraging [979, 837, 1080, 891, 1192]. LHAM [147]. libraries [201]. life [283]. lifecycle [375, 482]. lifted [751]. lifting [910, 1105]. lightweight [1053]. like [388]. limited [906, 233]. Lindex [600]. line [523, 385]. Lineage [223, 350, 1076]. LineageChain [931]. Linear [668, 231, 818, 1158, 704, 300]. Lines [70, 69]. link [240, 282]. linkage [1035]. linked [619]. LIST [1207]. lists [519, 489]. literature [422]. Llunatic [900]. Load [116, 547, 988]. Local [1086, 513, 671, 945, 896]. locality [725, 779, 928]. locality-aware [928]. locality-sensitive [779]. Localized [414]. locally [1232, 671]. Locating [200]. Location [852, 957, 1035, 806, 654, 243, 673, 707, 289]. Location- [957]. location-aware [1035, 707]. location-based [243, 289]. locations [658]. lock [701]. Locking [93, 945]. Locks [51]. Log [147, 525, 1130, 379]. Log-based [525, 1130]. Log-Structured [147]. logging [561, 802]. Logical [119, 832]. logs [890, 528]. long [1232, 882, 666]. longest [743]. longest-lasting [743]. looking [815]. lossless [1037]. lossy [1158]. lots [1018]. Low [1025, 532, 235]. low-cost [532]. Low-latency [1025]. lowest [766]. lowest-cost [766]. Lp [548]. Lp-norm [548]. LRU [68]. LSH [833, 940, 1033]. LSM [1140, 880, 946, 1181, 1170]. LSM-based [880, 1181, 1170].

M [133]. M-Trees [133]. machine

[962, 249, 818, 1021, 960, 903, 1128, 1113, 989, 966, 759, 1145, 1131]. machines [336, 760]. Main [1129, 753, 1029, 805, 1040, 1138, 701, 663]. main-memory [753, 1029, 805, 1138]. Maintaining [40, 347, 386, 54, 39]. Maintenance [117, 28, 636, 1213, 841, 1064, 474, 1153, 236, 244]. Make [895, 335]. Making [1037, 871, 1139, 773, 825]. malicious [534]. Management [44, 5, 1, 6, 88, 60, 30, 24, 8, 228, 428, 838, 417, 962, 507, 984, 1052, 163, 402, 286, 943, 302, 215, 283, 751, 293, 567, 195, 537, 810, 1009, 332, 443, 665, 727, 874, 1201, 547, 475, 466, 1159, 183, 855, 823, 192, 1217]. manager [252, 701]. Managing [962, 238, 217, 173, 718, 457, 363]. Manipulation [71]. manual [1139]. Many [816]. Many-query [816]. MapMerge [559]. Mapped [67]. Mapping [726, 354, 481, 188, 258]. Mapping-equivalence [726]. mappings [559, 586, 522, 1108]. MapReduce [643, 768, 848, 727, 576, 577]. Mariposa [76]. market [205, 831]. marketing [733]. markets [1151]. Markov [123]. Markov-Chain [123]. massive [847, 540, 961, 782, 1050, 941, 682, 861, 774]. massive-scale [941]. massively [945, 655]. Master [153, 560]. match [239, 554]. Matching [1239, 925, 708, 800, 523, 531, 853, 567, 1000, 1044, 400, 1049, 693, 311, 319, 314, 322, 1087, 926, 469, 1137, 1093, 196, 620, 647, 741, 584, 222, 1080, 687, 571, 605]. materialized [297, 426]. matrices [1176]. matrix [704]. max [1036, 276]. maximal [841, 310, 318, 1002, 362]. maximization [992, 923, 677, 1231]. Maximizing [548]. Maximum [1034, 668, 905, 854, 312, 320, 1175, 1234]. MDDE [1015]. meaning [479]. meaningful [355]. measure [854, 290]. measures [429, 1206, 437, 792, 868, 397, 793, 1115]. Measuring [580, 1230, 1112]. mechanism [1036, 704, 773, 825]. Mechanisms [109, 896]. Media [129, 137, 644, 834]. mediation [270]. mediator [272]. Mediators [273]. medoids [383]. meet [577]. meets [534, 1089]. memoriam [345]. Memory [959, 1240, 126, 963, 209, 208, 992, 1111, 401, 869, 753, 1081, 1029, 805, 1040, 1138, 1129, 169, 1168, 300, 974, 897, 845, 701, 945, 663, 958, 863, 1170, 1154, 968, 950, 781, 1033, 907, 916]. Memory-aware [963]. memory-efficient [992, 1168]. memory-resident [208]. memos [442]. MEMS [330]. merge [946]. message [346]. meta [1072]. meta-level [1072]. Metadata [141, 406, 402, 404]. metaphysical [351]. Method [115, 147, 569, 426, 849, 352, 328, 435, 289]. methodology [383, 249]. Methods [471, 1219, 208, 232, 446, 1133, 175, 513, 284, 335]. metric [1008, 775, 689, 785, 199, 397, 977]. Micro [958]. Micro-architectural [958]. microaggregation [306, 661]. microaggregation-based [661]. Microblogs [874]. microdata [487]. middleware [351, 525]. MIL [140]. Min [276]. Min/Max [276]. Mine [1228]. MiniCon [189]. Minimal [26, 786, 801, 307, 1203]. minimization [216, 873]. Minimum [1156, 297, 358, 780]. Mining [1228, 401, 596, 732, 150, 583, 524, 882, 423, 377, 283, 702, 313, 321, 540, 1180, 679, 1002, 1212, 303, 379, 255, 672, 362, 987, 384, 394]. MinJoin [1107]. mirrors [226]. Mis [953]. Mis-categorized [953]. MisMatch [692]. mitigate [962]. mitigating [404]. Mixed [1203, 805]. ML [1186]. MM [1129]. MM-DIRECT [1129]. Mobile [64, 65, 542, 537, 281, 393]. mobility [540, 1192]. modal [1048, 713]. Model [403, 50, 960, 1215, 95, 53, 30, 3, 228, 173, 965, 391, 206, 616, 955, 246, 341, 175, 1201, 106, 222, 999]. Model-independent [403]. Modeling [81, 45, 118, 608, 296, 882, 269, 674, 243, 395, 579]. Modelling [339]. models [456, 758, 1212, 1137, 251, 339, 454, 1195, 591, 1151, 1217]. Modern [736, 313, 321, 816, 1167]. monitoring [995, 414, 658]. Monotonic [194]. Morphtree [1138]. Morton [893]. most [733, 651, 895]. motif [1156]. motif-cut [1156]. Motifs [1218, 1120]. motion [496]. motion-aware [496]. movement [844]. Mover [573]. moving [299, 562, 468, 296, 499, 995, 1061, 484, 663, 477, 670, 899, 356, 574]. moving-object [663]. MSQL [785]. Multi [97, 810, 20, 142, 1062, 131, 755, 24, 441, 8, 1057, 391, 91, 360, 1079, 168, 1210, 614, 998, 341, 924, 1221, 1048, 1191, 172, 1137, 546, 844, 713, 1173, 308, 634, 1063, 1217, 851]. multi-ary [1173]. multi-attribute [91, 844]. multi-class [1079]. Multi-Clip [131]. Multi-constraint [1062]. Multi-dimensional [441, 168, 614, 1191]. multi-facet [1137]. Multi-Join [97]. multi-layer [851]. Multi-Level [24]. multi-modal [1048, 713]. Multi-objective [755]. multi-probe [1063]. Multi-Query [8]. multi-ranker [341]. multi-relational [308]. Multi-Resolution [142, 391]. Multi-schema-version [810]. multi-step [1221]. multi-tenant [1217]. multi-tier [546, 634]. multi-truth [1057]. multi-tuple [1210]. multi-valued [924]. multi-visual [172]. Multi-Way [20, 360, 998]. multicore [745, 1029, 561, 760]. Multidatabase [5, 1, 14, 99, 21, 251]. Multidatabases [12]. Multidimensional [44, 243, 154, 365, 1112, 206, 274, 565, 443, 633, 191, 288]. Multimedia [134, 162, 128, 132, 163, 174, 646]. Multimodal [506, 1183]. Multiple [70, 154, 357, 160, 231, 764, 69, 238, 624, 494]. multiplicities [444]. multisocket [561]. multisource [244]. multistore [973]. multitasking [1015]. multithreaded [945]. multivariate [306, 1115]. Multiversion [89, 217, 738]. MyBenchmark [667].

Nail [33]. Naïve [381]. native [215, 214]. Natural [884, 858, 173, 963, 1143, 1108]. navigation [160]. Navigational [109, 1190]. near [1096]. near-optimal [1096]. Nearest [299, 164, 553, 581, 463, 526, 770, 499, 1174, 445, 940, 493, 673, 1188, 543, 548, 757, 1063, 289, 774]. nearest-neighbor [463, 493, 673, 289]. negative [465, 942]. negotiation [182]. Neighbor [164, 299, 463, 526, 499, 1174, 445, 940, 493, 673, 1188, 548, 1063, 289]. neighborhood [715, 776]. neighborhood-centric [715]. neighbors [553, 581, 770, 543, 757]. Nested [113, 9, 1085, 1229, 411, 171]. nested-transaction [411]. Nesting [86]. Netherite [1199]. network [447, 656, 1053, 563, 611, 219, 495, 1022, 1174, 1121, 1135, 1067, 393, 352, 538, 504, 1088, 658, 997]. network-based [1135]. network-bound [219]. networked [82]. Networks [1225, 852, 553, 1035, 333, 359, 440, 1098, 296, 505, 1172, 685, 645, 677, 782, 857, 924, 1061, 986, 1106, 870, 533, 550, 1089, 1196, 1238, 602, 647, 804, 265, 325, 759, 414, 791, 767, 1118, 1007, 334]. neural [1053, 1121, 1135, 1067, 1196, 1165]. nexus [1019]. NN [391, 1069, 1061, 1033]. Node [1218, 951, 91, 896, 1164]. node-local [896]. nodes [1007]. noise [1130]. noise-aware [1130]. noisy [1036]. Non [623, 792, 1086, 1040, 527, 415]. Non-binary [792]. non-dedicated [415]. non-deterministic [527]. non-numeric [1086]. Non-uniformity [623]. non-volatile [1040]. nonuniform [903]. norm [779, 548]. normal [1197]. Normalization [522, 409, 348]. normalizing [1109]. NOT-twig [589]. Nothing [17, 94, 18]. Notification [114, 533]. novel [1205]. now-relative [197]. NScale [715]. NUMA [1040]. NUMA-aware [1040]. numeric [1086]. NVRAM [896]. NVRAM-centric [896].

O [948, 961, 940, 762, 682]. O-efficient [774]. Object [59, 57, 41, 108, 107, 29, 62, 86, 95, 125, 30, 122, 67, 275, 496, 63, 58, 726, 562, 100, 468, 744, 82, 232, 1182, 499, 138, 1178, 106, 663, 192, 585, 249]. object- [192]. object-based [82]. object-creating [726]. Object-Oriented [41, 108, 86, 125, 30, 275, 138]. Object-Process [249]. Object-Relational [122]. Objectbase [60]. ObjectGlobe [181]. objective [755]. Objects [112, 90, 38, 24, 299, 166, 937, 296, 281, 1061, 694, 670, 356, 574, 582]. Observations [1236]. obstructed [806]. offline [943]. offs [208, 1204]. oid [726]. oid-equivalence [726]. OLAP [315, 323, 973, 371, 805, 1168, 1161]. OLTP [1216, 805, 1040, 737, 958, 597]. OLTP/OLAP [805]. Omni [335]. Omni-family [335]. On-line [523, 385]. on-road [801]. One [168, 862, 463, 1093]. One-dimensional [168, 463]. One-pass [862]. one-to-one [1093]. ones [527]. Online [474, 292, 170, 413, 992, 184, 943, 739, 677, 1072, 646, 668, 777, 1007]. Ontological [1039, 242, 617, 702]. Ontologies [1239, 239, 238]. ontology [246]. ontology-based [246]. OODB [115, 110, 111]. OODBMS [81, 73]. Open [1162, 783]. operational [738]. operations [267]. Operator [262, 1228, 432, 339]. operators [945]. opinion [1231, 1151]. Opportunities [1029]. OPT [138]. Optimal [606, 89, 789, 131, 1007, 1096, 806, 440, 410, 1059, 1042, 376, 222, 672, 658]. optimality [771]. Optimisation [80]. optimism [1029]. Optimization [87, 367, 20, 98, 102, 783, 166, 301, 611, 522, 158, 1186, 724, 138, 409, 976, 1055, 1060, 815, 1205, 1235, 1121, 824, 473, 691, 433, 755, 1142, 576, 1063, 1242]. optimizations [1106]. Optimized [826, 740, 1040, 574]. optimizer [1148, 919, 1091]. Optimizers [1236, 1030, 983]. Optimizing [1110, 467, 113, 154, 169, 1190, 1170, 276, 1219, 175, 704, 287, 1203]. options [415]. Oracle [87]. oracles [1149]. Order [815, 92, 817, 819, 1146, 639, 719, 485, 1010, 1205, 978, 963, 811, 970, 793, 753]. order-preserving [92]. Ordered [51]. Ordering [131, 102, 981]. orders [556, 1146]. organizations [307]. Organizing [124, 240]. Oriented [41, 108, 86, 125, 30, 275, 72, 64, 306, 232, 138, 768, 1121, 369]. Orthogonally [58, 57]. OSAM*.KBMS [84]. outer [436, 311, 319]. outer-join [311, 319]. outlier [279, 1008, 1016]. Outliers [149]. outsourced [565, 426, 438, 486]. Outsourcing [613]. overhead [648]. overlap [906, 979, 771]. Overview [5, 387, 994]. OXPath [593].

P [1067]. P2P [481, 404]. PACE [794]. Package [817]. Page [68, 209]. PageRank [352, 863, 700]. Pair [164, 689, 1033]. Pair-Wise [164]. pairing [413]. pairwise [744, 1210, 939, 830, 843]. PANDA [759]. PANE [1088]. panorama [1214]. Papers [1084, 144, 516, 915, 557, 697, 883, 470, 812, 253, 427, 309, 317, 709, 590, 750, 749, 635, 1024, 894]. paradigm [180, 794]. Parallel [20, 1002, 805, 66, 116, 804, 244, 19, 1008, 432, 752, 94, 666, 655, 829, 987, 577]. parallel-friendly [1008]. Parallelism [17, 24, 22, 650, 981]. Parallelization [1154, 415]. Parallelizing [73, 863]. Parametric [847, 98, 755]. Pareto [1059]. parking [801, 1018]. PARROT [1065]. Parsimonious [564]. Parsing [864]. Part [4, 7]. Partial [112, 26, 819, 759]. partial-order-based [819]. Partially [426, 420]. participants [534]. partition [383, 790, 289]. partition-based [790]. partitioned [428, 332, 1065]. Partitioning [650, 116, 769, 167, 840, 1041, 232, 725, 835, 1156, 597]. partnership [220]. pass [862]. passing [1030]. Path [714, 764, 85, 329, 495, 115, 964, 1176, 809, 367, 795, 1062, 1089, 1152, 794, 767, 1118, 794]. Path-based [764]. PAth-CEntric [794]. Path-Method [115]. paths [766, 906, 997]. Pattern [52, 29, 603, 216, 374, 606, 853, 721, 313, 321, 1180, 469, 420, 498, 1065, 554]. Pattern-Based [29, 1065]. patterns [1095, 690, 1232, 401, 764, 679, 389, 233, 666, 987, 261]. pay [671]. Payment [1018]. PBS [640]. PC [144, 143]. PCA [1136]. PCCP [329]. peer [352, 325, 425]. peer-to-peer [352, 325, 425]. perception [1214]. Performance [530, 139, 23, 24, 67, 783, 208, 520, 859, 73, 948, 935, 235, 256, 1204, 868, 966, 657, 760]. Performant [1169]. Periodic [129, 437]. Permutation [137, 788]. Persistent [57, 380, 63, 58, 1111, 916]. personal [1052, 405]. personalization [184]. Personalized [647, 863, 1080, 684, 700]. perspective [210, 1070, 308]. perturbation [394]. pervasive [277]. phantoms [494]. Physical [79, 472, 609]. physics [1229]. physiological [597]. Picket [1014]. PicoDBMS [186]. PicShark [404]. Picture [662]. Piecewise [668]. PIM [1240]. PIM-tree [1240]. pipeline [1038, 1056]. pipelines [1021, 895, 1141]. Pivot [977]. place [901]. placement [665, 1060, 94]. plan [764]. planning [1085]. plans [814, 287]. platform [669, 578, 919, 1020]. platforms [183, 646]. PM [1033]. PM-LSH [1033]. PMG [115]. PMR [204]. POESIA [242]. Pointer [62]. policies [587, 293, 599, 946, 504]. polyglot [1082]. polymorphic [1138]. portable [745]. Possible [734]. postorder [606]. Power [71, 392, 610, 375]. Power-law [392]. powered [1166, 1127]. powerful [1177, 760]. Practical [1085, 284, 195, 828, 845]. PrDB [457]. pre [1048, 1179]. pre-trained [1179]. pre-training [1048]. Précis [343]. precision [972]. Precomputed [115]. Predicate [501, 75, 378, 368]. Predicate-based [501, 75]. predicates [1186, 396, 1030, 947]. Predictable [520]. Predicting [357]. Prediction [52, 852, 982, 523, 495, 1192, 1202, 620, 1018, 681]. prediction-enhanced [1192]. Predictions [123, 1144]. predictive [495]. Preference [517, 911, 939, 927]. preference-based [911]. preferences [1103]. preferred [766]. prefetch [166]. Prefetching [123, 329]. PrefixFPM [987]. preorder [606]. Prescriptive [850]. presence [401, 448, 233, 482, 420, 576]. Presentations [134]. preservation [493]. Preserving [258, 92, 374, 447, 305, 1052, 654, 302, 1015, 1067, 303, 490, 381, 1012, 1005]. pretrained [1217]. prevent [487]. prevention [634]. price [1007]. pricing [1202]. Primal [398]. primitives [140]. principle [358]. principled [735]. Principles [469]. prioritized [517]. Priority [74, 1045, 1100]. priority-aware [1100]. Privacy [447, 1000, 533, 504, 381, 308, 1005, 305, 1052, 370, 430, 656, 654, 1086, 302, 613, 1015, 704, 1067, 303, 490, 661, 1012]. Privacy-preserving [447, 381, 1005, 305, 1052, 654, 1015, 303, 490, 1012]. private [796, 1194, 625, 486, 394]. PrivPfC [796]. Probabilistic [514, 820, 456, 337, 619, 1069, 626, 453, 460, 462, 459, 521, 583, 758, 445, 886, 416, 458, 339, 457, 573, 483]. Probability [52]. probe [1063]. Problem [102, 692, 210, 735, 681]. Problems [136, 783, 410]. process [594, 528, 249]. processes [696, 1004]. Processing [87, 145, 101, 118, 107, 1220, 1240, 720, 663, 17, 1236, 1225, 19, 455, 705, 606, 984, 1224, 181, 568, 190, 529, 581, 463, 532, 638, 859, 167, 100, 391, 440, 608, 247, 643, 1116, 526, 650, 260, 1126, 1229, 360, 566, 710, 628, 1049, 366, 639, 368, 485, 1209, 445, 202, 886, 489, 395, 888, 607, 1155, 342, 549, 823, 18, 579, 1195, 1077, 1104, 1142, 450, 425, 1091, 670, 461, 899, 1076, 856, 802, 767, 1118, 1114, 622, 582, 629, 799]. Processing-in-Memory [1240]. processor [259, 432]. processors [821, 313, 321]. products [733, 651]. profile [1007]. Profiling [695, 1011]. programmable [948]. Programming [59, 61, 80, 84]. Progressive [171, 549, 327, 1069, 954]. project [387]. projections [231]. promises [398]. promising [733]. prompt4vis [1212]. prompting [1212]. Prone [1, 452]. Propagation [153, 285, 751]. Propagations [113]. properties [586, 454]. Property [1228, 1197]. proportional [744, 1168]. ProS [1069]. protection [370, 613, 1169, 493]. protein [285, 212].

Protocol [40, 39, 562]. Prototypes [32, 37]. Provenance [696, 798, 787, 813, 828, 931]. Provenance-based [696]. providers [533]. Providing [377, 509]. provisioning [642]. proximity [1008, 533, 793]. proxy [361, 384]. PSoup [229]. public [699]. publication [507, 656, 796, 625]. publish [909, 763, 707]. publish/subscribe [909, 763, 707]. publishing [654, 187]. PUG [828]. Purpose [370, 335, 987]. pushdown [1161]. Python [1195].

QFilter [527]. QoS [611]. QoS-aware [611]. QQL [340]. quadtree [204, 325]. quadtree-based [204]. qualified [988]. Qualitative [462, 46]. quality [518, 1069, 265, 1070, 605]. Quantifiable [150]. Quantifying [640, 1103]. Quantiles [729]. Quantitative [62, 1111]. quantization [1196]. quantization-aware [1196]. quasi [1002, 176]. quasi-cliques [1002]. quasi-sparse [176]. quaternary [364]. QUBLE [645]. Queries [97, 113, 70, 154, 131, 1058, 690, 496, 1176, 299, 726, 1223, 1052, 817, 367, 703, 229, 553, 581, 463, 775, 638, 157, 297, 444, 359, 552, 612, 798, 596, 1171, 853, 996, 1098, 1086, 611, 689, 390, 274, 542, 193, 725, 955, 995, 565, 245, 446, 515, 645, 889, 254, 488, 418, 967, 651, 69, 175, 233, 500, 704, 939, 445, 514, 747, 527, 1102, 517, 385, 742, 1190, 484, 1089, 489, 1039, 720, 980, 718, 189, 458, 865, 607, 343, 673, 630, 589, 491, 171, 420, 549, 1193, 863, 1152, 498, 827, 657, 899, 757, 414, 572, 1207, 441, 483, 574, 1203, 622]. queries [289, 629, 780, 554]. Query [87, 586, 301, 50, 107, 448, 1220, 12, 360, 20, 779, 98, 459, 815, 202, 80, 340, 18, 17, 659, 461, 8, 1236, 19, 216, 294, 705, 481, 890, 181, 821, 1213, 161, 520, 568, 190, 529, 188, 1148, 532, 167, 337, 965, 391, 440, 884, 247, 643, 814, 1025, 526, 158, 1186, 724, 446, 219, 138, 1030, 1049, 366, 976, 1055, 895, 368, 485, 355, 758, 1061, 1209, 824, 667, 816, 660, 886, 415, 917, 902, 439, 251, 395, 1082, 433, 376, 663, 983, 1143, 653, 1194, 342, 541, 755, 425, 873, 261, 765, 479, 767, 1063, 1073, 1118, 582]. query [1242, 629, 799, 652, 822, 999, 693]. Query-aware [779]. Query-By-Humming [693]. query-reuse [1082]. query-time [758]. Querying [278, 936, 706, 248, 1200, 993, 268, 140, 957, 382, 540, 296, 495, 383, 353, 272, 788, 1001, 861]. question [1143]. questions [747]. QuickStore [67]. Quo [1229].

R [345, 1181, 442, 653]. R-tree [1181]. R-trees [442, 653]. RailwayDB [716]. random [955, 474, 386, 963, 793]. Randomized [102]. range [775, 979, 1171, 390, 274, 565, 495, 693, 1188, 922]. rank [1148]. ranked [245]. ranker [341]. Ranking [515, 543, 521, 514, 1006, 433, 491, 705]. rankings [994]. rapid [892]. Raster [1182]. Rate [137]. RATEWeb [449]. Ratio [150]. RCBench [1117]. Rdb [87]. RDBMS [509]. RDF [993, 833, 1224, 912, 616, 913, 675, 475, 720, 1156, 990, 855, 901]. RDF-3X [475]. RDFFrames [989]. RDMA [1117]. RDMA-enabled [1117]. re [505, 227]. re-identification [505]. RE-tree [227]. reach [980]. Reachability [788, 638, 980, 602, 1152, 1001, 572, 861, 799]. Read [740, 1139]. Read/write [740]. Read/write-optimized [740]. Real [51, 563, 146, 11, 93, 74, 670, 834, 636, 506, 274, 539, 928, 657]. Real-Time [51, 146, 11, 74, 563, 670, 834, 636, 506, 539]. real-world [657]. Realization [62]. Realm [55]. Realm-Based [55]. Reasoning [690, 619]. recoding [513]. recognition [878]. recommendation [503, 1048, 508, 1124, 952, 777, 834]. recommendations [627, 542, 1046, 994]. recommender [1012]. Recommending [609]. reconciliation [269]. Reconciling [1185]. reconstruction [891]. record [531, 222]. Recoverable [114]. Recovery [104, 103, 91, 282, 1129, 525, 910, 1105, 802]. rectifying [925]. recursion [161, 610]. recursive [467, 448]. redesign [701]. REdistribution [512]. Reducing [436, 562, 1177, 731]. reduction [298, 328, 762, 799]. redundancy [348]. Reference [125, 397]. Reference-based [397]. Referential [85]. Refined [109]. refinement [348]. Refiner [1166, 1127]. region [829]. regional [584]. regions [944]. registers [1027]. regression [316, 324]. regret [873]. Regular [70, 1176, 227, 706]. Reinforcement [1215, 1172, 1122, 986, 971]. related [1036, 303]. relation [1103, 947]. Relational [83, 107, 12, 113, 38, 122, 695, 933, 230, 209, 268, 370, 627, 519, 753, 751, 885, 254, 712, 389, 1132, 967, 409, 816, 660, 478, 832, 339, 187, 589, 192, 655, 308, 330]. relationalize [1221]. relations [166, 467, 808, 837]. relationship [392]. Relationships [53, 25, 1097, 221]. relative [203, 197]. relaxation [327, 660]. Relaxing [81]. Relevance [115, 744]. relevant [694, 901]. Reliability [1144]. Reliable [126, 1166, 1127]. Remote [38]. RemusDB [592]. REncoder [1171]. reordering [170]. repair [587]. repairing [925, 776]. repairs [632, 1042]. replacement [293]. replica [665]. Replicated [14, 153]. replication [805, 393, 896]. report [535]. report-generation [535]. repositories [451, 200]. repository [211]. Representation [46, 455, 1110, 305, 1209, 1048, 1023, 668]. representations [846, 249, 990]. representative [1147, 662]. Representing [454]. Reputation [449]. research [783, 369]. resident [208, 489]. Resilience [1019]. resistant [1240]. Resisting [505]. Resolution [141, 142, 421, 819, 391, 1210, 1162, 1093, 465, 624, 631, 1179]. resolve [692]. resolving [406]. Resource [129, 7, 752, 1043]. Resource-aware [1043]. response [158]. responsible [1009, 1019]. Restricted [1220]. result [627]. result-driven [627]. results [884, 1068, 492, 1042, 595]. retention [599]. rethinking [1161]. Retrieval [112, 45, 246, 120, 17, 133, 291, 338, 1178, 416, 198, 339, 424, 18, 713, 585, 901]. Retrospective [264]. Return [8]. reusability [1215]. Reuse [49, 1082]. Reverse [1058, 757, 299, 553, 581, 770, 651, 445, 747, 365, 659, 548]. Revisited [77, 958, 1147, 1205, 649, 1168]. Revisiting [482]. rewriting [527, 461]. RFID [351, 450]. RHEEMix [919]. rich [1213, 457]. riding [334]. riding-aware [334]. Rights [493, 305]. Risk [795]. Risk-aware [795]. risky [1020]. RNE [997]. road [495, 801, 1061, 1089, 658, 767, 1118, 997]. Robust [1066, 809, 1133, 1040, 352]. robustness [1140]. ROLAP [375]. ROSE [55]. rotated [437]. rotations [1136]. Round [1225]. Route [1235, 801, 1048, 376]. routes [679]. routing [911, 898, 684]. RPQs [1110]. Rtree [451]. Rule [286, 247, 925, 523, 702, 1099, 1222, 290, 1228]. Rule-based [286, 247]. Rules [81, 1228, 150, 180, 560, 206, 462, 465, 631]. RUM [442]. RUM-tree [442]. rumours [1050]. running [815]. runtime [1082].

S [708, 159]. S.I. [1028]. SABRE [512]. safe [766, 81, 476]. sample [347]. samples [474, 386]. Sampling [632, 418, 623, 1189, 918, 928, 771, 942]. Sampling-based [418]. Sancus [1196, 1196]. satellite [372]. Scalability [561, 832]. Scalable [891, 124, 463, 768, 848, 1121, 926, 511, 597, 497, 708, 969, 1054, 817, 748, 184, 593, 1227, 710, 756, 366, 860, 1051, 1201, 385, 475, 849, 546, 189, 1011, 666, 963, 829, 1173, 1066, 1088, 572, 843, 1164]. Scale [1218, 933, 211, 852, 558, 951, 404, 619, 643, 451, 818, 1098, 1004, 270, 1172, 732, 575, 1064, 1034, 1099, 1222, 466, 715, 630, 985, 941, 1017, 1073]. ScaLeKB [748]. Scaling [382, 674, 760, 802, 805, 280, 186]. Scan [809]. Scans [68, 655, 576]. scarcity [404]. SCCs [682]. scenarios [314, 322]. Scheduled [700]. schedules [576]. Scheduling [27, 11, 535, 262, 129, 801]. Schema [81, 481, 270, 620, 559, 586, 403, 847, 594, 567, 522, 810, 400, 1003, 314, 322, 1060, 1135, 196, 1108, 258, 348, 355]. Schema-Free [355]. Schemas [3]. Schematic [90]. Scheme [75, 779, 426]. schemes [217]. science [1009, 1141]. sciences [283]. SCOPE [577]. scoring [290]. SD [451]. SD-Rtree [451]. SDS [35]. Search [164, 692, 618, 912, 503, 529, 876, 1053, 937, 501, 1069, 784, 914, 879, 499, 257, 685, 645, 961, 492, 536, 509, 723, 1051, 940, 785, 1034, 349, 205, 1175, 352, 511, 595, 422, 434, 1188, 365, 398, 646, 335, 904, 1173, 1165, 759, 573, 698, 486, 761, 791, 843, 683, 717, 797, 1063, 1033, 774, 851]. search-based [698]. searchable [1101]. searches [176, 790]. Searching [199, 26, 341, 1208, 742]. Seaweed [353]. Second [793, 864, 963]. Second-order [793, 963]. Secret [778, 608]. Section [64, 664, 72, 575]. Secure [146, 565, 304, 1160, 527, 534, 931, 1101, 1193, 827]. security [778]. Seeking [105]. segmentation [400]. segments [1158]. Selected [1084]. selecting [789]. Selection [124, 809, 210, 795, 246, 665, 752, 1205, 977, 700]. selectivities [396]. Selectivity [274, 28, 168, 312, 320, 591]. Self [443, 833, 392, 1238, 439]. self-adaptable [439]. self-clustering [833]. self-similarity [392]. self-supervised [1238]. Self-tuning [443]. Semantic [141, 157, 45, 249, 90, 944, 25, 192, 294, 839, 405, 269, 270, 400, 1003, 1135, 416, 1047, 901, 428, 239, 238, 241]. Semantically [26]. Semantically-Equivalent [26]. semantics [586, 608, 240, 583, 547, 706, 966]. semi [308, 1164]. semi-supervised [308, 1164]. semistructured [161, 580, 630, 342]. Sensitive [66, 444, 779, 512]. sensitivity [1086]. sensor [333, 359, 611, 265, 414]. sentences [173]. SEPIA [396]. sequence [423, 212, 743, 1238, 666, 584]. Sequenced [541, 376]. sequences [1046]. sequential [469, 956]. serializable [773, 825]. series [969, 1054, 1206, 1069, 680, 382, 795, 437, 1158, 860, 1160, 926, 292, 974, 956, 922, 1125, 1173, 1165, 1124, 1065, 746, 822, 288]. Server [49, 118, 114, 75, 103, 182, 525]. serverless [1199, 1113]. Servers [134, 256]. Service [142, 369, 533, 528, 1202, 718]. services [271, 180, 242, 537, 243, 411, 449, 178, 241, 183, 646, 287, 486, 363, 289]. Set [13, 26, 800, 263, 235, 1177, 1210, 1187, 811, 513, 837, 435, 807, 968, 950]. set-based [811]. set-expression [263]. set-valued [235, 513]. sets [429, 1046, 670]. settings [766]. shapelet [1160]. shapes [437, 198]. SHARC [518]. sharding [1092]. share [1027]. Shared [51, 17, 816, 94, 504, 18, 597, 863, 576, 1018]. shared-everything [597]. Shared-Nothing [17, 94]. Sharing [4, 778, 568, 82, 270]. shifted [479]. Shooting [545]. short [708]. shortest [367, 906, 613, 1062, 1089, 767, 1118, 997]. shot [1165]. shuffle [1145]. SI [887, 867, 546]. SI-Cache [546]. sideways [1030]. sight [913]. SIGMOD [72, 64]. signal [891]. Signature [124]. signed [1234]. significance [662]. SIMD [895, 917]. similar [490, 198, 1123]. similar-bicliques [1123]. Similarities [90]. Similarity [112, 607, 133, 1151, 891, 1057, 800, 951, 392, 1069, 569, 551, 580, 400, 614, 1107, 723, 785, 1187, 205, 804, 676, 646, 335, 176, 837, 1173, 573, 761, 843, 687, 683, 866, 622, 790]. Similarity-driven [1151]. Simple [964, 942, 1236, 335]. simplet [1189]. simplification [862]. SimRank [473, 830, 972, 791]. SimRank* [843].

Simultaneously [154]. single [726, 863, 972, 759, 248, 907]. single-dimensional [248]. single-function [726]. single-source [863, 972]. sites [160, 159, 361, 504]. situ [1043]. situation [252]. size [297, 744, 347, 623]. size- [744]. SKCompress [903]. sketch [1163, 1075]. Sketching [686]. Skew [20, 1240, 1219]. Skew-resistant [1240]. sky [479, 485]. SkyAlign [745]. Skyframe [425]. Skyline [1220, 865, 745, 995, 566, 485, 633, 924, 1235, 517, 425, 829, 873, 479]. skyline-based [1235]. skylines [662]. Sleepers [64, 65]. Sliding [1081, 488, 686, 970, 763]. Sliding-window [488, 686, 970]. slow [801]. small [857]. small-world [857]. smart [680]. smartcard [186]. Smooth [809]. snapshot [671, 581, 463, 721]. snapshots [1128]. Snorkel [892]. Social [503, 852, 507, 1035, 506, 505, 1000, 1192, 677, 986, 533, 504, 644, 777, 834, 780, 1007, 705]. Software [49, 314, 322, 1102]. SOLE [385]. solid [740]. solid-state [740]. Solution [136]. solutions [985, 363]. Solving [681, 826]. some [706]. Sort [568, 66, 556]. Sort-sharing-aware [568]. sortable [860]. sorted [489]. Sorting [550, 167, 731, 331]. source [1057, 620, 863, 972]. sourced [615]. sources [1112, 618, 1047, 273]. Space [1241, 46, 508, 497, 553, 806, 1171, 410, 1079, 1051, 1106, 990, 356, 248]. space-time [1171]. spaces [1008, 689, 785, 199, 397, 548, 977]. Span [1001]. Span-reachability [1001]. Spark [960, 813]. SPARQL [993, 890, 725, 720, 652]. sparse [893, 905, 1036, 903, 176, 1083]. sparsity [681]. Spatial [43, 55, 130, 120, 46, 203, 152, 875, 582, 1058, 208, 48, 912, 553, 530, 440, 937, 451, 853, 526, 1182, 42, 204, 1235, 383, 199, 647, 804, 412, 694, 1188, 855, 1193, 477, 1142, 763, 829, 585, 604, 899, 438, 757, 1207, 486, 276, 797, 1100, 452]. spatial-keyword [937, 763]. spatial-range-constrained [1188]. Spatio [800, 298, 654, 385, 197, 676, 326, 1207, 831]. spatio-bitemporal [197]. Spatio-temporal [298, 654, 385, 676, 326, 831]. Spatio-textual [800, 1207]. spatiotemporal [247, 295, 1192, 1048, 398, 541, 901, 730]. Special [516, 664, 915, 22, 37, 48, 63, 72, 57, 557, 64, 1131, 697, 736, 883, 984, 615, 470, 812, 1226, 283, 427, 43, 537, 309, 317, 453, 1009, 709, 575, 590, 749, 1167, 1084, 930, 635, 128, 1024, 32, 894, 17, 402, 338, 302, 260, 750]. specification [159, 195]. specified [420]. speech [1143, 1143]. speech-driven [1143]. Speech-to-SQL [1143]. Speed [510, 424, 1094]. Speeding [204, 1051]. spurious [492]. SQL [703, 448, 1119, 1186, 555, 1074, 734, 785, 1143, 192, 326]. SQL2 [85]. SQUID [1073]. SRX [855]. SSD [948]. SSDs [738]. STAIRS [544]. staleness [1196]. standard [751]. stars [545]. Start [967]. state [783, 740, 411, 946, 823]. state-of-the-art [946]. stateful [650]. static [1049, 860, 648, 830, 1237]. Statistical [466]. Status [1229]. Status-Quo [1229]. step [1221]. Stepwise [141]. Stochastic [1145, 898, 794]. Storage [109, 123, 103, 877, 332, 731, 880, 349, 547, 474, 416, 716, 932, 280, 330, 642, 907]. storage-centric [547]. Store [67, 781, 428]. stores [993, 955, 1184]. Storing [268]. story [636, 1026]. Strategies [64, 62, 153, 65]. Strategy [137, 423, 1156]. Stratosphere [669]. Stream [771, 228, 262, 1224, 859, 608, 1116, 432, 650, 260, 360, 255, 1077, 1104, 668, 261, 856, 1114]. Streaming [1094, 494, 636, 708, 229, 259, 563, 410, 860, 547, 1050, 598, 408]. streams [464, 909, 471, 263, 488, 278, 366, 928, 1163, 729, 1187, 688, 385, 478, 686, 511, 865, 706, 412, 603, 579, 498, 644]. string [92, 417, 569, 396, 1107, 918, 364, 741, 761]. strings [497]. strong [918, 655]. Structural [136, 409, 492, 161, 951, 505, 580, 685, 842]. Structure [47, 227, 464, 913, 1066, 790]. Structured [119, 109, 147, 615, 301, 343]. structured-document [301]. structures [235, 1169, 466, 896]. study [852, 890, 960, 235, 1230, 977]. Sub [1122, 300]. sub-linear [300]. Sub-trajectory [1122]. subgraph [636, 1000, 645, 1049, 768, 1102, 904, 1153, 687, 1075]. subgraphs [583, 988]. subjective [936]. subscribe [909, 763, 707]. subsequence [969, 1054, 693, 926, 571]. Subset [26, 418]. subset-based [418]. subspace [633, 549]. substring [168]. subtrajectory [1073]. Succinct [305]. suffix [982, 284]. summaries [744]. summarization [913, 292]. summarizations [860]. Summarizing [839]. Super [614]. Super-EGO [614]. supervised [1238, 308, 1164]. supervision [892]. Support [85, 10, 142, 2, 53, 962, 406, 392, 167, 184, 641, 195, 1227, 813, 336, 221, 192, 525, 707]. Supporting [14, 254, 437, 174, 375, 1101, 491, 405, 91, 753, 351, 1150, 442]. Suppressing [487]. surface [391]. surfing [357]. surprising [888]. surrounder [542]. Survey [1159, 1077, 695, 858, 993, 778, 962, 1224, 839, 876, 867, 957, 643, 914, 869, 879, 877, 1116, 850, 878, 193, 787, 566, 675, 1074, 1003, 976, 1055, 1191, 880, 874, 1184, 1183, 393, 1016, 871, 196, 888, 1155, 868, 823, 875, 598, 873, 1023, 1114, 977, 1104]. SVD [1136]. SW [428]. SW-Store [428]. sweeping [947]. sweeping-based [947]. switching [423]. Swizzling [62]. SWOOP [1187]. Swoosh [421]. SWORD [665]. Symbolic [132]. Sync [285]. Synchronization [103]. synchronous [862]. synopses [347, 419]. synopsis [410, 390, 854]. synthesis [1106]. synthesizing [1221]. synthetic [314, 322]. System [64, 1, 406, 33, 114, 60, 103, 34, 132, 76, 36, 72, 507, 229, 82, 943, 201, 215, 1126, 506, 710, 336, 1166, 1127, 931, 666, 221, 829, 1012, 941, 781, 971, 907]. System-oriented [64, 72]. systematic [1113, 845]. Systems [57, 40, 6, 14, 146, 148, 43, 11, 4, 7, 139, 142, 99, 88, 32, 116, 93, 17, 21, 380, 22, 37, 48, 63, 58, 262, 962, 39, 160, 1224, 1052, 370, 753, 1116, 751, 195, 42, 919, 665, 752, 178, 94, 592, 1159, 439, 872, 31, 701, 272, 823, 18, 192, 1077, 1104, 407, 425, 1145, 760].

Table [1210, 555]. tables [519, 1221].

Tabular [1106, 1212, 1014]. TADOC [938]. tagged [1188]. tagging [637]. tailoring [1147]. Taking [662]. targeted [733]. Task [691, 1151, 1100]. task-driven [1151]. tasks [786, 826]. Taxonomies [124, 803]. Taxonomy [660, 78, 273]. Taxonomy-based [660, 273]. Technique [40, 39, 680, 303]. Techniques [118, 10, 473, 880, 202, 1162, 393, 186, 251, 598, 280, 331]. technologies [850, 224, 369]. Technology [101, 35, 100]. Tee [1184]. Tee-based [1184]. template [206]. Temporal [1095, 83, 27, 130, 928, 162, 395, 1218, 1185, 846, 298, 654, 173, 965, 267, 564, 1227, 500, 1071, 385, 991, 949, 676, 399, 326, 407, 1152, 1001, 236, 1200, 861, 831]. Tempura [1091]. tenant [1217]. tensor [849]. tensors [732]. term [909]. termination [940]. Test [13]. testing [703, 480]. tests [316, 324]. Text [1232, 124, 938, 708, 231, 1074, 409, 536, 524, 544, 18]. text-to-SQL [1074]. textual [800, 909, 957, 1207]. Their [136, 4, 7, 54, 714, 1120, 607]. Thematic [838]. Thémis [61]. theoretic [580, 175]. theoretical [927]. Theory [21, 1223, 789, 870, 534, 1231]. THESUS [240]. theta [889]. thinker [988]. Third [1197]. Three [93, 1211, 548]. three-dimensional [548]. Threshold [1223, 483]. Threshold-based [483]. thresholds [379]. Tidy [967]. tier [546, 634]. Tiered [1219, 932]. Tiered-Indexing [1219]. TIGUKAT [60].

TIMBER [214]. Time [51, 1241, 146, 11, 137, 10, 70, 1071, 74, 922, 1124, 636, 992, 1206, 680, 1171, 382, 563, 506, 158, 410, 795, 437, 1158, 69, 539, 758, 801, 1061, 1160, 292, 898, 300, 956, 970, 1125, 603, 407, 670, 1165, 791, 834, 288]. time-dependent [1061]. time-series [1206, 680, 288]. Time-topology [1071]. time-varying [795, 898, 791]. timed [1095]. Timestamping [151, 1185]. timestamps [949]. TM [708]. tolerance [859, 902]. tolerances [693]. Tool [79, 1011]. Toolkit [6]. tools [989, 1139]. Top [909, 612, 446, 685, 500, 763, 901, 1058, 429, 401, 798, 937, 995, 254, 1068, 488, 651, 939, 545, 789, 747, 1163, 1034, 1187, 489, 694, 1188, 342, 873, 757, 441, 717, 483, 774]. Top- [909, 612, 446, 500, 763, 901, 685, 1058, 429, 798, 937, 995, 254, 1068, 488, 651, 939, 545, 789, 747, 1163, 1034, 1187, 489, 694, 1188, 342, 873, 757, 441, 717, 483, 774]. top-k [401]. Topic [124]. topics [831]. topology [1071, 759, 843]. topology-based [759]. TopX [342]. TR [1018]. TR-GCN [1018]. traceability [1076]. traces [552]. tracing [223]. Tracking [263, 798, 539]. trade [208, 1204]. trade-offs [208, 1204]. tradeoff [256]. tradeoffs [1216]. trained [1179]. training [1048, 1067, 1201, 1196, 892]. trajectories [499, 679, 543, 730]. trajectory [943, 540, 679, 539, 801, 1122, 862, 493, 538, 647, 804, 676, 868]. Transaction [5, 1, 101, 10, 2, 24, 1225, 100, 195, 628, 411, 407, 525, 802, 1117]. transaction-time [407]. Transactional [641, 125, 1214, 206, 739, 1155, 1114]. Transactions [27, 9, 671, 935, 1029]. transformation [1160]. transformations [223, 281, 1221, 607]. Transformer [1239]. transformers [1087]. transforms [303]. translation [403, 481, 188, 448]. transparent [592]. transportation [1048]. travel [795]. traversals [73]. treatment [1022, 1125]. treatment-adaptive [1022].

Tree

[216, 89, 164, 47, 383, 388, 227, 530, 740, 1181, 23, 420, 498, 464, 707, 356, 91, 1240, 442]. Tree-based [383]. tree-like [388]. tree-pattern [420, 498]. Trees [104, 133, 19, 982, 616, 1140, 282, 300, 653, 284, 442]. trends [877, 850]. triangle [1081, 928]. trichotomy [458]. Trie [569]. trie-based [569]. Trie-join [569]. Triejoin [1241]. tries [417, 497]. trillion [1073]. trillion-scale [1073]. triples [616]. Truss [1233, 961]. Trust [449]. Trustworthy [349]. truth [1057, 105]. truths [972]. Tunable [833]. Tunable-LSH [833]. tuning [472, 443, 314, 322, 1139, 1142, 971]. tuple [1185, 1210, 1129]. Tuples [967, 1006]. TurboLift [910, 1105]. Tutorial [4, 7]. TV [47]. TV-Tree [47]. twenty [810]. twenty-first [810]. twig [606, 469, 589]. Twitter [644]. Two [46, 1225, 1128, 854, 548]. two- [548]. Two-Dimensional [46, 854]. Two-Round [1225]. Type [81, 552, 110, 536, 111, 1047]. type-ahead [536]. Type-safe [81]. Types [55, 122, 1077, 1104]. typicality [446].

Ubiquitous [181]. ubiquity [888]. UDA [758]. UDF [1076]. ULISSE [926]. Umbra [967]. unbiased [1163]. Uncertain [1220, 315, 323, 463, 775, 770, 453, 795, 515, 488, 583, 445, 514, 545, 454, 491, 543, 579, 604, 683, 1078]. uncertainty [350, 431, 567, 898, 673]. Understanding [25, 479, 617]. undirected [588]. unfairness [962]. Unified [1048, 678, 521, 835, 761]. Uniform [60]. uniformity [623]. unifying [249]. units [524, 670]. univariate [956]. unleashed [1210]. unobstructed [806]. UnQL [161]. unscalable [628]. unstructured [343]. Unsupervised [969, 255, 1054]. untrusted [533]. Unveiling [540]. Update [153, 285, 263, 443, 547, 220, 1181, 663]. update-intensive [443, 1181]. Updates [14, 117, 636, 467, 889, 1132, 364, 442, 1203]. Updating [54, 562, 251]. UpStream [547]. usage [1081]. use [1019]. User [836, 191, 800, 1035, 462, 1214, 1080, 866, 777]. User-cognizant [191]. users [477]. Using [112, 10, 150, 135, 115, 120, 125, 132, 325, 833, 357, 1232, 1057, 268, 190, 297, 803, 476, 344, 619, 798, 596, 744, 551, 158, 674, 193, 808, 245, 336, 409, 281, 918, 314, 322, 621, 758, 801, 1169, 1221, 862, 1062, 785, 527, 303, 1050, 415, 1047, 386, 189, 754, 706, 1101, 442, 1108, 966, 1109, 670, 461, 494, 1203, 997, 777]. utility [661, 700]. Utility-based [700]. utilizing [829]. UV [604]. UV-diagram [604].

V* [484]. V*-kNN [484]. valid [362]. validation [786]. Value [378, 11, 1057, 1184, 781]. Value-Based [11]. valued [235, 924, 513]. Values [71, 26, 435]. Variable [137, 13]. Variable-Bit-Rate [137]. variants [873]. variety [973]. varying [795, 898, 791]. VBTree [827]. VDDA [712]. vector [1036, 310, 318, 336, 1159, 1027]. vectorized [918, 917]. vectors [724, 1188, 176]. vehicle [911]. verification [426]. Versatile [79, 1174, 342]. Version [1091, 65, 810, 885, 1214]. Versioning [30, 885]. vertex [1045]. vertical [232]. vertically [428]. vertices [1172]. Very [152, 250, 474, 386, 1011, 666, 176, 572]. via [231, 656, 1053, 859, 948, 627, 1086, 1049, 1189, 1051, 1174, 312, 320, 1201, 772, 661, 811, 762, 1023, 554]. video [247, 571, 777]. View [119, 163, 311, 319, 38, 1213, 210, 244]. View-Objects [38]. Views [211, 275, 107, 113, 467, 297, 193, 245, 639, 311, 319, 189, 434]. violations [632]. VIP [917]. virtual [307, 434]. visibility [935]. visible [526]. visual [645, 1043, 172, 1214, 904, 765, 249]. visualization [712, 1212, 871]. visualization-driven [712]. visualizing [941]. Visually [952]. ViSWeb [249]. VLDB [750, 144, 516, 887, 557, 1028, 697, 883, 470, 812, 867, 1226, 253, 427, 309, 317, 709, 590, 749, 1084, 930, 635, 143, 144]. VLL [701]. VOD [142]. VoIP [413]. volatile [1040]. VolcanoML [1051]. Voronoi [412, 604]. voting [1231]. vp [164]. vp-Tree [164]. vs [1236]. VUS [1206].

walk [963, 793]. walk-based [793]. wall [983]. warehouse [223]. warehouses [1185, 194]. warfare [718]. warping [382, 437, 603]. Watermarking [230]. WaveCluster [152]. Wavelet [390, 152, 419]. wavelets [190]. WavingSketch [1163]. Way [20, 360, 998, 742, 335]. weak [892]. wearing [731]. web [593, 361, 528, 352, 694, 287, 585, 428, 155, 271, 160, 618, 615, 157, 184, 518, 612, 239, 249, 519, 159, 242, 617, 240, 257, 256, 411, 238, 449, 241, 379, 363]. WebSemantics [200]. WebSources [158]. weight [636, 602]. weighted [925]. weighting [503]. Weights [115]. Which [556, 398]. While [54]. why-not [828, 747]. Wide [76]. Wide-Area [76]. wild [1223]. window [1119, 1081, 360, 488, 686, 970, 1077, 1104, 763]. window-based [1081]. Windowed [570]. windows [771]. wireless [414, 289, 452]. Wise [164, 257]. WISE-Integrator [257]. within [962]. without [809, 1221, 1145]. word [400]. work [745]. work-efficient [745]. Workaholics [64, 65]. workarounds [623]. worker [1005]. workflow [286, 242, 195, 183]. workflows [783, 1199]. workload [562, 665, 1156]. workload-aware [665, 1156]. workloads [535, 722, 390, 739, 805, 667, 1138, 663, 576, 822]. workstations [443]. world [140, 588, 857, 657]. worlds [455, 873]. worst [970, 653, 771]. worst-case [970, 653, 771]. wrapper [178]. write [587, 561]. write-access [587]. write-ahead [561]. write-optimized [740]. writes [731]. WWW [357].

xDBTagger [1108]. XML [387, 456, 211, 690, 570, 268, 692, 180, 481, 467, 587, 291, 218, 301, 217, 609, 215, 616, 567, 219, 214, 278, 459, 366, 341, 368, 492, 364, 527, 395, 187, 434, 183, 407, 498, 598, 348, 629]. XML-based [407]. XML-enabled [183]. XPath [218, 448, 630]. XQuery [259, 355, 261].

Yannakakis [889]. years [1128, 1211]. YmalDB [627]. York [144].

Z [485]. Z-order [485]. Z-SKY [485]. Zen [1040]. zero [648, 1165]. zero-overhead [648]. zero-shot [1165]. Zigzag [19]. zones [766, 1101].

References

Breitbart:1992:TMI

[1] Yuri Breitbart, Abraham Silberschatz, and Glenn R. Thompson. Transaction management issues in a failure-prone multidatabase system environment. VLDB Journal: Very Large Data Bases, 1(1):1– 39, July 1992. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/b/Breitbart:Yuri.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
s/Silberschatz:Abraham.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/t/Thompson:

Glenn_R=.html. Electronic edition.

Nodine:1992:CTH

[2] Marian H. Nodine and Stanley B.

Zdonik. Cooperative transaction hierarchies: Transaction support for design applications. VLDB Journal: Very Large Data Bases, 1 (1):41–80, July 1992. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/n/Nodine:
Marian_H=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/z/Zdonik:Stanley_B=.html.

Electronic edition.

Spaccapietra:1992:MIA

[3] Stefano Spaccapietra, Christine Parent, and Yann Dupont. Model independent assertions for integration of heterogeneous schemas. VLDB Journal: Very Large Data Bases, 1(1):81– 126, July 1992. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/d/Dupont:Yann.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
p/Parent:Christine.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/s/Spaccapietra:

Stefano.html. Electronic edition.

Hsiao:1992:FDSa

[4] David K. Hsiao. Federated databases and systems: Part I — a tutorial on their data sharing. VLDB Journal: Very Large Data Bases, 1 (1):127–179, July 1992. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/h/Hsiao:

David_K=.html. Electronic edition.

Breitbart:1992:OMT

[5] Yuri Breitbart, Hector Garcia-Molina, and Abraham Silberschatz. Overview of multidatabase transaction management. VLDB Journal: Very Large Data Bases, 1(2):181–240, October 1992. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/b/Breitbart:Yuri.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
g/Garcia=Molina:Hector.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/s/Silberschatz:

Abraham.html. Electronic edition.

Drew:1992:TII

[6] Pamela Drew, Roger King, and Dennis Heimbigner. A toolkit for the incremental implementation of heterogeneous database management systems. VLDB Journal: Very Large Data Bases, 1(2):241–284, October 1992. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/d/Drew:Pamela.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
h/Heimbigner:Dennis.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/k/King:Roger.

html. Electronic edition.

Hsiao:1992:FDSb

[7] David K. Hsiao. Federated databases and systems: Part II — a tutorial on their resource consolidation. VLDB Journal: Very Large Data Bases, 1 (2):285–310, October 1992. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/h/Hsiao:

David_K=.html. Electronic edition.

Yu:1993:BMB

[8] Philip S. Yu and Douglas W. Cornell. Buffer management based on return on consumption in a multiquery environment. VLDB Journal: Very Large Data Bases, 2(1):1–37, January 1993. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/c/Cornell:Douglas_
                                        W=.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/y/Yu:Philip_S=.html. Elec-

tronic edition.

Harder:1993:CCI

[9] Theo Härder and Kurt Rothermel. Concurrency control issues in nested transactions. VLDB Journal: Very Large Data Bases, 2(1):39–74, January 1993. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/h/H=auml=rder:
                                        Theo.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/r/Rothermel:Kurt.html. Elec-

tronic edition.

Jensen:1993:UDT

[10] Christian S. Jensen, Leo Mark, Nick Roussopoulos, and Timos K. Sellis. Using differential techniques to efficiently support transaction time. VLDB Journal: Very Large Data Bases, 2 (1):75–116, January 1993. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/j/Jensen:
Christian_S=.html;   http://ftp.
informatik.rwth-aachen.de/dblp/
db/indices/a-tree/m/Mark:Leo.html;
http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/r/Roussopoulos:
Nick.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/s/Sellis:Timos_K=.html. Elec-

tronic edition.

Haritsa:1993:VBS

[11] Jayant R. Haritsa, Michael J. Carey, and Miron Livny. Value-based scheduling in real-time database systems. VLDB Journal: Very Large Data Bases, 2(2):117–152, April 1993. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/c/Carey:Michael_
J=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/h/Haritsa:Jayant_R=.html;
http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/l/Livny:

Miron.html. Electronic edition.

Grant:1993:QLR

[12] John Grant, Witold Litwin, Nick Roussopoulos, and Timos K. Sellis. Query languages for relational multidatabases. VLDB Journal: Very Large Data Bases, 2(2):153– 171, April 1993. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/g/Grant:John.
html;   http://ftp.informatik.rwth-

aachen.de/dblp/db/indices/a-tree/
l/Litwin:Witold.html;   http://
ftp.informatik.rwth-aachen.de/dblp/
db/indices/a-tree/r/Roussopoulos:
Nick.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/s/Sellis:Timos_K=.html. Elec-

tronic edition.

Neufeld:1993:GCT

[13] Andrea Neufeld, Guido Moerkotte, and Peter C. Lockemann. Generating consistent test data for a variable set of general consistency constraints. VLDB Journal: Very Large Data Bases, 2(2):173–213, April 1993. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/l/Lockemann:Peter_
C=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/m/Moerkotte:Guido.html;
http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/n/Neufeld:

Andrea.html. Electronic edition.

Du:1993:SCU

[14] Weimin Du, Ahmed K. Elmagarmid, Won Kim, and Omran A. Bukhres. Supporting consistent updates in replicated multidatabase systems. VLDB Journal: Very Large Data Bases, 2(2):215–241, April 1993. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/b/Bukhres:Omran_
A=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/d/Du:Weimin.html;   http:
//ftp.informatik.rwth-aachen.de/

dblp/db/indices/a-tree/e/Elmagarmid:
Ahmed_K=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/k/Kim:Won.html. Electronic

edition.

Anonymous:1993:Ca

[15] Anonymous. Column. VLDB Journal: Very Large Data Bases, 2(2):??, April 1993. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Anonymous:1993:Cb

[16] Anonymous. Column. VLDB Journal: Very Large Data Bases, 2(2):??, April 1993. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Tomasic:1993:SIP

[17] Anthony Tomasic and Hector Garcia-Molina. Special issue in parallelism in database systems: Query processing and inverted indices in shared-nothing document information retrieval systems. VLDB Journal: Very Large Data Bases, 2 (3):243–275, July 1993. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/g/Garcia=
Molina:Hector.html;   http://ftp.
informatik.rwth-aachen.de/dblp/
db/indices/a-tree/t/Tomasic:Anthony.

html. Electronic edition.

Tomasic:1993:QPI

[18] Anthony Tomasic and Hector Garcia-Molina. Query processing and inverted indices in shared: nothing text document information retrieval systems. VLDB Journal: Very Large Data Bases, 2(3):243–276, July 1993.

CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ziane:1993:PQP

[19] Mikal Ziane, Mohamed Zaït, and Pascale Borla-Salamet. Parallel query processing with zigzag trees. VLDB Journal: Very Large Data Bases, 2 (3):277–301, July 1993. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/b/Borla=
Salamet:Pascale.html;   http://
ftp.informatik.rwth-aachen.de/dblp/
db/indices/a-tree/z/Za=iuml=t:Mohamed.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
z/Ziane:Mikal.html. Electronic edi-

tion.

Hua:1993:CDS

[20] Kien A. Hua, Yu lung Lo, and Honesty C. Young. Considering data skew factor in multi-way join query optimization for parallel execution. VLDB Journal: Very Large Data Bases, 2(3):303–330, July 1993. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/h/Hua:Kien_A=
.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/l/Lo:Yu=lung.html; http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/y/Young:

Honesty_C=.html. Electronic edition.

Zhang:1993:TGC

[21] Aidong Zhang and Ahmed K. Elmagarmid. A theory of global concurrency control in multidatabase systems. VLDB Journal: Very Large Data Bases, 2(3):331–360, July 1993. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/e/Elmagarmid:
Ahmed_K=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/z/Zhang:Aidong.html. Elec-

tronic edition.

Anonymous:1993:SIP

[22] Anonymous. Special issue in parallelism in database systems. VLDB Journal: Very Large Data Bases, 2 (3):??, July 1993. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Srinivasan:1993:PBT

[23] V. Srinivasan and Michael J. Carey. Performance of B^{+} tree concurrency control algorithms. VLDB Journal: Very Large Data Bases, 2(4): 361–406, October 1993. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/c/Carey:
Michael_J=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/s/Srinivasan:V=.html. Elec-

tronic edition.

Weikum:1993:MLT

[24] Gerhard Weikum and Christof Hasse. Multi-level transaction management for complex objects: Implementation, performance, parallelism. VLDB Journal: Very Large Data Bases, 2(4): 407–453, October 1993. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/h/Hasse:
Christof.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/w/Weikum:Gerhard.html. Elec-

tronic edition.

Storey:1993:USR

[25] Veda C. Storey. Understanding semantic relationships. VLDB Journal: Very Large Data Bases, 2(4): 455–488, October 1993. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/s/Storey:

Veda_C=.html. Electronic edition.

Tseng:1993:SMS

[26] Frank Shou-Cheng Tseng, Arbee L. P. Chen, and W.-P. Yang. Searching a minimal semantically-equivalent subset of a set of partial values. VLDB Journal: Very Large Data Bases, 2(4):489– 512, October 1993. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/c/Chen:Arbee_
L=_P=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/t/Tseng:Frank_Shou=Cheng.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
y/Yang:W==P=.html. Electronic edi-

tion.

Georgakopoulos:1994:CST

[27] Dimitrios Georgakopoulos, Marek Rusinkiewicz, and Witold Litwin. Chronological scheduling of transactions with temporal dependencies. VLDB Journal: Very Large Data Bases, 3(1):1–28, January 1994. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
g/Georgakopoulos:Dimitrios.html;
 http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/l/Litwin:
Witold.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/r/Rusinkiewicz:Marek.html.

Electronic edition.

Whang:1994:DMD

[28] Kyu Young Whang, Sang Wook Kim, and Gio Wiederhold. Dynamic maintenance of data distribution for selectivity estimation. VLDB Journal: Very Large Data Bases, 3(1):29–51, January 1994. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/k/Kim:Sang=Wook.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
w/Whang:Kyu=Young.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/w/Wiederhold:

Gio.html. Electronic edition.

Kamel:1994:PBO

[29] Nabil Kamel, Ping Wu, and Stanley Y. W. Su. A pattern-based object calculus. VLDB Journal: Very Large Data Bases, 3(1):53–76, January 1994. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/k/Kamel:Nabil.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
s/Su:Stanley_Y=_W=.html;   http:

                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/w/Wu:Ping.

html. Electronic edition.

Sciore:1994:VCM

[30] Edward Sciore. Versioning and configuration management in an objectoriented data model. VLDB Journal: Very Large Data Bases, 3(1): 77–106, January 1994. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/s/Sciore:

Edward.html. Electronic edition.

Ramamohanarao:1994:IDD

[31] Kotagiri Ramamohanarao and James Harland. An introduction to deductive database languages and systems. VLDB Journal: Very Large Data Bases, 3(2):107–122, April 1994. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ramamohanarao:1994:SIP

[32] Kotagiri Ramamohanarao and James Harland. Special issue on prototypes of deductive database systems: An introduction to deductive database languages and systems. VLDB Journal: Very Large Data Bases, 3(2):107– 122, April 1994. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/h/Harland:James.
                                        html;   http://ftp.informatik.rwth-
                                        aachen.de/dblp/db/indices/a-tree/
                                        r/Ramamohanarao:Kotagiri.html. Elec-

tronic edition.

Derr:1994:GND

[33] Marcia A. Derr, Shinichi Morishita, and Geoffrey Phipps. The glue-nail deductive database system: Design, implementation, and evaluation. VLDB Journal: Very Large Data Bases, 3(2):123–160, April 1994. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/d/Derr:Marcia_
A=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/m/Morishita:Shinichi.html;
http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/p/Phipps:

Geoffrey.html. Electronic edition.

Ramakrishnan:1994:CDS

[34] Raghu Ramakrishnan, Divesh Srivastava, S. Sudarshan, and Praveen Seshadri. The CORAL deductive system. VLDB Journal: Very Large Data Bases, 3(2):161–210, April 1994. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/r/Ramakrishnan:
Raghu.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/s/Seshadri:Praveen.html;
http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/s/Srivastava:
Divesh.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/s/Sudarshan:S=.html. Elec-

tronic edition.

Kiessling:1994:DSE

[35] Werner Kießling, Helmut Schmidt, Werner Strauß, and Gerhard Dünzinger.

DECLARE and SDS: Early efforts to commercialize deductive database technology. VLDB Journal: Very Large Data Bases, 3 (2):211–243, April 1994. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/d/D=uuml=
nzinger:Gerhard.html;   http://
ftp.informatik.rwth-aachen.de/dblp/
db/indices/a-tree/k/Kie=szlig=ling:
Werner.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/s/Schmidt:Helmut.html;
http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/s/Strau=
szlig=:Werner.html. Electronic edi-

tion.

Vaghani:1994:ADD

[36] Jayen Vaghani, Kotagiri Ramamohanarao, David B. Kemp, Zoltan Somogyi, Peter J. Stuckey, Tim S. Leask, and James Harland. The Aditi deductive database system. VLDB Journal: Very Large Data Bases, 3(2):245– 288, April 1994. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/h/Harland:James.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
k/Kemp:David_B=.html;   http://
ftp.informatik.rwth-aachen.de/dblp/
db/indices/a-tree/l/Leask:Tim_S=
.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/r/Ramamohanarao:Kotagiri.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
s/Somogyi:Zoltan.html;   http:/
/ftp.informatik.rwth-aachen.de/

dblp/db/indices/a-tree/s/Stuckey:
Peter_J=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/v/Vaghani:Jayen.html. Elec-

tronic edition.

Anonymous:1994:SIP

[37] Anonymous. Special issue on prototypes of deductive database systems. VLDB Journal: Very Large Data Bases, 3(2):??, April 1994. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lee:1994:EIV

[38] Byung Suk Lee and Gio Wiederhold. Efficiently instantiating view-objects from remote relational databases. VLDB Journal: Very Large Data Bases, 3(3):289–323, July 1994. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
l/Lee:Byung_Suk.html;   http://
ftp.informatik.rwth-aachen.de/dblp/
db/indices/a-tree/w/Wiederhold:

Gio.html. Electronic edition.

Barbara-Milla:1994:DPT

[39] Daniel Barbará-Millá and Hector Garcia-Molina. The demarcation protocol: a technique for maintaining constraints in distributed database systems. VLDB Journal: Very Large Data Bases, 3(3):325–353, July 1994. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Barbara:1994:DPT

[40] Daniel Barbará and Hector Garcia-Molina. The demarcation protocol: a technique for maintaining constraints in distributed database systems. VLDB Journal: Very Large Data Bases, 3 (3):325–353, July 1994. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/b/Barbar=
aacute=:Daniel.html;   http://ftp.
informatik.rwth-aachen.de/dblp/
db/indices/a-tree/g/Garcia=Molina:

Hector.html. Electronic edition.

Bertino:1994:ICO

[41] Elisa Bertino. Index configuration in object-oriented databases. VLDB Journal: Very Large Data Bases, 3 (3):355–399, July 1994. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/b/Bertino:

Elisa.html. Electronic edition.

Guting:1994:ISD

[42] Ralf Hartmut Güting. An introduction to spatial database systems. VLDB Journal: Very Large Data Bases, 3 (4):357–399, October 1994. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Guting:1994:SIS

[43] Ralf Hartmut Güting. Special issue on spatial database systems: An introduction to spatial database systems. VLDB Journal: Very Large Data Bases, 3(4):357–399, October 1994. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/g/G=uuml=ting:
Ralf_Hartmut.html. Electronic edi-

tion.

Baumann:1994:MMD

[44] Peter Baumann. Management of multidimensional discrete data. VLDB Journal: Very Large Data Bases, 3(4):401– 444, October 1994. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/b/Baumann:Peter.

html. Electronic edition.

Chu:1994:SMA

[45] Wesley W. Chu, Ion Tim Ieong, and Ricky K. Taira. A semantic modeling approach for image retrieval by content. VLDB Journal: Very Large Data Bases, 3(4):445–477, October 1994. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/c/Chu:Wesley_
W=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/i/Ieong:Ion_Tim.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/t/Taira:

Ricky_K=.html. Electronic edition.

Papadias:1994:QRS

[46] Dimitris Papadias and Timos K. Sellis. Qualitative representation of spatial knowledge in two-dimensional space. VLDB Journal: Very Large Data Bases, 3(4):479–516, October 1994. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/p/Papadias:Dimitris.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/

                                        s/Sellis:Timos_K=.html. Electronic

edition.

Lin:1994:TTI

[47] King Ip Lin, H. V. Jagadish, and Christos Faloutsos. The TVtree: An index structure for highdimensional data. VLDB Journal: Very Large Data Bases, 3(4):517–542, October 1994. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/f/Faloutsos:Christos.
                                        html;   http://ftp.informatik.rwth-
                                        aachen.de/dblp/db/indices/a-tree/
                                        j/Jagadish:H=_V=.html;   http:/
                                        /ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/l/Lin:King=

Ip.html. Electronic edition.

Anonymous:1994:SIS

[48] Anonymous. Special issue on spatial database systems. VLDB Journal: Very Large Data Bases, 3(4):??, October 1994. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Constantopoulos:1995:SIB

[49] Panos Constantopoulos, Matthias Jarke, John Mylopoulos, and Yannis Vassiliou. The software information base: a server for reuse. VLDB Journal: Very Large Data Bases, 4(1):1– 43, January 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/c/Constantopoulos:
                                        Panos.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/j/Jarke:Matthias.html;

http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/m/Mylopoulos:
John.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/v/Vassiliou:Yannis.html.

Electronic edition.

Clifton:1995:HDQ

[50] Chris Clifton, Hector Garcia-Molina, and David Bloom. HyperFile: a data and query model for documents. VLDB Journal: Very Large Data Bases, 4(1):45–86, January 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/b/Bloom:David.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
c/Clifton:Chris.html;   http://
ftp.informatik.rwth-aachen.de/dblp/
db/indices/a-tree/g/Garcia=Molina:

Hector.html. Electronic edition.

Agrawal:1995:OSL

[51] Divyakant Agrawal, Amr El Abbadi, Richard Jeffers, and Lijing Lin. Ordered shared locks for realtime databases. VLDB Journal: Very Large Data Bases, 4(1):87–126, January 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/a/Abbadi:Amr_
El.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/a/Agrawal:Divyakant.html;
 http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/j/Jeffers:
Richard.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/

                                        a-tree/l/Lin:Lijing.html. Elec-

tronic edition.

Dan:1995:CDA

[52] Asit Dan, Philip S. Yu, and Jen Yao Chung. Characterization of database access pattern for analytic prediction of buffer hit probability. VLDB Journal: Very Large Data Bases, 4 (1):127–154, January 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/c/Chung:
                                        Jen=Yao.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/d/Dan:Asit.html;   http:
                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/y/Yu:Philip_

S=.html. Electronic edition.

Peckham:1995:DME

[53] Joan Peckham, Bonnie MacKellar, and Michael Doherty. Data model for extensible support of explicit relationships in design databases. VLDB Journal: Very Large Data Bases, 4(2):157– 191, April 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/d/Doherty:Michael.
                                        html;   http://ftp.informatik.rwth-
                                        aachen.de/dblp/db/indices/a-tree/
                                        m/MacKellar:Bonnie.html;   http:
                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/p/Peckham:

Joan.html. Electronic edition.

Teniente:1995:UKB

[54] Ernest Teniente and Antoni Olivé. Updating knowledge bases while maintaining their consistency. VLDB Journal: Very Large Data Bases, 4(2):193– 241, April 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/o/Oliv=eacute=
:Antoni.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/t/Teniente:Ernest.html. Elec-

tronic edition.

Guting:1995:RBS

[55] Ralf Hartmut Güting and Markus Schneider. Realm-based spatial data types: The ROSE algebra. VLDB Journal: Very Large Data Bases, 4 (2):243–286, April 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/g/G=uuml=
ting:Ralf_Hartmut.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/s/Schneider:

Markus.html. Electronic edition.

Templeton:1995:IDC

[56] Marjorie Templeton, Herbert Henley, Edward Maros, and Darrel J. Van Buer. InterViso: Dealing with the complexity of federated database access. VLDB Journal: Very Large Data Bases, 4(2):287– 317, April 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/b/Buer:Darrel_
J=_Van.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/h/Henley:Herbert.html;
http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/m/Maros:
Edward.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/

                                        a-tree/t/Templeton:Marjorie.html.

Electronic edition.

Atkinson:1995:SIP

[57] Malcolm P. Atkinson and Ronald Morrison. Special issue on persistent object systems: Orthogonally persistent object systems. VLDB Journal: Very Large Data Bases, 4(3):319– 401, July 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/a/Atkinson:Malcolm_
                                        P=.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/m/Morrison:Ronald.html. Elec-

tronic edition.

Atkinson:1995:OPO

[58] Malcolm Atkinson and Ronald Morrison. Orthogonally persistent object systems. VLDB Journal: Very Large Data Bases, 4(3):319–402, July 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Albano:1995:FPL

[59] Antonio Albano, Giorgio Ghelli, and Renzo Orsini. Fibonacci: a programming language for object databases. VLDB Journal: Very Large Data Bases, 4(3):403–444, July 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/a/Albano:Antonio.
                                        html;   http://ftp.informatik.rwth-
                                        aachen.de/dblp/db/indices/a-tree/
                                        g/Ghelli:Giorgio.html;   http:/
                                        /ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/o/Orsini:

Renzo.html. Electronic edition.

Ozsu:1995:TUB

[60] M. Tamer Özsu, Randal J. Peters, Duane Szafron, Boman Irani, Anna Lipka, and Adriana Muñoz. TIGUKAT: a uniform behavioral objectbase management system. VLDB Journal: Very Large Data Bases, 4 (3):445–492, July 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/=/=Ouml=
zsu:M=_Tamer.html;   http://ftp.
informatik.rwth-aachen.de/dblp/
db/indices/a-tree/i/Irani:Boman.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
l/Lipka:Anna.html; http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/m/Mu=ntilde=oz:Adriana.html;
 http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/p/Peters:
Randal_J=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/s/Szafron:Duane.html. Elec-

tronic edition.

Benzaken:1995:TDP

[61] Véronique Benzaken and Anne Doucet. Thémis: a database programming language handling integrity constraints. VLDB Journal: Very Large Data Bases, 4(3):493–517, July 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/b/Benzaken:V=
eacute=ronique.html;   http://ftp.
informatik.rwth-aachen.de/dblp/
db/indices/a-tree/d/Doucet:Anne.

html. Electronic edition.

Kemper:1995:APS

[62] Alfons Kemper and Donald Kossmann. Adaptable pointer swizzling strategies in object bases: Design, realization, and quantitative analysis. VLDB Journal: Very Large Data Bases, 4 (3):519–566, July 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/k/Kemper:
                                        Alfons.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/k/Kossmann:Donald.html. Elec-

tronic edition.

Anonymous:1995:SIP

[63] Anonymous. Special issue on persistent object systems. VLDB Journal: Very Large Data Bases, 4(3):??, July 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Barbara:1995:SSO

[64] Daniel Barbará and Tomasz Imielinski. Special system-oriented section: The best of SIGMOD 1994: Sleepers and workaholics: Caching strategies in mobile environments. VLDB Journal: Very Large Data Bases, 4 (4):567–602, October 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/b/Barbar=
                                        aacute=:Daniel.html;   http://ftp.
                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/i/Imielinski:

Tomasz.html. Electronic edition.

Barbara:1995:SWC

[65] Daniel Barbará and Tomasz Imieliński. Sleepers and workaholics: caching strategies in mobile environments (extended version). VLDB Journal: Very Large Data Bases, 4(4):567–602, October 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Nyberg:1995:ACS

[66] Chris Nyberg, Tom Barclay, Zarka Cvetanovic, Jim Gray, and David B. Lomet. AlphaSort: a cache-sensitive parallel external sort. VLDB Journal: Very Large Data Bases, 4(4):603– 627, October 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/b/Barclay:Tom.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
c/Cvetanovic:Zarka.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/g/Gray:Jim.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
l/Lomet:David_B=.html;   http:/
/ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/n/Nyberg:

Chris.html. Electronic edition.

White:1995:QHP

[67] Seth J. White and David J. De-Witt. QuickStore: a high performance mapped object store. VLDB Journal: Very Large Data Bases, 4 (4):629–673, October 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/d/DeWitt:
David_J=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/w/White:Seth_J=.html. Elec-

tronic edition.

Swami:1995:EPF

[68] Arun N. Swami and K. Bernhard Schiefer. Estimating page fetches for index scans with finite LRU buffers. VLDB Journal: Very Large Data Bases, 4(4):675–701, October 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/s/Schiefer:K=_
                                        Bernhard.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/s/Swami:Arun_N=.html. Elec-

tronic edition.

Landau:1995:HQA

[69] Gad M. Landau, Jeanette P. Schmidt, and Vassilis J. Tsotras. Historical queries along multiple lines of time evolution. VLDB Journal: Very Large Data Bases, 4(4):703–726, October 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Landau:1995:RJA

[70] Gad M. Landau, Jeanette P. Schmidt, and Vassilis J. Tsotras. Regular journal articles: Historical queries along multiple lines of time evolution. VLDB Journal: Very Large Data Bases, 4(4):703– 726, October 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/l/Landau:Gad_
                                        M=.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/s/Schmidt:Jeanette_P=.html;
                                         http://ftp.informatik.rwth-aachen.
                                        de/dblp/db/indices/a-tree/t/Tsotras:

Vassilis_J=.html. Electronic edition.

Abiteboul:1995:PLM

[71] Serge Abiteboul and Catriel Beeri. The power of languages for the manipulation of complex values. VLDB Journal: Very Large Data Bases, 4(4):727– 794, October 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/a/Abiteboul:Serge.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
b/Beeri:Catriel.html. Electronic

edition.

Anonymous:1995:SSO

[72] Anonymous. Special system-oriented section: the best of SIGMOD ‘94. VLDB Journal: Very Large Data Bases, 4(4):??, October 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

DeWitt:1996:POT

[73] David J. De Witt, Jeffrey F. Naughton, John C. Shafer, and Shivakumar Venkataraman. Parallelizing OODBMS traversals: a performance evaluation. VLDB Journal: Very Large Data Bases, 5 (1):3–18, January 1996. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/d/DeWitt:
David_J=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/n/Naughton:Jeffrey_F=.html;
http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/s/Shafer:
John_C=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/v/Venkataraman:Shivakumar.

html;   http://link.springer.de/
link/service/journals/00778/bibs/
6005001/60050003.htm; http://link.
springer.de/link/service/journals/
00778/papers/6005001/60050003.pdf;
 http://link.springer.de/link/service/
journals/00778/papers/6005001/60050003.

ps.gz. Electronic edition.

Sivasankaran:1996:PAR

[74] Rajendran M. Sivasankaran, John A. Stankovic, Donald F. Towsley, Bhaskar Purimetla, and Krithi Ramamritham. Priority assignment in real-time active databases. VLDB Journal: Very Large Data Bases, 5(1):19–34, January 1996. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/p/Purimetla:Bhaskar.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
r/Ramamritham:Krithi.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/s/Sivasankaran:
Rajendran_M=.html;   http://ftp.
informatik.rwth-aachen.de/dblp/
db/indices/a-tree/s/Stankovic:John_
A=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/t/Towsley:Donald_F=.html;
 http://link.springer.de/link/service/
journals/00778/bibs/6005001/60050019.
htm; http://link.springer.de/link/
service/journals/00778/papers/6005001/
60050019.pdf;   http://link.springer.
de/link/service/journals/00778/
papers/6005001/60050019.ps.gz. Elec-

tronic edition.

Keller:1996:PBC

[75] Arthur M. Keller and Julie Basu. A predicate-based caching scheme for client-server database architectures. VLDB Journal: Very Large Data Bases, 5(1):35–47, January 1996. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/b/Basu:Julie.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
k/Keller:Arthur_M=.html;   http:
//link.springer.de/link/service/
journals/00778/bibs/6005001/60050035.
htm; http://link.springer.de/link/
service/journals/00778/papers/6005001/
60050035.pdf;   http://link.springer.
de/link/service/journals/00778/
papers/6005001/60050035.ps.gz. Elec-

tronic edition.

Stonebraker:1996:MWA

[76] Michael Stonebraker, Paul M. Aoki, Witold Litwin, Avi Pfeffer, Adam Sah, Jeff Sidell, Carl Staelin, and Andrew Yu. Mariposa: a wide-area distributed database system. VLDB Journal: Very Large Data Bases, 5(1):48– 63, January 1996. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/a/Aoki:Paul_M=
.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/l/Litwin:Witold.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/p/Pfeffer:
Avi.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/s/Sah:Adam.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/s/Sidell:
Jeff.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/

                                        a-tree/s/Staelin:Carl.html;   http:
                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/s/Stonebraker:
                                        Michael.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/y/Yu:Andrew.html;   http:
                                        //link.springer.de/link/service/
                                        journals/00778/bibs/6005001/60050048.
                                        htm; http://link.springer.de/link/
                                        service/journals/00778/papers/6005001/
                                        60050048.pdf;   http://link.springer.
                                        de/link/service/journals/00778/
                                        papers/6005001/60050048.ps.gz. Elec-

tronic edition.

Harris:1996:JAC

[77] Evan P. Harris and Kotagiri Ramamohanarao. Join algorithm costs revisited. VLDB Journal: Very Large Data Bases, 5(1):64–84, January 1996. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/h/Harris:Evan_
                                        P=.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/r/Ramamohanarao:Kotagiri.
                                        html;   http://link.springer.de/
                                        link/service/journals/00778/bibs/
                                        6005001/60050064.htm; http://link.
                                        springer.de/link/service/journals/
                                        00778/papers/6005001/60050064.pdf;
                                         http://link.springer.de/link/service/
                                        journals/00778/papers/6005001/60050064.

ps.gz. Electronic edition.

Ramamritham:1996:TCC

[78] Krithi Ramamritham and Panos K. Chrysanthis. A taxonomy of correctness criteria in database applications (*). VLDB Journal: Very Large Data Bases, 5(1):85–97, January 1996. CODEN VLDBFR.

ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/c/Chrysanthis:
Panos_K=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/r/Ramamritham:Krithi.html;
http://link.springer.de/link/service/
journals/00778/bibs/6005001/60050085.
htm; http://link.springer.de/link/
service/journals/00778/papers/6005001/
60050085.pdf;   http://link.springer.
de/link/service/journals/00778/
papers/6005001/60050085.ps.gz. Elec-

tronic edition.

Tsatalos:1996:GVT

[79] Odysseas G. Tsatalos, Marvin H. Solomon, and Yannis E. Ioannidis. The GMAP: a versatile tool for physical data independence. VLDB Journal: Very Large Data Bases, 5(2):101– 118, April 1996. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/i/Ioannidis:Yannis_
E=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/s/Solomon:Marvin_H=.html;
http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/t/Tsatalos:
Odysseas_G=.html;   http://link.
springer.de/link/service/journals/
00778/bibs/6005002/60050101.htm;
http://link.springer.de/link/service/
journals/00778/papers/6005002/60050101.
pdf;   http://link.springer.de/
link/service/journals/00778/papers/
6005002/60050101.ps.gz. Electronic

edition.

Poulovassilis:1996:AQO

[80] Alexandra Poulovassilis and Carol Small. Algebraic query optimisation for database programming languages. VLDB Journal: Very Large Data Bases, 5(2):119–132, April 1996. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/p/Poulovassilis:
Alexandra.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/s/Small:Carol.html;   http:
//link.springer.de/link/service/
journals/00778/bibs/6005002/60050119.
htm; http://link.springer.de/link/
service/journals/00778/papers/6005002/
60050119.pdf;   http://link.springer.
de/link/service/journals/00778/
papers/6005002/60050119.ps.gz. Elec-

tronic edition.

Amiel:1996:TSR

[81] Eric Amiel, Marie-Jo Bellosta, Eric Dujardin, and Eric Simon. Typesafe relaxing of schema consistency rules for flexible modeling in OODBMS. VLDB Journal: Very Large Data Bases, 5(2):133–150, April 1996. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/a/Amiel:Eric.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
b/Bellosta:Marie=Jo.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/d/Dujardin:
Eric.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/s/Simon:Eric.html; http:
//link.springer.de/link/service/
journals/00778/bibs/6005002/60050133.
htm;   http://link.springer.de/

link/service/journals/00778/papers/
6005002/60050133.pdf;   http://
link.springer.de/link/service/journals/
00778/papers/6005002/60050133.ps.

gz. Electronic edition.

Fang:1996:EOB

[82] Doug Fang, Shahram Ghandeharizadeh, and Dennis McLeod. An experimental object-based sharing system for networked databases. VLDB Journal: Very Large Data Bases, 5(2):151– 165, April 1996. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/f/Fang:Doug.html;
http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/g/Ghandeharizadeh:
Shahram.html;   http://link.springer.
de/link/service/journals/00778/
bibs/6005002/60050151.htm; http:
//link.springer.de/link/service/
journals/00778/papers/6005002/60050151.
pdf;   http://link.springer.de/
link/service/journals/00778/papers/
6005002/60050151.ps.gz;   http:
//link.springer.de/link/service/
journals/00778/tocs/mailto:helpdesk@

link.springer.de. Electronic edition.

Dey:1996:CTR

[83] Debabrata Dey, Terence M. Barron, and Veda C. Storey. A complete temporal relational algebra. VLDB Journal: Very Large Data Bases, 5 (3):167–180, August 1996. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/b/Barron:
Terence_M=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/

a-tree/d/Dey:Debabrata.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/s/Storey:
Veda_C=.html;   http://link.springer.
de/link/service/journals/00778/
bibs/6005003/60050167.htm; http:
//link.springer.de/link/service/
journals/00778/papers/6005003/60050167.
pdf;   http://link.springer.de/
link/service/journals/00778/papers/
6005003/60050167.ps.gz. Electronic

edition.

Shyy:1996:DIK

[84] Yuh-Ming Shyy, Javier Arroyo, Stanley Y. W. Su, and Herman Lam. The design and implementation of K: a highlevel knowledge-base programming language of OSAM*.KBMS. VLDB Journal: Very Large Data Bases, 5 (3):181–195, August 1996. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/a/Arroyo:
Javier.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/l/Lam:Herman.html; http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/s/Shyy:Yuh=
Ming.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/s/Su:Stanley_Y=_W=.html;
 http://link.springer.de/link/service/
journals/00778/bibs/6005003/60050181.
htm; http://link.springer.de/link/
service/journals/00778/papers/6005003/
60050181.pdf;   http://link.springer.
de/link/service/journals/00778/
papers/6005003/60050181.ps.gz. Elec-

tronic edition.

Harder:1996:APS

[85] Theo Härder and Joachim Reinert. Access path support for referential integrity in SQL2. VLDB Journal: Very Large Data Bases, 5(3):196–214, August 1996. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/h/H=auml=rder:
Theo.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/r/Reinert:Joachim.html;
http://link.springer.de/link/service/
journals/00778/bibs/6005003/60050196.
htm; http://link.springer.de/link/
service/journals/00778/papers/6005003/
60050196.pdf;   http://link.springer.
de/link/service/journals/00778/
papers/6005003/60050196.ps.gz. Elec-

tronic edition.

Ooi:1996:INE

[86] Beng Chin Ooi, Jiawei Han, Hongjun Lu, and Kian Lee Tan. Index nesting — an efficient approach to indexing in object-oriented databases. VLDB Journal: Very Large Data Bases, 5(3):215–228, August 1996. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/h/Han:Jiawei.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
l/Lu:Hongjun.html; http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/o/Ooi:Beng_Chin.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/t/Tan:Kian=
Lee.html;   http://link.springer.
de/link/service/journals/00778/
bibs/6005003/60050215.htm; http:
//link.springer.de/link/service/
journals/00778/papers/6005003/60050215.

                                        pdf;   http://link.springer.de/
                                        link/service/journals/00778/papers/
                                        6005003/60050215.ps.gz;   http:
                                        //link.springer.de/link/service/
                                        journals/00778/tocs/mailto:helpdesk@

link.springer.de. Electronic edition.

Antoshenkov:1996:QPO

[87] Gennady Antoshenkov and Mohamed Ziauddin. Query processing and optimization in Oracle Rdb. VLDB Journal: Very Large Data Bases, 5(4):229–237, December 1996. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/a/Antoshenkov:
                                        Gennady.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/z/Ziauddin:Mohamed.html;
                                         http://link.springer.de/link/service/
                                        journals/00778/bibs/6005004/60050229.
                                        htm; http://link.springer.de/link/
                                        service/journals/00778/papers/6005004/
                                        60050229.pdf;   http://link.springer.
                                        de/link/service/journals/00778/
                                        papers/6005004/60050229.ps.gz. Elec-

tronic edition.

Mylopoulos:1996:BKB

[88] John Mylopoulos, Vinay K. Chaudhri, Dimitris Plexousakis, Adel Shrufi, and Thodoros Topologlou. Building knowledge base management systems. VLDB Journal: Very Large Data Bases, 5(4):238–263, December 1996. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/c/Chaudhri:Vinay_
                                        K=.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/

a-tree/m/Mylopoulos:John.html;
http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/p/Plexousakis:
Dimitris.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/s/Shrufi:Adel.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/t/Topaloglou:
Thodoros.html;   http://link.springer.
de/link/service/journals/00778/
bibs/6005004/60050238.htm; http:
//link.springer.de/link/service/
journals/00778/papers/6005004/60050238.
pdf;   http://link.springer.de/
link/service/journals/00778/papers/
6005004/60050238.ps.gz. Electronic

edition.

Becker:1996:AOM

[89] Bruno Becker, Stephan Gschwind, Thomas Ohler, Bernhard Seeger, and Peter Widmayer. An asymptotically optimal multiversion B-tree. VLDB Journal: Very Large Data Bases, 5 (4):264–275, December 1996. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/b/Becker:
Bruno.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/g/Gschwind:Stephan.html;
 http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/o/Ohler:
Thomas.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/s/Seeger:Bernhard.html;
http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/w/Widmayer:
Peter.html;   http://link.springer.
de/link/service/journals/00778/
bibs/6005004/60050264.htm; http:
//link.springer.de/link/service/
journals/00778/papers/6005004/60050264.

                                        pdf;   http://link.springer.de/
                                        link/service/journals/00778/papers/
                                        6005004/60050264.ps.gz. Electronic

edition.

Kashyap:1996:SSS

[90] Vipul Kashyap and Amit P. Sheth. Semantic and schematic similarities between database objects: a contextbased approach. VLDB Journal: Very Large Data Bases, 5(4):276–304, December 1996. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/k/Kashyap:Vipul.
                                        html;   http://ftp.informatik.rwth-
                                        aachen.de/dblp/db/indices/a-tree/
                                        s/Sheth:Amit_P=.html;   http://
                                        link.springer.de/link/service/journals/
                                        00778/bibs/6005004/60050276.htm;
                                        http://link.springer.de/link/service/
                                        journals/00778/papers/6005004/60050276.
                                        pdf;   http://link.springer.de/
                                        link/service/journals/00778/papers/
                                        6005004/60050276.ps.gz;   http:
                                        //link.springer.de/link/service/
                                        journals/00778/tocs/mailto:helpdesk@

link.springer.de. Electronic edition.

Evangelidis:1997:HTM

[91] Georgios Evangelidis, David B. Lomet, and Betty Salzberg. The hB ^{Π}- tree: a multi-attribute index supporting concurrency, recovery and node consolidation. VLDB Journal: Very Large Data Bases, 6(1):1–25, February 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/e/Evangelidis:
                                        Georgios.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/

a-tree/l/Lomet:David_B=.html;
http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/s/Salzberg:
Betty.html;   http://link.springer.
de/link/service/journals/00778/
bibs/7006001/70060001.htm; http:
//link.springer.de/link/service/
journals/00778/papers/7006001/70060001.
pdf;   http://link.springer.de/
link/service/journals/00778/papers/
7006001/70060001.ps.gz. Electronic

edition.

Antoshenkov:1997:DBO

[92] Gennady Antoshenkov. Dictionarybased order-preserving string compression (*). VLDB Journal: Very Large Data Bases, 6(1):26–39, February 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/a/Antoshenkov:
Gennady.html;   http://link.springer.
de/link/service/journals/00778/
bibs/7006001/70060026.htm; http:
//link.springer.de/link/service/
journals/00778/papers/7006001/70060026.
pdf;   http://link.springer.de/
link/service/journals/00778/papers/
7006001/70060026.ps.gz. Electronic

edition.

Singhal:1997:ALB

[93] Vigyan Singhal and Alan Jay Smith. Analysis of locking behavior in three real database systems. VLDB Journal: Very Large Data Bases, 6(1):40– 52, February 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/s/Singhal:Vigyan.
html;   http://ftp.informatik.rwth-

                                        aachen.de/dblp/db/indices/a-tree/
                                        s/Smith:Alan_Jay.html;   http:/
                                        /link.springer.de/link/service/
                                        journals/00778/bibs/7006001/70060040.
                                        htm; http://link.springer.de/link/
                                        service/journals/00778/papers/7006001/
                                        70060040.pdf;   http://link.springer.
                                        de/link/service/journals/00778/
                                        papers/7006001/70060040.ps.gz. Elec-

tronic edition.

Mehta:1997:DPS

[94] Manish Mehta and David J. DeWitt. Data placement in shared-nothing parallel database systems (*). VLDB Journal: Very Large Data Bases, 6 (1):53–72, February 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/d/DeWitt:
                                        David_J=.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/m/Mehta:Manish.html;   http:
                                        //link.springer.de/link/service/
                                        journals/00778/bibs/7006001/70060053.
                                        htm; http://link.springer.de/link/
                                        service/journals/00778/papers/7006001/
                                        70060053.pdf;   http://link.springer.
                                        de/link/service/journals/00778/
                                        papers/7006001/70060053.ps.gz;
                                        http://link.springer.de/link/service/
                                        journals/00778/tocs/mailto:helpdesk@

link.springer.de. Electronic edition.

Papazoglou:1997:DMO

[95] Mike P. Papazoglou and Bernd J. Krämer. A database model for object dynamics. VLDB Journal: Very Large Data Bases, 6(2):73–96, August 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/

db/indices/a-tree/k/Kr=auml=mer:
Bernd_J=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/p/Papazoglou:Mike_P=.html;
http://link.springer.de/link/service/
journals/00778/bibs/7006002/70060073.
htm; http://link.springer.de/link/
service/journals/00778/papers/7006002/
70060073.pdf;   http://link.springer.
de/link/service/journals/00778/
papers/7006002/70060073.ps.gz. Elec-

tronic edition. See erratum [106].

Catarci:1997:GIH

[96] Tiziana Catarci, Giuseppe Santucci, and John Cardiff. Graphical interaction with heterogeneous databases (*). VLDB Journal: Very Large Data Bases, 6(2):97–120, August 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/c/Cardiff:John.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
c/Catarci:Tiziana.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/s/Santucci:
Giuseppe.html;   http://link.springer.
de/link/service/journals/00778/
bibs/7006002/70060097.htm; http:
//link.springer.de/link/service/
journals/00778/papers/7006002/70060097.
pdf;   http://link.springer.de/
link/service/journals/00778/papers/
7006002/70060097.ps.gz. Electronic

edition.

Chen:1997:AHF

[97] Ming-Syan Chen, Hui-I Hsiao, and Philip S. Yu. On applying hash filters to improving the execution of multi-join queries. VLDB Journal:

Very Large Data Bases, 6(2):121–131, August 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/c/Chen:Ming=Syan.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
h/Hsiao:Hui=I.html;   http://ftp.
informatik.rwth-aachen.de/dblp/
db/indices/a-tree/y/Yu:Philip_S=
.html;   http://link.springer.de/
link/service/journals/00778/bibs/
7006002/70060121.htm; http://link.
springer.de/link/service/journals/
00778/papers/7006002/70060121.pdf;
 http://link.springer.de/link/service/
journals/00778/papers/7006002/70060121.

ps.gz. Electronic edition.

Ioannidis:1997:PQO

[98] Yannis E. Ioannidis, Raymond T. Ng, Kyuseok Shim, and Timos K. Sellis. Parametric query optimization. VLDB Journal: Very Large Data Bases, 6(2):132–151, August 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/i/Ioannidis:Yannis_
E=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/n/Ng:Raymond_T=.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/s/Sellis:
Timos_K=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/s/Shim:Kyuseok.html;   http:
//link.springer.de/link/service/
journals/00778/bibs/7006002/70060132.
htm; http://link.springer.de/link/
service/journals/00778/papers/7006002/

70060132.pdf;   http://link.springer.
de/link/service/journals/00778/
papers/7006002/70060132.ps.gz. Elec-

tronic edition.

Mehrotra:1997:CCH

[99] Sharad Mehrotra, Henry F. Korth, and Avi Silberschatz. Concurrency control in hierarchical multidatabase systems. VLDB Journal: Very Large Data Bases, 6(2):152–172, August 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/k/Korth:Henry_
F=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/m/Mehrotra:Sharad.html;
http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/s/Silberschatz:
Abraham.html;   http://link.springer.
de/link/service/journals/00778/
bibs/7006002/70060152.htm; http:
//link.springer.de/link/service/
journals/00778/papers/7006002/70060152.
pdf;   http://link.springer.de/
link/service/journals/00778/papers/
7006002/70060152.ps.gz;   http:
//link.springer.de/link/service/
journals/00778/tocs/mailto:helpdesk@

link.springer.de. Electronic edition.

Cobb:1997:IOT

[100] Edward E. Cobb. The impact of object technology on commercial transaction processing. VLDB Journal: Very Large Data Bases, 6(3):173–190, August 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Cobb:1997:ITC

[101] Edward E. Cobb. The impact of technology on commercial transaction processing. VLDB Journal: Very Large Data Bases, 6(3):173–190, August 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/c/Cobb:Edward_
E=.html;   http://link.springer.
de/link/service/journals/00778/
bibs/7006003/70060173.htm. Elec-

tronic edition.

Steinbrunn:1997:HRO

[102] Michael Steinbrunn, Guido Moerkotte, and Alfons Kemper. Heuristic and randomized optimization for the join ordering problem. VLDB Journal: Very Large Data Bases, 6(3): 191–208, August 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/k/Kemper:
Alfons.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/m/Moerkotte:Guido.html;
http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/s/Steinbrunn:
Michael.html;   http://link.springer.
de/link/service/journals/00778/
bibs/7006003/70060191.htm. Elec-

tronic edition.

Panagos:1997:SRC

[103] Euthimios Panagos and Alexandros Biliris. Synchronization and recovery in a client-server storage system. VLDB Journal: Very Large Data Bases, 6(3):209–223, August 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/

db/indices/a-tree/b/Biliris:Alexandros.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
p/Panagos:Euthimios.html;   http:
//link.springer.de/link/service/
journals/00778/bibs/7006003/70060209.

htm. Electronic edition.

Lomet:1997:CRI

[104] David B. Lomet and Betty Salzberg. Concurrency and recovery for index trees. VLDB Journal: Very Large Data Bases, 6(3):224–240, August 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/l/Lomet:David_
B=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/s/Salzberg:Betty.html;
http://link.springer.de/link/service/
journals/00778/bibs/7006003/70060224.

htm. Electronic edition.

Haas:1997:STA

[105] Laura M. Haas, Michael J. Carey, Miron Livny, and Amit Shukla. Seeking the truth about ad hoc join costs. VLDB Journal: Very Large Data Bases, 6(3):241–256, August 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/c/Carey:Michael_
J=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/h/Haas:Laura_M=.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/l/Livny:
Miron.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/s/Shukla:Amit.html;   http:

                                        //link.springer.de/link/service/
                                        journals/00778/bibs/7006003/70060241.

htm. Electronic edition.

Papazoglou:1997:EDM

[106] Mike P. Papazoglou and Bernd J. Krämer. Erratum — A database model for object dynamics. VLDB Journal: Very Large Data Bases, 6(3):257– 260, August 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/k/Kr=auml=mer:
                                        Bernd_J=.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/p/Papazoglou:Mike_P=.html;
                                         http://link.springer.de/link/service/
                                        journals/00778/bibs/7006003/70060257.

htm. Electronic edition. See [95].

Fahl:1997:QPO

[107] Gustav Fahl and Tore Risch. Query processing over object views of relational data. VLDB Journal: Very Large Data Bases, 6(4):261–281, November 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/f/Fahl:Gustav.
                                        html;   http://ftp.informatik.rwth-
                                        aachen.de/dblp/db/indices/a-tree/
                                        r/Risch:Tore.html; http://link.
                                        springer.de/link/service/journals/
                                        00778/bibs/7006004/70060261.htm;
                                         http://link.springer.de/link/service/
                                        journals/00778/papers/7006004/70060261.

pdf. Electronic edition.

Diaz:1997:EEA

[108] Oscar Díaz and Arturo Jaime. EX-ACT: An extensible approach to active object-oriented databases. VLDB Journal: Very Large Data Bases, 6 (4):282–295, November 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/d/D=iacute=
az:Oscar.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/j/Jaime:Arturo.html;   http:
//link.springer.de/link/service/
journals/00778/bibs/7006004/70060282.
htm; http://link.springer.de/link/
service/journals/00778/papers/7006004/

70060282.pdf. Electronic edition.

Bohm:1997:SDS

[109] Klemens Böhm, Karl Aberer, Erich J. Neuhold, and Xiaoya Yang. Structured document storage and refined declarative and navigational access mechanisms in HyperStorM. VLDB Journal: Very Large Data Bases, 6(4):296–311, November 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/a/Aberer:Karl.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
b/B=ouml=hm:Klemens.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/n/Neuhold:
Erich_J=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/y/Yang:Xiaoya.html;   http:
//link.springer.de/link/service/
journals/00778/bibs/7006004/70060296.
htm; http://link.springer.de/link/
service/journals/00778/papers/7006004/

70060296.pdf. Electronic edition.

Muck:1997:CTH

[110] Thomas A. Mück and Martin L. Polaschek. A configurable type hierarchy index for OODB. VLDB Journal: Very Large Data Bases, 6(4): 312–332, November 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/m/M=uuml=
ck:Thomas_A=.html;   http://ftp.
informatik.rwth-aachen.de/dblp/
db/indices/a-tree/p/Polaschek:Martin_

L=.html. Electronic edition.

Mueck:1997:CTH

[111] Thomas A. Mueck and Martin L. Polaschek. A configurable type hierarchy index for OODB. VLDB Journal: Very Large Data Bases, 6(4): 312–332, November 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.de/link/service/
journals/00778/bibs/7006004/70060312.
htm;   http://link.springer.de/
link/service/journals/00778/papers/
7006004/70060312.pdf.

Berchtold:1997:UEF

[112] Stefan Berchtold, Daniel A. Keim, and Hans-Peter Kriegel. Using extended feature objects for partial similarity retrieval. VLDB Journal: Very Large Data Bases, 6(4):333–348, November 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/b/Berchtold:Stefan.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
k/Keim:Daniel_A=.html;   http:/
/ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/k/Kriegel:
Hans=Peter.html;   http://link.

springer.de/link/service/journals/
00778/bibs/7006004/70060333.htm;
http://link.springer.de/link/service/
journals/00778/papers/7006004/70060333.

pdf. Electronic edition.

Han:1998:ORQ

[113] Jia Liang Han. Optimizing relational queries in connection hypergraphs: Nested queries, views, and binding propagations. VLDB Journal: Very Large Data Bases, 7(1):1– 11, February 1998. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/h/Han:Jia_Liang.
html;   http://link.springer.de/
link/service/journals/00778/bibs/
8007001/80070001.htm; http://link.
springer.de/link/service/journals/
00778/papers/8007001/80070001.pdf.

Electronic edition.

Hanson:1998:FRC

[114] Eric N. Hanson, I.-Cheng Chen, Roxana Dastur, Kurt Engel, Vijay Ramaswamy, Wendy Tan, and Chun Xu. A flexible and recoverable client/server database event notification system. VLDB Journal: Very Large Data Bases, 7(1): 12–24, February 1998. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/c/Chen:I=
=Cheng.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/d/Dastur:Roxana.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/e/Engel:
Kurt.html; http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/

                                        a-tree/h/Hanson:Eric_N=.html;
                                        http://ftp.informatik.rwth-aachen.
                                        de/dblp/db/indices/a-tree/r/Ramaswamy:
                                        Vijay.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/t/Tan:Wendy.html;   http:
                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/x/Xu:Chun.
                                        html;   http://link.springer.de/
                                        link/service/journals/00778/bibs/
                                        8007001/80070012.htm; http://link.
                                        springer.de/link/service/journals/
                                        00778/papers/8007001/80070012.pdf.

Electronic edition.

Mehta:1998:OPM

[115] Ashish Mehta, James Geller, Yehoshua Perl, and Erich J. Neuhold. The OODB path-method generator (PMG) using access weights and precomputed access relevance. VLDB Journal: Very Large Data Bases, 7(1): 25–47, February 1998. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/g/Geller:
                                        James.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/m/Mehta:Ashish.html;   http:
                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/n/Neuhold:
                                        Erich_J=.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/p/Perl:Yehoshua.html;   http:
                                        //link.springer.de/link/service/
                                        journals/00778/bibs/8007001/80070025.
                                        htm; http://link.springer.de/link/
                                        service/journals/00778/papers/8007001/

80070025.pdf. Electronic edition.

Scheuermann:1998:DPL

[116] Peter Scheuermann, Gerhard Weikum, and Peter Zabback. Data partitioning and load balancing in parallel disk systems. VLDB Journal: Very Large Data Bases, 7(1):48–66, February 1998. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/s/Scheuermann:
Peter.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/w/Weikum:Gerhard.html;
http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/z/Zabback:
Peter.html;   http://link.springer.
de/link/service/journals/00778/
bibs/8007001/80070048.htm; http:
//link.springer.de/link/service/
journals/00778/papers/8007001/80070048.

pdf. Electronic edition.

Ishakbeyoglu:1998:MII

[117] Naci S. Ishakbeyoğlu and Z. Meral Özsoyoğlu. Maintenance of implication integrity constraints under updates to constraints. VLDB Journal: Very Large Data Bases, 7 (2):67–78, May 1998. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/=/=Ouml=
zsoyoglu:Z=_Meral.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/i/Ishakbeyoglu:
Naci.html;   http://link.springer.
de/link/service/journals/00778/
bibs/8007002/80070067.htm; http:
//link.springer.de/link/service/
journals/00778/papers/8007002/80070067.

pdf. Electronic edition.

Dessloch:1998:ADP

[118] Stefan Deßloch, Theo Härder, Nelson Mendonça Mattos, Bernhard Mitschang, and Joachim Thomas. Advanced data processing in KRISYS: Modeling concepts, implementation techniques, and client/server issues. VLDB Journal: Very Large Data Bases, 7(2):79–95, May 1998. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/d/De=szlig=
loch:Stefan.html;   http://ftp.
informatik.rwth-aachen.de/dblp/
db/indices/a-tree/h/H=auml=rder:
Theo.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/m/Mattos:Nelson_Mendon=ccedil=
a.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/m/Mitschang:Bernhard.html;
 http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/t/Thomas:
Joachim.html;   http://link.springer.
de/link/service/journals/00778/
bibs/8007002/80070079.htm; http:
//link.springer.de/link/service/
journals/00778/papers/8007002/80070079.

pdf. Electronic edition.

Abiteboul:1998:LVS

[119] Serge Abiteboul, Sophie Cluet, and Tova Milo. A logical view of structured files. VLDB Journal: Very Large Data Bases, 7(2):96– 114, May 1998. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/a/Abiteboul:Serge.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
c/Cluet:Sophie.html;   http://ftp.
informatik.rwth-aachen.de/dblp/

db/indices/a-tree/m/Milo:Tova.html;
 http://link.springer.de/link/service/
journals/00778/bibs/8007002/80070096.
htm; http://link.springer.de/link/
service/journals/00778/papers/8007002/

80070096.pdf. Electronic edition.

Ooi:1998:FIR

[120] Beng Chin Ooi, Kian-Lee Tan, Tat Seng Chua, and Wynne Hsu. Fast image retrieval using colorspatial information. VLDB Journal: Very Large Data Bases, 7(2):115– 128, May 1998. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/c/Chua:Tat=Seng.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
h/Hsu:Wynne.html;   http://ftp.
informatik.rwth-aachen.de/dblp/
db/indices/a-tree/o/Ooi:Beng_Chin.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
t/Tan:Kian=Lee.html;   http://link.
springer.de/link/service/journals/
00778/bibs/8007002/80070115.htm;
http://link.springer.de/link/service/
journals/00778/papers/8007002/80070115.

pdf. Electronic edition.

Jarke:1998:GE

[121] Matthias Jarke. Guest Editorial. VLDB Journal: Very Large Data Bases, 7(3):129, August 1998. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.de/link/service/
journals/00778/bibs/8007003/80070129.
htm;   http://link.springer.de/
link/service/journals/00778/papers/
8007003/80070129.pdf.

Seshadri:1998:EAD

[122] Praveen Seshadri. Enhanced abstract data types in object-relational databases. VLDB Journal: Very Large Data Bases, 7(3):130–140, August 1998. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/s/Seshadri:Praveen.
                                        html;   http://link.springer.de/
                                        link/service/journals/00778/bibs/
                                        8007003/80070130.htm; http://link.
                                        springer.de/link/service/journals/
                                        00778/papers/8007003/80070130.pdf.

Special Issue on VLDB 1997. Electronic

edition.

Kraiss:1998:IDC

[123] Achim Kraiss and Gerhard Weikum. Integrated document caching and prefetching in storage hierarchies based on Markov-chain predictions. VLDB Journal: Very Large Data Bases, 7 (3):141–162, August 1998. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/k/Kraiss:
                                        Achim.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/w/Weikum:Gerhard.html;
                                        http://link.springer.de/link/service/
                                        journals/00778/bibs/8007003/80070141.
                                        htm; http://link.springer.de/link/
                                        service/journals/00778/papers/8007003/

80070141.pdf. Electronic edition.

Chakrabarti:1998:SFS

[124] Soumen Chakrabarti, Byron Dom, Rakesh Agrawal, and Prabhakar Raghavan. Scalable feature selection, classification and signature generation for organizing large text databases into hierarchical topic taxonomies. VLDB Journal: Very Large Data Bases, 7(3):163–178, August 1998. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/a/Agrawal:Rakesh.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
c/Chakrabarti:Soumen.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/d/Dom:Byron.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
r/Raghavan:Prabhakar.html;   http:
//link.springer.de/link/service/
journals/00778/bibs/8007003/80070163.
htm; http://link.springer.de/link/
service/journals/00778/papers/8007003/

80070163.pdf. Electronic edition.

Roy:1998:GCO

[125] Prasan Roy, S. Seshadri, Abraham Silberschatz, S. Sudarshan, and S. Ashwin. Garbage collection in objectoriented databases using transactional cyclic reference counting. VLDB Journal: Very Large Data Bases, 7 (3):179–193, August 1998. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/a/Ashwin:
S=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/r/Roy:Prasan.html; http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/s/Seshadri:
S=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/s/Silberschatz:Abraham.html;

                                         http://ftp.informatik.rwth-aachen.
                                        de/dblp/db/indices/a-tree/s/Sudarshan:
                                        S=.html;   http://link.springer.
                                        de/link/service/journals/00778/
                                        bibs/8007003/80070179.htm; http:
                                        //link.springer.de/link/service/
                                        journals/00778/papers/8007003/80070179.

pdf. Electronic edition.

Ng:1998:IRM

[126] Wee Teck Ng and Peter M. Chen. Integrating reliable memory in databases. VLDB Journal: Very Large Data Bases, 7(3):194–204, August 1998. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

                                        URL http://ftp.informatik.rwth-
                                        aachen.de/dblp/db/indices/a-tree/
                                        c/Chen:Peter_M=.html;   http://
                                        ftp.informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/n/Ng:Wee_Teck.
                                        html;   http://link.springer.de/
                                        link/service/journals/00778/bibs/
                                        8007003/80070194.htm; http://link.
                                        springer.de/link/service/journals/
                                        00778/papers/8007003/80070194.pdf.

Electronic edition.

Ozsu:1998:I

[127] M. Tamer Özsu and Stavros Christodoulakis. Introduction. VLDB Journal: Very Large Data Bases, 7(4):205, December 1998. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ozsu:1998:SIM

[128] M. Tamer Özsu and Stavros Christodoulakis. Special issue on multimedia databases: Introduction. VLDB Journal: Very Large Data Bases, 7(4):205, December 1998. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/=/=Ouml=zsu:M=
_Tamer.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/c/Christodoulakis:Stavros.
html;   http://link.springer.de/
link/service/journals/00778/bibs/
8007004/80070205.htm; http://link.
springer.de/link/service/journals/
00778/papers/8007004/80070205.pdf.

Electronic edition.

Garofalakis:1998:PRS

[129] Minos N. Garofalakis, Banu Özden, and Avi Silberschatz. On periodic resource scheduling for continuousmedia databases. VLDB Journal: Very Large Data Bases, 7(4):206–225, December 1998. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/=/=Ouml=zden:
Banu.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/g/Garofalakis:Minos_N=.html;
http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/s/Silberschatz:
Abraham.html;   http://link.springer.
de/link/service/journals/00778/
bibs/8007004/80070206.htm; http:
//link.springer.de/link/service/
journals/00778/papers/8007004/80070206.

pdf. Electronic edition.

Jiang:1998:STC

[130] Haitao Jiang and Ahmed K. Elmagarmid. Spatial and temporal content-based access to hypervideo databases. VLDB Journal: Very Large Data Bases, 7(4):226–238, December 1998. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/e/Elmagarmid:
Ahmed_K=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/j/Jiang:Haitao.html;   http:
//link.springer.de/link/service/
journals/00778/bibs/8007004/80070226.
htm; http://link.springer.de/link/
service/journals/00778/papers/8007004/

80070226.pdf. Electronic edition.

Ng:1998:OCO

[131] Raymond T. Ng and Paul Shum. Optimal clip ordering for multi-clip queries. VLDB Journal: Very Large Data Bases, 7(4):239–252, December 1998. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/n/Ng:Raymond_
T=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/s/Shum:Paul.html;   http:
//link.springer.de/link/service/
journals/00778/bibs/8007004/80070239.
htm; http://link.springer.de/link/
service/journals/00778/papers/8007004/

80070239.pdf. Electronic edition.

Soffer:1998:ISI

[132] Aya Soffer and Hanan Samet. Integrating symbolic images into a multimedia database system using classification and abstraction approaches. VLDB Journal: Very Large Data Bases, 7(4):253–274, December 1998. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/s/Samet:Hanan.

html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
s/Soffer:Aya.html; http://link.
springer.de/link/service/journals/
00778/bibs/8007004/80070253.htm;
 http://link.springer.de/link/service/
journals/00778/papers/8007004/80070253.

pdf. Electronic edition.

Zezula:1998:ASR

[133] Pavel Zezula, Pasquale Savino, Giuseppe Amato, and Fausto Rabitti. Approximate similarity retrieval with M-trees. VLDB Journal: Very Large Data Bases, 7(4):275–293, December 1998. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/a/Amato:Giuseppe.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
r/Rabitti:Fausto.html;   http:/
/ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/s/Savino:
Pasquale.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/z/Zezula:Pavel.html;   http:
//link.springer.de/link/service/
journals/00778/bibs/8007004/80070275.
htm; http://link.springer.de/link/
service/journals/00778/papers/8007004/

80070275.pdf. Electronic edition.

Balkir:1998:DPM

[134] Nevzat Hurkan Balkir and Gultekin Özsoyoglu. Delivering presentations from multimedia servers. VLDB Journal: Very Large Data Bases, 7(4): 294–307, December 1998. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/

                                        dblp/db/indices/a-tree/=/=Ouml=
                                        zsoyoglu:Gultekin.html;   http:
                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/b/Balkir:
                                        Nevzat_Hurkan.html;   http://link.
                                        springer.de/link/service/journals/
                                        00778/bibs/8007004/80070294.htm;
                                        http://link.springer.de/link/service/
                                        journals/00778/papers/8007004/80070294.

pdf. Electronic edition.

Li:1999:FJU

[135] Zhe Li and Kenneth A. Ross. Fast joins using join indices. VLDB Journal: Very Large Data Bases, 8(1):1– 24, April 1999. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/l/Li:Zhe.html;
                                         http://ftp.informatik.rwth-aachen.
                                        de/dblp/db/indices/a-tree/r/Ross:
                                        Kenneth_A=.html;   http://link.
                                        springer.de/link/service/journals/
                                        00778/bibs/9008001/90080001.htm;
                                        http://link.springer.de/link/service/
                                        journals/00778/papers/9008001/90080001.

pdf. Electronic edition.

Harder:1999:IPS

[136] Theo Härder, Günter Sauter, and Joachim Thomas. The intrinsic problems of structural heterogeneity and an approach to their solution. VLDB Journal: Very Large Data Bases, 8(1): 25–43, April 1999. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/h/H=auml=rder:
                                        Theo.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/s/Sauter:G=uuml=nter.html;

http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/t/Thomas:
Joachim.html;   http://link.springer.
de/link/service/journals/00778/
bibs/9008001/90080025.htm; http:
//link.springer.de/link/service/
journals/00778/papers/9008001/90080025.

pdf. Electronic edition.

Huang:1999:CTP

[137] Yueh-Min Huang, Jen-Wen Ding, and Shiao-Li Tsao. Constant time permutation: An efficient block allocation strategy for variable-bit-rate continuous media data. VLDB Journal: Very Large Data Bases, 8(1):44– 54, April 1999. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/d/Ding:Jen=Wen.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
h/Huang:Yueh=Min.html;   http:/
/ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/t/Tsao:Shiao=
Li.html;   http://link.springer.
de/link/service/journals/00778/
bibs/9008001/90080044.htm; http:
//link.springer.de/link/service/
journals/00778/papers/9008001/90080044.

pdf. Electronic edition.

Kabra:1999:OOO

[138] Navin Kabra and David J. DeWitt. OPT++: an object-oriented implementation for extensible database query optimization. VLDB Journal: Very Large Data Bases, 8 (1):55–78, April 1999. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/

dblp/db/indices/a-tree/d/DeWitt:
David_J=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/k/Kabra:Navin.html;   http:
//link.springer.de/link/service/
journals/00778/bibs/9008001/90080055.
htm; http://link.springer.de/link/
service/journals/00778/papers/9008001/

90080055.pdf. Electronic edition.

Krivokapic:1999:DDD

[139] Natalija Krivokapić, Alfons Kemper, and Ehud Gudes. Deadlock detection in distributed database systems: a new algorithm and a comparative performance analysis. VLDB Journal: Very Large Data Bases, 8(2):79– 100, April 1999. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/g/Gudes:Ehud.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
k/Kemper:Alfons.html;   http://
ftp.informatik.rwth-aachen.de/dblp/
db/indices/a-tree/k/Krivokapic:
Natalija.html;   http://link.springer.
de/link/service/journals/00778/
bibs/9008002/90080079.htm; http:
//link.springer.de/link/service/
journals/00778/papers/9008002/90080079.

pdf. Electronic edition.

Boncz:1999:MPQ

[140] Peter A. Boncz and Martin L. Kersten. MIL primitives for querying a fragmented world. VLDB Journal: Very Large Data Bases, 8(2):101– 119, April 1999. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/

db/indices/a-tree/b/Boncz:Peter_
A=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/k/Kersten:Martin_L=.html;
 http://link.springer.de/link/service/
journals/00778/bibs/9008002/90080101.
htm; http://link.springer.de/link/
service/journals/00778/papers/9008002/

90080101.pdf. Electronic edition.

Aslan:1999:SHR

[141] Goksel Aslan and Dennis McLeod. Semantic heterogeneity resolution in federated databases by metadata implantation and stepwise evolution. VLDB Journal: Very Large Data Bases, 8 (2):120–132, April 1999. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/a/Aslan:
Goksel.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/m/McLeod:Dennis.html;   http:
//link.springer.de/link/service/
journals/00778/bibs/9008002/90080120.
htm; http://link.springer.de/link/
service/journals/00778/papers/9008002/

90080120.pdf. Electronic edition.

Law:1999:ESI

[142] Kelvin K. W. Law, John C. S. Lui, and Leana Golubchik. Efficient support for interactive service in multiresolution VOD systems. VLDB Journal: Very Large Data Bases, 8(2):133– 153, April 1999. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/g/Golubchik:Leana.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/

                                        l/Law:Kelvin_K=_W=.html;   http:
                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/l/Lui:John_
                                        C=_S=.html;   http://link.springer.
                                        de/link/service/journals/00778/
                                        bibs/9008002/90080133.htm; http:
                                        //link.springer.de/link/service/
                                        journals/00778/papers/9008002/90080133.

pdf. Electronic edition.

Shmueli:2000:FVP

[143] O. Shmueli and J. Widom. Foreword by the VLDB ‘98 PC Chairmen. VLDB Journal: Very Large Data Bases, 8 (3–4):155, February 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Widom:2000:BPV

[144] Jennifer Widom and Oded Shmueli. Best papers of VLDB ‘98, New York: Foreword by the VLDB ‘98 PC Chairmen: Best Papers of VLDB ‘98. VLDB Journal: Very Large Data Bases, 8(3–4):155, February 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/s/Shmueli:Oded.
                                        html;   http://ftp.informatik.rwth-
                                        aachen.de/dblp/db/indices/a-tree/
                                        w/Widom:Jennifer.html;   http:/
                                        /link.springer.de/link/service/
                                        journals/00778/bibs/0008003/00080155.
                                        htm; http://link.springer.de/link/
                                        service/journals/00778/papers/0008003/

00080155.pdf. Electronic edition.

Braumandl:2000:FJP

[145] Reinhard Braumandl, Jens Claußen, Alfons Kemper, and Donald Kossmann. Functional-join processing. VLDB Journal: Very Large Data Bases, 8(3–4):156–177, February 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/b/Braumandl:Reinhard.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
c/Clau=szlig=en:Jens.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/k/Kemper:
Alfons.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/k/Kossmann:Donald.html;
http://link.springer.de/link/service/
journals/00778/bibs/0008003/00080156.
htm; http://link.springer.de/link/
service/journals/00778/papers/0008003/

00080156.pdf. Electronic edition.

George:2000:SBF

[146] Binto George and Jayant R. Haritsa. Secure buffering in firm realtime database systems. VLDB Journal: Very Large Data Bases, 8(3– 4):178–198, February 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/g/George:
Binto.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/h/Haritsa:Jayant_R=.html;
http://link.springer.de/link/service/
journals/00778/bibs/0008003/00080178.
htm; http://link.springer.de/link/
service/journals/00778/papers/0008003/

00080178.pdf. Electronic edition.

Muth:2000:LLS

[147] Peter Muth, Patrick E. O’Neil, Achim Pick, and Gerhard Weikum. The LHAM log-structured history data access method. VLDB Journal: Very Large Data Bases, 8(3–4):199–221, February 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/m/Muth:Peter.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
o/O=Neil:Patrick_E=.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/p/Pick:Achim.
html; http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
w/Weikum:Gerhard.html;   http:/
/link.springer.de/link/service/
journals/00778/bibs/0008003/00080199.
htm; http://link.springer.de/link/
service/journals/00778/papers/0008003/

00080199.pdf. Electronic edition.

Gibson:2000:CCD

[148] David Gibson, Jon M. Kleinberg, and Prabhakar Raghavan. Clustering categorical data: An approach based on dynamical systems. VLDB Journal: Very Large Data Bases, 8(3– 4):222–236, February 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/g/Gibson:
David.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/k/Kleinberg:Jon_M=.html;
 http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/r/Raghavan:
Prabhakar.html;   http://link.springer.
de/link/service/journals/00778/
bibs/0008003/00080222.htm; http:
//link.springer.de/link/service/
journals/00778/papers/0008003/00080222.

pdf. Electronic edition.

Knorr:2000:DBO

[149] Edwin M. Knorr, Raymond T. Ng, and Vladimir Tucakov. Distancebased outliers: Algorithms and applications. VLDB Journal: Very Large Data Bases, 8(3–4):237–253, February 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/k/Knorr:Edwin_
M=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/n/Ng:Raymond_T=.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/t/Tucakov:
V=.html; http://link.springer.de/
link/service/journals/00778/bibs/
0008003/00080237.htm;   http://
link.springer.de/link/service/journals/
00778/papers/0008003/00080237.pdf.

Electronic edition.

Korn:2000:QDM

[150] Flip Korn, Alexandros Labrinidis, Yannis Kotidis, and Christos Faloutsos. Quantifiable data mining using ratio rules. VLDB Journal: Very Large Data Bases, 8(3–4):254–266, February 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/f/Faloutsos:Christos.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
k/Korn:Flip.html;   http://ftp.
informatik.rwth-aachen.de/dblp/
db/indices/a-tree/k/Kotidis:Yannis.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
l/Labrinidis:Alexandros.html;

                                        http://link.springer.de/link/service/
                                        journals/00778/bibs/0008003/00080254.
                                        htm; http://link.springer.de/link/
                                        service/journals/00778/papers/0008003/

00080254.pdf. Electronic edition.

Torp:2000:ETD

[151] Kristian Torp, Christian S. Jensen, and Richard Thomas Snodgrass. Effective timestamping in databases. VLDB Journal: Very Large Data Bases, 8(3– 4):267–288, February 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/j/Jensen:
                                        Christian_S=.html;   http://ftp.
                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/s/Snodgrass:Richard_
                                        T=.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/t/Torp:Kristian.html;   http:
                                        //link.springer.de/link/service/
                                        journals/00778/bibs/0008003/00080267.
                                        htm; http://link.springer.de/link/
                                        service/journals/00778/papers/0008003/

00080267.pdf. Electronic edition.

Sheikholeslami:2000:WWB

[152] Gholamhosein Sheikholeslami, Surojit Chatterjee, and Aidong Zhang. WaveCluster: a wavelet based clustering approach for spatial data in very large databases. VLDB Journal: Very Large Data Bases, 8(3–4):289–304, February 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/c/Chatterjee:
                                        Surojit.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/s/Sheikholeslami:Gholamhosein.

html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
z/Zhang:Aidong.html;   http://link.
springer.de/link/service/journals/
00778/bibs/0008003/00080289.htm;
http://link.springer.de/link/service/
journals/00778/papers/0008003/00080289.

pdf. Electronic edition.

Pacitti:2000:UPS

[153] Esther Pacitti and Eric Simon. Update propagation strategies to improve freshness in lazy master replicated databases. VLDB Journal: Very Large Data Bases, 8(3–4):305–318, February 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/p/Pacitti:Esther.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
s/Simon:Eric.html; http://link.
springer.de/link/service/journals/
00778/bibs/0008003/00080305.htm;
 http://link.springer.de/link/service/
journals/00778/papers/0008003/00080305.

pdf. Electronic edition.

Liang:2000:OMD

[154] Weifa Liang, Maria E. Orlowska, and Jeffrey X. Yu. Optimizing multiple dimensional queries simultaneously in multidimensional databases. VLDB Journal: Very Large Data Bases, 8(3–4):319–338, February 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/l/Liang:Weifa.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/

                                        o/Orlowska:Maria_E=.html;   http:
                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/y/Yu:Jeffrey_
                                        X=.html;   http://link.springer.
                                        de/link/service/journals/00778/
                                        bibs/0008003/00080319.htm; http:
                                        //link.springer.de/link/service/
                                        journals/00778/papers/0008003/00080319.

pdf. Electronic edition.

Atzeni:2000:DWG

[155] Paolo Atzeni and Alberto O. Mendelzon. Databases and the Web: Guest editorial: Databases and the Web. VLDB Journal: Very Large Data Bases, 9(1):1, March 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/a/Atzeni:
                                        Paolo.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/m/Mendelzon:Alberto_O=.html;
                                         http://link.springer.de/link/service/
                                        journals/00778/bibs/0009001/00090001.
                                        htm; http://link.springer.de/link/
                                        service/journals/00778/papers/0009001/

00090001.pdf. Electronic edition.

Atzeni:2000:GE

[156] Paolo Atzeni and Alberto O. Mendelzon. Guest editorial. VLDB Journal: Very Large Data Bases, 9(1):1, March 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chidlovskii:2000:SCW

[157] Boris Chidlovskii and Uwe M. Borghoff. Semantic caching of Web queries. VLDB Journal: Very Large Data Bases, 9(1):2–17, March 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/b/Borghoff:Uwe_
M=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/c/Chidlovskii:Boris.html;
 http://link.springer.de/link/service/
journals/00778/bibs/0009001/00090002.
htm; http://link.springer.de/link/
service/journals/00778/papers/0009001/

00090002.pdf. Electronic edition.

Gruser:2000:LRT

[158] Jean-Robert Gruser, Louiqa Raschid, Vladimir Zadorozhny, and Tao Zhan. Learning response time for WebSources using query feedback and application in query optimization. VLDB Journal: Very Large Data Bases, 9 (1):18–37, March 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/g/Gruser:
Jean=Robert.html;   http://ftp.
informatik.rwth-aachen.de/dblp/
db/indices/a-tree/r/Raschid:Louiqa.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
z/Zadorozhny:Vladimir.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/z/Zhan:Tao.
html;   http://link.springer.de/
link/service/journals/00778/bibs/
0009001/00090018.htm; http://link.
springer.de/link/service/journals/
00778/papers/0009001/00090018.pdf.

Electronic edition.

Fernandez:2000:DSW

[159] Mary Fernández, Daniela Florescu, Alon Levy, and Dan Suciu. Declarative specification of Web sites with S. VLDB Journal: Very Large Data Bases, 9(1):38–55, March 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

                                        URL http://ftp.informatik.rwth-
                                        aachen.de/dblp/db/indices/a-tree/
                                        f/Fernandez:Mary_F=.html;   http:
                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/f/Florescu:
                                        Daniela.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/l/Levy:Alon_Y=.html;   http:
                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/s/Suciu:
                                        Dan.html;   http://link.springer.
                                        de/link/service/journals/00778/
                                        bibs/0009001/00090038.htm; http:
                                        //link.springer.de/link/service/
                                        journals/00778/papers/0009001/00090038.

pdf. Electronic edition.

Berendt:2000:ANB

[160] Bettina Berendt and Myra Spiliopoulou. Analysis of navigation behaviour in Web sites integrating multiple information systems. VLDB Journal: Very Large Data Bases, 9(1):56–75, March 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/b/Berendt:Bettina.
                                        html;   http://ftp.informatik.rwth-
                                        aachen.de/dblp/db/indices/a-tree/
                                        s/Spiliopoulou:Myra.html;   http:
                                        //link.springer.de/link/service/
                                        journals/00778/bibs/0009001/00090056.
                                        htm; http://link.springer.de/link/
                                        service/journals/00778/papers/0009001/

00090056.pdf. Electronic edition.

Buneman:2000:UQL

[161] Peter Buneman, Mary F. Fernandez, and Dan Suciu. UnQL: a query language and algebra for semistructured data based on structural recursion. VLDB Journal: Very Large Data Bases, 9(1):76–110, March 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/b/Buneman:Peter.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
f/Fernandez:Mary_F=.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/s/Suciu:
Dan.html;   http://link.springer.
de/link/service/journals/00778/
bibs/0009001/00090076.htm; http:
//link.springer.de/link/service/
journals/00778/papers/0009001/00090076.

pdf. Electronic edition.

Mirbel:2000:CTI

[162] Isabelle Mirbel, Barbara Pernici, Timos K. Sellis, S. Tserkezoglou, and Michalis Vazirgiannis. Checking the temporal integrity of interactive multimedia documents. VLDB Journal: Very Large Data Bases, 9 (2):111–130, July 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/m/Mirbel:
Isabelle.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/p/Pernici:Barbara.html;
http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/s/Sellis:
Timos_K=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/t/Tserkezoglou:S=.html;
http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/v/Vazirgiannis:

                                        Michalis.html;   http://link.springer.
                                        de/link/service/journals/00778/
                                        bibs/0009002/00090111.htm; http:
                                        //link.springer.de/link/service/
                                        journals/00778/papers/0009002/00090111.

pdf. Electronic edition.

Candan:2000:VMM

[163] K. Selçuk Candan, Eric Lemar, and V. S. Subrahmanian. View management in multimedia databases. VLDB Journal: Very Large Data Bases, 9 (2):131–153, July 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/c/Candan:
                                        K=_Sel=ccedil=uk.html;   http:/
                                        /ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/l/Lemar:
                                        Eric.html; http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/s/Subrahmanian:V=_S=.html;
                                         http://link.springer.de/link/service/
                                        journals/00778/bibs/0009002/00090131.
                                        htm; http://link.springer.de/link/
                                        service/journals/00778/papers/0009002/

00090131.pdf. Electronic edition.

Fu:2000:DVT

[164] Ada Wai chee Fu, Polly Mei shuen Chan, Yin-Ling Cheung, and Yiu Sang Moon. Dynamic vp-tree indexing for n-nearest neighbor search given pair-wise distances. VLDB Journal: Very Large Data Bases, 9(2):154– 173, July 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/c/Chan:Polly_
                                        Mei=shuen.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/

a-tree/c/Cheung:Yin=Ling.html;
http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/f/Fu:
Ada_Wai=Chee.html;   http://ftp.
informatik.rwth-aachen.de/dblp/
db/indices/a-tree/m/Moon:Yiu_Sang.
html;   http://link.springer.de/
link/service/journals/00778/bibs/
0009002/00090154.htm; http://link.
springer.de/link/service/journals/
00778/papers/0009002/00090154.pdf.

Electronic edition.

Atkinson:2000:GE

[165] Malcolm P. Atkinson. Guest editorial. VLDB Journal: Very Large Data Bases, 9(3):175–176, December 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/0009003/00090175.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/0009003/00090175.
pdf.

Bernstein:2000:CBP

[166] Philip A. Bernstein, Shankar Pal, and David Shutt. Context-based prefetch

— an optimization for implementing objects on relations. VLDB Journal: Very Large Data Bases, 9(3): 177–189, December 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.de/link/service/
journals/00778/bibs/0009003/00090177.
htm;   http://link.springer.de/
link/service/journals/00778/papers/
0009003/00090177.pdf.

Claussen:2000:EES

[167] J. Claussen, A. Kemper, D. Kossmann, and C. Wiesner. Exploiting early sorting and early partitioning for decision support query processing. VLDB Journal: Very Large Data Bases, 9(3):190–213, December 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/0009003/00090190.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/0009003/00090190.
pdf.

Jagadish:2000:ODM

[168] H. V. Jagadish, Olga Kapitskaia, Raymond T. Ng, and Divesh Srivastava. One-dimensional and multidimensional substring selectivity estimation. VLDB Journal: Very Large Data Bases, 9(3):214–230, December 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/0009003/00090214.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/0009003/00090214.
pdf.

Manegold:2000:ODA

[169] Stefan Manegold, Peter A. Boncz, and Martin L. Kersten. Optimizing database architecture for the new bottleneck: memory access. VLDB Journal: Very Large Data Bases, 9(3):231–246, December 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/0009003/00090231.htm;   http:
//link.springer.de/link/service/

journals/00778/papers/0009003/00090231.
pdf.

Raman:2000:ODR

[170] Vijayshankar Raman, Bhaskaran Raman, and Joseph M. Hellerstein. Online dynamic reordering. VLDB Journal: Very Large Data Bases, 9(3):247–260, December 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/0009003/00090247.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/0009003/00090247.
pdf.

Tan:2000:PEN

[171] Kian-Lee Tan, Cheng Hian Goh, and Beng Chin Ooi. Progressive evaluation of nested aggregate queries. VLDB Journal: Very Large Data Bases, 9(3):261–278, December 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/0009003/00090261.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/0009003/00090261.
pdf.

Ngu:2001:CMV

[172] Anne H. H. Ngu, Quan Z. Sheng, Du Q. Huynh, and Ron Lei. Combining multivisual features for efficient indexing in a large image database. VLDB Journal: Very Large Data Bases, 9(4):279–293, April 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/

                                        bibs/1009004/10090279.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00778/papers/1009004/10090279.
                                        pdf.

Combi:2001:HTD

[173] Carlo Combi and Giuseppe Pozzi. HMAP — a temporal data model managing intervals with different granularities and indeterminacy from natural language sentences. VLDB Journal: Very Large Data Bases, 9(4):294–311, April 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/journals/00778/
                                        bibs/1009004/10090294.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00778/papers/1009004/10090294.
                                        pdf.

Li:2001:SEM

[174] Wen-Syan Li, K. Selçuk Candan, Kyoji Hirata, and Yoshinori Hara. Supporting efficient multimedia database exploration. VLDB Journal: Very Large Data Bases, 9(4):312–326, April 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/journals/00778/
                                        bibs/1009004/10090312.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00778/papers/1009004/10090312.
                                        pdf.

Lee:2001:GTM

[175] Chiang Lee, Chi-Sheng Shih, and Yaw-Huei Chen. A graph-theoretic model for optimizing queries involving methods. VLDB Journal: Very Large Data Bases, 9(4):327–343, April 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/1009004/10090327.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/1009004/10090327.
pdf.

Wang:2001:IVH

[176] Changzhou Wang and X. Sean Wang. Indexing very high-dimensional sparse and quasi-sparse vectors for similarity searches. VLDB Journal: Very Large Data Bases, 9(4):344–361, April 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/1009004/10090344.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/1009004/10090344.
pdf.

Casati:2001:GE

[177] Fabio Casati, Ming-Chien Shan, and Dimitrios Georgakopoulos. Guest editorial. VLDB Journal: Very Large Data Bases, 10(1):1, August 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/1010001/10100001.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/1010001/10100001.
pdf.

Mecella:2001:DWC

[178] Massimo Mecella and Barbara Pernici. Designing wrapper components for e-services in integrating heterogeneous systems. VLDB Journal: Very Large Data Bases, 10(1):2–15, August 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/journals/00778/
                                        bibs/1010001/10100002.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00778/papers/1010001/10100002.
                                        pdf.

Eyal:2001:ICH

[179] Anat Eyal and Tova Milo. Integrating and customizing heterogeneous ecommerce applications. VLDB Journal: Very Large Data Bases, 10 (1):16–38, August 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //link.springer.de/link/service/
                                        journals/00778/bibs/1010001/10100016.
                                        htm;   http://link.springer.de/
                                        link/service/journals/00778/papers/
                                        1010001/10100016.pdf.

Bonifati:2001:ARX

[180] Angela Bonifati, Stefano Ceri, and Stefano Paraboschi. Active rules for XML: a new paradigm for Eservices. VLDB Journal: Very Large Data Bases, 10(1):39–47, August 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/journals/00778/
                                        bibs/1010001/10100039.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00778/papers/1010001/10100039.
                                        pdf.

Braumandl:2001:OUQ

[181] R. Braumandl, M. Keidl, A. Kemper, D. Kossmann, A. Kreutz, S. Seltzsam, and K. Stocker. ObjectGlobe: Ubiquitous query processing on the Internet. VLDB Journal: Very Large Data Bases, 10(1):48–71, August 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/1010001/10100048.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/1010001/10100048.
pdf.

Su:2001:IBN

[182] Stanley Y. W. Su, Chunbo Huang, Joachim Hammer, Yihua Huang, Haifei Li, Liu Wang, Youzhong Liu, Charnyote Pluempitiwiriyawej, Minsoo Lee, and Herman Lam. An Internet-based negotiation server for e-commerce. VLDB Journal: Very Large Data Bases, 10(1):72–90, August 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/1010001/10100072.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/1010001/10100072.
pdf.

Shegalov:2001:XEW

[183] German Shegalov, Michael Gillmann, and Gerhard Weikum. XMLenabled workflow management for eservices across heterogeneous platforms. VLDB Journal: Very Large Data Bases, 10(1):91–103, August 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/1010001/10100091.htm;   http:
//link.springer.de/link/service/

                                        journals/00778/papers/1010001/10100091.
                                        pdf.

Datta:2001:ASS

[184] Anindya Datta, Kaushik Dutta, Debra VanderMeer, Krithi Ramamritham, and Shamkant B. Navathe. An architecture to support scalable online personalization on the Web. VLDB Journal: Very Large Data Bases, 10(1):104–117, August 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/journals/00778/
                                        bibs/1010001/10100104.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00778/papers/1010001/10100104.
                                        pdf.

ElAbbadi:2001:GE

[185] Amr El Abbadi, Gunter Schlageter, and Kyu-Young Whang. Guest editorial. VLDB Journal: Very Large Data Bases, 10(2–3):119, September 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/journals/00778/
                                        bibs/1010002/10100119.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00778/papers/1010002/10100119.
                                        pdf.

Pucheral:2001:PSD

[186] Philippe Pucheral, Luc Bouganim, Patrick Valduriez, and Christophe Bobineau. PicoDBMS: Scaling down database techniques for the smartcard. VLDB Journal: Very Large Data Bases, 10(2–3):120–132, September 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/1010002/10100120.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/1010002/10100120.
pdf.

Shanmugasundaram:2001:EPR

[187] Jayavel Shanmugasundaram, Eugene Shekita, Rimon Barr, Michael Carey, Bruce Lindsay, Hamid Pirahesh, and Berthold Reinwald. Efficiently publishing relational data as XML documents. VLDB Journal: Very Large Data Bases, 10(2–3):133–154, September 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/1010002/10100133.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/1010002/10100133.
pdf.

Chang:2001:AQM

[188] Kevin Chen-Chuan Chang and Héctor García-Molina. Approximate query mapping: Accounting for translation closeness. VLDB Journal: Very Large Data Bases, 10(2–3):155–181, September 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/1010002/10100155.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/1010002/10100155.
pdf.

Pottinger:2001:MSA

[189] Rachel Pottinger and Alon Halevy. MiniCon: a scalable algorithm for answering queries using views.

VLDB Journal: Very Large Data Bases, 10(2–3):182–198, September 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/journals/00778/
                                        bibs/1010002/10100182.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00778/papers/1010002/10100182.
                                        pdf.

Chakrabarti:2001:AQP

[190] Kaushik Chakrabarti, Minos Garofalakis, Rajeev Rastogi, and Kyuseok Shim. Approximate query processing using wavelets. VLDB Journal: Very Large Data Bases, 10(2–3): 199–223, September 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //link.springer.de/link/service/
                                        journals/00778/bibs/1010002/10100199.
                                        htm;   http://link.springer.de/
                                        link/service/journals/00778/papers/
                                        1010002/10100199.pdf.

Sarawagi:2001:UCM

[191] Sunita Sarawagi. User-cognizant multidimensional analysis. VLDB Journal: Very Large Data Bases, 10(2–3): 224–239, September 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //link.springer.de/link/service/
                                        journals/00778/bibs/1010002/10100224.
                                        htm;   http://link.springer.de/
                                        link/service/journals/00778/papers/
                                        1010002/10100224.pdf.

Turker:2001:SIS

[192] Can Türker and Michael Gertz. Semantic integrity support in SQL:1999 and commercial (object-)relational database management systems. VLDB Journal: Very Large Data Bases, 10 (4):241–269, December 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.de/link/service/
journals/00778/bibs/1010004/10100241.
htm;   http://link.springer.de/
link/service/journals/00778/papers/
1010004/10100241.pdf.

Halevy:2001:AQU

[193] Alon Y. Halevy. Answering queries using views: a survey. VLDB Journal: Very Large Data Bases, 10(4): 270–294, December 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.de/link/service/
journals/00778/bibs/1010004/10100270.
htm;   http://link.springer.de/
link/service/journals/00778/papers/
1010004/10100270.pdf.

Laurent:2001:MCI

[194] D. Laurent, J. Lechtenbörger, N. Spyratos, and G. Vossen. Monotonic complements for independent data warehouses. VLDB Journal: Very Large Data Bases, 10(4):295–315, December 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/1010004/10100295.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/1010004/10100295.
pdf.

Grefen:2001:GTS

[195] Paul Grefen, Jochem Vonk, and Peter Apers. Global transaction support for workflow management systems: from formal specification to practical implementation. VLDB Journal: Very Large Data Bases, 10(4):316–333, December 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/journals/00778/
                                        bibs/1010004/10100316.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00778/papers/1010004/10100316.
                                        pdf.

Rahm:2001:SAA

[196] Erhard Rahm and Philip A. Bernstein. A survey of approaches to automatic schema matching. VLDB Journal: Very Large Data Bases, 10(4): 334–350, December 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //link.springer.de/link/service/
                                        journals/00778/bibs/1010004/10100334.
                                        htm;   http://link.springer.de/
                                        link/service/journals/00778/papers/
                                        1010004/10100334.pdf.

Saltenis:2002:INR

[197] Simonas Šaltenis and Christian S. Jensen. Indexing of now-relative spatio-bitemporal data. VLDB Journal: Very Large Data Bases, 11 (1):1–16, August 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //link.springer.de/link/service/
                                        journals/00778/bibs/2011001/20110001.
                                        htm;   http://link.springer.de/
                                        link/service/journals/00778/papers/
                                        2011001/20110001.pdf.

Rafiei:2002:ERS

[198] Davood Rafiei and Alberto O. Mendelzon. Efficient retrieval of similar shapes. VLDB Journal: Very Large Data Bases, 11(1):17–27, August 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/2011001/20110017.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/2011001/20110017.
pdf.

Navarro:2002:SMS

[199] Gonzalo Navarro. Searching in metric spaces by spatial approximation. VLDB Journal: Very Large Data Bases, 11(1):28–46, August 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/2011001/20110028.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/2011001/20110028.
pdf.

Mihaila:2002:LAD

[200] George A. Mihaila, Louiqa Raschid, and Anthony Tomasic. Locating and accessing data repositories with Web-Semantics. VLDB Journal: Very Large Data Bases, 11(1):47–57, August 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/2011001/20110047.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/2011001/20110047.
pdf.

Ferrari:2002:ASD

[201] E. Ferrari, N. R. Adam, V. Atluri, E. Bertino, and U. Capuozzo. An authorization system for digital libraries. VLDB Journal: Very Large Data Bases, 11(1):58–67, August 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/2011001/20110058.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/2011001/20110058.
pdf.

Marathe:2002:QPT

[202] Arunprasad P. Marathe and Kenneth Salem. Query processing techniques for arrays. VLDB Journal: Very Large Data Bases, 11(1):68–91, August 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/2011001/20110068.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/2011001/20110068.
pdf.

Sakurai:2002:SIH

[203] Yasushi Sakurai, Masatoshi Yoshikawa, Shunsuke Uemura, and Haruhiko Kojima. Spatial indexing of highdimensional data based on relative approximation. VLDB Journal: Very Large Data Bases, 11(2):93–108, October 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/2011002/20110093.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/2011002/20110093.
pdf.

Hjaltason:2002:SCP

[204] Gisli R. Hjaltason and Hanan Samet. Speeding up construction of PMR quadtree-based spatial indexes. VLDB Journal: Very Large Data Bases, 11 (2):109–137, October 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.de/link/service/
journals/00778/bibs/2011002/20110109.
htm;   http://link.springer.de/
link/service/journals/00778/papers/
2011002/20110109.pdf.

Nanopoulos:2002:ESS

[205] Alexandros Nanopoulos and Yannis Manolopoulos. Efficient similarity search for market basket data. VLDB Journal: Very Large Data Bases, 11(2):138–152, October 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/2011002/20110138.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/2011002/20110138.
pdf.

Feng:2002:TMM

[206] Ling Feng, Jeffrey Xu Yu, Hongjun Lu, and Jiawei Han. A template model for multidimensional inter-transactional association rules. VLDB Journal: Very Large Data Bases, 11(2):153–175, October 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/2011002/20110153.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/2011002/20110153.
pdf.

Apers:2002:E

[207] Peter Apers, Stefano Ceri, and Richard Snodgrass. Editorial. VLDB Journal: Very Large Data Bases, 11(3): 177–178, November 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //link.springer.de/link/service/
                                        journals/00778/bibs/2011003/20110177.
                                        htm;   http://link.springer.de/
                                        link/service/journals/00778/papers/
                                        2011003/20110177.pdf. Special issue

VLDB best papers 2001.

An:2002:EPT

[208] Ning An, Sudhanva Gurumurthi, Anand Sivasubramaniam, Narayanan Vijaykrishnan, Mahmut Kandemir, and Mary Jane Irwin. Energyperformance trade-offs for spatial access methods on memory-resident data. VLDB Journal: Very Large Data Bases, 11(3):179–197, November 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/journals/00778/
                                        bibs/2011003/20110179.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00778/papers/2011003/20110179.

pdf. Special issue VLDB best papers

2001.

Ailamaki:2002:DPL

[209] Anastassia Ailamaki, David J. DeWitt, and Mark D. Hill. Data page layouts for relational databases on deep memory hierarchies. VLDB Journal: Very Large Data Bases, 11(3): 198–215, November 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //link.springer.de/link/service/

journals/00778/bibs/2011003/20110198.
htm;   http://link.springer.de/
link/service/journals/00778/papers/
2011003/20110198.pdf. Special issue

VLDB best papers 2001.

Chirkova:2002:FPV

[210] Rada Chirkova, Alon Y. Halevy, and Dan Suciu. A formal perspective on the view selection problem. VLDB Journal: Very Large Data Bases, 11(3):216–237, November 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/2011003/20110216.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/2011003/20110216.

pdf. Special issue VLDB best papers 2001.

Aguilera:2002:VLS

[211] Vincent Aguilera, Sophie Cluet, Tova Milo, Pierangelo Veltri, and Dan Vodislav. Views in a large-scale XML repository. VLDB Journal: Very Large Data Bases, 11(3):238–255, November 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/2011003/20110238.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/2011003/20110238.

pdf. Special issue VLDB best papers 2001.

Hunt:2002:DIL

[212] Ela Hunt, Malcolm P. Atkinson, and Robert W. Irving. Database indexing for large DNA and protein sequence collections. VLDB Journal: Very Large Data Bases, 11(3):256–271, November 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/2011003/20110256.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/2011003/20110256.

pdf. Special issue VLDB best papers 2001.

Halevy:2002:GE

[213] Alon Y. Halevy. Guest editorial. VLDB Journal: Very Large Data Bases, 11(4):273, December 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/2011004/20110273.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/2011004/20110273.
pdf.

Jagadish:2002:TNX

[214] H. V. Jagadish, S. Al-Khalifa, A. Chapman, L. V. S. Lakshmanan, A. Nierman, S. Paparizos, J. M. Patel, D. Srivastava, N. Wiwatwattana, Y. Wu, and C. Yu. TIMBER: a native XML database. VLDB Journal: Very Large Data Bases, 11(4):274–291, December 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/2011004/20110274.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/2011004/20110274.
pdf.

Fiebig:2002:ANX

[215] T. Fiebig, S. Helmer, C.-C. Kanne, G. Moerkotte, J. Neumann, R. Schiele, and T. Westmann. Anatomy of a native XML base management system. VLDB Journal: Very Large Data Bases, 11(4):292–314, December 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/2011004/20110292.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/2011004/20110292.
pdf.

Amer-Yahia:2002:TPQ

[216] S. Amer-Yahia, S. Cho, L. V. S. Lakshmanan, and D. Srivastava. Tree pattern query minimization. VLDB Journal: Very Large Data Bases, 11(4): 315–331, December 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.de/link/service/
journals/00778/bibs/2011004/20110315.
htm;   http://link.springer.de/
link/service/journals/00778/papers/
2011004/20110315.pdf.

Chien:2002:ESM

[217] S.-Y. Chien, V. J. Tsotras, and C. Zaniolo. Efficient schemes for managing multiversion XML documents. VLDB Journal: Very Large Data Bases, 11(4):332–353, December 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/2011004/20110332.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/2011004/20110332.
pdf.

Chan:2002:EFX

[218] C.-Y. Chan, P. Felber, M. Garofalakis, and R. Rastogi. Efficient filtering of XML documents with XPath expressions. VLDB Journal: Very Large Data Bases, 11(4):354–379, December 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/journals/00778/
                                        bibs/2011004/20110354.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00778/papers/2011004/20110354.
                                        pdf.

Ives:2002:XQE

[219] Zachary G. Ives, A. Y. Halevy, and D. S. Weld. An XML query engine for network-bound data. VLDB Journal: Very Large Data Bases, 11(4):380–402, December 2002. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/journals/00778/
                                        bibs/2011004/20110380.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00778/papers/2011004/20110380.
                                        pdf.

Ozsu:2003:NPA

[220] M. Tamer Özsu. New partnership with ACM and update on the journal. VLDB Journal: Very Large Data Bases, 12(1):1, May 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/journals/00778/
                                        bibs/3012001/30120001.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00778/papers/3012001/30120001.
                                        pdf.

Sheth:2003:CRK

[221] A. Sheth, S. Thacker, and S. Patel. Complex relationships and knowledge discovery support in the Info-Quilt system. VLDB Journal: Very Large Data Bases, 12(1):2–27, May 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/3012001/30120002.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/3012001/30120002.
pdf.

Verykios:2003:BDM

[222] V. S. Verykios, G. V. Moustakides, and M. G. Elfeky. A Bayesian decision model for cost optimal record matching. VLDB Journal: Very Large Data Bases, 12(1):28–40, May 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/3012001/30120028.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/3012001/30120028.
pdf.

Cui:2003:LTG

[223] Y. Cui and J. Widom. Lineage tracing for general data warehouse transformations. VLDB Journal: Very Large Data Bases, 12(1):41–58, May 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
de/link/service/journals/00778/
bibs/3012001/30120041.htm;   http:
//link.springer.de/link/service/
journals/00778/papers/3012001/30120041.
pdf.

Medjahed:2003:BBI

[224] B. Medjahed, B. Benatallah, A. Bouguettaya, A. H. H. Ngu, and A. K. Elmagarmid. Business-to-business interactions: issues and enabling technologies. VLDB Journal: Very Large Data Bases, 12(1):59–85, May 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/journals/00778/
                                        bibs/3012001/30120059.htm;   http:
                                        //link.springer.de/link/service/
                                        journals/00778/papers/3012001/30120059.
                                        pdf.

Bernstein:2003:GE

[225] Philip A. Bernstein, Yannis Ioannidis, and Raghu Ramakrishnan. Guest editorial. VLDB Journal: Very Large Data Bases, 12(2):87–88, August 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ramamurthy:2003:CFM

[226] Ravishankar Ramamurthy, David J. DeWitt, and Qi Su. A case for fractured mirrors. VLDB Journal: Very Large Data Bases, 12(2):89–101, August 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chan:2003:RTE

[227] Chee-Yong Chan, Minos Garofalakis, and Rajeev Rastogi. RE-tree: an efficient index structure for regular expressions. VLDB Journal: Very Large Data Bases, 12(2):102–119, August 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Abadi:2003:ANM

[228] Daniel J. Abadi, Don Carney, Ugur Çetintemel, Mitch Cherniack, Christian Convey, Sangdon Lee, Michael Stonebraker, Nesime Tatbul, and Stan Zdonik. Aurora: a new model and architecture for data stream management. VLDB Journal: Very Large Data Bases, 12(2):120–139, August 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chandrasekaran:2003:PSS

[229] Sirish Chandrasekaran and Michael J. Franklin. PSoup: a system for streaming queries over streaming data. VLDB Journal: Very Large Data Bases, 12 (2):140–156, August 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Agrawal:2003:WRD

[230] Rakesh Agrawal, Peter J. Haas, and Jerry Kiernan. Watermarking relational data: framework, algorithms and analysis. VLDB Journal: Very Large Data Bases, 12(2):157–169, August 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chakrabarti:2003:FAT

[231] Soumen Chakrabarti, Shourya Roy, and Mahesh V. Soundalgekar. Fast and accurate text classification via multiple linear discriminant projections. VLDB Journal: Very Large Data Bases, 12 (2):170–185, August 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Fung:2003:CDV

[232] Chi-Wai Fung, Kamalakar Karlapalem, and Qing Li. Cost-driven vertical class partitioning for methods in object oriented databases. VLDB Journal: Very Large Data Bases, 12(3):187–210, October 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Li:2003:CCA

[233] Chen Li. Computing complete answers to queries in the presence of limited access patterns. VLDB Journal: Very Large Data Bases, 12(3):211–227, October 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chua:2003:IBA

[234] Cecil Eng H. Chua, Roger H. L. Chiang, and Ee-Peng Lim. Instance-based attribute identification in database integration. VLDB Journal: Very Large Data Bases, 12(3):228–243, October 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Helmer:2003:PSF

[235] Sven Helmer and Guido Moerkotte. A performance study of four index structures for set-valued attributes of low cardinality. VLDB Journal: Very Large Data Bases, 12(3):244–261, October 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yang:2003:ICM

[236] Jun Yang and Jennifer Widom. Incremental computation and maintenance of temporal aggregates. VLDB Journal: Very Large Data Bases, 12 (3):262–283, October 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Atluri:2003:GE

[237] Vijay Atluri, Anupam Joshi, and Yelena Yesha. Guest editorial. VLDB Journal: Very Large Data Bases, 12 (4):285, November 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Maedche:2003:MMD

[238] A. Maedche, B. Motik, and L. Stojanovic. Managing multiple and distributed ontologies on the Semantic Web. VLDB Journal: Very Large Data Bases, 12(4):286–302, November 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Doan:2003:LMO

[239] AnHai Doan, Jayant Madhavan, Robin Dhamankar, Pedro Domingos, and Alon Halevy. Learning to match ontologies on the Semantic Web. VLDB Journal: Very Large Data Bases, 12 (4):303–319, November 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Halkidi:2003:TOW

[240] Maria Halkidi, Benjamin Nguyen, Iraklis Varlamis, and Michalis Vazirgiannis. THESUS: Organizing Web document collections based on link semantics. VLDB Journal: Very Large Data Bases, 12(4):320–332, November 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Medjahed:2003:CWS

[241] Brahim Medjahed, Athman Bouguettaya, and Ahmed K. Elmagarmid. Composing Web services on the Semantic Web. VLDB Journal: Very Large Data Bases, 12(4):333–351, November 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Fileto:2003:POW

[242] Renato Fileto, Ling Liu, Calton Pu, Eduardo Delgado Assad, and Claudia Bauzer Medeiros. POESIA: an ontological workflow approach for composing Web services in agriculture. VLDB Journal: Very Large Data Bases, 12(4):352–367, November 2003. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Jensen:2004:MDM

[243] Christian S. Jensen, Augustas Kligys, Torben Bach Pedersen, and Igor Timko. Multidimensional data modeling for location-based services. VLDB Journal: Very Large Data Bases, 13 (1):1–21, January 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhang:2004:PMV

[244] Xin Zhang, Lingli Ding, and Elke A. Rundensteiner. Parallel multisource view maintenance. VLDB Journal: Very Large Data Bases, 13(1):22–48, January 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Hristidis:2004:AAA

[245] Vagelis Hristidis and Yannis Papakonstantinou. Algorithms and applications for answering ranked queries using ranked views. VLDB Journal: Very Large Data Bases, 13(1):49–70, January 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Khan:2004:REO

[246] Latifur Khan, Dennis McLeod, and Eduard Hovy. Retrieval effectiveness of an ontology-based model for information selection. VLDB Journal: Very Large Data Bases, 13(1):71–85, January 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Donderler:2004:RBS

[247] Mehmet Emin Dönderler, Özgür Ulusoy, and Ugur Güdükbay. Rule-based spatiotemporal query processing for video databases. VLDB Journal: Very Large Data Bases, 13(1):86–103, January 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yu:2004:QHD

[248] Cui Yu, Stéphane Bressan, Beng Chin Ooi, and Kian-Lee Tan. Querying highdimensional data in single-dimensional space. VLDB Journal: Very Large Data Bases, 13(2):105–119, May 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Dori:2004:VVS

[249] Dov Dori. ViSWeb — the Visual Semantic Web: unifying human and machine knowledge representations with Object-Process Methodology. VLDB Journal: Very Large Data Bases, 13 (2):120–147, May 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Fu:2004:EHA

[250] Lixin Fu and Sanguthevar Rajasekaran. Evaluating holistic aggregators efficiently for very large datasets. VLDB Journal: Very Large Data Bases, 13 (2):148–161, May 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Rahal:2004:ETU

[251] Amira Rahal, Qiang Zhu, and Per-Åke Larson. Evolutionary techniques for updating query cost models in a dynamic multidatabase environment. VLDB Journal: Very Large Data Bases, 13(2):162–176, May 2004. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Adi:2004:ASM

[252] Asaf Adi and Opher Etzion. Amit — the situation manager. VLDB Journal: Very Large Data Bases, 13(2):177–203, May 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Freytag:2004:BPV

[253] Johann-Christoph Freytag, Serge Abiteboul, and Mike Carey. Best papers of VLDB 2003. VLDB Journal: Very Large Data Bases, 13(3): 205–206, September 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ilyas:2004:STJ

[254] Ihab F. Ilyas, Walid G. Aref, and Ahmed K. Elmagarmid. Supporting top-k join queries in relational databases. VLDB Journal: Very Large Data Bases, 13(3):207–221, September 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Papadimitriou:2004:AUS

[255] Spiros Papadimitriou, Anthony Brockwell, and Christos Faloutsos. Adaptive, unsupervised stream mining. VLDB Journal: Very Large Data Bases, 13 (3):222–239, September 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Labrinidis:2004:ETB

[256] Alexandros Labrinidis and Nick Roussopoulos. Exploring the tradeoff between performance and data freshness in database-driven Web servers. VLDB Journal: Very Large Data Bases, 13 (3):240–255, September 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

He:2004:AIW

[257] Hai He, Weiyi Meng, Clement Yu, and Zonghuan Wu. Automatic integration of Web search interfaces with WISEintegrator. VLDB Journal: Very Large Data Bases, 13(3):256–273, September 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Velegrakis:2004:PMC

[258] Yannis Velegrakis, René J. Miller, and Lucian Popa. Preserving mapping consistency under schema changes. VLDB Journal: Very Large Data Bases, 13 (3):274–293, September 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Florescu:2004:BSX

[259] Daniela Florescu, Chris Hillery, Donald Kossmann, Paul Lucas, Fabio Riccardi, Till Westmann, J. Carey, and Arvind Sundararajan. The BEA streaming XQuery processor. VLDB Journal: Very Large Data Bases, 13(3):294–315, September 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gehrke:2004:GES

[260] Johannes Gehrke and M. Hellerstein. Guest Editorial to the special issue on data stream processing. VLDB Journal: Very Large Data Bases, 13 (4):317, December 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yang:2004:FHQ

[261] Huai Yang, Li Lee, and Wynne Hsu. Finding hot query patterns over an XQuery stream. VLDB Journal: Very Large Data Bases, 13(4):318–332, December 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Babcock:2004:OSD

[262] Brian Babcock, Shivnath Babu, Mayur Datar, Rajeev Motwani, and Dilys Thomas. Operator scheduling in data stream systems. VLDB Journal: Very Large Data Bases, 13(4):333–353, December 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ganguly:2004:TSE

[263] Sumit Ganguly, Minos Garofalakis, and Rajeev Rastogi. Tracking setexpression cardinalities over continuous update streams. VLDB Journal: Very Large Data Bases, 13(4): 354–369, December 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Balakrishnan:2004:RA

[264] Hari Balakrishnan, Magdalena Balazinska, Don Carney, Uḡur Çetintemel, Mitch Cherniack, Christian Convey, Eddie Galvez, Jon Salz, Michael Stonebraker, Nesime Tatbul, Richard Tibbetts, and Stan Zdonik. Retrospective on Aurora. VLDB Journal: Very Large Data Bases, 13(4):370–383, December 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Sharaf:2004:BEE

[265] A. Sharaf, Jonathan Beaver, Alexandros Labrinidis, and K. Chrysanthis. Balancing energy efficiency and quality of aggregate data in sensor networks. VLDB Journal: Very Large Data Bases, 13(4):384–403, December 2004. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ozsu:2005:E

[266] Tamer Özsu. Editorial. VLDB Journal: Very Large Data Bases, 14(1):1, March 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gao:2005:JOT

[267] Dengfeng Gao, S. Jensen, T. Snodgrass, and D. Soo. Join operations in temporal databases. VLDB Journal: Very Large Data Bases, 14(1):2–29, March 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Balmin:2005:SQX

[268] Andrey Balmin and Yannis Papakonstantinou. Storing and querying XML data using denormalized relational databases. VLDB Journal: Very Large Data Bases, 14(1):30–49, March 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gal:2005:FME

[269] Avigdor Gal, Ateret Anaby-Tavor, Alberto Trombetta, and Danilo Montesi. A framework for modeling and evaluating automatic semantic reconciliation. VLDB Journal: Very Large Data Bases, 14(1):50–67, March 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Halevy:2005:SML

[270] Y. Halevy, G. Ives, Dan Suciu, and Igor Tatarinov. Schema mediation for largescale semantic data sharing. VLDB Journal: Very Large Data Bases, 14 (1):68–83, March 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Benatallah:2005:AWS

[271] Boualem Benatallah, Mohand-Said Hacid, Alain Leger, Christophe Rey, and Farouk Toumani. On automating Web services discovery. VLDB Journal: Very Large Data Bases, 14(1):84–96, March 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Sattler:2005:CBQ

[272] Kai-Uwe Sattler, Ingolf Geist, and Eike Schallehn. Concept-based querying in mediator systems. VLDB Journal: Very Large Data Bases, 14(1):97–111, March 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Tzitzikas:2005:MTB

[273] Yannis Tzitzikas, Nicolas Spyratos, and Panos Constantopoulos. Mediators over taxonomy-based information sources. VLDB Journal: Very Large Data Bases, 14(1):112–136, March 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gunopulos:2005:SEM

[274] Dimitrios Gunopulos, George Kollios, J. Tsotras, and Carlotta Domeniconi. Selectivity estimators for multidimensional range queries over real attributes. VLDB Journal: Very Large Data Bases, 14(2):137–154, April 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Alhajj:2005:VFC

[275] Reda Alhajj, Faruk Polat, and Cem Yílmaz. Views as first-class citizens in object-oriented databases. VLDB Journal: Very Large Data Bases, 14 (2):155–169, April 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhang:2005:OSM

[276] Donghui Zhang and J. Tsotras. Optimizing spatial Min/Max aggregations. VLDB Journal: Very Large Data Bases, 14(2):170–181, April 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Perich:2005:CJP

[277] Filip Perich, Anupam Joshi, Yelena Yesha, and Tim Finin. Collaborative joins in a pervasive computing environment. VLDB Journal: Very Large Data Bases, 14(2):182–196, April 2005.

CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Josifovski:2005:QXS

[278] Vanja Josifovski, Marcus Fontoura, and Attila Barta. Querying XML streams. VLDB Journal: Very Large Data Bases, 14(2):197–210, April 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Aggarwal:2005:EEA

[279] C. Aggarwal and S. Yu. An effective and efficient algorithm for highdimensional outlier detection. VLDB Journal: Very Large Data Bases, 14 (2):211–221, April 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yao:2005:HBL

[280] D. Yao, Cyrus Shahabi, and Per-Åke Larson. Hash-based labeling techniques for storage scaling. VLDB Journal: Very Large Data Bases, 14(2):222–237, April 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Kollios:2005:IMO

[281] George Kollios, Dimitris Papadopoulos, Dimitrios Gunopulos, and J. Tsotras. Indexing mobile objects using dual transformations. VLDB Journal: Very Large Data Bases, 14(2):238–256, April 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Jaluta:2005:CCR

[282] Ibrahim Jaluta, Seppo Sippu, and Eljas Soisalon-Soininen. Concurrency control and recovery for balanced B-link trees. VLDB Journal: Very Large Data Bases, 14(2):257–277, April 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gaasterland:2005:SID

[283] Terry Gaasterland, H. V. Jagadish, and Louiqa Raschid. Special issue on data management, analysis, and mining for the life sciences. VLDB Journal: Very Large Data Bases, 14(3): 279–280, September 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0938-1287&
volume=14&issue=3&spage=279.

Tian:2005:PMC

[284] Yuanyuan Tian, Sandeep Tata, Richard A. Hankins, and Jignesh M. Patel. Practical methods for constructing suffix trees. VLDB Journal: Very Large Data Bases, 14(3):281– 299, September 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0938-1287&
volume=14&issue=3&spage=281.

Claypool:2005:SYD

[285] Kajal T. Claypool and Elke A. Rundensteiner. Sync your data: update propagation for heterogeneous protein databases. VLDB Journal: Very Large Data Bases, 14(3): 300–317, September 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0938-1287&
volume=14&issue=3&spage=300.

Conery:2005:RBW

[286] John S. Conery, Julian M. Catchen, and Michael Lynch. Rule-based workflow management for bioinformatics. VLDB Journal: Very Large Data Bases, 14(3):318–329, September 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://www.

                                        springerlink.com/openurl.asp?genre=
                                        article&issn=0938-1287&volume=14&
                                        issue=3&spage=318.

Thakkar:2005:COE

[287] Snehal Thakkar, José Luis Ambite, and Craig A. Knoblock. Composing, optimizing, and executing plans for bioinformatics web services. VLDB Journal: Very Large Data Bases, 14(3): 330–353, September 2005. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=0938-1287&
                                        volume=14&issue=3&spage=330.

Vlachos:2006:IMT

[288] Michail Vlachos, Marios Hadjieleftheriou, Dimitrios Gunopulos, and Eamonn Keogh. Indexing Multidimensional Time-Series. VLDB Journal: Very Large Data Bases, 15(1):1–20, January 2006. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zheng:2006:GPI

[289] Baihua Zheng, Jianliang Xu, Wang-Chien Lee, and Lun Lee. Grid-partition index: a hybrid method for nearestneighbor queries in wireless locationbased services. VLDB Journal: Very Large Data Bases, 15(1):21–39, January 2006. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Tamir:2006:CGM

[290] Raz Tamir and Yehuda Singer. On a confidence gain measure for association rule discovery and scoring. VLDB Journal: Very Large Data Bases, 15 (1):40–52, January 2006. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Bremer:2006:IDD

[291] Jan-Marco Bremer and Michael Gertz. Integrating document and data retrieval based on XML. VLDB Journal: Very Large Data Bases, 15(1):53– 83, January 2006. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ogras:2006:OSD

[292] Y. Ogras and Hakan Ferhatosmanoglu. Online summarization of dynamic time series data. VLDB Journal: Very Large Data Bases, 15(1):84–98, January 2006. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Goh:2006:DBM

[293] Leng Goh, Yanfeng Shu, Zhiyong Huang, and Chin Ooi. Dynamic buffer management with extensible replacement policies. VLDB Journal: Very Large Data Bases, 15(2):99–120, June 2006. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Arasu:2006:CCQ

[294] Arvind Arasu, Shivnath Babu, and Jennifer Widom. The CQL continuous query language: semantic foundations and query execution. VLDB Journal: Very Large Data Bases, 15(2):121–142, June 2006. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Hadjieleftheriou:2006:ISA

[295] Marios Hadjieleftheriou, George Kollios, J. Tsotras, and Dimitrios Gunopulos. Indexing spatiotemporal archives. VLDB Journal: Very Large Data Bases, 15(2):143–164, June 2006. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Guting:2006:MQM

[296] Hartmut Güting, Teixeira de Almeida, and Zhiming Ding. Modeling and querying moving objects in networks. VLDB Journal: Very Large Data Bases, 15(2):165–190, June 2006. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chirkova:1999:AQU

[297] Rada Chirkova, Chen Li, and Jia Li. Answering queries using materialized views with minimum size. VLDB Journal: Very Large Data Bases, 15 (3):191–210, April 1999. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Cao:1999:STD

[298] Hu Cao, Ouri Wolfson, and Goce Trajcevski. Spatio-temporal data reduction with deterministic error bounds. VLDB Journal: Very Large Data Bases, 15(3):211–228, April 1999. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Benetis:1999:NRN

[299] Rimantas Benetis, S. Jensen, Gytis Karçiauskas, and Simonas S̊altenis. Nearest and reverse nearest neighbor queries for moving objects. VLDB Journal: Very Large Data Bases, 15 (3):229–249, April 1999. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Pelleg:1999:DTS

[300] Dan Pelleg and Andrew Moore. Dependency trees in sub-linear time and bounded memory. VLDB Journal: Very Large Data Bases, 15(3):250–262, April 1999. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Che:1999:QOX

[301] Dunren Che, Karl Aberer, and Tamer Özsu. Query optimization in XML structured-document databases. VLDB Journal: Very Large Data Bases, 15(3):263–289, April 1999. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ferrari:2006:GES

[302] Elena Ferrari and Bhavani Thuraisingham. Guest editorial: special issue on privacy preserving data management. VLDB Journal: Very Large Data Bases, 15(4):291–292, November 2006. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Mukherjee:2006:PPT

[303] Shibnath Mukherjee, Zhiyuan Chen, and Aryya Gangopadhyay. A privacypreserving technique for Euclidean distance-based mining algorithms using Fourier-related transforms. VLDB Journal: Very Large Data Bases, 15 (4):293–315, November 2006. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Jiang:2006:SDF

[304] Wei Jiang and Chris Clifton. A secure distributed framework for achieving kanonymity. VLDB Journal: Very Large Data Bases, 15(4):316–333, November 2006. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Blanton:2006:SRF

[305] Marina Blanton and Mikhail Atallah. Succinct representation of flexible and privacy-preserving access rights. VLDB Journal: Very Large Data Bases, 15 (4):334–354, November 2006. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Domingo-Ferrer:2006:EMD

[306] Josep Domingo-Ferrer, Antoni Martínez-Ballesté, Josep Maria Mateo-Sanz, and Francesc Sebé. Efficient multivariate data-oriented microaggregation. VLDB Journal: Very Large Data Bases, 15 (4):355–369, November 2006. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Massacci:2006:HHD

[307] Fabio Massacci, John Mylopoulos, and Nicola Zannone. Hierarchical Hippocratic databases with minimal disclosure for virtual organizations. VLDB Journal: Very Large Data Bases, 15 (4):370–387, November 2006. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Xiong:2006:PLM

[308] Hui Xiong, Michael Steinbach, and Vipin Kumar. Privacy leakage in multirelational databases: a semi-supervised learning perspective. VLDB Journal: Very Large Data Bases, 15(4): 388–402, November 2006. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Haas:2006:SIB

[309] Laura M. Haas, Christian S. Jensen, and Martin L. Kersten. Special issue: best papers of VLDB 2005. VLDB Journal: Very Large Data Bases, 16(1): 1–3, October 2006. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Godfrey:2006:AAM

[310] Parke Godfrey, Ryan Shipley, and Jarek Gryz. Algorithms and analyses for maximal vector computation. VLDB Journal: Very Large Data Bases, 16(1):5–28, October 2006. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Larson:2006:VMO

[311] Per-Åke Larson and Jingren Zhou. View matching for outer-join views. VLDB Journal: Very Large Data Bases, 16(1):29–53, October 2006. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Markl:2006:CSE

[312] V. Markl, P. J. Haas, M. Kutsch, N. Megiddo, U. Srivastava, and T. M. Tran. Consistent selectivity estimation via maximum entropy. VLDB Journal: Very Large Data Bases, 16 (1):55–76, October 2006. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ghoting:2006:CCF

[313] Amol Ghoting, Gregory Buehrer, Srinivasan Parthasarathy, Daehyun Kim, Anthony Nguyen, Yen-Kuang Chen, and Pradeep Dubey. Cache-conscious frequent pattern mining on modern and emerging processors. VLDB Journal: Very Large Data Bases, 16(1):77– 96, October 2006. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lee:2006:ETS

[314] Yoonkyong Lee, Mayssam Sayyadian, AnHai Doan, and Arnon S. Rosenthal. eTuner: tuning schema matching software using synthetic scenarios. VLDB Journal: Very Large Data Bases, 16 (1):97–122, October 2006. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Burdick:2006:OUI

[315] Doug Burdick, Prasad M. Deshpande, T. S. Jayram, Raghu Ramakrishnan, and Shivakumar Vaithyanathan. OLAP over uncertain and imprecise data. VLDB Journal: Very Large Data Bases, 16(1):123–144, October 2006. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Haftmann:2006:FER

[316] Florian Haftmann, Donald Kossmann, and Eric Lo. A framework for efficient regression tests on database applications. VLDB Journal: Very Large Data Bases, 16(1):145–164, October 2006. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Haas:2007:SIB

[317] Laura M. Haas, Christian S. Jensen, and Martin L. Kersten. Special issue: best papers of VLDB 2005. VLDB Journal: Very Large Data Bases, 16(1): 1–3, January 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Godfrey:2007:AAM

[318] Parke Godfrey, Ryan Shipley, and Jarek Gryz. Algorithms and analyses for maximal vector computation. VLDB Journal: Very Large Data Bases, 16(1):5–28, January 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Larson:2007:VMO

[319] Per-Åke Larson and Jingren Zhou. View matching for outer-join views. VLDB Journal: Very Large Data Bases, 16(1):29–53, January 2007. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Markl:2007:CSE

[320] V. Markl, P. J. Haas, M. Kutsch, N. Megiddo, U. Srivastava, and T. M. Tran. Consistent selectivity estimation via maximum entropy. VLDB Journal: Very Large Data Bases, 16 (1):55–76, January 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ghoting:2007:CCF

[321] Amol Ghoting, Gregory Buehrer, Srinivasan Parthasarathy, Daehyun Kim, Anthony Nguyen, Yen-Kuang Chen, and Pradeep Dubey. Cache-conscious frequent pattern mining on modern and emerging processors. VLDB Journal: Very Large Data Bases, 16(1):77– 96, January 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lee:2007:ETS

[322] Yoonkyong Lee, Mayssam Sayyadian, AnHai Doan, and Arnon S. Rosenthal. eTuner: tuning schema matching software using synthetic scenarios. VLDB Journal: Very Large Data Bases, 16 (1):97–122, January 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Burdick:2007:OUI

[323] Doug Burdick, Prasad M. Deshpande, T. S. Jayram, Raghu Ramakrishnan, and Shivakumar Vaithyanathan. OLAP over uncertain and imprecise data. VLDB Journal: Very Large Data Bases, 16(1):123–144, January 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Haftmann:2007:FER

[324] Florian Haftmann, Donald Kossmann, and Eric Lo. A framework for efficient regression tests on database applications. VLDB Journal: Very Large Data Bases, 16(1):145–164, January 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Tanin:2007:UDQ

[325] Egemen Tanin, Aaron Harwood, and Hanan Samet. Using a distributed quadtree index in peer-to-peer networks. VLDB Journal: Very Large Data Bases, 16(2):165–178, April 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Viqueira:2007:SES

[326] Jose R. Rios Viqueira and Nikos A. Lorentzos. SQL extension for spatiotemporal data. VLDB Journal: Very Large Data Bases, 16(2):179–200, April 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Dai:2007:CDC

[327] Bi-Ru Dai, Cheng-Ru Lin, and Ming-Syan Chen. Constrained data clustering by depth control and progressive constraint relaxation. VLDB Journal: Very Large Data Bases, 16(2):201–217, April 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Shen:2007:ADD

[328] Heng Tao Shen, Xiaofang Zhou, and Aoying Zhou. An adaptive and dynamic dimensionality reduction method for high-dimensional indexing. VLDB Journal: Very Large Data Bases, 16(2):219–234, April 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

He:2007:PCC

[329] Zhen He and Alonso Marquez. Path and cache conscious prefetching (PCCP). VLDB Journal: Very Large Data Bases, 16(2):235–249, April 2007.

CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yu:2007:MBS

[330] Hailing Yu, Divyakant Agrawal, and Amr El Abbadi. MEMS based storage architecture for relational databases. VLDB Journal: Very Large Data Bases, 16(2):251–268, April 2007. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yiannis:2007:CTF

[331] John Yiannis and Justin Zobel. Compression techniques for fast external sorting. VLDB Journal: Very Large Data Bases, 16(2):269–291, April 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Jermaine:2007:PEF

[332] Christopher Jermaine, Edward Omiecinski, and Wai Gen Yee. The partitioned exponential file for database storage management. VLDB Journal: Very Large Data Bases, 16(4):417–437, October 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Deligiannakis:2007:DCH

[333] Antonios Deligiannakis, Yannis Kotidis, and Nick Roussopoulos. Dissemination of compressed historical information in sensor networks. VLDB Journal: Very Large Data Bases, 16 (4):439–461, October 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Bohm:2007:FRA

[334] Klemens Böhm and Erik Buchmann. Free riding-aware forwarding in Content-Addressable Networks. VLDB Journal: Very Large Data Bases, 16 (4):463–482, October 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Traina:2007:OFA

[335] Caetano Traina, Jr., Roberto F. Filho, Agma J. Traina, Marcos R. Vieira, and Christos Faloutsos. The Omni-family of all-purpose access methods: a simple and effective way to make similarity search more efficient. VLDB Journal: Very Large Data Bases, 16 (4):483–505, October 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Khan:2007:NID

[336] Latifur Khan, Mamoun Awad, and Bhavani Thuraisingham. A new intrusion detection system using support vector machines and hierarchical clustering. VLDB Journal: Very Large Data Bases, 16(4):507–521, October 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Dalvi:2007:EQE

[337] Nilesh Dalvi and Dan Suciu. Efficient query evaluation on probabilistic databases. VLDB Journal: Very Large Data Bases, 16(4):523–544, October 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Croft:2008:ISI

[338] W. Bruce Croft and Hans-J. Schek. Introduction to the special issue on database and information retrieval integration. VLDB Journal: Very Large Data Bases, 17(1):1–3, January 2008.

CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Roelleke:2008:MRM

[339] Thomas Roelleke, Hengzhi Wu, Jun Wang, and Hany Azzam. Modelling retrieval models in a probabilistic relational algebra with a new operator: the relational Bayes. VLDB Journal: Very Large Data Bases, 17(1):5–37, January 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Schmitt:2008:QDQ

[340] Ingo Schmitt. QQL: A DB&IR Query Language. VLDB Journal: Very Large Data Bases, 17(1):39–56, January 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lau:2008:MRM

[341] Ho Lam Lau and Wilfred Ng. A multiranker model for adaptive XML searching. VLDB Journal: Very Large Data Bases, 17(1):57–80, January 2008. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Theobald:2008:TEV

[342] Martin Theobald, Holger Bast, Debapriyo Majumdar, Ralf Schenkel, and Gerhard Weikum. TopX: efficient and versatile top-k query processing for semistructured data. VLDB Journal: Very Large Data Bases, 17(1):81–115, January 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Simitsis:2008:PUK

[343] Alkis Simitsis, Georgia Koutrika, and Yannis Ioannidis. Précis: from unstructured keywords as queries to structured databases as answers. VLDB Journal: Very Large Data Bases, 17 (1):117–149, January 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Cornacchia:2008:FEI

[344] Roberto Cornacchia, Sándor Héman, Marcin Zukowski, Arjen P. Vries, and Peter Boncz. Flexible and efficient IR using array databases. VLDB Journal: Very Large Data Bases, 17 (1):151–168, January 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lockemann:2008:MKR

[345] Peter C. Lockemann. In memoriam Klaus R. Dittrich (1950—2007). VLDB Journal: Very Large Data Bases, 17 (2):169–170, March 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Alonso:2008:GEM

[346] Gustavo Alonso, David Lomet, and Umesh Dayal. Guest Editors’ message. VLDB Journal: Very Large Data Bases, 17(2):171–172, March 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gemulla:2008:MBS

[347] Rainer Gemulla, Wolfgang Lehner, and Peter J. Haas. Maintaining bounded-size sample synopses of evolving datasets. VLDB Journal: Very Large Data Bases, 17(2):173–201, March 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yu:2008:XSR

[348] Cong Yu and H. V. Jagadish. XML schema refinement through redundancy detection and normalization. VLDB Journal: Very Large Data Bases, 17 (2):203–223, March 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Mitra:2008:TKS

[349] Soumyadeb Mitra, Marianne Winslett, Windsor W. Hsu, and Kevin Chen-Chuan Chang. Trustworthy keyword search for compliance storage. VLDB Journal: Very Large Data Bases, 17 (2):225–242, March 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Benjelloun:2008:DUL

[350] Omar Benjelloun, Anish Das Sarma, Alon Halevy, Martin Theobald, and Jennifer Widom. Databases with uncertainty and lineage. VLDB Journal: Very Large Data Bases, 17(2):243–264, March 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Jeffery:2008:ARM

[351] Shawn R. Jeffery, Michael J. Franklin, and Minos Garofalakis. An adaptive RFID middleware for supporting metaphysical data independence. VLDB Journal: Very Large Data Bases, 17 (2):265–289, March 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Parreira:2008:JAP

[352] Josiane Xavier Parreira, Carlos Castillo, Debora Donato, Sebastian Michel, and Gerhard Weikum. The Juxtaposed approximate PageRank method for robust PageRank approximation in a peer-to-peer web search network. VLDB Journal: Very Large Data Bases, 17(2):291–313, March 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Narayanan:2008:DAQ

[353] Dushyanth Narayanan, Austin Donnelly, Richard Mortier, and Antony Rowstron. Delay aware querying with Seaweed. VLDB Journal: Very Large Data Bases, 17(2):315–331, March 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Bernstein:2008:IMC

[354] Philip A. Bernstein, Todd J. Green, Sergey Melnik, and Alan Nash. Implementing mapping composition. VLDB Journal: Very Large Data Bases, 17 (2):333–353, March 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Li:2008:ESF

[355] Yunyao Li, Cong Yu, and H. V. Jagadish. Enabling Schema-Free XQuery with meaningful query focus. VLDB Journal: Very Large Data Bases, 17 (3):355–377, May 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yiu:2008:BTI

[356] Man Lung Yiu, Yufei Tao, and Nikos Mamoulis. The Bdual-Tree: indexing moving objects by space filling curves in the dual space. VLDB Journal: Very Large Data Bases, 17(3):379–400, May 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Awad:2008:PWS

[357] Mamoun Awad, Latifur Khan, and Bhavani Thuraisingham. Predicting WWW surfing using multiple evidence combination. VLDB Journal: Very Large Data Bases, 17(3):401–417, May 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wang:2008:HBM

[358] Hai Wang and Kenneth C. Sevcik. Histograms based on the minimum description length principle. VLDB Journal: Very Large Data Bases, 17(3):419– 442, May 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Deligiannakis:2008:BCQ

[359] Antonios Deligiannakis, Yannis Kotidis, and Nick Roussopoulos. Bandwidthconstrained queries in sensor networks. VLDB Journal: Very Large Data Bases, 17(3):443–467, May 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Hammad:2008:QPM

[360] Moustafa A. Hammad, Walid G. Aref, and Ahmed K. Elmagarmid. Query processing of multi-way stream window joins. VLDB Journal: Very Large Data Bases, 17(3):469–488, May 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Luo:2008:FBP

[361] Qiong Luo, Jeffrey F. Naughton, and Wenwei Xue. Form-based proxy caching for database-backed web sites: keywords and functions. VLDB Journal: Very Large Data Bases, 17(3):489– 513, May 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wang:2008:EAM

[362] Yida Wang, Ee-Peng Lim, and San-Yih Hwang. Efficient algorithms for mining maximal valid groups. VLDB Journal: Very Large Data Bases, 17(3):515–535, May 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yu:2008:DMW

[363] Qi Yu, Xumin Liu, Athman Bouguettaya, and Brahim Medjahed. Deploying and managing Web services: issues, solutions, and directions. VLDB Journal: Very Large Data Bases, 17(3):537–572, May 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Li:2008:EUD

[364] Changqing Li, Tok Wang Ling, and Min Hu. Efficient updates in dynamic XML data: from binary string to quaternary string. VLDB Journal: Very Large Data Bases, 17(3):573–601, May 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Tao:2007:MRK

[365] Yufei Tao, Dimitris Papadias, Xiang Lian, and Xiaokui Xiao. Multidimensional reverse kNN search. VLDB Journal: Very Large Data Bases, 16(3):293– 316, July 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Koch:2007:AGS

[366] Christoph Koch and Stefanie Scherzinger. Attribute grammars for scalable query processing on XML streams. VLDB Journal: Very Large Data Bases, 16 (3):317–342, July 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chan:2007:OES

[367] Edward P. Chan and Heechul Lim. Optimization and evaluation of shortest path queries. VLDB Journal: Very Large Data Bases, 16(3):343–369, July 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lee:2007:DPI

[368] Jae-Gil Lee, Kyu-Young Whang, Wook-Shin Han, and Il-Yeol Song. The dynamic predicate: integrating access control with query processing in XML databases. VLDB Journal: Very Large Data Bases, 16(3):371–387, July 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Papazoglou:2007:SOA

[369] Mike P. Papazoglou and Willem-Jan Heuvel. Service oriented architectures: approaches, technologies and research issues. VLDB Journal: Very Large Data Bases, 16(3):389–415, July 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Byun:2008:PBA

[370] Ji-Won Byun and Ninghui Li. Purpose based access control for privacy protection in relational database systems. VLDB Journal: Very Large Data Bases, 17(4):603–619, July 2008.

CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Karayannidis:2008:HCO

[371] Nikos Karayannidis and Timos Sellis. Hierarchical clustering for OLAP: the CUBE File approach. VLDB Journal: Very Large Data Bases, 17(4):621–655, July 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Plattner:2008:EDS

[372] Christian Plattner, Gustavo Alonso, and M. Tamer Özsu. Extending DBMSs with satellite databases. VLDB Journal: Very Large Data Bases, 17(4):657–682, July 2008. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Hsieh:2008:DEF

[373] Ming-Jyh Hsieh, Wei-Guang Teng, Ming-Syan Chen, and Philip S. Yu. DAWN: an efficient framework of DCT for data with error estimation. VLDB Journal: Very Large Data Bases, 17 (4):683–702, July 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Atzori:2008:APP

[374] Maurizio Atzori, Francesco Bonchi, Fosca Giannotti, and Dino Pedreschi. Anonymity preserving pattern discovery. VLDB Journal: Very Large Data Bases, 17(4):703–727, July 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Morfonios:2008:SDC

[375] Konstantinos Morfonios and Yannis Ioannidis. Supporting the data cube lifecycle: the power of ROLAP. VLDB Journal: Very Large Data Bases, 17 (4):729–764, July 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Sharifzadeh:2008:OSR

[376] Mehdi Sharifzadeh, Mohammad Kolahdouzan, and Cyrus Shahabi. The optimal sequenced route query. VLDB Journal: Very Large Data Bases, 17 (4):765–787, July 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Friedman:2008:PAD

[377] Arik Friedman, Ran Wolff, and Assaf Schuster. Providing k-anonymity in data mining. VLDB Journal: Very Large Data Bases, 17(4):789–804, July 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Harder:2008:VCC

[378] Theo Härder and Andreas Bühmann. Value complete, column complete, predicate complete. VLDB Journal: Very Large Data Bases, 17(4):805–826, July 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ou:2008:EAI

[379] Jian-Chih Ou, Chang-Hung Lee, and Ming-Syan Chen. Efficient algorithms for incremental Web log mining with dynamic thresholds. VLDB Journal: Very Large Data Bases, 17(4):827–845, July 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Alagic:2008:GJP

[380] Suad Alagić and Mark Royer. Genericity in Java: persistent and database systems implications. VLDB Journal: Very Large Data Bases, 17(4):847–878, July 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Vaidya:2008:PPN

[381] Jaideep Vaidya, Murat Kantarcıoğlu, and Chris Clifton. Privacy-preserving Naïve Bayes classification. VLDB Journal: Very Large Data Bases, 17(4):879– 898, July 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Fu:2008:STW

[382] Ada Wai-Chee Fu, Eamonn Keogh, Leo Yung Lau, Chotirat Ann Ratanamahatana, and Raymond Chi-Wing Wong. Scaling and time warping in time series querying. VLDB Journal: Very Large Data Bases, 17(4):899–921, July 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Mouratidis:2008:TBP

[383] Kyriakos Mouratidis, Dimitris Papadias, and Spiros Papadimitriou. Treebased partition querying: a methodology for computing medoids in large spatial datasets. VLDB Journal: Very Large Data Bases, 17(4):923–945, July 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yu:2008:DMP

[384] Jeffrey Xu Yu, Zhiheng Li, and Guimei Liu. A data mining proxy approach for efficient frequent itemset mining. VLDB Journal: Very Large Data Bases, 17(4):947–970, July 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Mokbel:2008:SSL

[385] Mohamed F. Mokbel and Walid G. Aref. SOLE: scalable on-line execution of continuous queries on spatiotemporal data streams. VLDB Journal: Very Large Data Bases, 17(5):971–995, August 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Pol:2008:MVL

[386] Abhijit Pol, Christopher Jermaine, and Subramanian Arumugam. Maintaining very large random samples using the geometric file. VLDB Journal: Very Large Data Bases, 17(5):997–1018, August 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Abiteboul:2008:AXP

[387] Serge Abiteboul, Omar Benjelloun, and Tova Milo. The Active XML project: an overview. VLDB Journal: Very Large Data Bases, 17(5):1019–1040, August 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Buccafurri:2008:EHT

[388] Francesco Buccafurri, Gianluca Lax, Domenico Saccà, Luigi Pontieri, and Domenico Rosaci. Enhancing histograms by tree-like bucket indices. VLDB Journal: Very Large Data Bases, 17(5):1041–1061, August 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Kamra:2008:DAA

[389] Ashish Kamra, Evimaria Terzi, and Elisa Bertino. Detecting anomalous access patterns in relational databases. VLDB Journal: Very Large Data Bases, 17(5):1063–1077, August 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Guha:2008:WSH

[390] Sudipto Guha, Hyoungmin Park, and Kyuseok Shim. Wavelet synopsis for hierarchical range queries with workloads. VLDB Journal: Very Large Data Bases, 17(5):1079–1099, August 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Deng:2008:MRS

[391] Ke Deng, Xiaofang Zhou, Heng Tao Shen, Qing Liu, Kai Xu, and Xuemin Lin. A multi-resolution surface distance model for k-NN query processing. VLDB Journal: Very Large Data Bases, 17(5):1101–1119, August 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chuang:2008:PLR

[392] Kun-Ta Chuang, Jiun-Long Huang, and Ming-Syan Chen. Power-law relationship and self-similarity in the itemset support distribution: analysis and applications. VLDB Journal: Very Large Data Bases, 17(5): 1121–1141, August 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Padmanabhan:2008:SDR

[393] Prasanna Padmanabhan, Le Gruenwald, Anita Vallur, and Mohammed Atiquzzaman. A survey of data replication techniques for mobile ad hoc network databases. VLDB Journal: Very Large Data Bases, 17(5):1143–1164, August 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhong:2008:GPT

[394] Sheng Zhong and Zhiqiang Yang. Guided perturbation: towards private and accurate mining. VLDB Journal: Very Large Data Bases, 17(5): 1165–1177, August 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Rizzolo:2008:TXM

[395] Flavio Rizzolo and Alejandro A. Vaisman. Temporal XML: modeling, indexing, and query processing. VLDB Journal: Very Large Data Bases, 17 (5):1179–1212, August 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Jin:2008:SES

[396] Liang Jin, Chen Li, and Rares Vernica. SEPIA: estimating selectivities of approximate string predicates in large databases. VLDB Journal: Very Large Data Bases, 17(5):1213–1229, August 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Venkateswaran:2008:RBI

[397] Jayendra Venkateswaran, Tamer Kahveci, Christopher Jermaine, and Deepak Lachwani. Reference-based indexing for metric spaces with costly distance measures. VLDB Journal: Very Large Data Bases, 17(5):1231–1251, August 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Tao:2008:PDW

[398] Yufei Tao and Xiaokui Xiao. Primal or dual: which promises faster spatiotemporal search? VLDB Journal: Very Large Data Bases, 17(5): 1253–1270, August 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Tao:2008:ETC

[399] Yufei Tao and Xiaokui Xiao. Efficient temporal counting with bounded error. VLDB Journal: Very Large Data Bases, 17(5):1271–1292, August 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Islam:2008:ACB

[400] Aminul Islam, Diana Inkpen, and Iluju Kiringa. Applications of corpusbased semantic similarity and word segmentation to database schema matching. VLDB Journal: Very Large Data Bases, 17(5):1293–1320, August 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chuang:2008:MTK

[401] Kun-Ta Chuang, Jiun-Long Huang, and Ming-Syan Chen. Mining top-k frequent patterns in the presence of the memory constraint. VLDB Journal: Very Large Data Bases, 17(5): 1321–1344, August 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Catarci:2008:GES

[402] Tiziana Catarci and René J. Miller. Guest editorial: special issue on metadata management. VLDB Journal: Very Large Data Bases, 17(6): 1345–1346, November 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Atzeni:2008:MIS

[403] Paolo Atzeni, Paolo Cappellari, Riccardo Torlone, Philip A. Bernstein, and Giorgio Gianforme. Modelindependent schema translation. VLDB Journal: Very Large Data Bases, 17(6): 1347–1370, November 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Cudre-Mauroux:2008:PMM

[404] Philippe Cudré-Mauroux, Adriana Budura, Manfred Hauswirth, and Karl Aberer. PicShark: mitigating metadata scarcity through large-scale P2P collaboration. VLDB Journal: Very Large Data Bases, 17(6):1371–1384, November 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Cruz:2008:LFS

[405] Isabel F. Cruz and Huiyong Xiao. A layered framework supporting personal information integration and application design for the semantic desktop. VLDB Journal: Very Large Data Bases, 17(6):1385–1406, November 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Candan:2008:SSE

[406] K. Selçuk Candan, Huiping Cao, Yan Qi, and Maria Luisa Sapino. System support for exploration and expert feedback in resolving conflicts during integration of metadata. VLDB Journal: Very Large Data Bases, 17(6): 1407–1444, November 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wang:2008:AXB

[407] Fusheng Wang, Carlo Zaniolo, and Xin Zhou. ArchIS: an XML-based approach to transaction-time temporal database systems. VLDB Journal: Very Large Data Bases, 17(6): 1445–1463, November 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhou:2008:DSD

[408] Yongluan Zhou, Beng Chin Ooi, and Kian-Lee Tan. Disseminating streaming data in a dynamic environment: an adaptive and cost-based approach. VLDB Journal: Very Large Data Bases, 17(6):1465–1483, November 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Kim:2008:SOF

[409] Min-Soo Kim, Kyu-Young Whang, Jae-Gil Lee, and Min-Jae Lee. Structural optimization of a full-text ngram index using relational normalization. VLDB Journal: Very Large Data Bases, 17(6):1485–1507, November 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Guha:2008:STO

[410] Sudipto Guha. On the space–time of optimal, approximate and streaming algorithms for synopsis construction problems. VLDB Journal: Very Large Data Bases, 17(6):1509–1535, November 2008. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lakhal:2009:FFE

[411] Neila Ben Lakhal, Takashi Kobayashi, and Haruo Yokota. FENECIA: failure endurable nested-transaction based execution of composite Web services with incorporated state analysis. VLDB Journal: Very Large Data Bases, 18 (1):1–56, January 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Sharifzadeh:2009:AVC

[412] Mehdi Sharifzadeh and Cyrus Shahabi. Approximate Voronoi cell computation on spatial data streams. VLDB Journal: Very Large Data Bases, 18 (1):57–75, January 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Vlachos:2009:OPV

[413] Michail Vlachos, Aris Anagnostopoulos, Olivier Verscheure, and Philip S. Yu. Online pairing of VoIP conversations. VLDB Journal: Very Large Data Bases, 18(1):77–98, January 2009. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yao:2009:LMK

[414] Yuxia Yao, Xueyan Tang, and Ee-Peng Lim. Localized monitoring of kNN queries in wireless sensor networks. VLDB Journal: Very Large Data Bases, 18(1):99–117, January 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Paton:2009:AQP

[415] Norman W. Paton, Jorge Buenabad-Chavez, Mengsong Chen, Vijayshankar Raman, Garret Swart, Inderpal Narang, Daniel M. Yellin, and Alvaro A. Fernandes. Autonomic query parallelization using non-dedicated computers: an evaluation of adaptivity options. VLDB Journal: Very Large Data Bases, 18(1):119–140, January 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Park:2009:ESR

[416] Laurence A. Park and Kotagiri Ramamohanarao. Efficient storage and retrieval of probabilistic latent semantic information for information retrieval. VLDB Journal: Very Large Data Bases, 18(1):141–155, January 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Askitis:2009:BTD

[417] Nikolas Askitis and Justin Zobel. Btries for disk-based string management. VLDB Journal: Very Large Data Bases, 18(1):157–179, January 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Joshi:2009:SBE

[418] Shantanu Joshi and Christopher Jermaine. Sampling-based estimators for subset-based queries. VLDB Journal: Very Large Data Bases, 18(1):181–202, January 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Sacharidis:2009:HCW

[419] Dimitris Sacharidis, Antonios Deligiannakis, and Timos Sellis. Hierarchically compressed wavelet synopses. VLDB Journal: Very Large Data Bases, 18 (1):203–231, January 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Theodoratos:2009:CPS

[420] Dimitri Theodoratos, Pawel Placek, Theodore Dalamagas, Stefanos Souldatos, and Timos Sellis. Containment of partially specified tree-pattern queries in the presence of dimension graphs. VLDB Journal: Very Large Data Bases, 18(1):233–254, January 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Benjelloun:2009:SGA

[421] Omar Benjelloun, Hector Garcia-Molina, David Menestrina, Qi Su, Steven Euijong Whang, and Jennifer Widom. Swoosh: a generic approach to entity resolution. VLDB Journal: Very Large Data Bases, 18(1):255–276, January 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ratprasartporn:2009:CBL

[422] Nattakarn Ratprasartporn, Jonathan Po, Ali Cakmak, Sulieman Bani-Ahmad, and Gultekin Ozsoyoglu. Context-based literature digital collection search. VLDB Journal: Very Large Data Bases, 18(1):277–301, January 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chiu:2009:EFS

[423] Ding-Ying Chiu, Yi-Hung Wu, and Arbee L. Chen. Efficient frequent sequence mining by a dynamic strategy switching algorithm. VLDB Journal: Very Large Data Bases, 18(1):303– 327, January 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Shen:2009:SII

[424] Heng Tao Shen, Shouxu Jiang, Kian-Lee Tan, Zi Huang, and Xiaofang Zhou. Speed up interactive image retrieval. VLDB Journal: Very Large Data Bases, 18(1):329–343, January 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wang:2009:SFS

[425] Shiyuan Wang, Quang Hieu Vu, Beng Chin Ooi, Anthony K. Tung, and Lizhen Xu. Skyframe: a framework for skyline query processing in peer-to-peer systems. VLDB Journal: Very Large Data Bases, 18(1):345–362, January 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Mouratidis:2009:PMD

[426] Kyriakos Mouratidis, Dimitris Sacharidis, and Hweehwa Pang. Partially materialized digest scheme: an efficient verification method for outsourced databases. VLDB Journal: Very Large Data Bases, 18(1):363–381, January 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Garofalakis:2009:SIB

[427] Minos Garofalakis, Johannes Gehrke, and Divesh Srivastava. Special issue: best papers of VLDB 2007. VLDB Journal: Very Large Data Bases, 18 (2):383–384, April 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Abadi:2009:SSV

[428] Daniel J. Abadi, Adam Marcus, Samuel R. Madden, and Kate Hollenbach. SW-Store: a vertically partitioned DBMS for Semantic Web data management. VLDB Journal: Very Large Data Bases, 18(2):385–406, April 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Arai:2009:AMT

[429] Benjamin Arai, Gautam Das, Dimitrios Gunopulos, and Nick Koudas. Anytime measures for top-k algorithms on exact and fuzzy data sets. VLDB Journal: Very Large Data Bases, 18 (2):407–427, April 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chen:2009:AKD

[430] Bee-Chung Chen, Kristen Lefevre, and Raghu Ramakrishnan. Adversarialknowledge dimensions in data privacy. VLDB Journal: Very Large Data Bases, 18(2):429–467, April 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Dong:2009:DIU

[431] Xin Luna Dong, Alon Halevy, and Cong Yu. Data integration with uncertainty. VLDB Journal: Very Large Data Bases, 18(2):469–500, April 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gedik:2009:CPS

[432] Buğra Gedik, Rajesh R. Bordawekar, and Philip S. Yu. CellJoin: a parallel stream join operator for the Cell processor. VLDB Journal: Very Large Data Bases, 18(2):501–519, April 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Schnaitter:2009:DER

[433] Karl Schnaitter, Joshua Spiegel, and Neoklis Polyzotis. Depth estimation for ranking query optimization. VLDB Journal: Very Large Data Bases, 18 (2):521–542, April 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Shao:2009:EKS

[434] Feng Shao, Lin Guo, Chavdar Botev, Anand Bhaskar, Muthiah Chettiar, Fan Yang, and Jayavel Shanmugasundaram. Efficient keyword search over virtual XML views. VLDB Journal: Very Large Data Bases, 18(2):543–570, April 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wu:2009:GEV

[435] Mingxi Wu and Chris Jermaine. Guessing the extreme values in a data set: a Bayesian method and its applications. VLDB Journal: Very Large Data Bases, 18(2):571–597, April 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Hill:2009:ROJ

[436] Gerhard Hill and Andrew Ross. Reducing outer joins. VLDB Journal: Very Large Data Bases, 18(3):599–610, June 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Keogh:2009:SEI

[437] Eamonn Keogh, Li Wei, Xiaopeng Xi, Michail Vlachos, Sang-Hee Lee, and Pavlos Protopapas. Supporting exact indexing of arbitrarily rotated shapes and periodic time series under Euclidean and warping distance measures. VLDB Journal: Very Large Data Bases, 18(3):611–630, June 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yang:2009:AIO

[438] Yin Yang, Stavros Papadopoulos, Dimitris Papadias, and George Kollios. Authenticated indexing for outsourced spatial databases. VLDB Journal: Very Large Data Bases, 18(3):631–648, June 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Quiane-Ruiz:2009:SAQ

[439] Jorge-Arnulfo Quiané-Ruiz, Philippe Lamarre, and Patrick Valduriez. A self-adaptable query allocation framework for distributed information systems. VLDB Journal: Very Large Data Bases, 18(3):649–674, June 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Deng:2009:IOQ

[440] Ke Deng, Xiaofang Zhou, Heng Tao Shen, Shazia Sadiq, and Xue Li. Instance optimal query processing in spatial networks. VLDB Journal: Very Large Data Bases, 18(3):675–693, June 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yiu:2009:MDT

[441] Man Lung Yiu and Nikos Mamoulis. Multi-dimensional top-k dominating queries. VLDB Journal: Very Large Data Bases, 18(3):695–718, June 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Silva:2009:RTS

[442] Yasin N. Silva, Xiaopeng Xiong, and Walid G. Aref. The RUM-tree: supporting frequent updates in R-trees using memos. VLDB Journal: Very Large Data Bases, 18(3):719–738, June 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Kriakov:2009:STM

[443] Vassil Kriakov, George Kollios, and Alex Delis. Self-tuning management of update-intensive multidimensional data in clusters of workstations. VLDB Journal: Very Large Data Bases, 18 (3):739–764, June 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Cohen:2009:EQS

[444] Sara Cohen. Equivalence of queries that are sensitive to multiplicities. VLDB Journal: Very Large Data Bases, 18(3):765–785, June 2009. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lian:2009:EPP

[445] Xiang Lian and Lei Chen. Efficient processing of probabilistic reverse nearest neighbor queries over uncertain data. VLDB Journal: Very Large Data Bases, 18(3):787–808, June 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Hua:2009:TTQ

[446] Ming Hua, Jian Pei, Ada W. Fu, Xuemin Lin, and Ho-Fung Leung. Top-k typicality queries and efficient query answering methods on large databases. VLDB Journal: Very Large Data Bases, 18(3):809–835, June 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Bawa:2009:PPI

[447] Mayank Bawa, Roberto J. Bayardo, Jr., Rakesh Agrawal, and Jaideep Vaidya. Privacy-preserving indexing of documents on the network. VLDB Journal: Very Large Data Bases, 18 (4):837–856, August 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Fan:2009:QTX

[448] Wenfei Fan, Jeffrey Xu Yu, Jianzhong Li, Bolin Ding, and Lu Qin. Query translation from XPath to SQL in the presence of recursive DTDs. VLDB Journal: Very Large Data Bases, 18 (4):857–883, August 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Malik:2009:RRA

[449] Zaki Malik and Athman Bouguettaya. RATEWeb: Reputation Assessment for Trust Establishment among Web services. VLDB Journal: Very Large Data Bases, 18(4):885–911, August 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wang:2009:CRE

[450] Fusheng Wang, Shaorong Liu, and Peiya Liu. Complex RFID event processing. VLDB Journal: Very Large Data Bases, 18(4):913–931, August 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

DuMouza:2009:LSI

[451] Cédric Du Mouza, Witold Litwin, and Philippe Rigaux. Large-scale indexing of spatial data in distributed repositories: the SD-rtree. VLDB Journal: Very Large Data Bases, 18(4):933–958, August 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zheng:2009:DSI

[452] Baihua Zheng, Wang-Chien Lee, Ken C. Lee, Dik Lun Lee, and Min Shao. A distributed spatial index for error-prone wireless data broadcast. VLDB Journal: Very Large Data Bases, 18(4):959–986, August 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Haas:2009:SIU

[453] Peter J. Haas and Dan Suciu. Special issue on uncertain and probabilistic databases. VLDB Journal: Very Large Data Bases, 18(5):987–988, October 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Sarma:2009:RUD

[454] Anish Das Sarma, Omar Benjelloun, Alon Halevy, Shubha Nabar, and Jennifer Widom. Representing uncertain data: models, properties, and algorithms. VLDB Journal: Very Large Data Bases, 18(5):989–1019, October 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Antova:2009:WBE

[455] Lyublena Antova, Christoph Koch, and Dan Olteanu. 10^{(10^{6})} worlds and beyond: efficient representation and processing of incomplete information. VLDB Journal: Very Large Data Bases, 18(5):1021–1040, October 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Abiteboul:2009:EPX

[456] Serge Abiteboul, Benny Kimelfeld, Yehoshua Sagiv, and Pierre Senellart. On the expressiveness of probabilistic XML models. VLDB Journal: Very Large Data Bases, 18(5): 1041–1064, October 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Sen:2009:PME

[457] Prithviraj Sen, Amol Deshpande, and Lise Getoor. PrDB: managing and exploiting rich correlations in probabilistic databases. VLDB Journal: Very Large Data Bases, 18(5):1065–1090, October 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Re:2009:THQ

[458] Christopher Ré and Dan Suciu. The trichotomy of HAVING queries on a probabilistic database. VLDB Journal: Very Large Data Bases, 18(5): 1091–1116, October 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Kimelfeld:2009:QEP

[459] Benny Kimelfeld, Yuri Kosharovsky, and Yehoshua Sagiv. Query evaluation over probabilistic XML. VLDB Journal: Very Large Data Bases, 18 (5):1117–1140, October 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Hassanzadeh:2009:CPD

[460] Oktie Hassanzadeh and Renée J. Miller. Creating probabilistic databases from duplicated data. VLDB Journal: Very Large Data Bases, 18(5): 1141–1166, October 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wolf:2009:QPI

[461] Garrett Wolf, Aravind Kalavagattu, Hemal Khatri, Raju Balakrishnan, Bhaumik Chokshi, Jianchun Fan, Yi Chen, and Subbarao Kambhampati. Query processing over incomplete autonomous databases: query rewriting using learned data dependencies. VLDB Journal: Very Large Data Bases, 18(5):1167–1190, October 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Keulen:2009:QEK

[462] Maurice Keulen and Ander Keijzer. Qualitative effects of knowledge rules and user feedback in probabilistic data integration. VLDB Journal: Very Large Data Bases, 18(5):1191–1217, October 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chen:2009:SPS

[463] Jinchuan Chen, Reynold Cheng, Mohamed Mokbel, and Chi-Yin Chow. Scalable processing of snapshot and continuous nearest-neighbor queries over one-dimensional uncertain data. VLDB Journal: Very Large Data Bases, 18(5):1219–1240, October 2009. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chen:2010:TFD

[464] Keke Chen and Ling Liu. HE-Tree: a framework for detecting changes in clustering structure for categorical data streams. VLDB Journal: Very Large Data Bases, 18(6):1241–1260, December 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Whang:2010:GER

[465] Steven Euijong Whang, Omar Benjelloun, and Hector Garcia-Molina. Generic entity resolution with negative rules. VLDB Journal: Very Large Data Bases, 18(6):1261–1277, December 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ntarmos:2010:SSI

[466] Nikos Ntarmos, Peter Triantafillou, and Gerhard Weikum. Statistical structures for Internet-scale data management. VLDB Journal: Very Large Data Bases, 18(6):1279–1312, December 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Bramandia:2010:OUR

[467] Ramadhana Bramandia, Jiefeng Cheng, Byron Choi, and Jeffrey Xu Yu. Optimizing updates of recursive XML views of relations. VLDB Journal: Very Large Data Bases, 18(6):1313–1333, December 2010. CODEN VLDBFR.

ISSN 1066-8888 (print), 0949-877X (electronic).

Duntgen:2010:BBM

[468] Christian Düntgen, Thomas Behr, and Ralf Hartmut Güting. Berlin-MOD: a benchmark for moving object databases. VLDB Journal: Very Large Data Bases, 18(6):1335–1368, December 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Mandreoli:2010:PHS

[469] Federica Mandreoli, Riccardo Martoglia, and Pavel Zezula. Principles of Holism for sequential twig pattern matching. VLDB Journal: Very Large Data Bases, 18(6):1369–1392, December 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Buneman:2010:SIB

[470] Peter Buneman, Volker Markl, Beng Chin Ooi, and Kenneth Ross. Special issue: best papers of VLDB 2008. VLDB Journal: Very Large Data Bases, 19 (1):1–2, February 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Cormode:2010:MFF

[471] Graham Cormode and Marios Hadjieleftheriou. Methods for finding frequent items in data streams. VLDB Journal: Very Large Data Bases, 19 (1):3–20, February 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Bruno:2010:CPD

[472] Nicolas Bruno and Surajit Chaudhuri. Constrained physical design tuning. VLDB Journal: Very Large Data Bases, 19(1):21–44, February 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lizorkin:2010:AEO

[473] Dmitry Lizorkin, Pavel Velikhov, Maxim Grinev, and Denis Turdakov. Accuracy estimate and optimization techniques for SimRank computation. VLDB Journal: Very Large Data Bases, 19(1):45–66, February 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Nath:2010:OMV

[474] Suman Nath and Phillip B. Gibbons. Online maintenance of very large random samples on flash storage. VLDB Journal: Very Large Data Bases, 19 (1):67–90, February 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Neumann:2010:RES

[475] Thomas Neumann and Gerhard Weikum. The RDF-3X engine for scalable management of RDF data. VLDB Journal: Very Large Data Bases, 19 (1):91–113, February 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Cormode:2010:ABG

[476] Graham Cormode, Divesh Srivastava, Ting Yu, and Qing Zhang. Anonymizing bipartite graph data using safe groupings. VLDB Journal: Very Large Data Bases, 19(1):115–139, February 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

U:2010:CSA

[477] Leong Hou U., Kyriakos Mouratidis, and Nikos Mamoulis. Continuous spatial assignment of moving users. VLDB Journal: Very Large Data Bases, 19 (2):141–160, April 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Papadopoulos:2010:CAR

[478] Stavros Papadopoulos, Yin Yang, and Dimitris Papadias. Continuous authentication on relational streams. VLDB Journal: Very Large Data Bases, 19 (2):161–180, April 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhang:2010:UMS

[479] Zhenjie Zhang, Hua Lu, Beng Chin Ooi, and Anthony K. Tung. Understanding the meaning of a shifted sky: a general framework on extending skyline query. VLDB Journal: Very Large Data Bases, 19(2):181–201, April 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lo:2010:FTD

[480] Eric Lo, Carsten Binnig, Donald Kossmann, M. Tamer Özsu, and Wing-Kai Hon. A framework for testing DBMS features. VLDB Journal: Very Large Data Bases, 19(2):203–230, April 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Bonifati:2010:SMQ

[481] Angela Bonifati, Elaine Chang, Terence Ho, Laks V. Lakshmanan, Rachel Pottinger, and Yongik Chung. Schema mapping and query translation in heterogeneous P2P XML databases. VLDB Journal: Very Large Data Bases, 19(2):231–256, April 2010. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Morfonios:2010:RCL

[482] Konstantinos Morfonios and Yannis Ioannidis. Revisiting the cube lifecycle in the presence of hierarchies. VLDB Journal: Very Large Data Bases, 19 (2):257–282, April 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhang:2010:TBP

[483] Wenjie Zhang, Xuemin Lin, Ying Zhang, Jian Pei, and Wei Wang. Threshold-based probabilistic top-k dominating queries. VLDB Journal: Very Large Data Bases, 19(2):283–305, April 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Nutanong:2010:AEV

[484] Sarana Nutanong, Rui Zhang, Egemen Tanin, and Lars Kulik. Analysis and evaluation of V*-kNN: an efficient algorithm for moving kNN queries. VLDB Journal: Very Large Data Bases, 19(3):307–332, June 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lee:2010:ZSE

[485] Ken C. Lee, Wang-Chien Lee, Baihua Zheng, Huajing Li, and Yuan Tian. Z-SKY: an efficient skyline query processing framework based on Z-order. VLDB Journal: Very Large Data Bases, 19(3):333–362, June 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yiu:2010:ESS

[486] Man Lung Yiu, Gabriel Ghinita, Christian S. Jensen, and Panos Kalnis. Enabling search services on outsourced private spatial data. VLDB Journal: Very Large Data Bases, 19(3):363–384, June 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Hintoglu:2010:SMP

[487] Ayça Azgin Hintoglu and Yücel Saygın. Suppressing microdata to prevent classification based inference. VLDB Journal: Very Large Data Bases, 19(3):385– 410, June 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Jin:2010:SWT

[488] Cheqing Jin, Ke Yi, Lei Chen, Jeffrey Xu Yu, and Xuemin Lin. Slidingwindow top-k queries on uncertain streams. VLDB Journal: Very Large Data Bases, 19(3):411–435, June 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Pang:2010:EPE

[489] Hweehwa Pang, Xuhua Ding, and Baihua Zheng. Efficient processing of exact top-k queries over disk-resident sorted lists. VLDB Journal: Very Large Data Bases, 19(3):437–456, June 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Murugesan:2010:EPP

[490] Mummoorthy Murugesan, Wei Jiang, Chris Clifton, Luo Si, and Jaideep Vaidya. Efficient privacy-preserving similar document detection. VLDB Journal: Very Large Data Bases, 19 (4):457–475, August 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Soliman:2010:SRQ

[491] Mohamed A. Soliman, Ihab F. Ilyas, and Shalev Ben-David. Supporting ranking queries on uncertain and incomplete data. VLDB Journal: Very Large Data Bases, 19(4):477–501, August 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lee:2010:SCE

[492] Ki-Hoon Lee, Kyu-Young Whang, Wook-Shin Han, and Min-Soo Kim. Structural consistency: enabling XML keyword search to eliminate spurious results consistently. VLDB Journal: Very Large Data Bases, 19(4):503–529, August 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lucchese:2010:RPT

[493] Claudio Lucchese, Michail Vlachos, Deepak Rajan, and Philip S. Yu. Rights protection of trajectory datasets with nearest-neighbor preservation. VLDB Journal: Very Large Data Bases, 19(4):531–556, August 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhang:2010:SMA

[494] Rui Zhang, Nick Koudas, Beng Chin Ooi, Divesh Srivastava, and Pu Zhou. Streaming multiple aggregations using phantoms. VLDB Journal: Very Large Data Bases, 19(4):557–583, August 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Jeung:2010:PPP

[495] Hoyoung Jeung, Man Lung Yiu, Xiaofang Zhou, and Christian S. Jensen. Path prediction and predictive range querying in road network databases. VLDB Journal: Very Large Data Bases, 19(4):585–602, August 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ali:2010:MAA

[496] Mohammed Eunus Ali, Egemen Tanin, Rui Zhang, and Lars Kulik. A motionaware approach for efficient evaluation of continuous queries on 3D object databases. VLDB Journal: Very Large Data Bases, 19(5):603–632, October 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Askitis:2010:ESC

[497] Nikolas Askitis and Ranjan Sinha. Engineering scalable, cache and space efficient tries for strings. VLDB Journal: Very Large Data Bases, 19(5):633– 660, October 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wu:2010:EEG

[498] Xiaoying Wu, Dimitri Theodoratos, and Calisto Zuzarte. Efficient evaluation of generalized tree-pattern queries on XML streams. VLDB Journal: Very Large Data Bases, 19(5):661–686, October 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Guting:2010:ENN

[499] Ralf Hartmut Güting, Thomas Behr, and Jianqiu Xu. Efficient k-nearest neighbor search on moving object trajectories. VLDB Journal: Very Large Data Bases, 19(5):687–714, October 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Li:2010:TQT

[500] Feifei Li, Ke Yi, and Wangchao Le. Top-k queries on temporal data. VLDB Journal: Very Large Data Bases, 19 (5):715–733, October 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Duda:2010:PBI

[501] Cristian Duda, Donald Kossmann, and Chong Zhou. Predicate-based indexing for desktop search. VLDB Journal: Very Large Data Bases, 19(5):735– 758, October 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Bohm:2010:F

[502] Klemens Böhm and Laks V. Lakshmanan. Foreword. VLDB Journal: Very Large Data Bases, 19(6): 759–760, December 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Carmel:2010:SBW

[503] David Carmel, Haggai Roitman, and Elad Yom-Tov. Social bookmark weighting for search and recommendation. VLDB Journal: Very Large Data Bases, 19(6):761–775, December 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Squicciarini:2010:PPS

[504] Anna C. Squicciarini, Mohamed Shehab, and Joshua Wede. Privacy policies for shared content in social network sites. VLDB Journal: Very Large Data Bases, 19(6):777–796, December 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Hay:2010:RSR

[505] Michael Hay, Gerome Miklau, David Jensen, Don Towsley, and Chao Li. Resisting structural re-identification in anonymized social networks. VLDB Journal: Very Large Data Bases, 19 (6):797–823, December 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gruhl:2010:MSI

[506] Daniel Gruhl, Meena Nagarajan, Jan Pieper, Christine Robson, and Amit Sheth. Multimodal social intelligence in a real-time dashboard system. VLDB Journal: Very Large Data Bases, 19 (6):825–848, December 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Benz:2010:SBP

[507] Dominik Benz, Andreas Hotho, Robert Jäschke, Beate Krause, Folke Mitzlaff, Christoph Schmitz, and Gerd Stumme. The social bookmark and publication management system bibsonomy. VLDB Journal: Very Large Data Bases, 19 (6):849–875, December 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Roy:2010:SEG

[508] Senjuti Basu Roy, Sihem Amer-Yahia, Ashish Chawla, Gautam Das, and Cong Yu. Space efficiency in group recommendation. VLDB Journal: Very Large Data Bases, 19(6):877–900, December 2010. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Li:2011:PBK

[509] Guoliang Li, Jianhua Feng, Xiaofang Zhou, and Jianyong Wang. Providing built-in keyword search capabilities in RDBMS. VLDB Journal: Very Large Data Bases, 20(1):1–19, February 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Cai:2011:SKD

[510] Deng Cai, Xiaofei He, and Jiawei Han. Speed up kernel discriminant analysis. VLDB Journal: Very Large Data Bases, 20(1):21–33, February 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Qin:2011:SKS

[511] Lu Qin, Jeffrey Xu Yu, and Lijun Chang. Scalable keyword search on large data streams. VLDB Journal: Very Large Data Bases, 20(1):35–57, February 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Cao:2011:SSA

[512] Jianneng Cao, Panagiotis Karras, Panos Kalnis, and Kian-Lee Tan.

SABRE: a Sensitive Attribute Bucketization and REdistribution framework for t-closeness. VLDB Journal: Very Large Data Bases, 20(1):59–81, February 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Terrovitis:2011:LGR

[513] Manolis Terrovitis, Nikos Mamoulis, and Panos Kalnis. Local and global recoding methods for anonymizing setvalued data. VLDB Journal: Very Large Data Bases, 20(1):83–106, February 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lian:2011:PIR

[514] Xiang Lian and Lei Chen. Probabilistic inverse ranking queries in uncertain databases. VLDB Journal: Very Large Data Bases, 20(1):107–127, February 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Hua:2011:RQU

[515] Ming Hua, Jian Pei, and Xuemin Lin. Ranking queries on uncertain data. VLDB Journal: Very Large Data Bases, 20(1):129–153, February 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Abiteboul:2011:SIB

[516] Serge Abiteboul, Volker Markl, Tova Milo, and Jignesh Patel. Special issue: best papers of VLDB 2009. VLDB Journal: Very Large Data Bases, 20 (2):155–156, April 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Mindolin:2011:PEP

[517] Denis Mindolin and Jan Chomicki. Preference elicitation in prioritized skyline queries. VLDB Journal: Very Large Data Bases, 20(2):157–182, April 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Denev:2011:SFD

[518] Dimitar Denev, Arturas Mazeika, Marc Spaniol, and Gerhard Weikum. The SHARC framework for data quality in Web archiving. VLDB Journal: Very Large Data Bases, 20(2):183–207, April 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Elmeleegy:2011:HRT

[519] Hazem Elmeleegy, Jayant Madhavan, and Alon Halevy. Harvesting relational tables from lists on the Web. VLDB Journal: Very Large Data Bases, 20 (2):209–226, April 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Candea:2011:PPH

[520] George Candea, Neoklis Polyzotis, and Radek Vingralek. Predictable performance and high query concurrency for data analytics. VLDB Journal: Very Large Data Bases, 20(2):227–248, April 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Li:2011:UAR

[521] Jian Li, Barna Saha, and Amol Deshpande. A unified approach to ranking in probabilistic databases. VLDB Journal: Very Large Data Bases, 20 (2):249–275, April 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gottlob:2011:NOS

[522] Georg Gottlob, Reinhard Pichler, and Vadim Savenkov. Normalization and optimization of schema mappings. VLDB Journal: Very Large Data Bases, 20(2):277–302, April 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Cho:2011:LRM

[523] Chung-Wen Cho, Yi-Hung Wu, Show-Jane Yen, Ying Zheng, and Arbee L. Chen. On-line rule matching for event prediction. VLDB Journal: Very Large Data Bases, 20(3):303–334, June 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Liu:2011:MLD

[524] Jun Liu, Lu Jiang, Zhaohui Wu, Qinghua Zheng, and Yanan Qian. Mining learning-dependency between knowledge units from text. VLDB Journal: Very Large Data Bases, 20 (3):335–345, June 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wang:2011:LBM

[525] Rui Wang, Betty Salzberg, and David Lomet. Log-based middleware server recovery with transaction support. VLDB Journal: Very Large Data Bases, 20(3):347–370, June 2011. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gao:2011:CVN

[526] Yunjun Gao, Baihua Zheng, Gencai Chen, Qing Li, and Xiaofa Guo. Continuous visible nearest neighbor query processing in spatial databases. VLDB Journal: Very Large Data Bases, 20 (3):371–396, June 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Luo:2011:QRI

[527] Bo Luo, Dongwon Lee, Wang-Chien Lee, and Peng Liu. QFilter: rewriting insecure XML queries to secure ones using non-deterministic finite automata. VLDB Journal: Very Large Data Bases, 20(3):397–415, June 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Motahari-Nezhad:2011:ECP

[528] Hamid Reza Motahari-Nezhad, Regis Saint-Paul, Fabio Casati, and Boualem Benatallah. Event correlation for process discovery from web service interaction logs. VLDB Journal: Very Large Data Bases, 20(3):417–444, June 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chakrabarti:2011:IDQ

[529] Soumen Chakrabarti, Amit Pathak, and Manish Gupta. Index design and query processing for graph conductance search. VLDB Journal: Very Large Data Bases, 20(3):445–470, June 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chen:2011:PAD

[530] Shaoping Chen, Yi-Cheng Tu, and Yuni Xia. Performance analysis of a dual-tree algorithm for computing spatial distance histograms. VLDB Journal: Very Large Data Bases, 20 (4):471–494, August 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Fan:2011:DCR

[531] Wenfei Fan, Hong Gao, Xibei Jia, Jianzhong Li, and Shuai Ma. Dynamic constraints for record matching. VLDB Journal: Very Large Data Bases, 20 (4):495–520, August 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Cheng:2011:FGQ

[532] James Cheng, Yiping Ke, Ada Wai-Chee Fu, and Jeffrey Xu Yu. Fast graph query processing with a low-cost index. VLDB Journal: Very Large Data Bases, 20(4):521–539, August 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Mascetti:2011:PGS

[533] Sergio Mascetti, Dario Freni, Claudio Bettini, X. Sean Wang, and Sushil Jajodia. Privacy in geo-social networks: proximity notification with untrusted service providers and curious buddies. VLDB Journal: Very Large Data Bases, 20(4):541–566, August 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Mohammed:2011:AMG

[534] Noman Mohammed, Benjamin C. Fung, and Mourad Debbabi. Anonymity meets game theory: secure data integration with malicious participants. VLDB Journal: Very Large Data Bases, 20(4):567–588, August 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ahmad:2011:IAS

[535] Mumtaz Ahmad, Ashraf Aboulnaga, Shivnath Babu, and Kamesh Munagala. Interaction-aware scheduling of report-generation workloads. VLDB Journal: Very Large Data Bases, 20 (4):589–615, August 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Li:2011:EFF

[536] Guoliang Li, Shengyue Ji, Chen Li, and Jianhua Feng. Efficient fuzzy full-text type-ahead search. VLDB Journal: Very Large Data Bases, 20 (4):617–640, August 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Guting:2011:SID

[537] Ralf Hartmut Güting and Nikos Mamoulis. Special issue on data management for mobile services. VLDB Journal: Very Large Data Bases, 20 (5):641–642, October 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Popa:2011:INT

[538] Iulian Sandu Popa, Karine Zeitouni, Vincent Oria, Dominique Barth, and Sandrine Vial. Indexing in-network trajectory flows. VLDB Journal: Very Large Data Bases, 20(5):643–669, October 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lange:2011:ERT

[539] Ralph Lange, Frank Dürr, and Kurt Rothermel. Efficient real-time trajectory tracking. VLDB Journal: Very Large Data Bases, 20(5):671–694, October 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Giannotti:2011:UCH

[540] Fosca Giannotti, Mirco Nanni, Dino Pedreschi, Fabio Pinelli, Chiara Renso, Salvatore Rinzivillo, and Roberto Trasarti. Unveiling the complexity of human mobility by querying and mining massive trajectory data. VLDB Journal: Very Large Data Bases, 20 (5):695–719, October 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Timko:2011:SSA

[541] Igor Timko, Michael Böhlen, and Johann Gamper. Sequenced spatiotemporal aggregation for coarse query granularities. VLDB Journal: Very Large Data Bases, 20(5):721–741, October 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Guo:2011:DBS

[542] Xi Guo, Baihua Zheng, Yoshiharu Ishikawa, and Yunjun Gao. Directionbased surrounder queries for mobile recommendations. VLDB Journal: Very Large Data Bases, 20(5):743–766, October 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Trajcevski:2011:RCN

[543] Goce Trajcevski, Roberto Tamassia, Isabel F. Cruz, Peter Scheuermann, David Hartglass, and Christopher Zamierowski. Ranking continuous nearest neighbors for uncertain trajectories. VLDB Journal: Very Large Data Bases, 20(5):767–791, October 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Rao:2011:STE

[544] Weixiong Rao, Lei Chen, and Ada Wai-Chee Fu. STAIRS: Towards efficient full-text filtering and dissemination in DHT environments. VLDB Journal: Very Large Data Bases, 20 (6):793–817, December 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lian:2011:STS

[545] Xiang Lian and Lei Chen. Shooting top-k stars in uncertain databases. VLDB Journal: Very Large Data Bases, 20(6):819–840, December 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Perez-Sorrosal:2011:ESC

[546] Francisco Perez-Sorrosal, Marta Patiño-Martinez, Ricardo Jimenez-Peris, and Bettina Kemme. Elastic SI-Cache: consistent and scalable caching in multitier architectures. VLDB Journal: Very Large Data Bases, 20(6):841–865, December 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Moga:2011:USC

[547] Alexandru Moga, Irina Botan, and Nesime Tatbul. UpStream: storagecentric load management for streaming applications with update semantics. VLDB Journal: Very Large Data Bases, 20(6):867–892, December 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wong:2011:MBR

[548] Raymond Chi-Wing Wong, M. Tamer Özsu, Ada Wai-Chee Fu, Philip S. Yu, Lian Liu, and Yubao Liu. Maximizing bichromatic reverse nearest neighbor for Lp-norm in two- and threedimensional spaces. VLDB Journal: Very Large Data Bases, 20(6):893–919, December 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Tiakas:2011:PPS

[549] Eleftherios Tiakas, Apostolos N. Papadopoulos, and Yannis Manolopoulos. Progressive processing of subspace dominating queries. VLDB Journal: Very Large Data Bases, 20(6): 921–948, December 2011. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Mueller:2012:SNF

[550] Rene Mueller, Jens Teubner, and Gustavo Alonso. Sorting networks on FP-GAs. VLDB Journal: Very Large Data Bases, 21(1):1–23, February 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Georgoulas:2012:DSE

[551] Konstantinos Georgoulas and Yannis Kotidis. Distributed similarity estimation using derived dimensions. VLDB Journal: Very Large Data Bases, 21 (1):25–50, February 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Deutch:2012:TIT

[552] Daniel Deutch and Tova Milo. Type inference and type checking for queries over execution traces. VLDB Journal: Very Large Data Bases, 21(1):51– 68, February 2012. CODEN VLDBFR.

ISSN 1066-8888 (print), 0949-877X (electronic).

Cheema:2012:CRN

[553] Muhammad Aamir Cheema, Wenjie Zhang, Xuemin Lin, Ying Zhang, and Xuefei Li. Continuous reverse k nearest neighbors queries in Euclidean space and in spatial networks. VLDB Journal: Very Large Data Bases, 21(1):69– 95, February 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zou:2012:APM

[554] Lei Zou, Lei Chen, M. Tamer Özsu, and Dongyan Zhao. Answering pattern match queries in large graph databases via graph embedding. VLDB Journal: Very Large Data Bases, 21 (1):97–120, February 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Hartmann:2012:DES

[555] Sven Hartmann, Markus Kirchberg, and Sebastian Link. Design by example for SQL table definitions with functional dependencies. VLDB Journal: Very Large Data Bases, 21(1): 121–144, February 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Guravannavar:2012:WSO

[556] Ravindra Guravannavar, S. Sudarshan, Ajit A. Diwan, and Ch. Sobhan Babu. Which sort orders are interesting? VLDB Journal: Very Large Data Bases, 21(1):145–165, February 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Atzeni:2012:SIB

[557] Paolo Atzeni, Elisa Bertino, Masaru Kitsuregawa, and Kian-Lee Tan. Special issue: best papers of VLDB 2010. VLDB Journal: Very Large Data Bases, 21(2):167–168, April 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Bu:2012:HAL

[558] Yingyi Bu, Bill Howe, Magdalena Balazinska, and Michael D. Ernst. The HaLoop approach to large-scale iterative data analysis. VLDB Journal: Very Large Data Bases, 21(2):169–190, April 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Alexe:2012:MCI

[559] Bogdan Alexe, Mauricio Hernández, Lucian Popa, and Wang-Chiew Tan. MapMerge: correlating independent schema mappings. VLDB Journal: Very Large Data Bases, 21(2):191–211, April 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Fan:2012:TCF

[560] Wenfei Fan, Jianzhong Li, Shuai Ma, Nan Tang, and Wenyuan Yu. Towards certain fixes with editing rules and master data. VLDB Journal: Very Large Data Bases, 21(2):213–238, April 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Johnson:2012:SWA

[561] Ryan Johnson, Ippokratis Pandis, Radu Stoica, Manos Athanassoulis, and Anastasia Ailamaki. Scalability of write-ahead logging on multicore and multisocket hardware. VLDB Journal: Very Large Data Bases, 21(2):239–263, April 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chen:2012:AUP

[562] Su Chen, Beng Chin Ooi, and Zhenjie Zhang. An adaptive updating protocol for reducing moving object database workload. VLDB Journal: Very Large Data Bases, 21(2):265–286, April 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Fusco:2012:RTC

[563] Francesco Fusco, Michail Vlachos, and Marc Ph. Stoecklin. Real-time creation of bitmap indexes on streaming network data. VLDB Journal: Very Large Data Bases, 21(3):287–307, June 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gordevicus:2012:PTA

[564] Juozas Gordevičus, Johann Gamper, and Michael Böhlen. Parsimonious temporal aggregation. VLDB Journal: Very Large Data Bases, 21(3):309–332, June 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Hore:2012:SMR

[565] Bijit Hore, Sharad Mehrotra, Mustafa Canim, and Murat Kantarcioglu. Secure multidimensional range queries over outsourced data. VLDB Journal: Very Large Data Bases, 21(3):333–358, June 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Hose:2012:SSP

[566] Katja Hose and Akrivi Vlachou. A survey of skyline processing in highly distributed environments. VLDB Journal: Very Large Data Bases, 21(3):359–384, June 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gong:2012:EMU

[567] Jian Gong, Reynold Cheng, and David W. Cheung. Efficient management of uncertainty in XML schema matching. VLDB Journal: Very Large Data Bases, 21(3):385–409, June 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Cao:2012:SSA

[568] Yu Cao, Ramadhana Bramandia, Chee-Yong Chan, and Kian-Lee Tan. Sort-sharing-aware query processing. VLDB Journal: Very Large Data Bases, 21(3):411–436, June 2012. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Feng:2012:TJT

[569] Jianhua Feng, Jiannan Wang, and Guoliang Li. Trie-join: a trie-based method for efficient string similarity joins. VLDB Journal: Very Large Data Bases, 21(4):437–461, August 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Augsten:2012:WGA

[570] Nikolaus Augsten, Michael Böhlen, Curtis Dyreson, and Johann Gamper. Windowed pq-grams for approximate joins of data-centric XML. VLDB Journal: Very Large Data Bases, 21 (4):463–488, August 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhou:2012:ESM

[571] Xiangmin Zhou, Xiaofang Zhou, Lei Chen, and Athman Bouguettaya. Efficient subsequence matching over large video databases. VLDB Journal: Very Large Data Bases, 21(4):489–508, August 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yildirim:2012:GSI

[572] Hilmi Yildirim, Vineet Chaoji, and Mohammed J. Zaki. GRAIL: a scalable index for reachability queries in very large graphs. VLDB Journal: Very Large Data Bases, 21(4):509–534, August 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Xu:2012:EES

[573] Jia Xu, Zhenjie Zhang, Anthony K. Tung, and Ge Yu. Efficient and effective similarity search over probabilistic data based on Earth Mover’s Distance. VLDB Journal: Very Large Data Bases, 21(4):535–559, August 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhang:2012:HOA

[574] Rui Zhang, Jianzhong Qi, Dan Lin, Wei Wang, and Raymond Chi-Wing Wong. A highly optimized algorithm for continuous intersection join queries over moving objects. VLDB Journal: Very Large Data Bases, 21(4):561–586, August 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lehner:2012:SSL

[575] Wolfgang Lehner and Michael J. Franklin. Special section on large-scale analytics. VLDB Journal: Very Large Data Bases, 21(5):587–588, October 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wolf:2012:OSM

[576] Joel Wolf, Andrey Balmin, Deepak Rajan, Kirsten Hildrum, Rohit Khandekar, Sujay Parekh, Kun-Lung Wu, and Rares Vernica. On the optimization of schedules for MapReduce workloads in the presence of shared scans. VLDB Journal: Very Large Data Bases, 21(5):589–609, October 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhou:2012:SPD

[577] Jingren Zhou, Nicolas Bruno, Ming-Chuan Wu, Per-Ake Larson, Ronnie Chaiken, and Darren Shakib. SCOPE: parallel databases meet MapReduce. VLDB Journal: Very Large Data Bases, 21(5):611–636, October 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Kang:2012:GEA

[578] U. Kang, Hanghang Tong, Jimeng Sun, Ching-Yung Lin, and Christos Faloutsos. gbase: an efficient analysis platform for large graphs. VLDB Journal: Very Large Data Bases, 21 (5):637–650, October 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Tran:2012:CMP

[579] Thanh T. Tran, Liping Peng, Yanlei Diao, Andrew Mcgregor, and Anna Liu. CLARO: modeling and processing uncertain data streams. VLDB Journal: Very Large Data Bases, 21 (5):651–676, October 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Helmer:2012:MSS

[580] Sven Helmer, Nikolaus Augsten, and Michael Böhlen. Measuring structural similarity of semistructured data based on information-theoretic approaches. VLDB Journal: Very Large Data Bases, 21(5):677–702, October 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Cheema:2012:EPS

[581] Muhammad Aamir Cheema, Wenjie Zhang, Xuemin Lin, and Ying Zhang. Efficiently processing snapshot and continuous reverse k nearest neighbors queries. VLDB Journal: Very Large Data Bases, 21(5):703–728, October 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zheng:2012:SQP

[582] Kai Zheng, Xiaofang Zhou, Pui Cheong Fung, and Kexin Xie. Spatial query processing for fuzzy objects. VLDB Journal: Very Large Data Bases, 21 (5):729–751, October 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Li:2012:MFS

[583] Jianzhong Li, Zhaonian Zou, and Hong Gao. Mining frequent subgraphs over uncertain graph databases under probabilistic semantics. VLDB Journal: Very Large Data Bases, 21(6):753–777, December 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Vergoulis:2012:ARS

[584] Thanasis Vergoulis, Theodore Dalamagas, Dimitris Sacharidis, and Timos Sellis. Approximate regional sequence matching for genomic databases. VLDB Journal: Very Large Data Bases, 21(6):779–795, December 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wu:2012:FES

[585] Dingming Wu, Gao Cong, and Christian S. Jensen. A framework for efficient spatial web object retrieval. VLDB Journal: Very Large Data Bases, 21(6):797–822, December 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Arenas:2012:QLB

[586] Marcelo Arenas, Jorge Pérez, Juan Reutter, and Cristian Riveros. Query language-based inverses of schema mappings: semantics, computation, and closure properties. VLDB Journal: Very Large Data Bases, 21(6): 823–842, December 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Bravo:2012:CRX

[587] Loreto Bravo, James Cheney, Irini Fundulaki, and Ricardo Segovia. Consistency and repair for XML writeaccess control policies. VLDB Journal: Very Large Data Bases, 21(6):

843–867, December 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chang:2012:EDD

[588] Lijun Chang, Jeffrey Xu Yu, Lu Qin, Hong Cheng, and Miao Qiao. The exact distance to destination in undirected world. VLDB Journal: Very Large Data Bases, 21(6):869–888, December 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Soh:2012:AEE

[589] Kheng Hong Soh, Ba Quan Truong, and Sourav S. Bhowmick. ANDES: efficient evaluation of NOT-twig queries in relational databases. VLDB Journal: Very Large Data Bases, 21(6): 889–914, December 2012. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lehner:2013:SIB

[590] Wolfgang Lehner and Sunita Sarawagi. Special issue on best papers of VLDB 2011. VLDB Journal: Very Large Data Bases, 22(1):1–2, February 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Tzoumas:2013:EAG

[591] Kostas Tzoumas, Amol Deshpande, and Christian S. Jensen. Efficiently adapting graphical models for selectivity estimation. VLDB Journal: Very Large Data Bases, 22(1):3–27, February 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Minhas:2013:RTH

[592] Umar Farooq Minhas, Shriram Rajagopalan, Brendan Cully, Ashraf Aboulnaga, Kenneth Salem, and Andrew Warfield. RemusDB: transparent high availability for database systems. VLDB Journal: Very Large Data Bases, 22(1):29–45, February 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Furche:2013:OLS

[593] Tim Furche, Georg Gottlob, Giovanni Grasso, Christian Schallhart, and Andrew Sellers. OXPath: a language for scalable data extraction, automation, and crawling on the deep web. VLDB Journal: Very Large Data Bases, 22 (1):47–72, February 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Curino:2013:ADS

[594] Carlo Curino, Hyun Jin Moon, Alin Deutsch, and Carlo Zaniolo. Automating the database schema evolution process. VLDB Journal: Very Large Data Bases, 22(1):73–98, February 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ramesh:2013:KSF

[595] Aditya Ramesh, S. Sudarshan, Purva Joshi, and Manisha Naik Gaonkar. Keyword search on form results. VLDB Journal: Very Large Data Bases, 22 (1):99–123, February 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Dieng:2013:MFC

[596] Cheikh Tidiane Dieng, Tao-Yuan Jen, Dominique Laurent, and Nicolas Spyratos. Mining frequent conjunctive queries using functional and inclusion dependencies. VLDB Journal: Very Large Data Bases, 22(2):125–150, April 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Tozun:2013:SDB

[597] Pinar Tözün, Ippokratis Pandis, Ryan Johnson, and Anastasia Ailamaki. Scalable and dynamically balanced shared-everything OLTP with physiological partitioning. VLDB Journal: Very Large Data Bases, 22(2):151–175, April 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wu:2013:SXS

[598] Xiaoying Wu and Dimitri Theodoratos. A survey on XML streaming evaluation techniques. VLDB Journal: Very Large Data Bases, 22(2):177–202, April 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lu:2013:ADU

[599] Wentian Lu, Gerome Miklau, and Neil Immerman. Auditing a database under retention policies. VLDB Journal: Very Large Data Bases, 22(2):203–228, April 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yuan:2013:LLB

[600] Dayu Yuan and Prasenjit Mitra. Lindex: a lattice-based index for graph databases. VLDB Journal: Very Large Data Bases, 22(2):229–252, April 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Song:2013:CDH

[601] Shaoxu Song, Lei Chen, and Philip S. Yu. Comparable dependencies over heterogeneous data. VLDB Journal: Very Large Data Bases, 22(2):253–274, April 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Qiao:2013:CWC

[602] Miao Qiao, Hong Cheng, Lu Qin, Jeffrey Xu Yu, Philip S. Yu, and Lijun Chang. Computing weight constraint reachability in large networks. VLDB Journal: Very Large Data Bases, 22 (3):275–294, June 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Toyoda:2013:PDD

[603] Machiko Toyoda, Yasushi Sakurai, and Yoshiharu Ishikawa. Pattern discovery in data streams under the time warping distance. VLDB Journal: Very Large Data Bases, 22(3):295–318, June 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Xie:2013:UDV

[604] Xike Xie, Reynold Cheng, Man Lung Yiu, Liwen Sun, and Jinchuan Chen. UV-diagram: a Voronoi diagram for uncertain spatial databases. VLDB Journal: Very Large Data Bases, 22 (3):319–344, June 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhu:2013:HEQ

[605] Yuanyuan Zhu, Lu Qin, Jeffrey Xu Yu, Yiping Ke, and Xuemin Lin. High efficiency and quality: large graphs matching. VLDB Journal: Very Large Data Bases, 22(3):345–368, June 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Baca:2013:OEG

[606] Radim Baca, Michal Krátký, Tok Wang Ling, and Jiaheng Lu. Optimal and efficient generalized twig pattern processing: a combination of preorder and postorder filterings. VLDB Journal: Very Large Data Bases, 22(3):369–393, June 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Silva:2013:SQT

[607] Yasin N. Silva, Walid G. Aref, Per-Ake Larson, Spencer S. Pearson, and Mohamed H. Ali. Similarity queries: their conceptual evaluation, transformations, and processing. VLDB Journal: Very Large Data Bases, 22(3):395– 420, June 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Dindar:2013:MES

[608] Nihal Dindar, Nesime Tatbul, Renée J. Miller, Laura M. Haas, and Irina Botan. Modeling the execution semantics of stream processing engines with SECRET. VLDB Journal: Very Large Data Bases, 22(4):421–446, August 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Elghandour:2013:RXP

[609] Iman Elghandour, Ashraf Aboulnaga, Daniel C. Zilio, and Calisto Zuzarte. Recommending XML physical designs for XML databases. VLDB Journal: Very Large Data Bases, 22(4):447–470, August 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Mazuran:2013:EPD

[610] Mirjana Mazuran, Edoardo Serra, and Carlo Zaniolo. Extending the power of datalog recursion. VLDB Journal: Very Large Data Bases, 22(4):471–493, August 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Galpin:2013:QAO

[611] Ixent Galpin, Alvaro A. Fernandes, and Norman W. Paton. QoS-aware optimization of sensor network queries. VLDB Journal: Very Large Data Bases, 22(4):495–517, August 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Deutch:2013:TQW

[612] Daniel Deutch, Tova Milo, and Neoklis Polyzotis. Top-k queries over Web applications. VLDB Journal: Very Large Data Bases, 22(4):519–542, August 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gao:2013:OSD

[613] Jun Gao, Jeffrey Xu Yu, Ruoming Jin, Jiashuai Zhou, Tengjiao Wang, and Dongqing Yang. Outsourcing shortest distance computing with privacy protection. VLDB Journal: Very Large Data Bases, 22(4):543–559, August 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Kalashnikov:2013:SEF

[614] Dmitri V. Kalashnikov. Super-EGO: fast multi-dimensional similarity join. VLDB Journal: Very Large Data Bases, 22(4):561–585, August 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Brambilla:2013:SIS

[615] Marco Brambilla, Stefano Ceri, and Alon Halevy. Special issue on structured and crowd-sourced data on the Web. VLDB Journal: Very Large Data Bases, 22(5):587–588, October 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Goasdoue:2013:GTT

[616] François Goasdoué, Konstantinos Karanasos, Yannis Katsis, Julien Leblay, Ioana Manolescu, and Stamatis Zampetakis. Growing triples on trees: an XML–RDF hybrid model for annotated documents. VLDB Journal: Very Large Data Bases, 22(5):589–613, October 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Furche:2013:OKA

[617] Tim Furche, Georg Gottlob, Giovanni Grasso, Xiaonan Guo, Giorgio Orsi, and Christian Schallhart. The ontological key: automatically understanding and integrating forms to access the deep Web. VLDB Journal: Very Large Data Bases, 22(5):615–640, October 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Bozzon:2013:ESF

[618] Alessandro Bozzon, Marco Brambilla, Stefano Ceri, and Davide Mazza. Exploratory search framework for Web data sources. VLDB Journal: Very Large Data Bases, 22(5):641–663, October 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Demartini:2013:LSL

[619] Gianluca Demartini, Djellel Eddine Difallah, and Philippe Cudré-Mauroux. Large-scale linked data integration using probabilistic reasoning and crowdsourcing. VLDB Journal: Very Large Data Bases, 22(5):665–687, October 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Sagi:2013:SMP

[620] Tomer Sagi and Avigdor Gal. Schema matching prediction with applications to data source discovery and dynamic ensembling. VLDB Journal: Very Large Data Bases, 22(5):689–710, October 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lee:2013:HEC

[621] Jongwuk Lee, Hyunsouk Cho, Jin-Woo Park, Young-Rok Cha, Seung-Won Hwang, Zaiqing Nie, and Ji-Rong Wen. Hybrid entity clustering using crowds and data. VLDB Journal: Very Large Data Bases, 22 (5):711–726, October 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhao:2013:EPG

[622] Xiang Zhao, Chuan Xiao, Xuemin Lin, Wei Wang, and Yoshiharu Ishikawa. Efficient processing of graph similarity queries with edit distance constraints. VLDB Journal: Very Large Data Bases, 22(6):727–752, December 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gemulla:2013:NUI

[623] Rainer Gemulla, Peter J. Haas, and Wolfgang Lehner. Non-uniformity issues and workarounds in bounded-size sampling. VLDB Journal: Very Large Data Bases, 22(6):753–772, December 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Whang:2013:JER

[624] Steven Euijong Whang and Hector Garcia-Molina. Joint entity resolution on multiple datasets. VLDB Journal: Very Large Data Bases, 22(6): 773–795, December 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Xu:2013:DPH

[625] Jia Xu, Zhenjie Zhang, Xiaokui Xiao, Yin Yang, Ge Yu, and Marianne Winslett. Differentially private histogram publication. VLDB Journal: Very Large Data Bases, 22(6):797–822, December 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Fink:2013:AAP

[626] Robert Fink, Jiewen Huang, and Dan Olteanu. Anytime approximation in probabilistic databases. VLDB Journal: Very Large Data Bases, 22(6): 823–848, December 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Drosou:2013:YER

[627] Marina Drosou and Evaggelia Pitoura. YmalDB: exploring relational databases via result-driven recommendations. VLDB Journal: Very Large Data Bases, 22(6):849–874, December 2013. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Johnson:2014:EUC

[628] Ryan Johnson, Ippokratis Pandis, and Anastasia Ailamaki. Eliminating unscalable communication in transaction processing. VLDB Journal: Very Large Data Bases, 23(1):1–23, February 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhou:2014:EQP

[629] Junfeng Zhou, Zhifeng Bao, Wei Wang, Jinjia Zhao, and Xiaofeng Meng. Efficient query processing for XML keyword queries based on the IDList index. VLDB Journal: Very Large Data Bases, 23(1):25–50, February 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Slavov:2014:GBA

[630] Vasil Slavov and Praveen Rao. A gossip-based approach for Internetscale cardinality estimation of XPath queries over distributed semistructured data. VLDB Journal: Very Large Data Bases, 23(1):51–76, February 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Whang:2014:IER

[631] Steven Euijong Whang and Hector Garcia-Molina. Incremental entity resolution on rules and data. VLDB Journal: Very Large Data Bases, 23 (1):77–102, February 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Beskales:2014:SRC

[632] George Beskales, Ihab F. Ilyas, Lukasz Golab, and Artur Galiullin. Sampling from repairs of conditional functional dependency violations. VLDB Journal: Very Large Data Bases, 23 (1):103–128, February 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lee:2014:TEM

[633] Jongwuk Lee and Seung-Won Hwang. Toward efficient multidimensional subspace skyline computation. VLDB Journal: Very Large Data Bases, 23 (1):129–145, February 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zellag:2014:CAM

[634] Kamal Zellag and Bettina Kemme. Consistency anomalies in multi-tier architectures: automatic detection and prevention. VLDB Journal: Very Large Data Bases, 23(1):147–172, February 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ozsoyoglu:2014:SIB

[635] Z. Meral Özsoyoğlu, Uğur Çetintemel, Nilesh Dalvi, Hank Korth, and Anthony Tung. Special issue on best papers of VLDB 2012. VLDB Journal: Very Large Data Bases, 23(2):173–174, April 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-014-
0356-z;   http://link.springer.
com/content/pdf/10.1007/s00778-
014-0356-z.pdf.

Angel:2014:DSM

[636] Albert Angel, Nick Koudas, Nikos Sarkas, Divesh Srivastava, Michael Svendsen, and Srikanta Tirthapura. Dense subgraph maintenance under streaming edge weight updates for real-time story identification. VLDB Journal: Very Large Data Bases, 23 (2):175–199, April 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.com/article/10.
1007/s00778-013-0340-z.

Das:2014:EFE

[637] Mahashweta Das, Saravanan Thirumuruganathan, Sihem Amer-Yahia, Gautam Das, and Cong Yu. An expressive framework and efficient algorithms for the analysis of collaborative tagging. VLDB Journal: Very Large Data Bases, 23(2):201–226, April 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-013-
0341-y.

Cheng:2014:EPH

[638] James Cheng, Zechao Shang, Hong Cheng, Haixun Wang, and Jeffrey Xu Yu. Efficient processing of k-hop reachability queries. VLDB Journal: Very Large Data Bases, 23(2):227–252, April 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-013-
0346-6.

Koch:2014:DHO

[639] Christoph Koch, Yanif Ahmad, Oliver Kennedy, Milos Nikolic, Andres Nötzli, Daniel Lupei, and Amir Shaikhha. DBToaster: higher-order delta processing for dynamic, frequently fresh views. VLDB Journal: Very Large Data Bases, 23(2):253–278, April 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-013-
0348-4.

Bailis:2014:QEC

[640] Peter Bailis, Shivaram Venkataraman, Michael J. Franklin, Joseph M. Hellerstein, and Ion Stoica. Quantifying eventual consistency with PBS. VLDB Journal: Very Large Data Bases, 23 (2):279–302, April 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.com/article/10.
1007/s00778-013-0330-1.

Graefe:2014:TSA

[641] Goetz Graefe, Felix Halim, Stratos Idreos, Harumi Kuno, Stefan Manegold, and Bernhard Seeger. Transactional support for adaptive indexing. VLDB Journal: Very Large Data Bases, 23(2):303–328, April 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-013-
0345-7.

Zhang:2014:TCE

[642] Ning Zhang, Junichi Tatemura, Jignesh M. Patel, and Hakan Hacigumus. Toward cost-effective storage provisioning for DBMSs. VLDB Journal: Very Large Data Bases, 23(2):329–354, April 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-013-
0334-x.

Doulkeridis:2014:SLS

[643] Christos Doulkeridis and Kjetil Nørvåg. A survey of large-scale analytical query processing in MapReduce. VLDB Journal: Very Large Data Bases, 23(3):355– 380, June 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhou:2014:EDT

[644] Xiangmin Zhou and Lei Chen. Event detection over Twitter social media streams. VLDB Journal: Very Large Data Bases, 23(3):381–400, June 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Hung:2014:QTB

[645] Ho Hoang Hung, Sourav S. Bhowmick, Ba Quan Truong, Byron Choi, and Shuigeng Zhou. QUBLE: towards blending interactive visual subgraph search queries on large networks. VLDB Journal: Very Large Data Bases, 23(3):401–426, June 2014. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Teodoro:2014:ASS

[646] George Teodoro, Eduardo Valle, Nathan Mariano, Ricardo Torres, Wagner Meira, Jr., and Joel H. Saltz. Approximate similarity search for online multimedia services on distributed CPU–GPU platforms. VLDB Journal: Very Large Data Bases, 23(3):427–448, June 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Shang:2014:PTM

[647] Shuo Shang, Ruogu Ding, Kai Zheng, Christian S. Jensen, Panos Kalnis, and Xiaofang Zhou. Personalized trajectory matching in spatial networks. VLDB Journal: Very Large Data Bases, 23 (3):449–468, June 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Richter:2014:TZO

[648] Stefan Richter, Jorge-Arnulfo Quiané-Ruiz, Stefan Schuh, and Jens Dittrich. Towards zero-overhead static and adaptive indexing in Hadoop. VLDB Journal: Very Large Data Bases, 23 (3):469–494, June 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Meier:2014:BR

[649] Michael Meier. The backchase revisited. VLDB Journal: Very Large Data Bases, 23(3):495–516, June 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gedik:2014:PFS

[650] Bugra Gedik. Partitioning functions for stateful data parallelism in stream processing. VLDB Journal: Very Large Data Bases, 23(4):517–539, August 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Koh:2014:FKM

[651] Jia-Ling Koh, Chen-Yi Lin, and Arbee L. Chen. Finding kk most favorite products based on reverse toptt queries. VLDB Journal: Very Large Data Bases, 23(4):541–564, August 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zou:2014:GGB

[652] Lei Zou, M. Tamer Özsu, Lei Chen, Xuchuan Shen, Ruizhe Huang, and Dongyan Zhao. gStore: a graphbased SPARQL query engine. VLDB Journal: Very Large Data Bases, 23 (4):565–590, August 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Tao:2014:ILW

[653] Yufei Tao, Yi Yang, Xiaocheng Hu, Cheng Sheng, and Shuigeng Zhou. Instance-level worst-case query bounds on R-trees. VLDB Journal: Very Large Data Bases, 23(4):591–607, August 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Cicek:2014:ELD

[654] A. Ercument Cicek, Mehmet Ercan Nergiz, and Yucel Saygin. Ensuring location diversity in privacypreserving spatio-temporal data publishing. VLDB Journal: Very Large Data Bases, 23(4):609–625, August 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Unterbrunner:2014:HAE

[655] Philipp Unterbrunner, Gustavo Alonso, and Donald Kossmann. High availability, elasticity, and strong consistency for massively parallel scans over relational data. VLDB Journal: Very Large Data Bases, 23(4):627–652, August 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chen:2014:CND

[656] Rui Chen, Benjamin C. Fung, Philip S. Yu, and Bipin C. Desai. Correlated network data publication via differential privacy. VLDB Journal: Very Large Data Bases, 23(4):653–676, August 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Xiang:2014:AED

[657] Yang Xiang. Answering exact distance queries on real-world graphs with bounded performance guarantees. VLDB Journal: Very Large Data Bases, 23(5):677–695, October 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yao:2014:DMO

[658] Bin Yao, Xiaokui Xiao, Feifei Li, and Yifan Wu. Dynamic monitoring of optimal locations in road network databases. VLDB Journal: Very Large Data Bases, 23(5):697–720, October 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Tran:2014:QRE

[659] Quoc Trung Tran, Chee-Yong Chan, and Srinivasan Parthasarathy. Query reverse engineering. VLDB Journal: Very Large Data Bases, 23(5):721–746, October 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Martinenghi:2014:TBR

[660] Davide Martinenghi and Riccardo Torlone. Taxonomy-based relaxation of query answering in relational databases. VLDB Journal: Very Large Data Bases, 23(5):747–769, October 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Soria-Comas:2014:EDU

[661] Jordi Soria-Comas, Josep Domingo-Ferrer, David Sánchez, and Sergio Martínez. Enhancing data utility in differential privacy via microaggregationbased k-anonymity. VLDB Journal: Very Large Data Bases, 23(5):771–794, October 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Magnani:2014:TBP

[662] Matteo Magnani, Ira Assent, and Michael L. Mortensen. Taking the Big Picture: representative skylines based on significance and diversity. VLDB Journal: Very Large Data Bases, 23 (5):795–815, October 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Sidlauskas:2014:PEM

[663] Darius Šidlauskas, Simonas Šaltenis, and Christian S. Jensen. Processing of extreme moving-object update and query workloads in main memory. VLDB Journal: Very Large Data Bases, 23(5):817–841, October 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Aboulnaga:2014:SSD

[664] Ashraf Aboulnaga, Beng Chin Ooi, and Patrick Valduriez. Special section on data-intensive cloud infrastructure. VLDB Journal: Very Large Data Bases, 23(6):843, December 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Kumar:2014:SWA

[665] K. Ashwin Kumar, Abdul Quamar, Amol Deshpande, and Samir Khuller. SWORD: workload-aware data placement and replica selection for cloud data management systems. VLDB Journal: Very Large Data Bases, 23 (6):845–870, December 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Sahli:2014:ASP

[666] Majed Sahli, Essam Mansour, and Panos Kalnis. ACME: a scalable parallel system for extracting frequent patterns from a very long sequence. VLDB Journal: Very Large Data Bases, 23 (6):871–893, December 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lo:2014:MGD

[667] Eric Lo, Nick Cheng, Wilfred W. Lin, Wing-Kai Hon, and Byron Choi. My-Benchmark: generating databases for query workloads. VLDB Journal: Very Large Data Bases, 23(6):895–913, December 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Xie:2014:MEB

[668] Qing Xie, Chaoyi Pang, Xiaofang Zhou, Xiangliang Zhang, and Ke Deng. Maximum error-bounded Piecewise Linear Representation for online stream approximation. VLDB Journal: Very Large Data Bases, 23(6): 915–937, December 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Alexandrov:2014:SPB

[669] Alexander Alexandrov, Rico Bergmann, Stephan Ewen, Johann-Christoph Freytag, Fabian Hueske, Arvid Heise, Odej Kao, Marcus Leich, Ulf Leser, Volker Markl, Felix Naumann, Mathias Peters, Astrid Rheinländer, Matthias J. Sax, Sebastian Schelter, Mareike Höger, Kostas Tzoumas, and Daniel Warneke. The Stratosphere platform for big data analytics. VLDB Journal: Very Large Data Bases, 23(6): 939–964, December 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ward:2014:RTC

[670] Phillip G. Ward, Zhen He, Rui Zhang, and Jianzhong Qi. Real-time continuous intersection joins over large sets of moving objects using graphic processing units. VLDB Journal: Very Large Data Bases, 23(6):965–985, December 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Binnig:2014:DSI

[671] Carsten Binnig, Stefan Hildenbrand, Franz Färber, Donald Kossmann, Juchang Lee, and Norman May. Distributed snapshot isolation: global transactions pay globally, local transactions pay locally. VLDB Journal: Very Large Data Bases, 23(6):987–1011, December 2014. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Vlachos:2015:CMF

[672] Michail Vlachos, Nikolaos M. Freris, and Anastasios Kyrillidis. Compressive mining: fast and optimal data mining in the compressed domain. VLDB Journal: Very Large Data Bases, 24 (1):1–24, February 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Sistla:2015:CNN

[673] A. Prasad Sistla, Ouri Wolfson, and Bo Xu. Continuous nearest-neighbor queries with location uncertainty. VLDB Journal: Very Large Data Bases, 24(1):25–50, February 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gur:2015:SFA

[674] Izzeddin Gür, Mehmet Güvercin, and Hakan Ferhatosmanoglu. Scaling forecasting algorithms using clustered modeling. VLDB Journal: Very Large Data Bases, 24(1):51–65, February 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Kaoudi:2015:RCS

[675] Zoi Kaoudi and Ioana Manolescu. RDF in the clouds: a survey. VLDB Journal: Very Large Data Bases, 24 (1):67–91, February 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Su:2015:CTD

[676] Han Su, Kai Zheng, Jiamin Huang, Haozhou Wang, and Xiaofang Zhou. Calibrating trajectory data for spatiotemporal similarity analysis. VLDB Journal: Very Large Data Bases, 24 (1):93–116, February 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Li:2015:CAI

[677] Hui Li, Sourav S. Bhowmick, Aixin Sun, and Jiangtao Cui. Conformityaware influence maximization in online social networks. VLDB Journal: Very Large Data Bases, 24(1): 117–141, February 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Deng:2015:UFA

[678] Dong Deng, Guoliang Li, Jianhua Feng, Yi Duan, and Zhiguo Gong. A unified framework for approximate dictionarybased entity extraction. VLDB Journal: Very Large Data Bases, 24(1): 143–167, February 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Hung:2015:CAC

[679] Chih-Chieh Hung, Wen-Chih Peng, and Wang-Chien Lee. Clustering and aggregating clues of trajectories for mining trajectory patterns and routes. VLDB Journal: Very Large Data Bases, 24(2):169–192, April 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Eichinger:2015:TSC

[680] Frank Eichinger, Pavel Efros, Stamatis Karnouskos, and Klemens Böhm. A time-series compression technique and its application to the smart grid. VLDB Journal: Very Large Data Bases, 24 (2):193–218, April 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Xue:2015:SDS

[681] Andy Yuan Xue, Jianzhong Qi, Xing Xie, Rui Zhang, Jin Huang, and Yuan Li. Solving the data sparsity problem in destination prediction. VLDB Journal: Very Large Data Bases, 24 (2):219–243, April 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhang:2015:ECS

[682] Zhiwei Zhang, Jeffrey Xu Yu, Lu Qin, Lijun Chang, and Xuemin Lin. I/O efficient: computing SCCs in massive graphs. VLDB Journal: Very Large Data Bases, 24(2):245–270, April 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yuan:2015:GSS

[683] Ye Yuan, Guoren Wang, Lei Chen, and Haixun Wang. Graph similarity search on large uncertain graph databases. VLDB Journal: Very Large Data Bases, 24(2):271–296, April 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yang:2015:TPC

[684] Bin Yang, Chenjuan Guo, Yu Ma, and Christian S. Jensen. Toward personalized, context-aware routing. VLDB Journal: Very Large Data Bases, 24 (2):297–318, April 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Huang:2015:TKS

[685] Xin Huang, Hong Cheng, Rong-Hua Li, Lu Qin, and Jeffrey Xu Yu. Top-K structural diversity search in large networks. VLDB Journal: Very Large Data Bases, 24(3):319–343, June 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Papapetrou:2015:SDS

[686] Odysseas Papapetrou, Minos Garofalakis, and Antonios Deligiannakis. Sketching distributed sliding-window data streams. VLDB Journal: Very Large Data Bases, 24(3):345–368, June 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yuan:2015:EDS

[687] Ye Yuan, Guoren Wang, Jeffery Yu Xu, and Lei Chen. Efficient distributed subgraph similarity matching. VLDB Journal: Very Large Data Bases, 24 (3):369–394, June 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Mirylenka:2015:CHH

[688] Katsiaryna Mirylenka, Graham Cormode, Themis Palpanas, and Divesh Srivastava. Conditional heavy hitters: detecting interesting correlations in data streams. VLDB Journal: Very Large Data Bases, 24(3):395–414, June 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gao:2015:ECP

[689] Yunjun Gao, Lu Chen, Xinhan Li, Bin Yao, and Gang Chen. Efficient kkclosest pair queries in general metric spaces. VLDB Journal: Very Large Data Bases, 24(3):415–439, June 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Aksoy:2015:RPE

[690] Cem Aksoy, Aggeliki Dimitriou, and Dimitri Theodoratos. Reasoning with patterns to effectively answer XML keyword queries. VLDB Journal: Very Large Data Bases, 24(3):441–465, June 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Roy:2015:TAO

[691] Senjuti Basu Roy, Ioanna Lykourentzou, Saravanan Thirumuruganathan, Sihem Amer-Yahia, and Gautam Das. Task assignment optimization in knowledge-intensive crowdsourcing. VLDB Journal: Very Large Data Bases, 24(4):467–491, August 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Bao:2015:GFR

[692] Zhifeng Bao, Yong Zeng, Tok Wang Ling, Dongxiang Zhang, Guoliang Li, and H. V. Jagadish. A general framework to resolve the MisMatch problem in XML keyword search. VLDB Journal: Very Large Data Bases, 24 (4):493–518, August 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Kotsifakos:2015:EBS

[693] Alexios Kotsifakos, Isak Karlsson, Panagiotis Papapetrou, Vassilis Athitsos, and Dimitrios Gunopulos. Embeddingbased subsequence matching with gaps — range — tolerances: a Query-By-Humming application. VLDB Journal: Very Large Data Bases, 24(4):519–536, August 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Skovsgaard:2015:FTR

[694] Anders Skovsgaard and Christian S. Jensen. Finding top-k relevant groups of spatial web objects. VLDB Journal: Very Large Data Bases, 24(4):537–555, August 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Abedjan:2015:PRD

[695] Ziawasch Abedjan, Lukasz Golab, and Felix Naumann. Profiling relational data: a survey. VLDB Journal: Very Large Data Bases, 24(4):557–581, August 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Deutch:2015:PBA

[696] Daniel Deutch, Yuval Moskovitch, and Val Tannen. Provenance-based analysis of data-centric processes. VLDB Journal: Very Large Data Bases, 24 (4):583–607, August 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Bohlen:2015:SIB

[697] Michael H. Böhlen and Christoph Koch. Special issue on best papers of VLDB 2013. VLDB Journal: Very Large Data Bases, 24(5):609–610, October 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yan:2015:ALK

[698] Zhepeng Yan, Nan Zheng, Zachary G. Ives, Partha Pratim Talukdar, and Cong Yu. Active learning in keyword search-based data integration. VLDB Journal: Very Large Data Bases, 24 (5):611–631, October 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zou:2015:CDA

[699] Tao Zou, Ronan Bras, Marcos Vaz Salles, Alan Demers, and Johannes Gehrke. ClouDiA: a deployment advisor for public clouds. VLDB Journal: Very Large Data Bases, 24(5):633– 653, October 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhu:2015:SAP

[700] Fanwei Zhu, Yuan Fang, Kevin Chen-Chuan Chang, and Jing Ying. Scheduled approximation for Personalized PageRank with Utility-based Hub Selection. VLDB Journal: Very Large Data Bases, 24(5):655–679, October 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ren:2015:VLM

[701] Kun Ren, Alexander Thomson, and Daniel J. Abadi. VLL: a lock manager redesign for main memory database systems. VLDB Journal: Very Large Data Bases, 24(5):681–705, October 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Galarraga:2015:FRM

[702] Luis Galárraga, Christina Teflioudi, Katja Hose, and Fabian M. Suchanek. Fast rule mining in ontological knowledge bases with AMIE++. VLDB Journal: Very Large Data Bases, 24 (6):707–730, December 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chandra:2015:DGT

[703] Bikash Chandra, Bhupesh Chawda, Biplab Kar, K. V. Reddy, Shetal Shah, and S. Sudarshan. Data generation for testing and grading SQL queries. VLDB Journal: Very Large Data Bases, 24(6):731–755, December 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Li:2015:MMO

[704] Chao Li, Gerome Miklau, Michael Hay, Andrew Mcgregor, and Vibhor Rastogi. The matrix mechanism: optimizing linear counting queries under differential privacy. VLDB Journal: Very Large Data Bases, 24(6):757–781, December 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Armenatzoglou:2015:GSR

[705] Nikos Armenatzoglou, Ritesh Ahuja, and Dimitris Papadias. Geo-Social Ranking: functions and query processing. VLDB Journal: Very Large Data Bases, 24(6):783–799, December 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Santini:2015:QSU

[706] Simone Santini. Querying streams using regular expressions: some semantics, decidability, and efficiency issues. VLDB Journal: Very Large Data Bases, 24(6):801–821, December 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wang:2015:ATE

[707] Xiang Wang, Ying Zhang, Wenjie Zhang, Xuemin Lin, and Wei Wang. AP-Tree: efficiently support locationaware Publish/Subscribe. VLDB Journal: Very Large Data Bases, 24(6): 823–848, December 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Basik:2015:STS

[708] Fuat Basık, Buğra Gedik, Hakan Ferhatosmanoğlu, and Mert Emin Kalender. S^{33}-TM: scalable streaming short text matching. VLDB Journal: Very Large Data Bases, 24(6):849–866, December 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Jagadish:2016:SIB

[709] H. V. Jagadish and Aoying Zhou. Special issue on best papers of VLDB 2014. VLDB Journal: Very Large Data Bases, 25(1):1–2, February 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Jiang:2016:EES

[710] Dawei Jiang, Sai Wu, Gang Chen, Beng Chin Ooi, Kian-Lee Tan, and Jun Xu. epiC: an extensible and scalable system for processing Big Data. VLDB Journal: Very Large Data Bases, 25 (1):3–26, February 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Schuhknecht:2016:EEA

[711] Felix Martin Schuhknecht, Alekh Jindal, and Jens Dittrich. An experimental evaluation and analysis of database cracking. VLDB Journal: Very Large Data Bases, 25(1):27–52, February 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Jugel:2016:VAV

[712] Uwe Jugel, Zbigniew Jerzak, Gregor Hackenbroich, and Volker Markl.

VDDA: automatic visualization-driven data aggregation in relational databases. VLDB Journal: Very Large Data Bases, 25(1):53–77, February 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wang:2016:EDL

[713] Wei Wang, Xiaoyan Yang, Beng Chin Ooi, Dongxiang Zhang, and Yueting Zhuang. Effective deep learningbased multi-modal retrieval. VLDB Journal: Very Large Data Bases, 25 (1):79–101, February 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Funke:2016:KPC

[714] Stefan Funke, André Nusser, and Sabine Storandt. On k-Path Covers and their applications. VLDB Journal: Very Large Data Bases, 25 (1):103–123, February 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Quamar:2016:NNC

[715] Abdul Quamar, Amol Deshpande, and Jimmy Lin. NScale: neighborhoodcentric large-scale graph analytics in the cloud. VLDB Journal: Very Large Data Bases, 25(2):125–150, April 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Soule:2016:RAS

[716] Robert Soulé and Bugra Gedik. RailwayDB: adaptive storage of interaction graphs. VLDB Journal: Very Large Data Bases, 25(2):151–169, April 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yuan:2016:DTK

[717] Long Yuan, Lu Qin, Xuemin Lin, Lijun Chang, and Wenjie Zhang. Diversified top-k clique search. VLDB Journal: Very Large Data Bases, 25 (2):171–196, April 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Pham:2016:ACW

[718] Thao N. Pham, Panos K. Chrysanthis, and Alexandros Labrinidis. Avoiding class warfare: managing continuous queries with differentiated classes of service. VLDB Journal: Very Large Data Bases, 25(2):197–221, April 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Langer:2016:EOD

[719] Philipp Langer and Felix Naumann. Efficient order dependency detection. VLDB Journal: Very Large Data Bases, 25(2):223–241, April 2016. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Peng:2016:PSQ

[720] Peng Peng, Lei Zou, M. Tamer Özsu, Lei Chen, and Dongyan Zhao. Processing SPARQL queries over distributed RDF graphs. VLDB Journal: Very Large Data Bases, 25(2):243–268, April 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gao:2016:TCP

[721] Jun Gao, Chang Zhou, and Jeffrey Xu Yu. Toward continuous pattern detection over evolving large graph with snapshot isolation. VLDB Journal: Very Large Data Bases, 25(2):269–290, April 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Baumann:2016:BDC

[722] Stephan Baumann, Peter Boncz, and Kai-Uwe Sattler. Bitwise dimensional co-clustering for analytical workloads. VLDB Journal: Very Large Data Bases, 25(3):291–316, June 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Li:2016:EAF

[723] Feifei Li, Ke Yi, Yufei Tao, Bin Yao, Yang Li, Dong Xie, and Min Wang. Exact and approximate flexible aggregate similarity search. VLDB Journal: Very Large Data Bases, 25(3):317–338, June 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Guzun:2016:HQO

[724] Gheorghi Guzun and Guadalupe Canahuate. Hybrid query optimization for hard-to-compress bit-vectors. VLDB Journal: Very Large Data Bases, 25(3):339–354, June 2016. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Harbi:2016:ASQ

[725] Razen Harbi, Ibrahim Abdelaziz, Panos Kalnis, Nikos Mamoulis, Yasser Ebrahim, and Majed Sahli. Accelerating SPARQL queries by exploiting hash-based locality and adaptive partitioning. VLDB Journal: Very Large Data Bases, 25(3):355–380, June 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Bonifati:2016:MEO

[726] Angela Bonifati, Werner Nutt, Riccardo Torlone, and Jan Van Den Bussche. Mapping-equivalence and oidequivalence of single-function objectcreating conjunctive queries. VLDB Journal: Very Large Data Bases, 25 (3):381–397, June 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lu:2016:DCE

[727] Yue Lu, Yuguan Li, and Mohamed Y. Eltabakh. Decorating the cloud: enabling annotation management in MapReduce. VLDB Journal: Very Large Data Bases, 25(3):399–424, June 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Sariyuce:2016:IKC

[728] Ahmet Erdem Sariyüce, Bugra Gedik, Gabriela Jacques-Silva, Kun-Lung Wu, and Ümit V. Çatalyürek. Incremental k-core decomposition: algorithms and evaluation. VLDB Journal: Very Large Data Bases, 25(3):425–447, June 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Luo:2016:QDS

[729] Ge Luo, Lu Wang, Ke Yi, and Graham Cormode. Quantiles over data streams: experimental comparisons, new analyses, and further improvements. VLDB Journal: Very Large Data Bases, 25 (4):449–472, August 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Xie:2016:EEI

[730] Xike Xie, Benjin Mei, Jinchuan Chen, Xiaoyong Du, and Christian S. Jensen.

Elite: an elastic infrastructure for big spatiotemporal trajectories. VLDB Journal: Very Large Data Bases, 25 (4):473–493, August 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Kanza:2016:ESF

[731] Yaron Kanza and Hadas Yaari. External sorting on flash storage: reducing cell wearing and increasing efficiency by avoiding intermediate writes. VLDB Journal: Very Large Data Bases, 25 (4):495–518, August 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Jeon:2016:MBS

[732] Inah Jeon, Evangelos E. Papalexakis, Christos Faloutsos, Lee Sael, and U. Kang. Mining billion-scale tensors: algorithms and discoveries. VLDB Journal: Very Large Data Bases, 25 (4):519–544, August 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Islam:2016:KYC

[733] Md. Saiful Islam and Chengfei Liu. Know your customer: computing kmost promising products for targeted marketing. VLDB Journal: Very Large Data Bases, 25(4):545–570, August 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Kohler:2016:PCK

[734] Henning Köhler, Uwe Leck, Sebastian Link, and Xiaofang Zhou. Possible and certain keys for SQL. VLDB Journal: Very Large Data Bases, 25 (4):571–596, August 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Mottin:2016:HPA

[735] Davide Mottin, Alice Marascu, Senjuti Basu Roy, Gautam Das, Themis Palpanas, and Yannis Velegrakis. A holistic and principled approach for the empty-answer problem. VLDB Journal: Very Large Data Bases, 25 (4):597–622, August 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Boncz:2016:SIM

[736] Peter Boncz, Wolfgang Lehner, and Thomas Neumann. Special issue: Modern hardware. VLDB Journal: Very Large Data Bases, 25(5):623–624, October 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Porobic:2016:CIH

[737] Danica Porobic, Ippokratis Pandis, Miguel Branco, Pinar Tözün, and Anastasia Ailamaki. Characterization of the impact of hardware islands on OLTP. VLDB Journal: Very Large Data Bases, 25(5):625–650, October 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Sadoghi:2016:ESO

[738] Mohammad Sadoghi, Kenneth A. Ross, Mustafa Canim, and Bishwaranjan Bhattacharjee. Exploiting SSDs in operational multiversion databases. VLDB Journal: Very Large Data Bases, 25(5):651–672, October 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Kang:2016:FCE

[739] Woon-Hak Kang, Sang-Won Lee, and Bongki Moon. Flash as cache extension for online transactional workloads. VLDB Journal: Very Large Data Bases, 25(5):673–694, October 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Jin:2016:RWO

[740] Peiquan Jin, Chengcheng Yang, Christian S. Jensen, Puyuan Yang, and Lihua Yue. Read/write-optimized tree indexing for solid-state drives. VLDB Journal: Very Large Data Bases, 25 (5):695–717, October 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Sitaridi:2016:GAS

[741] Evangelia A. Sitaridi and Kenneth A. Ross. GPU-accelerated string matching for database applications. VLDB Journal: Very Large Data Bases, 25 (5):719–740, October 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Mottin:2016:EQN

[742] Davide Mottin, Matteo Lissandrini, Yannis Velegrakis, and Themis Palpanas. Exemplar queries: a new way of searching. VLDB Journal: Very Large Data Bases, 25(6):741–765, December 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Li:2016:EDL

[743] Yuhong Li, Leong Hou U., Man Lung Yiu, and Zhiguo Gong. Efficient discovery of longest-lasting correlation in sequence databases. VLDB Journal: Very Large Data Bases, 25(6): 767–790, December 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Fakas:2016:DPS

[744] Georgios J. Fakas, Zhi Cai, and Nikos Mamoulis. Diverse and proportional size-l object summaries using pairwise relevance. VLDB Journal: Very Large Data Bases, 25(6):791–816, December 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

BOgh:2016:SPW

[745] Kenneth S. BØgh, Sean Chester, and Ira Assent. SkyAlign: a portable, workefficient skyline algorithm for multicore and GPU architectures. VLDB Journal: Very Large Data Bases, 25 (6):817–841, December 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zoumpatianos:2016:AAD

[746] Kostas Zoumpatianos, Stratos Idreos, and Themis Palpanas. ADS: the adaptive data series index. VLDB Journal: Very Large Data Bases, 25(6): 843–866, December 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Liu:2016:AWW

[747] Qing Liu, Yunjun Gao, Gang Chen, Baihua Zheng, and Linlin Zhou. Answering why-not and why questions on reverse top-k queries. VLDB Journal: Very Large Data Bases, 25(6): 867–892, December 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chen:2016:SSL

[748] Yang Chen, Daisy Zhe Wang, and Sean Goldberg. ScaLeKB: scalable learning and inference over large knowledge bases. VLDB Journal: Very Large Data Bases, 25(6):893–918, December 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Li:2017:SIB

[749] Chen Li and Volker Markl. Special issue on best papers of VLDB 2015. VLDB Journal: Very Large Data Bases, 26(1):1–2, February 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). See erratum [750].

Li:2017:ESI

[750] Chen Li and Volker Markl. Erratum to: Special issue on best papers of VLDB 2015. VLDB Journal: Very Large Data Bases, 26(1):3, February 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). See [749].

Gatterbauer:2017:DPA

[751] Wolfgang Gatterbauer and Dan Suciu. Dissociation and propagation for approximate lifted inference with standard relational database management systems. VLDB Journal: Very Large Data Bases, 26(1):5–30, February 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Li:2017:RBR

[752] Jiexing Li, Jeffrey F. Naughton, and Rimma V. Nehme. Resource bricolage and resource selection for parallel database systems. VLDB Journal:

Very Large Data Bases, 26(1):31–54, February 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Finis:2017:OIS

[753] Jan Finis, Robert Brunel, Alfons Kemper, Thomas Neumann, Norman May, and Franz Faerber. Order Indexes: supporting highly dynamic hierarchical data in relational main-memory database systems. VLDB Journal: Very Large Data Bases, 26(1):55–80, February 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Sa:2017:IKB

[754] Christopher Sa, Alex Ratner, Christopher Ré, Jaeho Shin, Feiran Wang, Sen Wu, and Ce Zhang. Incremental knowledge base construction using DeepDive. VLDB Journal: Very Large Data Bases, 26(1):81–105, February 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Trummer:2017:MOP

[755] Immanuel Trummer and Christoph Koch. Multi-objective parametric query optimization. VLDB Journal: Very Large Data Bases, 26(1):107–124, February 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Khayyat:2017:FSI

[756] Zuhair Khayyat, William Lucia, Meghna Singh, Mourad Ouzzani, Paolo Papotti, Jorge-Arnulfo Quiané-Ruiz, Nan Tang, and Panos Kalnis. Fast and scalable inequality joins. VLDB Journal: Very Large Data Bases, 26 (1):125–150, February 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yang:2017:RKN

[757] Shiyu Yang, Muhammad Aamir Cheema, Xuemin Lin, Ying Zhang, and Wenjie Zhang. Reverse k nearest neighbors queries and spatial reverse top-k queries. VLDB Journal: Very Large Data Bases, 26(2):151–176, April 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Li:2017:DBQ

[758] Kun Li, Xiaofeng Zhou, Daisy Zhe Wang, Christan Grant, Alin Dobra, and Christopher Dudley. In-database batch and query-time inference over probabilistic graphical models using UDA — GIST. VLDB Journal: Very Large Data Bases, 26(2):177–201, April 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Xie:2017:PTP

[759] Miao Xie, Sourav S. Bhowmick, Gao Cong, and Qing Wang. PANDA: toward partial topology-based search on large networks in a single machine. VLDB Journal: Very Large Data Bases, 26(2):203–228, April 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yang:2017:SPM

[760] Mohan Yang, Alexander Shkapsky, and Carlo Zaniolo. Scaling up the performance of more powerful Datalog systems on multicore machines. VLDB Journal: Very Large Data Bases, 26 (2):229–248, April 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yu:2017:UFS

[761] Minghe Yu, Jin Wang, Guoliang Li, Yong Zhang, Dong Deng, and Jianhua Feng. A unified framework for string similarity search with edit-distance constraint. VLDB Journal: Very Large Data Bases, 26(2):249–274, April 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yuan:2017:EEG

[762] Long Yuan, Lu Qin, Xuemin Lin, Lijun Chang, and Wenjie Zhang. I/O efficient ECC graph decomposition via graph reduction. VLDB Journal: Very Large Data Bases, 26(2):275–300, April 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wang:2017:TKS

[763] Xiang Wang, Wenjie Zhang, Ying Zhang, Xuemin Lin, and Zengfeng Huang. Top-k spatial-keyword publish/ subscribe over sliding window. VLDB Journal: Very Large Data Bases, 26 (3):301–326, June 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gao:2017:PBH

[764] Jun Gao, Yuqiong Liu, Chang Zhou, and Jeffrey Xu Yu. Path-based holistic detection plan for multiple patterns in distributed graph frameworks. VLDB Journal: Very Large Data Bases, 26 (3):327–345, June 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yi:2017:AVQ

[765] Peipei Yi, Byron Choi, Sourav S. Bhowmick, and Jianliang Xu. AutoG: a visual query autocompletion framework for graph databases. VLDB Journal: Very Large Data Bases, 26 (3):347–372, June 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Aljubayrin:2017:FLC

[766] Saad Aljubayrin, Jianzhong Qi, Christian S. Jensen, Rui Zhang, Zhen He, and Yuan Li. Finding lowest-cost paths in settings with safe and preferred zones. VLDB Journal: Very Large Data Bases, 26(3):373–397, June 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhang:2017:DSP

[767] Dongxiang Zhang, Dingyu Yang, Yuan Wang, Kian-Lee Tan, Jian Cao, and Heng Tao Shen. Distributed shortest path query processing on dynamic road networks. VLDB Journal: Very Large Data Bases, 26(3):399–419, June 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lai:2017:SSE

[768] Longbin Lai, Lu Qin, Xuemin Lin, and Lijun Chang. Scalable subgraph enumeration in MapReduce: a costoriented approach. VLDB Journal: Very Large Data Bases, 26(3):421–446, June 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Cafagna:2017:DIP

[769] Francesco Cafagna and Michael H. Böhlen. Disjoint interval partitioning. VLDB Journal: Very Large Data Bases, 26(3):447–466, June 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gao:2017:EFR

[770] Yunjun Gao, Xiaoye Miao, Gang Chen, Baihua Zheng, Deng Cai, and Huiyong Cui. On efficiently finding reverse k-nearest neighbors over uncertain graphs. VLDB Journal: Very Large Data Bases, 26(4):467–492, August 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Tao:2017:SSW

[771] Yufei Tao, Xiaocheng Hu, and Miao Qiao. Stream sampling over windows with worst-case optimality and ℓoverlap independence. VLDB Journal: Very Large Data Bases, 26(4):493–510, August 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Nguyen:2017:ADC

[772] Quoc Viet Nguyen, Chi Thang Duong, Thanh Tam Nguyen, Matthias Weidlich, Karl Aberer, Hongzhi Yin, and Xiaofang Zhou. Argument discovery via crowdsourcing. VLDB Journal: Very Large Data Bases, 26(4):511–535, August 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wang:2017:EMA

[773] Tianzheng Wang, Ryan Johnson, Alan Fekete, and Ippokratis Pandis. Efficiently making (almost) any concurrency control mechanism serializable. VLDB Journal: Very Large Data Bases, 26(4):537–562, August 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). See erratum [825].

Zhu:2017:EAT

[774] Qiankun Zhu, Hong Cheng, and Xin Huang. I/O-efficient algorithms for top-k nearest keyword search in massive graphs. VLDB Journal: Very Large Data Bases, 26(4):563–583, August 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chen:2017:IMU

[775] Lu Chen, Yunjun Gao, Aoxiao Zhong, Christian S. Jensen, Gang Chen, and Baihua Zheng. Indexing metric uncertain data for range queries and range joins. VLDB Journal: Very Large Data Bases, 26(4):585–610, August 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Song:2017:GRU

[776] Shaoxu Song, Boge Liu, Hong Cheng, Jeffrey Xu Yu, and Lei Chen. Graph repairing under neighborhood constraints. VLDB Journal: Very Large Data Bases, 26(5):611–635, October 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhou:2017:EOV

[777] Xiangmin Zhou, Lei Chen, Yanchun Zhang, Dong Qin, Longbing Cao, Guangyan Huang, and Chen Wang. Enhancing online video recommendation using social user interactions. VLDB Journal: Very Large Data Bases, 26(5):637–656, October 2017.

CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Attasena:2017:SSC

[778] Varunya Attasena, Jérôme Darmont, and Nouria Harbi. Secret sharing for cloud data security: a survey. VLDB Journal: Very Large Data Bases, 26 (5):657–681, October 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Huang:2017:QAL

[779] Qiang Huang, Jianlin Feng, Qiong Fang, Wilfred Ng, and Wei Wang. Query-aware locality-sensitive hashing scheme for l_{p} norm. VLDB Journal: Very Large Data Bases, 26(5):683– 708, October 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhu:2017:GSG

[780] Qijun Zhu, Haibo Hu, Cheng Xu, Jianliang Xu, and Wang-Chien Lee. Geo-social group queries with minimum acquaintance constraints. VLDB Journal: Very Large Data Bases, 26 (5):709–727, October 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhang:2017:DMK

[781] Kai Zhang, Kaibo Wang, Yuan Yuan, Lei Guo, Rubao Li, Xiaodong Zhang, Bingsheng He, Jiayu Hu, and Bei Hua. A distributed in-memory keyvalue store system on heterogeneous CPU–GPU cluster. VLDB Journal: Very Large Data Bases, 26(5):729–750, October 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Li:2017:FIC

[782] Rong-Hua Li, Lu Qin, Jeffrey Xu Yu, and Rui Mao. Finding influential communities in massive networks. VLDB Journal: Very Large Data Bases, 26 (6):751–776, December 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ali:2017:CDP

[783] Syed Muhammad Ali and Robert Wrembel. From conceptual design to performance optimization of ETL workflows: current state of research and open problems. VLDB Journal: Very Large Data Bases, 26(6): 777–801, December 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Fang:2017:EEA

[784] Yixiang Fang, Reynold Cheng, Yankai Chen, Siqiang Luo, and Jiafeng Hu. Effective and efficient attributed community search. VLDB Journal: Very Large Data Bases, 26(6):803–828, December 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lu:2017:MES

[785] Wei Lu, Jiajia Hou, Ying Yan, Meihui Zhang, Xiaoyong Du, and Thomas Moscibroda. MSQL: efficient similarity search in metric spaces using SQL. VLDB Journal: Very Large Data Bases, 26(6):829–854, December 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Hung:2017:AVG

[786] Nguyen Quoc Hung, Duong Chi Thang, Nguyen Thanh Tam, Matthias Weidlich, Karl Aberer, Hongzhi Yin, and Xiaofang Zhou. Answer validation for generic crowdsourcing tasks with minimal efforts. VLDB Journal: Very Large Data Bases, 26(6):855–880, December 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Herschel:2017:SPW

[787] Melanie Herschel, Ralf Diestelkämper, and Houssem Ben Lahmar. A survey on provenance: What for? What form? What from? VLDB Journal: Very Large Data Bases, 26(6): 881–906, December 2017. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wei:2018:RQI

[788] Hao Wei, Jeffrey Xu Yu, Can Lu, and Ruoming Jin. Reachability querying: an independent permutation labeling approach. VLDB Journal: Very Large Data Bases, 27(1):1–26, February 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lin:2018:OAS

[789] Chunbin Lin, Jiaheng Lu, Zhewei Wei, Jianguo Wang, and Xiaokui Xiao. Optimal algorithms for selecting top-k combinations of attributes: theory and applications. VLDB Journal: Very Large Data Bases, 27(1):27–52, February 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhao:2018:ESS

[790] Xiang Zhao, Chuan Xiao, Xuemin Lin, Wenjie Zhang, and Yang Wang. Efficient structure similarity searches: a partition-based approach. VLDB Journal: Very Large Data Bases, 27 (1):53–78, February 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yu:2018:DSS

[791] Weiren Yu, Xuemin Lin, Wenjie Zhang, and Julie A. Mccann. Dynamical SimRank search on time-varying networks. VLDB Journal: Very Large Data Bases, 27(1):79–104, February 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Sagi:2018:NBE

[792] Tomer Sagi and Avigdor Gal. Nonbinary evaluation measures for big data integration. VLDB Journal: Very Large Data Bases, 27(1):105–126, February 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wu:2018:SOR

[793] Yubao Wu, Xiang Zhang, Yuchen Bian, Zhipeng Cai, Xiang Lian, Xueting Liao, and Fengpan Zhao. Second-order random walk-based proximity measures in graph analysis: formulations and algorithms. VLDB Journal: Very Large Data Bases, 27(1):127–152, February 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yang:2018:PPC

[794] Bin Yang, Jian Dai, Chenjuan Guo, Christian S. Jensen, and Jilin Hu. PACE: a PAth-CEntric paradigm for stochastic path finding. VLDB Journal: Very Large Data Bases, 27(2):153– 178, April 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Hu:2018:RAP

[795] Jilin Hu, Bin Yang, Chenjuan Guo, and Christian S. Jensen. Risk-aware path selection with time-varying, uncertain travel costs: a time series approach. VLDB Journal: Very Large Data Bases, 27(2):179–200, April 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Su:2018:PDP

[796] Dong Su, Jianneng Cao, Ninghui Li, and Min Lyu. PrivPfC: differentially private data publication for classification. VLDB Journal: Very Large Data Bases, 27(2):201–223, April 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhang:2018:AKS

[797] Dongxiang Zhang, Yuchen Li, Xin Cao, Jie Shao, and Heng Tao Shen. Augmented keyword search on spatial entity databases. VLDB Journal: Very Large Data Bases, 27(2):225–244, April 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Deutch:2018:EPT

[798] Daniel Deutch, Amir Gilad, and Yuval Moskovitch. Efficient provenance tracking for datalog using topk queries. VLDB Journal: Very Large Data Bases, 27(2):245–269, April 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhou:2018:ARQ

[799] Junfeng Zhou, Jeffrey Xu Yu, Na Li, Hao Wei, Ziyang Chen, and Xian Tang. Accelerating reachability query processing based on D⃗AG reduction.

VLDB Journal: Very Large Data Bases, 27(2):271–296, April 2018. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Belesiotis:2018:STU

[800] Alexandros Belesiotis, Dimitrios Skoutas, Christodoulos Efstathiades, Vassilis Kaffes, and Dieter Pfoser. Spatiotextual user matching and clustering based on set similarity joins. VLDB Journal: Very Large Data Bases, 27 (3):297–320, June 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Li:2018:GSG

[801] Lei Li, Kai Zheng, Sibo Wang, Wen Hua, and Xiaofang Zhou. Go slow to go fast: minimal on-road time route scheduling with parking facilities using historical trajectory. VLDB Journal: Very Large Data Bases, 27(3):321–345, June 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yao:2018:SDT

[802] Chang Yao, Meihui Zhang, Qian Lin, Beng Chin Ooi, and Jiatao Xu. Scaling distributed transaction processing and recovery based on dependency logging. VLDB Journal: Very Large Data Bases, 27(3):347–368, June 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chodpathumwan:2018:CEC

[803] Yodsawalai Chodpathumwan, Ali Vakilian, Arash Termehchy, and Amir Nayyeri. Cost-effective conceptual design using taxonomies. VLDB Journal: Very Large Data Bases, 27(3):369–394, June 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Shang:2018:PTS

[804] Shuo Shang, Lisi Chen, Zhewei Wei, Christian S. Jensen, Kai Zheng, and Panos Kalnis. Parallel trajectory similarity joins in spatial networks. VLDB Journal: Very Large Data Bases, 27 (3):395–420, June 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lee:2018:PRA

[805] Juchang Lee, Wook-Shin Han, Hyoung Jun Na, Chang Gyoo Park, Kyu Hwan Kim, Deok Hoe Kim, Joo Yeon Lee, Sang Kyun Cha, and Seunghyun Moon. Parallel replication across formats for scaling out mixed OLTP/OLAP workloads in main-memory databases. VLDB Journal: Very Large Data Bases, 27(3):421– 444, June 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Choudhury:2018:FOL

[806] Farhana Murtaza Choudhury, J. Shane Culpepper, Zhifeng Bao, and Timos Sellis. Finding the optimal location and keywords in obstructed and unobstructed space. VLDB Journal: Very Large Data Bases, 27(4):445–470, August 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yang:2018:ESC

[807] Jianye Yang, Wenjie Zhang, Shiyu Yang, Ying Zhang, Xuemin Lin, and Long Yuan. Efficient set containment join. VLDB Journal: Very Large Data Bases, 27(4):471–495, August 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Hao:2018:DRU

[808] Shuang Hao, Nan Tang, Guoliang Li, Jian Li, and Jianhua Feng. Distilling relations using knowledge bases. VLDB Journal: Very Large Data Bases, 27 (4):497–519, August 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Borovica-Gajic:2018:SSR

[809] Renata Borovica-Gajic, Stratos Idreos, Anastasia Ailamaki, Marcin Zukowski, and Campbell Fraser. Smooth Scan: robust access path selection without cardinality estimation. VLDB Journal: Very Large Data Bases, 27(4):521–545, August 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Herrmann:2018:MSV

[810] Kai Herrmann, Hannes Voigt, Torben Bach Pedersen, and Wolfgang Lehner. Multi-schema-version data management: data independence in the twenty-first century. VLDB Journal: Very Large Data Bases, 27(4):547–571, August 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Szlichta:2018:ECD

[811] Jaroslaw Szlichta, Parke Godfrey, Lukasz Golab, Mehdi Kargar, and Divesh Srivastava. Effective and complete discovery of bidirectional order dependencies via set-based axioms. VLDB Journal: Very Large Data Bases, 27 (4):573–591, August 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chaudhuri:2018:SIB

[812] Surajit Chaudhuri and Jayant R. Haritsa. Special issue on best papers of VLDB 2016. VLDB Journal: Very Large Data Bases, 27(5):593–594, October 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Interlandi:2018:ADP

[813] Matteo Interlandi, Ari Ekmekji, Kshitij Shah, Muhammad Ali Gulzar, Sai Deep Tetali, Miryung Kim, Todd Millstein, and Tyson Condie. Adding data provenance support to Apache Spark. VLDB Journal: Very Large Data Bases, 27 (5):595–615, October 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Eich:2018:EGQ

[814] Marius Eich, Pit Fender, and Guido Moerkotte. Efficient generation of query plans containing group-by, join, and groupjoin. VLDB Journal: Very Large Data Bases, 27(5):617–641, October 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Leis:2018:QOT

[815] Viktor Leis, Bernhard Radke, Andrey Gubichev, Atanas Mirchev, Peter Boncz, Alfons Kemper, and Thomas Neumann. Query optimization through the looking glass, and what we found running the Join Order Benchmark. VLDB Journal: Very Large Data Bases, 27(5):643–668, October 2018.

CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Makreshanski:2018:MQJ

[816] Darko Makreshanski, Georgios Giannikis, Gustavo Alonso, and Donald Kossmann. Many-query join: efficient shared execution of relational joins on modern hardware. VLDB Journal: Very Large Data Bases, 27 (5):669–692, October 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Brucato:2018:PQE

[817] Matteo Brucato, Azza Abouzied, and Alexandra Meliou. Package queries: efficient and scalable computation of high-order constraints. VLDB Journal: Very Large Data Bases, 27(5):693– 718, October 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Elgohary:2018:CLA

[818] Ahmed Elgohary, Matthias Boehm, Peter J. Haas, Frederick R. Reiss, and Berthold Reinwald. Compressed linear algebra for large-scale machine learning. VLDB Journal: Very Large Data Bases, 27(5):719–744, October 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Chai:2018:POB

[819] Chengliang Chai, Guoliang Li, Jian Li, Dong Deng, and Jianhua Feng. A partial-order-based framework for costeffective crowdsourced entity resolution. VLDB Journal: Very Large Data Bases, 27(6):745–770, December 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Roblot:2018:PCC

[820] Tania Roblot, Miika Hannula, and Sebastian Link. Probabilistic cardinality constraints. VLDB Journal: Very Large Data Bases, 27(6):771–795, December 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Bress:2018:GCC

[821] Sebastian Breß, Bastian Köcher, Henning Funke, Steffen Zeuch, Tilmann Rabl, and Volker Markl. Generating custom code for efficient query execution on heterogeneous processors. VLDB Journal: Very Large Data Bases, 27(6):797–822, December 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zoumpatianos:2018:GDS

[822] Kostas Zoumpatianos, Yin Lou, Ioana Ileana, Themis Palpanas, and Johannes Gehrke. Generating data series query workloads. VLDB Journal: Very Large Data Bases, 27(6):823–846, December 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

To:2018:SSM

[823] Quoc-Cuong To, Juan Soto, and Volker Markl. A survey of state management in big data processing systems. VLDB Journal: Very Large Data Bases, 27 (6):847–872, December 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Liu:2018:ACE

[824] Yuchen Liu, Hai Liu, Dongqing Xiao, and Mohamed Y. Eltabakh. Adaptive correlation exploitation in big data query optimization. VLDB Journal: Very Large Data Bases, 27(6):873–898, December 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wang:2018:EEM

[825] Tianzheng Wang, Ryan Johnson, Alan Fekete, and Ippokratis Pandis. Erratum to: Efficiently making (almost) any concurrency control mechanism serializable. VLDB Journal: Very Large Data Bases, 27(6):899–900, December 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). See [773].

Rahman:2019:OGF

[826] Habibur Rahman, Senjuti Basu Roy, Saravanan Thirumuruganathan, Sihem Amer-Yahia, and Gautam Das. Optimized group formation for solving collaborative tasks. VLDB Journal: Very Large Data Bases, 28(1):1–23, February 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wu:2019:VFS

[827] Zhiqiang Wu and Kenli Li. VBTree: forward secure conjunctive queries over encrypted data for cloud computing. VLDB Journal: Very Large Data Bases, 28(1):25–46, February 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Lee:2019:PFP

[828] Seokki Lee, Bertram Ludäscher, and Boris Glavic. PUG: a framework and practical implementation for why and why-not provenance. VLDB Journal: Very Large Data Bases, 28(1):47–71, February 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wang:2019:SSS

[829] Wenlu Wang, Ji Zhang, Min-Te Sun, and Wei-Shinn Ku. A scalable spatial skyline evaluation system utilizing parallel independent region groups. VLDB Journal: Very Large Data Bases, 28 (1):73–98, February 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wang:2019:APS

[830] Yue Wang, Lei Chen, Yulin Che, and Qiong Luo. Accelerating pairwise Sim-Rank estimation over static and dynamic graphs. VLDB Journal: Very Large Data Bases, 28(1):99–122, February 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhao:2019:EMC

[831] Kaiqi Zhao, Gao Cong, Jin-Yao Chin, and Rong Wen. Exploring market competition over topics in spatio-temporal document collections. VLDB Journal: Very Large Data Bases, 28(1): 123–145, February 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Picado:2019:LSE

[832] Jose Picado, Arash Termehchy, Alan Fern, and Parisa Ataei. Logical scalability and efficiency of relational learning algorithms. VLDB Journal: Very Large Data Bases, 28(2):147–171, April 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Aluc:2019:BSC

[833] Günes Aluç, M. Tamer Özsu, and Khuzaima Daudjee. Building selfclustering RDF databases using Tunable-LSH. VLDB Journal: Very Large Data Bases, 28(2):173–195, April 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Zhou:2019:RTC

[834] Xiangmin Zhou, Dong Qin, Lei Chen, and Yanchun Zhang. Real-time context-aware social media recommendation. VLDB Journal: Very Large Data Bases, 28(2):197–219, April 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Ntaflos:2019:UAB

[835] Lefteris Ntaflos, George Trimponias, and Dimitris Papadias. A unified agent-based framework for constrained graph partitioning. VLDB Journal: Very Large Data Bases, 28(2):221–241, April 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Omidvar-Tehrani:2019:UGA

[836] Behrooz Omidvar-Tehrani, Sihem Amer-Yahia, and Ria Mae Borromeo. User group analytics: hypothesis generation and exploratory analysis of user data. VLDB Journal: Very Large Data Bases, 28(2):243–266, April 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wang:2019:LSR

[837] Xubo Wang, Lu Qin, Xuemin Lin, Ying Zhang, and Lijun Chang. Leveraging set relations in exact and dynamic set similarity join. VLDB Journal: Very Large Data Bases, 28(2):267–292, April 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Amer-Yahia:2019:TID

[838] Sihem Amer-Yahia, Lei Chen, and Renée J. Miller. Thematic issue on data management for graphs. VLDB Journal: Very Large Data Bases, 28 (3):293–294, June 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Cebiric:2019:SSG

[839] Šejla Cebirić, François Goasdoué, Haridimos Kondylakis, Dimitris Kotzinos, Ioana Manolescu, Georgia Troullinou, and Mussab Zneika. Summarizing semantic graphs: a survey. VLDB Journal: Very Large Data Bases, 28 (3):295–327, June 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Demirci:2019:CAP

[840] Gunduz Vehbi Demirci, Hakan Ferhatosmanoglu, and Cevdet Aykanat. Cascade-aware partitioning of large graph databases. VLDB Journal: Very Large Data Bases, 28(3):329–350, June 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Das:2019:IMM

[841] Apurba Das, Michael Svendsen, and Srikanta Tirthapura. Incremental maintenance of maximal cliques in a dynamic graph. VLDB Journal: Very Large Data Bases, 28(3):351–375, June 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Wen:2019:ESG

[842] Dong Wen, Lu Qin, Ying Zhang, Lijun Chang, and Xuemin Lin. Efficient structural graph clustering: an indexbased approach. VLDB Journal: Very Large Data Bases, 28(3):377–399, June 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yu:2019:SES

[843] Weiren Yu, Xuemin Lin, Wenjie Zhang, Jian Pei, and Julie A. Mccann. Sim-Rank*: effective and scalable pairwise similarity search based on graph topology. VLDB Journal: Very Large Data Bases, 28(3):401–426, June 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Valdes:2019:FEM

[844] Fabio Valdés and Ralf Hartmut Güting. A framework for efficient multi-attribute movement data analysis. VLDB Journal: Very Large Data Bases, 28(4):427–449, August 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-018-
0525-6.

Psaropoulos:2019:ICS

[845] Georgios Psaropoulos, Thomas Legler, Norman May, and Anastasia Ailamaki. Interleaving with coroutines: a systematic and practical approach to hide memory latency in index joins. VLDB Journal: Very Large Data Bases, 28 (4):451–471, August 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.com/article/10.
1007/s00778-018-0533-6.

Bronselaer:2019:CRT

[846] Antoon Bronselaer, Christophe Billiet, Robin De Mol, Joachim Nielandt, and Guy De Tré. Compact representations of temporal databases. VLDB Journal: Very Large Data Bases, 28 (4):473–496, August 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00778-018-0535-4.

Baazizi:2019:PSI

[847] Mohamed-Amine Baazizi, Dario Colazzo, Giorgio Ghelli, and Carlo Sartiani. Parametric schema inference for massive JSON datasets. VLDB Journal: Very Large Data Bases, 28 (4):497–521, August 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00778-018-0532-7.

Li:2019:SCG

[848] Yuan Li, Ahmed Eldawy, Jie Xue, Nadezda Knorozova, Mohamed F. Mokbel, and Ravi Janardan. Scalable computational geometry in MapReduce. VLDB Journal: Very Large Data Bases, 28(4):523–548, August 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00778-018-
                                        0534-5.

Park:2019:FSM

[849] Namyong Park, Sejoon Oh, and U Kang. Fast and scalable method for distributed Boolean tensor factorization. VLDB Journal: Very Large Data Bases, 28(4):549–574, August 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-019-
00538-z.

Frazzetto:2019:PAS

[850] Davide Frazzetto, Thomas Dyhre Nielsen, Torben Bach Pedersen, and Laurynas Šikšnys. Prescriptive analytics: a survey of emerging trends and technologies. VLDB Journal: Very Large Data Bases, 28(4):575–595, August 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-019-
00539-y.

Zhu:2019:FDC

[851] Rong Zhu, Zhaonian Zou, and Jianzhong Li. Fast diversified coherent core search on multi-layer graphs. VLDB Journal: Very Large Data Bases, 28(4):597–622, August 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-019-
00542-3.

AlHasanHaldar:2019:LPL

[852] Nur Al Hasan Haldar, Jianxin Li, Mark Reynolds, Timos Sellis, and Jeffrey Xu Yu. Location prediction in large-scale social networks: an in-depth benchmarking study. VLDB Journal: Very Large Data Bases, 28(5):623–648, October 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-019-
00553-0.

Fang:2019:EPM

[853] Yixiang Fang, Yun Li, Reynold Cheng, Nikos Mamoulis, and Gao Cong. Evaluating pattern matching queries for spatial databases. VLDB Journal: Very Large Data Bases, 28(5):649–673, October 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00778-019-
                                        00550-3.

Kim:2019:ETD

[854] Jinhyun Kim, Jun-Ki Min, and Kyuseok Shim. Efficient twodimensional Haar^{+} synopsis construction for the maximum absolute error measure. VLDB Journal: Very Large Data Bases, 28(5):675–701, October 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00778-019-
                                        00551-2.

Theocharidis:2019:SEM

[855] Konstantinos Theocharidis, John Liagouris, Nikos Mamoulis, Panagiotis Bouros, and Manolis Terrovitis. SRX: efficient management of spatial RDF data. VLDB Journal: Very Large Data Bases, 28(5):703–733, October 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00778-019-
                                        00554-z.

Yang:2019:FAS

[856] Tong Yang, Jie Jiang, Yang Zhou, Long He, Jinyang Li, Bin Cui, Steve Uhlig, and Xiaoming Li. Fast and accurate stream processing by filtering the cold. VLDB Journal: Very Large Data Bases, 28(5):735–763, October 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-019-
00560-1.

Li:2019:ESW

[857] Wentao Li, Miao Qiao, Lu Qin, Ying Zhang, Lijun Chang, and Xuemin Lin. Eccentricities on small-world networks. VLDB Journal: Very Large Data Bases, 28(5):765–792, October 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-019-
00566-9.

Affolter:2019:CSR

[858] Katrin Affolter, Kurt Stockinger, and Abraham Bernstein. A comparative survey of recent natural language interfaces for databases. VLDB Journal: Very Large Data Bases, 28(5): 793–819, October 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.com/article/10.
1007/s00778-019-00567-8;   http:
//link.springer.com/content/pdf/
10.1007/s00778-019-00567-8.pdf.

Cheng:2019:PCD

[859] Zhinan Cheng, Qun Huang, and Patrick P. C. Lee. On the performance and convergence of distributed stream processing via approximate fault tolerance. VLDB Journal: Very Large Data Bases, 28(5):821–846, October 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00778-019-
                                        00565-w.

Kondylakis:2019:CSS

[860] Haridimos Kondylakis, Niv Dayan, Kostas Zoumpatianos, and Themis Palpanas. Coconut: sortable summarizations for scalable indexes over static and streaming data series. VLDB Journal: Very Large Data Bases, 28(6): 847–869, December 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00778-019-00573-w.

Zhang:2019:EDR

[861] Tianming Zhang, Yunjun Gao, Lu Chen, Wei Guo, Shiliang Pu, Baihua Zheng, and Christian S. Jensen. Efficient distributed reachability querying of massive temporal graphs. VLDB Journal: Very Large Data Bases, 28(6): 871–896, December 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00778-019-00572-x.

Lin:2019:OPT

[862] Xuelian Lin, Jiahao Jiang, Shuai Ma, Yimeng Zuo, and Chunming Hu. One-pass trajectory simplification using the synchronous Euclidean distance. VLDB Journal: Very Large Data Bases, 28(6):897–921, December 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00778-019-
                                        00575-8.

Wang:2019:PAS

[863] Runhui Wang, Sibo Wang, and Xiaofang Zhou. Parallelizing approximate single-source personalized PageRank queries on shared memory. VLDB Journal: Very Large Data Bases, 28 (6):923–940, December 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.com/article/10.
1007/s00778-019-00576-7.

Langdale:2019:PGJ

[864] Geoff Langdale and Daniel Lemire. Parsing gigabytes of JSON per second. VLDB Journal: Very Large Data Bases, 28(6):941–960, December 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-019-
00578-5.

Ren:2019:SQI

[865] Weilong Ren, Xiang Lian, and Kambiz Ghazinour. Skyline queries over incomplete data streams. VLDB Journal: Very Large Data Bases, 28(6): 961–985, December 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.com/article/10.
1007/s00778-019-00577-6.

Zhang:2019:ECD

[866] Fan Zhang, Xuemin Lin, Ying Zhang, Lu Qin, and Wenjie Zhang. Efficient community discovery with user engagement and similarity. VLDB Journal: Very Large Data Bases, 28(6): 987–1012, December 2019. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.com/article/10.
1007/s00778-019-00579-4.

Chen:2020:VSS

[867] Lei Chen and Sihem Amer-Yahia. VLDB SI survey editorial. VLDB Journal: Very Large Data Bases, 29(1):1–2, January 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.com/article/10.
1007/s00778-019-00598-1;   http:
//link.springer.com/content/pdf/
10.1007/s00778-019-00598-1.pdf.

Su:2020:STD

[868] Han Su, Shuncheng Liu, Bolong Zheng, Xiaofang Zhou, and Kai Zheng. A survey of trajectory distance measures and performance evaluation. VLDB Journal: Very Large Data Bases, 29 (1):3–32, January 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.com/article/10.
1007/s00778-019-00574-9.

Fang:2020:MDA

[869] Jian Fang, Yvo T. B. Mulder, Jan Hidders, Jinho Lee, and H. Peter Hofstee. In-memory database acceleration on FPGAs: a survey. VLDB Journal: Very Large Data Bases, 29 (1):33–59, January 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.com/article/10.
1007/s00778-019-00581-w;   http:
//link.springer.com/content/pdf/
10.1007/s00778-019-00581-w.pdf.

Malliaros:2020:CDN

[870] Fragkiskos D. Malliaros, Christos Giatsidis, Apostolos N. Papadopoulos, and Michalis Vazirgiannis. The core decomposition of networks: theory, algorithms and applications. VLDB Journal: Very Large Data Bases, 29 (1):61–92, January 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.com/article/10.
1007/s00778-019-00587-4.

Qin:2020:MDV

[871] Xuedi Qin, Yuyu Luo, Nan Tang, and Guoliang Li. Making data visualization more efficient and effective: a survey. VLDB Journal: Very Large Data Bases, 29(1):93–117, January 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-019-
00588-3.

Rahman:2020:EID

[872] Protiva Rahman, Lilong Jiang, and Arnab Nandi. Evaluating interactive data systems. VLDB Journal: Very Large Data Bases, 29(1):119–146, January 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-019-
00589-2.

Xie:2020:ESR

[873] Min Xie, Raymond Chi-Wing Wong, and Ashwin Lall. An experimental survey of regret minimization query and variants: bridging the best worlds between top-k query and skyline query. VLDB Journal: Very Large Data Bases, 29(1):147–175, January 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-019-
00570-z.

Magdy:2020:MDM

[874] Amr Magdy, Laila Abdelhafeez, Yunfan Kang, Eric Ong, and Mohamed F. Mokbel. Microblogs data management: a survey. VLDB Journal: Very Large Data Bases, 29(1):177–216, January 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-019-
00569-6.

Tong:2020:SCS

[875] Yongxin Tong, Zimu Zhou, Yuxiang Zeng, Lei Chen, and Cyrus Shahabi. Spatial crowdsourcing: a survey. VLDB Journal: Very Large Data Bases, 29(1):217–250, January 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-019-
00568-7.

Chapman:2020:DSS

[876] Adriane Chapman, Elena Simperl, Laura Koesten, George Konstantinidis, Luis-Daniel Ibáñez, Emilia Kacprzak, and Paul Groth. Dataset search: a survey. VLDB Journal: Very Large Data Bases, 29(1):251–272, January 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-019-
00564-x;   http://link.springer.

com/content/pdf/10.1007/s00778-
019-00564-x.pdf.

Fevgas:2020:IFS

[877] Athanasios Fevgas, Leonidas Akritidis, Panayiotis Bozanis, and Yannis Manolopoulos. Indexing in flash storage devices: a survey on challenges, current approaches, and future trends. VLDB Journal: Very Large Data Bases, 29(1):273–311, January 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-019-
00559-8.

Giatrakos:2020:CER

[878] Nikos Giatrakos, Elias Alevizos, Alexander Artikis, Antonios Deligiannakis, and Minos Garofalakis. Complex event recognition in the Big Data era: a survey. VLDB Journal: Very Large Data Bases, 29(1):313–352, January 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-019-
00557-w.

Fang:2020:SCS

[879] Yixiang Fang, Xin Huang, Lu Qin, Ying Zhang, Wenjie Zhang, Reynold Cheng, and Xuemin Lin. A survey of community search over big graphs. VLDB Journal: Very Large Data Bases, 29(1):353–392, January 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-019-

00556-x. See correction [914].

Luo:2020:LBS

[880] Chen Luo and Michael J. Carey. LSM-based storage techniques: a survey. VLDB Journal: Very Large Data Bases, 29(1):393–418, January 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00778-019-
                                        00555-y.

Blumenthal:2020:CHG

[881] David B. Blumenthal, Nicolas Boria, Johann Gamper, Sébastien Bougleux, and Luc Brun. Comparing heuristics for graph edit distance computation. VLDB Journal: Very Large Data Bases, 29(1):419–458, January 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00778-019-
                                        00544-1.

Chen:2020:EMM

[882] Xinhong Chen and Qing Li. Event modeling and mining: a long journey toward explainable events. VLDB Journal: Very Large Data Bases, 29 (1):459–482, January 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00778-019-00545-0.

Boncz:2020:SIB

[883] Peter Boncz and Kenneth Salem. Special issue on best papers of VLDB 2017. VLDB Journal: Very Large Data Bases, 29(1):483–484, January 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00778-019-
00600-w;   http://link.springer.
com/content/pdf/10.1007/s00778-
019-00600-w.pdf.

Deutch:2020:ENL

[884] Daniel Deutch, Nave Frost, and Amir Gilad. Explaining natural language query results. VLDB Journal: Very Large Data Bases, 29(1):485–508, January 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-019-
00584-7.

Huang:2020:VOD

[885] Silu Huang, Liqi Xu, Jialin Liu, Aaron J. Elmore, and Aditya Parameswaran. OrpheusDB: bolt-on versioning for relational databases (extended version). VLDB Journal: Very Large Data Bases, 29(1):509–538, January 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-019-
00594-5.

Orr:2020:EPA

[886] Laurel Orr, Magdalena Balazinska, and Dan Suciu. EntropyDB: a probabilistic approach to approximate query processing. VLDB Journal: Very Large Data Bases, 29(1):539–567, January 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-019-
00582-9.

Amer-Yahia:2020:VSE

[887] Sihem Amer-Yahia and Jian Pei. VLDB SI 2018 editorial. VLDB Journal: Very Large Data Bases, 29 (2–3):593–594, May 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.com/article/10.
1007/s00778-019-00599-0;   http:
//link.springer.com/content/pdf/
10.1007/s00778-019-00599-0.pdf.

Sahu:2020:ULG

[888] Siddhartha Sahu, Amine Mhedhbi, Semih Salihoglu, Jimmy Lin, and M. Tamer Özsu. The ubiquity of large graphs and surprising challenges of graph processing: extended survey. VLDB Journal: Very Large Data Bases, 29(2–3):595–618, May 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-019-
00548-x.

Idris:2020:GDY

[889] Muhammad Idris, Martín Ugarte, Stijn Vansummeren, Hannes Voigt, and Wolfgang Lehner. General dynamic Yannakakis: conjunctive queries with theta joins under updates. VLDB Journal: Very Large Data Bases, 29 (2–3):619–653, May 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.com/article/10.
1007/s00778-019-00590-9.

Bonifati:2020:ASL

[890] Angela Bonifati, Wim Martens, and Thomas Timm. An analytical study of large SPARQL query logs. VLDB Journal: Very Large Data Bases, 29 (2–3):655–679, May 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.com/article/10.
1007/s00778-019-00558-9.

Asudeh:2020:SAS

[891] Abolfazl Asudeh, Jees Augustine, Azade Nazi, Saravanan Thirumuruganathan, Nan Zhang, Gautam Das, and Divesh Srivastava. Scalable algorithms for signal reconstruction by leveraging similarity joins. VLDB Journal: Very Large Data Bases, 29 (2–3):681–707, May 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.com/article/10.
1007/s00778-019-00562-z.

Ratner:2020:SRT

[892] Alexander Ratner, Stephen H. Bach, Henry Ehrenberg, Jason Fries, Sen Wu, and Christopher Ré. Snorkel: rapid training data creation with weak supervision. VLDB Journal: Very Large Data Bases, 29(2–3):709–730, May 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-019-
00552-1;   http://link.springer.
com/content/pdf/10.1007/s00778-
019-00552-1.pdf.

Breslow:2020:MFF

[893] Alex D. Breslow and Nuwan S. Jayasena. Morton filters: fast, compressed sparse cuckoo filters. VLDB Journal: Very Large Data Bases, 29 (2–3):731–754, May 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.com/article/10.
1007/s00778-019-00561-0.

Salem:2020:SIB

[894] Kenneth Salem. Special issue on best papers of DaMoN 2018. VLDB Journal: Very Large Data Bases, 29(2–3):755, May 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00778-019-00597-2;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00778-019-00597-2.pdf.

Lang:2020:MMY

[895] Harald Lang, Linnea Passing, Andreas Kipf, Peter Boncz, Thomas Neumann, and Alfons Kemper. Make the most out of your SIMD investments: counter control flow divergence in compiled query pipelines. VLDB Journal: Very Large Data Bases, 29 (2–3):757–774, May 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00778-019-00547-y;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00778-019-00547-y.pdf.

Zarubin:2020:ECN

[896] Mikhail Zarubin, Thomas Kissinger, Dirk Habich, Thomas Willhalm, and Wolfgang Lehner. Efficient compute node-local replication mechanisms for NVRAM-centric data structures. VLDB Journal: Very Large Data Bases, 29(2–3):775–795, May 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00778-019-
                                        00549-w.

Pohl:2020:JHB

[897] Constantin Pohl, Kai-Uwe Sattler, and Goetz Graefe. Joins on high-bandwidth memory: a new level in the memory hierarchy. VLDB Journal: Very Large Data Bases, 29(2–3):797–817, May 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00778-019-
00546-z.

Pedersen:2020:FSR

[898] Simon Aagaard Pedersen, Bin Yang, and Christian S. Jensen. Fast stochastic routing under time-varying uncertainty. VLDB Journal: Very Large Data Bases, 29(4):819–839, July 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-019-00585-
6.

Xu:2020:EPM

[899] Hongfei Xu, Yu Gu, and Rui Zhang. Efficient processing of moving collective spatial keyword queries. VLDB Journal: Very Large Data Bases, 29 (4):841–865, July 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-019-00583-8.

Geerts:2020:CDL

[900] Floris Geerts, Giansalvatore Mecca, and Donatello Santoro. Cleaning data with Llunatic. VLDB Journal: Very Large Data Bases, 29 (4):867–892, July 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-019-00586-5.

Wu:2020:TRS

[901] Dingming Wu, Hao Zhou, and Nikos Mamoulis. Top-k relevant semantic place retrieval on spatiotemporal RDF data. VLDB Journal: Very Large Data Bases, 29 (4):893–917, July 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-019-00591-8.

Qin:2020:EQA

[902] Jianbin Qin, Chuan Xiao, and Kunihiko Sadakane. Efficient query autocompletion with edit distancebased error tolerance. VLDB Journal: Very Large Data Bases, 29 (4):919–943, July 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-019-00595-4.

Jiang:2020:SCS

[903] Jiawei Jiang, Fangcheng Fu, and Bin Cui. SKCompress: compressing sparse and nonuniform gradient in distributed machine learning. VLDB Journal: Very Large Data Bases, 29 (5):945–972, September 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-019-00596-3.

Wang:2020:FEF

[904] Chaohui Wang, Miao Xie, and Shuigeng Zhou. FERRARI: an efficient framework for visual exploratory subgraph search in graph databases. VLDB Journal: Very Large Data Bases, 29(5):973–998, September 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-020-00601-
0.

Chang:2020:EMC

[905] Lijun Chang. Efficient maximum clique computation and enumeration over large sparse graphs. VLDB Journal: Very Large Data Bases, 29(5): 999–1022, September 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00602-z.

Chondrogiannis:2020:FSP

[906] Theodoros Chondrogiannis, Panagiotis Bouros, and David B. Blumenthal. Finding k-shortest paths with limited overlap. VLDB Journal: Very Large Data Bases, 29(5):1023– 1047, September 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00604-x.

Zou:2020:ADS

[907] Jia Zou, Arun Iyengar, and Chris Jermaine. Architecture of a distributed storage that combines file system, memory and computation in a single layer. VLDB Journal: Very Large Data Bases, 29(5):1049– 1073, September 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00605-w.

Liu:2020:ECC

[908] Boge Liu, Long Yuan, and Jingren Zhou. Efficient (α, β)-core computation in bipartite graphs. VLDB Journal: Very Large Data Bases, 29(5): 1075–1099, September 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-020-00606-9.

Chen:2020:TTP

[909] Lisi Chen, Shuo Shang, and Ling Shao. Top-k term publish/subscribe for geotextual data streams. VLDB Journal: Very Large Data Bases, 29(5): 1101–1128, September 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-020-00607-8.

Yang:2020:TFA

[910] Fan Yang, Faisal M. Almutairi, and Vladimir Zadorozhny. Turbo-Lift: fast accuracy lifting for historical data recovery. VLDB Journal: Very Large Data Bases, 29(5): 1129–1148, September 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-020-00609-6. See cor-

rection [1105].

Guo:2020:CAP

[911] Chenjuan Guo, Bin Yang, and Lu Chen. Context-aware, preferencebased vehicle routing. VLDB Journal: Very Large Data Bases, 29(5): 1149–1170, September 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00608-7.

Cai:2020:DSK

[912] Zhi Cai, Georgios Kalamatianos, and Dimitris Papadias. Diversified spatial keyword search on RDF data. VLDB Journal: Very Large Data Bases, 29(5): 1171–1189, September 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00610-z.

Goasdoue:2020:RGS

[913] François Goasdoué, Paweł Guzewicz, and Ioana Manolescu. RDF graph summarization for first-sight structure discovery. VLDB Journal: Very Large Data Bases, 29(5):1191– 1218, September 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00611-y.

Fang:2020:CSC

[914] Yixiang Fang, Xin Huang, and Xuemin Lin. Correction: A survey of community search over big graphs. VLDB Journal: Very Large Data Bases, 29 (5):1219, September 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-019-00592-7. See [879].

Aboulnaga:2020:SIB

[915] Ashraf Aboulnaga. Special issue on the best papers of DaMoN 2019. VLDB Journal: Very Large Data Bases, 29 (6):1221, November 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-020-00629-2.

vanRenen:2020:BBP

[916] Alexander van Renen, Lukas Vogel, and Alfons Kemper. Building blocks for persistent memory. VLDB Journal: Very Large Data Bases, 29(6): 1223–1241, November 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-020-00622-9.

Polychroniou:2020:VSV

[917] Orestis Polychroniou and Kenneth A. Ross. VIP: a SIMD vectorized analytical query engine. VLDB Journal: Very Large Data Bases, 29(6): 1243–1261, November 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-020-00621-w.

Lasch:2020:FSS

[918] Robert Lasch, Ismail Oukid, and Kai-Uwe Sattler. Faster & strong: string dictionary compression using sampling and fast vectorized decompression. VLDB Journal: Very Large Data Bases, 29(6):1263– 1285, November 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-020-00620-x.

Kruse:2020:RDJ

[919] Sebastian Kruse, Zoi Kaoudi, and Jorge-Arnulfo Quiané-Ruiz. RHEEMix in the data jungle: a cost-based optimizer for cross-platform systems. VLDB Journal: Very Large Data Bases, 29(6):1287–1310, November 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-020-00612-
x.

Yang:2020:GBF

[920] Jingru Yang, Ju Fan, and Xiaoyong Du. A game-based framework for crowdsourced data labeling. VLDB Journal: Very Large Data Bases, 29(6): 1311–1336, November 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00613-w.

Jacobs:2020:BBB

[921] Steven Jacobs, Xikui Wang, and Md Yusuf Sarwar Uddin. BAD to the bone: Big Active Data at its core. VLDB Journal: Very Large Data Bases, 29(6):1337– 1364, November 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00616-7.

Sun:2020:TSI

[922] Tao Sun, Hongbo Liu, and Xindong Wu. Time series indexing by dynamic covering with crossrange constraints. VLDB Journal: Very Large Data Bases, 29(6):1365– 1384, November 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00614-9.

Huang:2020:EAA

[923] Keke Huang, Jing Tang, and Andrew Lim. Efficient approximation algorithms for adaptive influence maximization. VLDB Journal: Very Large Data Bases, 29(6): 1385–1406, November 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-020-00615-8.

Li:2020:FSC

[924] Rong-Hua Li, Lu Qin, and Zibin Zheng. Finding skyline communities in multi-valued networks. VLDB Journal: Very Large Data Bases, 29(6): 1407–1432, November 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-020-00618-5.

Ahmad:2020:AWM

[925] Hiba Abu Ahmad and Hongzhi Wang. Automatic weighted matching rectifying rule discovery for data repairing. VLDB Journal: Very Large Data Bases, 29(6):1433– 1447, November 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-020-00617-6.

Linardi:2020:SDS

[926] Michele Linardi and Themis Palpanas. Scalable data series subsequence matching with ULISSE. VLDB Journal: Very Large Data Bases, 29(6): 1449–1474, November 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00619-4.

Song:2020:IPA

[927] Liangjun Song, Junhao Gan, and Timos Sellis. Incremental preference adjustment: a graph-theoretical approach. VLDB Journal: Very Large Data Bases, 29(6):1475–1500, November 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https://link.

springer.com/article/10.1007/s00778-
020-00623-8.

Lee:2020:TLA

[928] Dongjin Lee, Kijung Shin, and Christos Faloutsos. Temporal locality-aware sampling for accurate triangle counting in real graph streams. VLDB Journal: Very Large Data Bases, 29(6): 1501–1525, November 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00624-7.

Omidvar-Tehrani:2020:CAE

[929] Behrooz Omidvar-Tehrani, Sihem Amer-Yahia, and Laks V. S. Lakshmanan. Cohort analytics: efficiency and applicability. VLDB Journal: Very Large Data Bases, 29(6): 1527–1550, November 2020. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00625-6.

Ozcan:2021:GES

[930] Fatma Özcan and Lei Chen. Guest editorial: Special issue on VLDB 2019. VLDB Journal: Very Large Data Bases, 30(1):1–2, January 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-020-00630-
9.

Ruan:2021:LFG

[931] Pingcheng Ruan, Tien Tuan Anh Dinh, and Beng Chin Ooi. LineageChain: a fine-grained, secure and efficient data provenance system for blockchains. VLDB Journal: Very Large Data Bases, 30 (1):3–24, January 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00646-1.

Wu:2021:ATC

[932] Chenggang Wu, Vikram Sreekanti, and Joseph M. Hellerstein. Autoscaling tiered cloud storage in Anna. VLDB Journal: Very Large Data Bases, 30 (1):25–43, January 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00632-7.

Abuzaid:2021:DRI

[933] Firas Abuzaid, Peter Kraft, and Matei Zaharia. DIFF: a relational interface for large-scale data explanation. VLDB Journal: Very Large Data Bases, 30 (1):45–70, January 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00633-6.

Whittaker:2021:ICC

[934] Michael Whittaker and Joseph M. Hellerstein. Interactive checks for coordination avoidance. VLDB Journal: Very Large Data Bases, 30 (1):71–92, January 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00628-3.

Fan:2021:GBV

[935] Hua Fan and Wojciech Golab. Gossipbased visibility control for highperformance geo-distributed transactions. VLDB Journal: Very Large Data Bases, 30(1):93–114, January 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-020-00626-
5.

Li:2021:QSD

[936] Yuliang Li, Aaron Feng, and Wang-Chiew Tan. Querying subjective data. VLDB Journal: Very Large Data Bases, 30(1):115–140, January 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-020-00634-
5.

Dong:2021:CTS

[937] Yuyang Dong, Chuan Xiao, and Hiroyuki Kitagawa. Continuous topk spatial-keyword search on dynamic objects. VLDB Journal: Very Large Data Bases, 30(2): 141–161, March 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-020-00627-4.

Zhang:2021:TTA

[938] Feng Zhang, Jidong Zhai, and Xiaoyong Du. TADOC: Text analytics directly on compression. VLDB Journal: Very Large Data Bases, 30 (2):163–188, March 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-020-00636-3.

Li:2021:CTQ

[939] Yan Li, Hao Wang, and Zhiguo Gong. Crowdsourced top-k queries by pairwise preference judgments with confidence and budget control. VLDB Journal: Very Large Data Bases, 30 (2):189–213, March 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-020-00631-8.

Liu:2021:LET

[940] Wanqi Liu, Hanchen Wang, and Xuemin Lin. EI-LSH: an earlytermination driven I/O efficient incremental c-approximate nearest neighbor search. VLDB Journal: Very Large Data Bases, 30 (2):215–235, March 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-020-00635-4.

Yu:2021:GCC

[941] Jia Yu and Mohamed Sarwat. GeoSparkViz: a cluster computing system for visualizing massive-scale geospatial data.

VLDB Journal: Very Large Data Bases, 30(2):237–258, March 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-020-00645-
2.

Zhang:2021:SAN

[942] Yongqi Zhang, Quanming Yao, and Lei Chen. Simple and automated negative sampling for knowledge graph embedding. VLDB Journal: Very Large Data Bases, 30 (2):259–285, March 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00640-7.

Fang:2021:DHE

[943] Ziquan Fang, Lu Chen, and Christian S. Jensen. Dragoon: a hybrid and efficient big trajectory management system for offline and online analytics. VLDB Journal: Very Large Data Bases, 30 (2):287–310, March 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00652-x.

Paul:2021:SER

[944] Debjyoti Paul, Feifei Li, and Jeff M. Phillips. Semantic embedding for regions of interest. VLDB Journal: Very Large Data Bases, 30 (3):311–331, May 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00647-0.

Romanous:2021:ELL

[945] Bashar Romanous, Skyler Windh, and Vassilis Tsotras. Efficient local locking for massively multithreaded inmemory hash-based operators. VLDB Journal: Very Large Data Bases, 30 (3):333–359, May 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-020-00642-5.

Mao:2021:CES

[946] Qizhong Mao, Steven Jacobs, and Neal E. Young. Comparison and evaluation of state-of-the-art LSM merge policies. VLDB Journal: Very Large Data Bases, 30(3): 361–378, May 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-020-00638-1.

Piatov:2021:CES

[947] Danila Piatov, Sven Helmer, and Fabio Persia. Cache-efficient sweepingbased interval joins for extended Allen relation predicates. VLDB Journal: Very Large Data Bases, 30 (3):379–402, May 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-020-00650-5.

Do:2021:BDC

[948] Jaeyoung Do, Ivan Luiz Picoli, and Philippe Bonnet. Better database cost/performance via batched I/O on programmable SSD. VLDB Journal: Very Large Data Bases, 30 (3):403–424, May 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00648-z.

Song:2021:CTT

[949] Shaoxu Song, Ruihong Huang, and Jianmin Wang. Cleaning timestamps with temporal constraints. VLDB Journal: Very Large Data Bases, 30 (3):425–446, May 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00641-6.

Yang:2021:IEM

[950] Chengcheng Yang, Dong Deng, and Ling Shao. Internal and external memory set containment join. VLDB Journal: Very Large Data Bases, 30 (3):447–470, May 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00644-3. See cor-

rection [968].

Chen:2021:ESN

[951] Xiaoshuang Chen, Longbin Lai, and Xuemin Lin. Efficient structural node similarity computation on billion-scale graphs. VLDB Journal: Very Large Data Bases, 30 (3):471–493, May 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00654-9.

Yu:2021:VAR

[952] Wenhui Yu, Xiangnan He, and Zheng Qin. Visually aware recommendation with aesthetic features. VLDB Journal: Very Large Data Bases, 30 (4):495–513, July 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-021-00651-y.

Hao:2021:MCE

[953] Shuang Hao, Nan Tang, and Ning Wang. Mis-categorized entities detection. VLDB Journal: Very Large Data Bases, 30(4):515–536, July 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00778-021-00653-
                                        w.

Galhotra:2021:EEP

[954] Sainyam Galhotra, Donatella Firmani, and Divesh Srivastava. Efficient and effective ER with progressive blocking. VLDB Journal: Very Large Data Bases, 30 (4):537–557, July 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-021-00656-7.

Hewasinghage:2021:CMR

[955] Moditha Hewasinghage, Alberto Abelló, and Esteban Zimányi. A cost model for random access queries in document stores. VLDB Journal: Very Large Data Bases, 30 (4):559–578, July 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-021-00660-x.

Schneider:2021:DDS

[956] Johannes Schneider, Phillip Wenig, and Thorsten Papenbrock. Distributed detection of sequential anomalies in univariate time series. VLDB Journal: Very Large Data Bases, 30 (4):579–602, July 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00657-6.

Chen:2021:LKB

[957] Zhida Chen, Lisi Chen, and Christian S. Jensen. Location- and keyword-based querying of geo-textual data: a survey. VLDB Journal: Very Large Data Bases, 30 (4):603–640, July 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00661-w.

Sirin:2021:MAA

[958] Utku Sirin, Pinar Tözün, and Anastasia Ailamaki. Micro-architectural analysis of in-memory OLTP: Revisited. VLDB Journal: Very Large Data Bases, 30(4):641–665, July 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-021-00663-
8.

Bouros:2021:MIJ

[959] Panagiotis Bouros, Nikos Mamoulis, and Manolis Terrovitis. In-memory interval joins. VLDB Journal: Very Large Data Bases, 30(4): 667–691, July 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00639-0.

Guo:2021:MAD

[960] Yunyan Guo, Zhipeng Zhang, and Jianzhong Li. Model averaging in distributed machine learning: a case study with Apache Spark. VLDB Journal: Very Large Data Bases, 30 (4):693–712, July 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00664-7.

Jiang:2021:EEK

[961] Yuli Jiang, Xin Huang, and Hong Cheng. I/O efficient k-truss community search in massive graphs. VLDB Journal: Very Large Data Bases, 30 (5):713–738, September 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00649-y.

Balayn:2021:MBU

[962] Agathe Balayn, Christoph Lofi, and Geert-Jan Houben. Managing bias and unfairness in data for decision support: a survey of machine learning and data engineering approaches to identify and mitigate bias and unfairness within data management and analytics systems. VLDB Journal: Very Large Data Bases, 30(5): 739–768, September 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00671-8.

Shao:2021:MAF

[963] Yingxia Shao, Shiyue Huang, and Lei Chen. Memory-aware framework for fast and scalable secondorder random walk over billion-edge natural graphs. VLDB Journal: Very Large Data Bases, 30(5):769– 797, September 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00669-2.

Peng:2021:EHC

[964] You Peng, Xuemin Lin, and Jingren Zhou. Efficient hop-constrained s–t simple path enumeration. VLDB Journal: Very Large Data Bases, 30(5): 799–823, September 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00674-5.

Debrouvier:2021:MQL

[965] Ariel Debrouvier, Eliseo Parodi, and Alejandro Vaisman. A model and query language for temporal graph databases. VLDB Journal: Very Large Data Bases, 30(5):825– 858, September 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00675-4.

Wang:2021:FSH

[966] Jin Wang, Jiacheng Wu, and Carlo Zaniolo. Formal semantics and high performance in declarative machine learning using Datalog. VLDB Journal: Very Large Data Bases, 30(5): 859–881, September 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00665-6.

Kersten:2021:TTF

[967] Timo Kersten, Viktor Leis, and Thomas Neumann. Tidy tuples and flying start: fast compilation and fast execution of relational queries in Umbra. VLDB Journal: Very Large Data Bases, 30(5): 883–905, September 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-020-00643-4.

Yang:2021:CIE

[968] Chengcheng Yang, Dong Deng, and Ling Shao. Correction to: Internal and external memory set containment join. VLDB Journal: Very Large Data Bases, 30 (5):907, September 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00662-9. See [950].

Boniol:2021:USS

[969] Paul Boniol, Michele Linardi, and Emmanuel Remy. Unsupervised and scalable subsequence anomaly detection in large data series. VLDB Journal: Very Large Data Bases, 30(6): 909–931, November 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00655-8. See cor-

rection [1054].

Tangwongsan:2021:OSW

[970] Kanat Tangwongsan, Martin Hirzel, and Scott Schneider. In-order sliding-window aggregation in worstcase constant time. VLDB Journal: Very Large Data Bases, 30(6): 933–957, November 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00668-3.

Zhang:2021:HCE

[971] Ji Zhang, Ke Zhou, and Jiashu Xing. CDBTune^{+}: an efficient deep reinforcement learning-based automatic cloud database tuning system. VLDB Journal: Very Large Data Bases, 30(6): 959–987, November 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00670-9.

Wang:2021:EBS

[972] Hanzhi Wang, Zhewei Wei, and Ji-Rong Wen. ExactSim: benchmarking single-source SimRank algorithms with high-precision ground truths. VLDB Journal: Very Large Data Bases, 30 (6):989–1015, November 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00672-7.

Forresi:2021:DBF

[973] Chiara Forresi, Enrico Gallinucci, and Hamdi Ben Hamadou. A dataspacebased framework for OLAP analyses in a high-variety multistore. VLDB Journal: Very Large Data Bases, 30(6): 1017–1040, November 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-021-00682-5.

Peng:2021:FDS

[974] Botao Peng, Panagiota Fatourou, and Themis Palpanas. Fast data series indexing for in-memory data. VLDB Journal: Very Large Data Bases, 30(6): 1041–1067, November 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-021-00677-2.

Wei:2021:ADE

[975] Ziheng Wei, Sven Hartmann, and Sebastian Link. Algorithms for the discovery of embedded functional dependencies. VLDB Journal: Very Large Data Bases, 30(6): 1069–1093, November 2021. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-021-00684-3.

Kossmann:2022:DDQ

[976] Jan Kossmann, Thorsten Papenbrock, and Felix Naumann. Data dependencies for query optimization: a survey. VLDB Journal: Very Large Data Bases, 31 (1):1–22, January 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-021-00676-3. See cor-

rection [1055].

Zhu:2022:PSA

[977] Yifan Zhu, Lu Chen, and Christian S. Jensen. Pivot selection algorithms in metric spaces: a survey and experimental study. VLDB Journal: Very Large Data Bases, 31 (1):23–47, January 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00691-4.

Schmidl:2022:EDD

[978] Sebastian Schmidl and Thorsten Papenbrock. Efficient distributed discovery of bidirectional order dependencies. VLDB Journal: Very Large Data Bases, 31(1):49–74, January 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-021-00683-
4.

Dignos:2022:LRJ

[979] Anton Dignös, Michael H. Böhlen, and Peter Moser. Leveraging range joins for the computation of overlap joins. VLDB Journal: Very Large Data Bases, 31 (1):75–99, January 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00692-3.

Peng:2022:ARE

[980] You Peng, Xuemin Lin, and Lu Qin. Answering reachability and K-reach queries on large graphs with label constraints. VLDB Journal: Very Large Data Bases, 31(1): 101–127, January 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00695-0.

Li:2022:DLP

[981] Wentao Li, Miao Qiao, and Xuemin Lin. Distance labeling: on parallelism, compression, and ordering. VLDB Journal: Very Large Data Bases, 31 (1):129–155, January 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-021-00694-1.

Alevizos:2022:CEF

[982] Elias Alevizos, Alexander Artikis, and Georgios Paliouras. Complex event forecasting with prediction suffix trees. VLDB Journal: Very Large Data Bases, 31(1): 157–180, January 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-021-00698-x.

Snodgrass:2022:QOH

[983] Richard T. Snodgrass, Sabah Currim, and Young-Kyoon Suh. Have query optimizers hit the wall? VLDB Journal: Very Large Data Bases, 31 (1):181–200, January 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-021-00689-y.

Bonifati:2022:SIB

[984] Angela Bonifati and Hannes Voigt. Special issue on big graph data management and processing. VLDB Journal: Very Large Data Bases, 31 (2):201–202, March 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00732-6.

Wang:2022:TES

[985] Kai Wang, Xuemin Lin, and Ying Zhang. Towards efficient solutions of bitruss decomposition for largescale bipartite graphs. VLDB Journal: Very Large Data Bases, 31 (2):203–226, March 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00658-5.

Linghu:2022:ACE

[986] Qingyuan Linghu, Fan Zhang, and Ying Zhang. Anchored coreness: efficient reinforcement of social networks. VLDB Journal: Very Large Data Bases, 31(2):227–252, March 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-021-00673-
6.

Yan:2022:PPF

[987] Da Yan, Wenwen Qu, and Yang Zhou. PrefixFPM: a parallel framework for general-purpose mining of frequent and closed patterns. VLDB Journal: Very Large Data Bases, 31 (2):253–286, March 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00687-0.

Yan:2022:GTG

[988] Da Yan, Guimu Guo, and John C. S. Lui. G-thinker: a general distributed framework for finding qualified subgraphs in a big graph with load balancing. VLDB Journal: Very Large Data Bases, 31 (2):287–320, March 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00688-z.

Mohamed:2022:RKG

[989] Aisha Mohamed, Ghadeer Abuoda, and Ashraf Aboulnaga. RDFFrames: knowledge graph access for machine learning tools. VLDB Journal: Very Large Data Bases, 31 (2):321–346, March 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00690-5.

Sagi:2022:DSR

[990] Tomer Sagi, Matteo Lissandrini, and Katja Hose. A design space for RDF data representations. VLDB Journal: Very Large Data Bases, 31 (2):347–373, March 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00725-x.

Rost:2022:DTG

[991] Christopher Rost, Kevin Gomez, and Erhard Rahm. Distributed temporal graph analytics with GRADOOP. VLDB Journal: Very Large Data Bases, 31(2):375–401, March 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-021-00667-

Bevilacqua:2022:FME

[992] Glenn S. Bevilacqua and Laks V. S. Lakshmanan. A fractional memory-efficient approach for online continuous-time influence maximization. VLDB Journal: Very Large Data Bases, 31(2):403–429, March 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-021-00679-
0.

Ali:2022:SRS

[993] Waqas Ali, Muhammad Saleem, and Axel-Cyrille Ngonga Ngomo. A survey of RDF stores & SPARQL engines for querying knowledge graphs. VLDB Journal: Very Large Data Bases, 31(3):1–26, May 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00711-3.

Pitoura:2022:FRR

[994] Evaggelia Pitoura, Kostas Stefanidis, and Georgia Koutrika. Fairness in rankings and recommendations: an overview. VLDB Journal: Very Large Data Bases, 31 (3):431–458, May 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00697-y.

Hidayat:2022:CMM

[995] Arif Hidayat, Muhammad Aamir Cheema, and Ying Zhang. Continuous monitoring of moving skyline and top-k queries. VLDB Journal: Very Large Data Bases, 31 (3):459–482, May 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-021-00702-4.

Farhan:2022:FFD

[996] Muhammad Farhan, Qing Wang, and Brendan McKay. Fast fully dynamic labelling for distance queries. VLDB Journal: Very Large Data Bases, 31 (3):483–506, May 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-021-00707-z.

Zhao:2022:RCS

[997] Tianyu Zhao, Shuai Huang, and Guoliang Li. RNE: computing shortest paths using road network embedding. VLDB Journal: Very Large Data Bases, 31(3):507–528, May 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00778-021-00705-
                                        1.

Lai:2022:AMW

[998] Zhuohang Lai, Xibo Sun, and Xiaolong Xie. Accelerating multi-way joins on the GPU. VLDB Journal: Very Large Data Bases, 31 (3):529–553, May 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-021-00708-y.

diVimercati:2022:AMQ

[999] Sabrina De Capitani di Vimercati, Sara Foresti, and Pierangela Samarati. An authorization model for query execution in the cloud. VLDB Journal: Very Large Data Bases, 31 (3):555–579, May 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00709-x.

Huang:2022:PEG

[1000] Kai Huang, Haibo Hu, and Xiaofang Zhou. Privacy and efficiency guaranteed social subgraph matching. VLDB Journal: Very Large Data Bases, 31 (3):581–602, May 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00706-0.

Wen:2022:SRQ

[1001] Dong Wen, Bohua Yang, and Wenjie Zhang. Span-reachability querying in large temporal graphs. VLDB Journal: Very Large Data Bases, 31 (4):629–647, July 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00715-z.

Khalil:2022:PML

[1002] Jalal Khalil, Da Yan, and Lyuheng Yuan. Parallel mining of large maximal quasi-cliques. VLDB Journal: Very Large Data Bases, 31 (4):649–674, July 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00712-2.

Kellou-Menouer:2022:SSS

[1003] Kenza Kellou-Menouer, Nikolaos Kardoulakis, and Haridimos Kondylakis. A survey on semantic schema discovery. VLDB Journal: Very Large Data Bases, 31 (4):675–710, July 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00717-x.

Fritz:2022:EEC

[1004] Manuel Fritz, Michael Behringer, and Holger Schwarz. Efficient exploratory clustering analyses in large-scale exploration processes. VLDB Journal: Very Large Data Bases, 31 (4):711–732, July 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00716-y.

Zheng:2022:PPW

[1005] Libin Zheng, Lei Chen, and Peng Cheng. Privacy-preserving worker allocation in crowdsourcing. VLDB Journal: Very Large Data Bases, 31 (4):733–751, July 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00713-1.

Qin:2022:IDR

[1006] Xuedi Qin, Chengliang Chai, and Mourad Ouzzani. Interactively discovering and ranking desired tuples by data exploration. VLDB Journal: Very Large Data Bases, 31 (4):753–777, July 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00714-0.

Zhu:2022:OPP

[1007] Yuqing Zhu, Jing Tang, and Xueyan Tang. Optimal price profile for influential nodes in online social networks. VLDB Journal: Very Large Data Bases, 31(4):779–795, July 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-021-00727-
9.

Amagata:2022:FEP

[1008] Daichi Amagata, Makoto Onizuka, and Takahiro Hara. Fast, exact, and parallel-friendly outlier detection algorithms with proximity graph in metric spaces. VLDB Journal: Very Large Data Bases, 31 (4):797–821, July 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00729-1.

Huang:2022:SIR

[1009] Zi Huang, Yanyan Shen, and Divesh Srivastava. Special issue on responsible data management and data science. VLDB Journal: Very Large Data Bases, 31 (5):823, September 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00761-1.

Li:2022:AOD

[1010] Pei Li, Jaroslaw Szlichta, and Divesh Srivastava. ABC of order dependencies. VLDB Journal: Very Large Data Bases, 31(5): 825–849, September 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-021-00696-z.

Principe:2022:AHS

[1011] Renzo Arturo Alva Principe, Andrea Maurino, and Blerina Spahiu. ABSTAT-HD: a scalable tool for profiling very large knowledge graphs. VLDB Journal: Very Large Data Bases, 31(5):851–876, September 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00778-021-00704-
                                        2.

Wang:2022:FAP

[1012] Qinyong Wang, Hongzhi Yin, and Xiangliang Zhang. Fast-adapting and privacy-preserving federated recommender system. VLDB Journal: Very Large Data Bases, 31(5): 877–896, September 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-021-00700-6.

Xiang:2022:GGG

[1013] Sheng Xiang, Dong Wen, and Xuemin Lin. General graph generators: experiments, analyses, and improvements. VLDB Journal: Very Large Data Bases, 31(5):897– 925, September 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-021-00701-5.

Liu:2022:PGA

[1014] Zifan Liu, Zhechun Zhou, and Theodoros Rekatsinas. Picket: guarding against corrupted data in tabular data during learning and inference. VLDB Journal: Very Large Data Bases, 31(5):927– 955, September 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00699-w.

Ge:2022:MMD

[1015] Yong-Feng Ge, Maria Orlowska, and Yanchun Zhang. MDDE: multitasking distributed differential evolution for privacy-preserving database fragmentation. VLDB Journal: Very Large Data Bases, 31(5):957– 975, September 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00718-w.

Panjei:2022:SOE

[1016] Egawati Panjei, Le Gruenwald, and Shejuti Silvia. A survey on outlier explanations. VLDB Journal: Very Large Data Bases, 31(5):977– 1008, September 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00721-1.

Zeng:2022:EAS

[1017] Weixin Zeng, Xiang Zhao, and Wei Wang. On entity alignment at scale. VLDB Journal: Very Large Data Bases, 31(5):1009– 1033, September 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-021-00703-3.

Xu:2022:PBP

[1018] Qingyu Xu, Feng Zhang, and Xiaoyong Du. Payment behavior prediction on shared parking lots with TR-GCN. VLDB Journal: Very Large Data Bases, 31(5):1035– 1058, September 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-021-00722-0.

Sadiq:2022:IRN

[1019] Shazia Sadiq, Amir Aryani, and Xiaofang Zhou. Information resilience: the nexus of responsible and agile approaches to information use. VLDB Journal: Very Large Data Bases, 31(5): 1059–1084, September 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-021-00720-2.

Liu:2022:ECR

[1020] Fanzhen Liu, Zhao Li, and Quan Z. Sheng. eRiskCom: an e-commerce risky community detection platform. VLDB Journal: Very Large Data Bases, 31(5):1085–1101, September 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00778-021-00723-
                                        z.

Grafberger:2022:DDD

[1021] Stefan Grafberger, Paul Groth, and Sebastian Schelter. Data distribution debugging in machine learning pipelines. VLDB Journal: Very Large Data Bases, 31(5):1103– 1126, September 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00726-w.

Li:2022:DTA

[1022] Qian Li, Zhichao Wang, and Guandong Xu. Deep treatment-adaptive network for causal inference. VLDB Journal: Very Large Data Bases, 31(5): 1127–1142, September 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00724-y.

Zhang:2022:BCS

[1023] Rui Zhang, Bayu Distiawan Trisedya, and Jianzhong Qi. A benchmark and comprehensive survey on knowledge graph entity alignment via representation learning. VLDB Journal: Very Large Data Bases, 31(5): 1143–1168, September 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00747-z.

Porobic:2022:SIB

[1024] Danica Porobic. Special issue on the best papers of DaMoN 2020. VLDB Journal: Very Large Data Bases, 31 (6):1169, November 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00766-w.

Funke:2022:LLQ

[1025] Henning Funke, Jan Mühlig, and Jens Teubner. Low-latency query compilation. VLDB Journal: Very Large Data Bases, 31(6):1171–1184, November 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00778-022-00741-
                                        5.

Bang:2022:FSC

[1026] Tiemo Bang, Norman May, Ilia Petrov, and Carsten Binnig. The full story of 1000 cores. VLDB Journal: Very Large Data Bases, 31(6): 1185–1213, November 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-022-00742-4.

Pietrzyk:2022:SSV

[1027] Johannes Pietrzyk, Alexander Krause, Dirk Habich, and Wolfgang Lehner. To share or not to share vector registers? VLDB Journal: Very Large Data Bases, 31(6):1215– 1236, November 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-022-00744-2.

Balazinska:2022:EV

[1028] Magdalena Balazinska and Xiaofang Zhou. Editorial for S.I.: VLDB 2020. VLDB Journal: Very Large Data Bases, 31(6):1237– 1238, November 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00734-4.

Huang:2022:OOC

[1029] Yihe Huang, William Qian, Eddie Kohler, Barbara Liskov, and Liuba Shrira. Opportunities for optimism in contended main-memory multicore transactions. VLDB Journal: Very Large Data Bases, 31(6): 1239–1261, November 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00719-9.

Kandula:2022:DIP

[1030] Srikanth Kandula, Laurel Orr, and Surajit Chaudhuri. Data-induced predicates for sideways information passing in query optimizers. VLDB Journal: Very Large Data Bases, 31(6): 1263–1290, November 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00693-2.

Herlihy:2022:CCD

[1031] Maurice Herlihy, Barbara Liskov, and Liuba Shrira. Cross-chain deals and adversarial commerce. VLDB Journal: Very Large Data Bases, 31(6): 1291–1309, November 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00686-1.

Li:2022:AAC

[1032] Yuanbing Li, Xian Wu, Yifei Jin, Jian Li, Guoliang Li, and Jianhua Feng. Adapative algorithms for crowdaided categorization. VLDB Journal: Very Large Data Bases, 31(6): 1311–1337, November 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00685-2.

Zheng:2022:PLF

[1033] Bolong Zheng, Xi Zhao, Lianggui Weng, Quoc Viet Hung Nguyen, Hang Liu, and Christian S. Jensen. PM-LSH: a fast and accurate in-memory framework for high-dimensional approximate NN and closest pair search. VLDB Journal: Very Large Data Bases, 31(6): 1339–1363, November 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00680-7.

Lyu:2022:MTE

[1034] Bingqing Lyu, Lu Qin, Xuemin Lin, Ying Zhang, Zhengping Qian, and Jingren Zhou. Maximum and top-k diversified biclique search at scale. VLDB Journal: Very Large Data Bases, 31(6): 1365–1389, November 2022. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-021-00681-6.

Chen:2023:HHF

[1035] Wei Chen, Weiqing Wang, Hongzhi Yin, Lei Zhao, and Xiaofang Zhou. HFUL: a hybrid framework for user account linkage across locationaware social networks. VLDB Journal: Very Large Data Bases, 32 (1):1–22, January 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00730-8.

Ding:2023:FGE

[1036] Zeyu Ding, Yuxin Wang, Yingtai Xiao, Guanhong Wang, Danfeng Zhang, and Daniel Kifer. Free gap estimates from the exponential mechanism, sparse vector, noisy max and related algorithms. VLDB Journal: Very Large Data Bases, 32 (1):23–48, January 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00728-2.

Fan:2023:MGC

[1037] Wenfei Fan, Yuanhao Li, Muyang Liu, and Can Lu. Making graphs compact by lossless contraction. VLDB Journal: Very Large Data Bases, 32 (1):49–73, January 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00731-7.

Lourenco:2023:BID

[1038] Raoni Lourenço, Juliana Freire, Eric Simon, Gabriel Weber, and Dennis Shasha. BugDoc: Iterative debugging and explanation of pipeline. VLDB Journal: Very Large Data Bases, 32 (1):75–101, January 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00733-5. See cor-

rection [1056].

Pankowski:2023:ODF

[1039] Tadeusz Pankowski. Ontological databases with faceted queries. VLDB Journal: Very Large Data Bases, 32 (1):103–121, January 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-022-00735-3.

Liu:2023:ZRN

[1040] Gang Liu, Leying Chen, and Shimin Chen. Zen+: a robust NUMAaware OLTP engine optimized for nonvolatile main memory. VLDB Journal: Very Large Data Bases, 32(1): 123–148, January 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-022-00737-1.

Fan:2023:ADG

[1041] Wenfei Fan, Ruiqi Xu, Qiang Yin, Wenyuan Yu, and Jingren Zhou. Application-driven graph partitioning. VLDB Journal: Very Large Data Bases, 32(1):149–172, January 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00778-022-00736-
                                        2.

Miao:2023:AIR

[1042] Dongjing Miao, Pengfei Zhang, Jianzhong Li, Ye Wang, and Zhipeng Cai. Approximation and inapproximability results on computing optimal repairs. VLDB Journal: Very Large Data Bases, 32(1): 173–197, January 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00738-0.

Maroulis:2023:RAA

[1043] Stavros Maroulis, Nikos Bikakis, George Papastefanatos, Panos Vassiliadis, and Yannis Vassiliou. Resourceaware adaptive indexing for in situ visual exploration and analytics. VLDB Journal: Very Large Data Bases, 32 (1):199–227, January 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00739-z.

Huang:2023:DEM

[1044] Jiacheng Huang, Wei Hu, Zhifeng Bao, Qijin Chen, and Yuzhong Qu. Deep entity matching with adversarial active learning. VLDB Journal: Very Large Data Bases, 32(1): 229–255, January 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00745-1.

Wang:2023:ABC

[1045] Kai Wang, Xuemin Lin, Lu Qin, Wenjie Zhang, and Ying Zhang. Accelerated butterfly counting with vertex priority on bipartite graphs. VLDB Journal: Very Large Data Bases, 32 (2):257–281, March 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00746-0.

Nikookar:2023:DRS

[1046] Sepideh Nikookar, Mohammadreza Esfandiari, Ria Mae Borromeo, Paras Sakharkar, Sihem Amer-Yahia, and Senjuti Basu Roy. Diversifying recommendations on sequences of sets. VLDB Journal: Very Large Data Bases, 32(2):283–304, March 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-022-00740-
6.

Piai:2023:FGS

[1047] Federico Piai, Paolo Atzeni, Paolo Merialdo, and Divesh Srivastava. Fine-grained semantic type discovery for heterogeneous sources using clustering. VLDB Journal: Very Large Data Bases, 32(2): 305–324, March 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00743-3.

Liu:2023:URR

[1048] Hao Liu, Jindong Han, Yanjie Fu, Yanyan Li, Kai Chen, and Hui Xiong. Unified route representation learning for multi-modal transportation recommendation with spatiotemporal pre-training. VLDB Journal: Very Large Data Bases, 32 (2):325–342, March 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00748-y.

Kim:2023:FSQ

[1049] Hyunjoon Kim, Yunyoung Choi, Kunsoo Park, Xuemin Lin, Seok-Hee Hong, and Wook-Shin Han. Fast subgraph query processing and subgraph matching via static and dynamic equivalences. VLDB Journal: Very Large Data Bases, 32 (2):343–368, March 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00749-x.

Nguyen:2023:DRL

[1050] Thanh Tam Nguyen, Thanh Trung Huynh, Hongzhi Yin, Matthias Weidlich, Thanh Thi Nguyen, Thai Son Mai, and Quoc Viet Hung Nguyen. Detecting rumours with latency guarantees using massive streaming data. VLDB Journal: Very Large Data Bases, 32(2):369–387, March 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-022-00750-
4.

Li:2023:VSE

[1051] Yang Li, Yu Shen, Wentao Zhang, Ce Zhang, and Bin Cui. VolcanoML: speeding up end-to-end AutoML via scalable search space decomposition. VLDB Journal: Very Large Data Bases, 32(2): 389–413, March 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00752-2.

Bouganim:2023:HDP

[1052] Luc Bouganim, Julien Loudet, and Iulian Sandu Popa. Highly distributed and privacy-preserving queries on personal data management systems. VLDB Journal: Very Large Data Bases, 32(2):415–445, March 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00778-022-00753-
                                        1.

Chen:2023:IGL

[1053] Jiazun Chen, Jun Gao, and Bin Cui. ICS-GNN^{+}: lightweight interactive community search via graph neural network. VLDB Journal: Very Large Data Bases, 32 (2):447–467, March 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-022-00754-0.

Boniol:2023:CUS

[1054] Paul Boniol, Michele Linardi, Federico Roncallo, Themis Palpanas, Mohammed Meftah, and Emmanuel Remy. Correction to: Unsupervised and scalable subsequence anomaly detection in large data series. VLDB Journal: Very Large Data Bases, 32(2):469, March 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-021-00678-1. See [969].

Kossmann:2023:CDD

[1055] Jan Kossmann, Thorsten Papenbrock, and Felix Naumann. Correction to: Data dependencies for query optimization: a survey. VLDB Journal: Very Large Data Bases, 32(2):471, March 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-021-00710-4. See [976].

Lourenco:2023:CBI

[1056] Raoni Lourenço, Juliana Freire, Eric Simon, Gabriel Weber, and Dennis Shasha. Correction to: Bug-Doc iterative debugging and explanation of pipeline executions. VLDB Journal: Very Large Data Bases, 32(2):473, March 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00751-3.  See

[1038].

Azzalini:2023:EDA

[1057] Fabio Azzalini, Davide Piantella, Emanuele Rabosio, and Letizia Tanca. Enhancing domain-aware multi-truth data fusion using copy-based source authority and value similarity. VLDB Journal: Very Large Data Bases, 32 (3):475–500, May 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00757-x.

Ahmed:2023:RST

[1058] Pritom Ahmed, Ahmed Eldawy, Vagelis Hristidis, and Vassilis J. Tsotras. Reverse spatial top-k keyword queries. VLDB Journal: Very Large Data Bases, 32 (3):501–524, May 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00759-9.

Li:2023:POB

[1059] Meng Li, Rongbiao Xie, Deyi Chen, Haipeng Dai, Rong Gu, He Huang, Wanchun Dou, and Guihai Chen. A Pareto optimal Bloom filter family with hash adaptivity. VLDB Journal: Very Large Data Bases, 32 (3):525–548, May 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00755-z.

Lei:2023:HDP

[1060] Chuan Lei, Abdul Quamar, Vasilis Efthymiou, Fatma Özcan, and Rana Alotaibi. HERMES: data placement and schema optimization for enterprise knowledge bases. VLDB Journal: Very Large Data Bases, 32 (3):549–574, May 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00756-y.

Li:2023:ENQ

[1061] Jiajia Li, Cancan Ni, Dan He, Lei Li, Xiufeng Xia, and Xiaofang Zhou. Efficient k NN query for moving objects on time-dependent road networks. VLDB Journal: Very Large Data Bases, 32 (3):575–594, May 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00758-w.

Liu:2023:MCS

[1062] Ziyi Liu, Lei Li, Mengxuan Zhang, Wen Hua, and Xiaofang Zhou. Multiconstraint shortest path using forest hop labeling. VLDB Journal: Very Large Data Bases, 32 (3):595–621, May 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00760-2.

Zhang:2023:LBQ

[1063] Pengcheng Zhang, Bin Yao, Chao Gao, Bin Wu, Xiao He, Feifei Li, Yuanfei Lu, Chaoqun Zhan, and Feilong Tang. Learning-based query optimization for multi-probe approximate nearest neighbor search. VLDB Journal: Very Large Data Bases, 32 (3):623–645, May 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00762-0.

Luo:2023:TMH

[1064] Qi Luo, Dongxiao Yu, Zhipeng Cai, Xuemin Lin, Guanghui Wang, and Xiuzhen Cheng. Toward maintenance of hypercores in large-scale dynamic hypergraphs. VLDB Journal: Very Large Data Bases, 32 (3):647–664, May 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00763-z.

Zhang:2023:PPB

[1065] Liang Zhang, Noura Alghamdi, Huayi Zhang, Mohamed Y. Eltabakh, and Elke A. Rundensteiner. PARROT: pattern-based correlation exploitation in big partitioned data series. VLDB Journal: Very Large Data Bases, 32 (3):665–688, May 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00767-9.

Wellenzohn:2023:RSC

[1066] Kevin Wellenzohn, Michael H. Böhlen, Sven Helmer, Antoine Pietri, and Stefano Zacchiroli. Robust and scalable content-and-structure indexing. VLDB Journal: Very Large Data Bases, 32 (4):689–715, July 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-022-00764-y.

Miao:2023:PPP

[1067] Xupeng Miao, Wentao Zhang, Yuezihan Jiang, Fangcheng Fu, Yingxia Shao, Lei Chen, Yangyu Tao, Gang Cao, and Bin Cui. P^{2}CG: a privacy preserving collaborative graph neural network training framework. VLDB Journal: Very Large Data Bases, 32 (4):717–736, July 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-022-00768-8.

Islam:2023:GFE

[1068] Md Mouinul Islam, Mahsa Asadi, Sihem Amer-Yahia, and Senjuti Basu Roy. A generic framework for efficient computation of top-k diverse results. VLDB Journal: Very Large Data Bases, 32(4): 737–761, July 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-022-00770-0.

Echihabi:2023:PDS

[1069] Karima Echihabi, Theophanis Tsandilas, Anna Gogolou, Anastasia Bezerianos, and Themis Palpanas. ProS: data series progressive k-NN similarity search and classification with probabilistic quality guarantees. VLDB Journal: Very Large Data Bases, 32 (4):763–789, July 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00771-z.

Whang:2023:DCQ

[1070] Steven Euijong Whang, Yuji Roh, Hwanjun Song, and Jae-Gil Lee. Data collection and quality challenges in deep learning: a datacentric AI perspective. VLDB Journal: Very Large Data Bases, 32 (4):791–813, July 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00775-9.

Lou:2023:TTA

[1071] Yunkai Lou, Chaokun Wang, Tiankai Gu, Hao Feng, Jun Chen, and Jeffrey Xu Yu. Time-topology analysis on temporal graphs. VLDB Journal: Very Large Data Bases, 32 (4):815–843, July 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00772-y.

Ntroumpogiannis:2023:MLA

[1072] Antonios Ntroumpogiannis, Michail Giannoulis, Nikolaos Myrtakis, Vassilis Christophides, Eric Simon, and Ioannis Tsamardinos. A meta-level analysis of online anomaly detectors. VLDB Journal: Very Large Data Bases, 32 (4):845–886, July 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-022-00773-x.

Zhang:2023:SSQ

[1073] Dongxiang Zhang, Zhihao Chang, Dingyu Yang, Dongsheng Li, Kian-Lee Tan, Ke Chen, and Gang Chen. SQUID: subtrajectory query in trillionscale GPS database. VLDB Journal: Very Large Data Bases, 32 (4):887–904, July 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-022-00777-7.

Katsogiannis-Meimarakis:2023:SDL

[1074] George Katsogiannis-Meimarakis and Georgia Koutrika. A survey on deep learning approaches for text-to-SQL. VLDB Journal: Very Large Data Bases, 32(4):905–936, July 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00778-022-00776-
                                        8.

Zhao:2023:LSS

[1075] Kangfei Zhao, Jeffrey Xu Yu, Qiyan Li, Hao Zhang, and Yu Rong. Learned sketch for subgraph counting: a holistic approach. VLDB Journal: Very Large Data Bases, 32(5): 937–962, September 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-023-00781-5.

Yamada:2023:ALT

[1076] Masaya Yamada, Hiroyuki Kitagawa, Toshiyuki Amagasa, and Akiyoshi Matono. Augmented lineage: traceability of data analysis including complex UDF processing. VLDB Journal: Very Large Data Bases, 32(5): 963–983, September 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00769-7.

Verwiebe:2023:SWT

[1077] Juliane Verwiebe, Philipp M. Grulich, Jonas Traub, and Volker Markl. Survey of window types for aggregation in stream processing systems. VLDB Journal: Very Large Data Bases, 32(5): 985–1011, September 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00778-6.  See

[1104].

Zhou:2023:BCB

[1078] Alexander Zhou, Yue Wang, and Lei Chen. Butterfly counting and bitruss decomposition on uncertain bipartite graphs. VLDB Journal: Very Large Data Bases, 32(5):1013– 1036, September 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00782-4.

Hirsch:2023:EDK

[1079] Vitali Hirsch, Peter Reimann, Dennis Treder-Tschechlov, Holger Schwarz, and Bernhard Mitschang. Exploiting domain knowledge to address class imbalance and a heterogeneous feature space in multiclass classification. VLDB Journal: Very Large Data Bases, 32(5): 1037–1064, September 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00780-6.

Xu:2023:LUI

[1080] Jia Xu, Zulong Chen, Wanjie Tao, Ziyi Wang, Detao Lv, Yao Yu, and Chuanfei Xu. Leveraging user itinerary to improve personalized deep matching at Fliggy. VLDB Journal: Very Large Data Bases, 32(5): 1065–1086, September 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00787-z.

Gou:2023:SWB

[1081] Xiangyang Gou and Lei Zou. Sliding window-based approximate triangle counting with bounded memory usage. VLDB Journal: Very Large Data Bases, 32(5):1087–1110, September 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-023-00783-
3.

Schiavio:2023:DDQ

[1082] Filippo Schiavio, Daniele Bonetta, and Walter Binder. DynQ: a dynamic query engine with queryreuse capabilities embedded in a polyglot runtime. VLDB Journal: Very Large Data Bases, 32(5):1111– 1135, September 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00784-2.

Yang:2023:BCE

[1083] Jianye Yang, Yun Peng, Dian Ouyang, Wenjie Zhang, Xuemin Lin, and Xiang Zhao. (p, q)-biclique counting and enumeration for large sparse bipartite graphs. VLDB Journal: Very Large Data Bases, 32(5):1137– 1161, September 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00786-0.

Naumann:2023:ESI

[1084] Felix Naumann and Xin Luna Dong.

Editorial: Special issue for selected papers of VLDB 2021. VLDB Journal: Very Large Data Bases, 32 (6):1163, November 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00792-2.

Fent:2023:PPE

[1085] Philipp Fent, Altan Birler, and Thomas Neumann. Practical planning and execution of groupjoin and nested aggregates. VLDB Journal: Very Large Data Bases, 32(6): 1165–1190, November 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00765-x.

Farias:2023:LDD

[1086] Victor A. E. Farias, Felipe T. Brito, Cheryl Flynn, Javam C. Machado, Subhabrata Majumdar, and Divesh Srivastava. Local dampening: differential privacy for non-numeric queries via local sensitivity. VLDB Journal: Very Large Data Bases, 32(6): 1191–1214, November 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-022-00774-w.

Li:2023:EEM

[1087] Yuliang Li, Jinfeng Li, Yoshi Suhara, AnHai Doan, and Wang-Chiew Tan. Effective entity matching with transformers. VLDB Journal: Very Large Data Bases, 32(6):1215–1235, November 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https://link.

springer.com/article/10.1007/s00778-
023-00779-z.

Yang:2023:PSE

[1088] Renchi Yang, Jieming Shi, Xiaokui Xiao, Yin Yang, Sourav S. Bhowmick, and Juncheng Liu. PANE: scalable and effective attributed network embedding. VLDB Journal: Very Large Data Bases, 32(6):1237– 1262, November 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00790-4.

Ouyang:2023:WHM

[1089] Dian Ouyang, Dong Wen, Lu Qin, Lijun Chang, Xuemin Lin, and Ying Zhang. When hierarchy meets 2-hop-labeling: efficient shortest distance and path queries on road networks. VLDB Journal: Very Large Data Bases, 32(6): 1263–1287, November 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00789-x.

Qian:2023:IDD

[1090] Chaoqin Qian, Menglu Li, Zijing Tan, Ai Ran, and Shuai Ma. Incremental discovery of denial constraints. VLDB Journal: Very Large Data Bases, 32(6): 1289–1313, November 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00788-y.

Wang:2023:TGC

[1091] Zuozhi Wang, Kai Zeng, Botong Huang, Wei Chen, Xiaozong Cui, Bo Wang, Ji Liu, Liya Fan, Dachuan Qu, Zhenyu Hou, Tao Guan, Chen Li, and Jingren Zhou. Tempura: a general cost-based optimizer framework for incremental data processing (journal version). VLDB Journal: Very Large Data Bases, 32(6): 1315–1342, November 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00785-1.

Hellings:2023:BSB

[1092] Jelle Hellings and Mohammad Sadoghi.

ByShard: sharding in a Byzantine environment. VLDB Journal: Very Large Data Bases, 32(6):1343– 1367, November 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00794-0.

Papadakis:2023:AOO

[1093] George Papadakis, Vasilis Efthymiou, Emmanouil Thanos, Oktie Hassanzadeh, and Peter Christen. An analysis of one-to-one matching algorithms for entity resolution. VLDB Journal: Very Large Data Bases, 32(6): 1369–1400, November 2023. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-023-00791-3.

Wang:2024:SDC

[1094] Haoyu Wang, Aoqian Zhang, Shaoxu Song, and Jianmin Wang. Streaming data cleaning based on speed change. VLDB Journal: Very Large Data Bases, 33(1):1–24, January 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00778-023-00796-
                                        y.

Aghasadeghi:2024:TGP

[1095] Amir Aghasadeghi, Jan Van den Bussche, and Julia Stoyanovich. Temporal graph patterns by timed automata. VLDB Journal: Very Large Data Bases, 33(1):25–47, January 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00778-023-00795-
                                        z.

Chang:2024:NOA

[1096] Lijun Chang and Zhiyi Wang. A nearoptimal approach to edge connectivitybased hierarchical graph decomposition. VLDB Journal: Very Large Data Bases, 33(1):49–71, January 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-023-00797-
x.

Christodoulou:2024:HHI

[1097] George Christodoulou, Panagiotis Bouros, and Nikos Mamoulis. HINT: a hierarchical interval index for Allen relationships. VLDB Journal: Very Large Data Bases, 33 (1):73–100, January 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00798-w.

Farhan:2024:BBD

[1098] Muhammad Farhan, Henning Koehler, and Qing Wang. BatchHL^{+}: batch dynamic labelling for distance queries on large-scale networks. VLDB Journal: Very Large Data Bases, 33(1): 101–129, January 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00799-9.

Meilicke:2024:ABR

[1099] Christian Meilicke, Melisachew Wudage Chekol, Patrick Betz, Manuel Fink, and Heiner Stuckeschmidt. Anytime bottom-up rule learning for large-scale knowledge graph completion. VLDB Journal: Very Large Data Bases, 33 (1):131–161, January 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00800-5. See cor-

rection [1222].

Zhao:2024:CBT

[1100] Yan Zhao, Kai Zheng, Ziwei Wang, Liwei Deng, Bin Yang, Torben Bach Pedersen, Christian S. Jensen, and Xiaofang Zhou. Coalition-based task assignment with priority-aware fairness in spatial crowdsourcing. VLDB Journal: Very Large Data Bases, 33 (1):163–184, January 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-023-00802-3.

Shaham:2024:SSD

[1101] Sina Shaham, Gabriel Ghinita, and Cyrus Shahabi. Supporting secure dynamic alert zones using searchable encryption and graph embedding. VLDB Journal: Very Large Data Bases, 33 (1):185–206, January 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-023-00803-2.

Ma:2024:ADD

[1102] Chenhao Ma, Yixiang Fang, Reynold Cheng, Laks V. S. Lakshmanan, Xiaolin Han, and Xiaodong Li. Accelerating directed densest subgraph queries with software and hardware approaches. VLDB Journal: Very Large Data Bases, 33(1): 207–230, January 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-023-00805-0.

Mouratidis:2024:QCD

[1103] Kyriakos Mouratidis, Keming Li, and Bo Tang. Quantifying the competitiveness of a dataset in relation to general preferences. VLDB Journal: Very Large Data Bases, 33(1): 231–250, January 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00804-1.

Verwiebe:2024:CSW

[1104] Juliane Verwiebe, Philipp M. Grulich, Jonas Traub, and Volker Markl. Correction to: Survey of window types for aggregation in stream processing systems. VLDB Journal: Very Large Data Bases, 33 (1):251, January 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00793-1.  See

[1077].

Yang:2024:CTF

[1105] Fan Yang, Faisal M. Almutairi, Hyun Ah Song, Christos Faloutsos, Nicholas D. Sidiropoulos, and Vladimir Zadorozhny. Correction to: Turbo-Lift: fast accuracy lifting for historical data recovery. VLDB Journal: Very Large Data Bases, 33 (1):253, January 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00801-4. See [910].

Liu:2024:TDS

[1106] Tongyu Liu, Ju Fan, Guoliang Li, Nan Tang, and Xiaoyong Du. Tabular data synthesis with generative adversarial networks: design space and optimizations. VLDB Journal: Very Large Data Bases, 33 (2):255–280, March 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-023-00807-y.

Karpov:2024:MFA

[1107] Nikolai Karpov, Haoyu Zhang, and Qin Zhang. MinJoin++: a fast algorithm for string similarity joins under edit distance. VLDB Journal: Very Large Data Bases, 33 (2):281–299, March 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-023-00806-z.

Usta:2024:XEN

[1108] Arif Usta, Akifhan Karakayali, and Özgür Ulusoy. xDBTagger: explainable natural language interface to databases using keyword mappings and schema graph. VLDB Journal: Very Large Data Bases, 33 (2):301–321, March 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-023-00809-w.

Wang:2024:CEU

[1109] Jiayi Wang, Chengliang Chai, Jiabin Liu, and Guoliang Li. Cardinality estimation using normalizing flow. VLDB Journal: Very Large Data Bases, 33 (2):323–348, March 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-023-00808-x.

Arroyuelo:2024:ORC

[1110] Diego Arroyuelo, Adrián Gómez-Brandón, Aidan Hogan, Gonzalo Navarro, and Javiel Rojas-Ledesma. Optimizing RPQs over a compact graph representation. VLDB Journal: Very Large Data Bases, 33 (2):349–374, March 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00811-2.

Chen:2024:QEP

[1111] Zhiwen Chen, Daokun Hu, Wenkui Che, Jianhua Sun, and Hao Chen. A quantitative evaluation of persistent memory hash indexes. VLDB Journal: Very Large Data Bases, 33 (2):375–397, March 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00812-1.

Abello:2024:EEM

[1112] Alberto Abelló and James Cheney.

Eris: efficiently measuring discord in multidimensional sources. VLDB Journal: Very Large Data Bases, 33 (2):399–423, March 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00810-3.

Jiang:2024:SEM

[1113] Jiawei Jiang, Shaoduo Gan, Bo Du, Gustavo Alonso, Ana Klimovic, Ankit Singla, Wentao Wu, Sheng Wang, and Ce Zhang. A systematic evaluation of machine learning on serverless infrastructure. VLDB Journal: Very Large Data Bases, 33 (2):425–449, March 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-023-00813-0.

Zhang:2024:STS

[1114] Shuhao Zhang, Juan Soto, and Volker Markl. A survey on transactional stream processing. VLDB Journal: Very Large Data Bases, 33 (2):451–479, March 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-023-00814-z.

dHondt:2024:EDM

[1115] Jens E. d’Hondt, Koen Minartz, and Odysseas Papapetrou. Efficient detection of multivariate correlations with different correlation measures. VLDB Journal: Very Large Data Bases, 33 (2):481–505, March 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-023-00815-y.

Fragkoulis:2024:SES

[1116] Marios Fragkoulis, Paris Carbone, Vasiliki Kalavri, and Asterios Katsifodimos. A survey on the evolution of stream processing systems. VLDB Journal: Very Large Data Bases, 33 (2):507–541, March 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-023-00819-8.

Zhao:2024:RRE

[1117] Hongyao Zhao, Jingyao Li, Wei Lu, Qian Zhang, Wanqing Yang, Jiajia Zhong, Meihui Zhang, Haixiang Li, Xiaoyong Du, and Anqun Pan. RCBench: an RDMA-enabled transaction framework for analyzing concurrency control algorithms. VLDB Journal: Very Large Data Bases, 33 (2):543–567, March 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00821-0.

Zhang:2024:LCS

[1118] Junhua Zhang, Long Yuan, Wentao Li, Lu Qin, Ying Zhang, and Wenjie Zhang. Label-constrained shortest path query processing on road networks. VLDB Journal: Very Large Data Bases, 33(3): 569–593, May 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00825-w.

Fang:2024:NWC

[1119] James Fang, Dmitry Lychagin, Michael J.

Carey, and Vassilis J. Tsotras. A new window clause for SQL++. VLDB Journal: Very Large Data Bases, 33 (3):595–623, May 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00830-z.

Lee:2024:HMT

[1120] Geon Lee, Seokbum Yoon, Jihoon Ko, Hyunju Kim, and Kijung Shin.

Hypergraph motifs and their extensions beyond binary. VLDB Journal: Very Large Data Bases, 33 (3):625–665, May 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00827-8.

Liao:2024:SDG

[1121] Ningyi Liao, Dingheng Mo, Siqiang Luo, Xiang Li, and Pengcheng Yin. Scalable decoupling graph neural network with feature-oriented optimization. VLDB Journal: Very Large Data Bases, 33(3):667–683, May 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-023-00829-
6.

Liang:2024:STC

[1122] Anqi Liang, Bin Yao, Bo Wang, Yinpei Liu, Zhida Chen, Jiong Xie, and Feifei Li. Sub-trajectory clustering with deep reinforcement learning. VLDB Journal: Very Large Data Bases, 33 (3):685–702, May 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00833-w.

Yao:2024:ISB

[1123] Kai Yao, Lijun Chang, and Jeffrey Xu Yu. Identifying similarbicliques in bipartite graphs. VLDB Journal: Very Large Data Bases, 33 (3):703–726, May 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00834-9.

Xia:2024:TSD

[1124] Tianrui Xia, Jinzhao Xiao, Yuxiang Huang, Changyu Hu, Shaoxu Song, Xiangdong Huang, and Jianmin Wang. Time series data encoding in Apache IoTDB: comparative analysis and recommendation. VLDB Journal: Very Large Data Bases, 33 (3):727–752, May 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00840-5.

Ting:2024:NDT

[1125] Kai Ming Ting, Zongyou Liu, Lei Gong, Hang Zhang, and Ye Zhu. A new distributional treatment for time series anomaly detection. VLDB Journal: Very Large Data Bases, 33 (3):753–780, May 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00832-x.

Gong:2024:IAI

[1126] Shufeng Gong, Chao Tian, Qiang Yin, Zhengdong Wang, Song Yu, Yanfeng Zhang, Wenyuan Yu, Liang Geng, Chong Fu, Ge Yu, and Jingren Zhou. Ingress: an automated incremental graph processing system. VLDB Journal: Very Large Data Bases, 33 (3):781–806, May 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00838-z.

Lin:2024:RRE

[1127] Hong Lin, Ke Chen, Dawei Jiang, Lidan Shou, and Gang Chen. Refiner: a reliable and efficient incentivedriven federated learning system powered by blockchain. VLDB Journal: Very Large Data Bases, 33 (3):807–831, May 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00839-y. See cor-

rection [1166].

Jiang:2024:HGM

[1128] Jiawei Jiang, Yi Wei, Yu Liu, Wentao Wu, Chuang Hu, Zhigao Zheng, Ziyi Zhang, Yingxia Shao, and Ce Zhang. How good are machine learning clouds? Benchmarking two snapshots over 5 years. VLDB Journal: Very Large Data Bases, 33 (3):833–857, May 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00842-3.

Magalhaes:2024:MDM

[1129] Arlino Magalhaes, Angelo Brayner, and Jose Maria Monteiro. MM-DIRECT: Main memory database instant recovery with tuple consistent checkpoint. VLDB Journal: Very Large Data Bases, 33 (3):859–882, May 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00846-z.

Jia:2024:HNA

[1130] Tong Jia, Ying Li, Yong Yang, and Gang Huang. Hilogx: noiseaware log-based anomaly detection with human feedback. VLDB Journal: Very Large Data Bases, 33 (3):883–900, May 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00843-2.

Boehm:2024:SIM

[1131] Matthias Boehm and Nesime Tatbul. Special issue on “Machine learning and databases”. VLDB Journal: Very Large Data Bases, 33(4):901, July 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00848-x.

Kara:2024:FIA

[1132] Ahmet Kara, Milos Nikolic, Dan Olteanu, and Haozhe Zhang. F-IVM: analytics over relational databases under updates. VLDB Journal: Very Large Data Bases, 33(4): 903–929, July 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00817-w.

Huang:2024:ERA

[1133] Enhui Huang, Yanlei Diao, Anna Liu, Liping Peng, and Luciano Di Palma. Efficient and robust active learning methods for interactive database exploration. VLDB Journal: Very Large Data Bases, 33 (4):931–956, July 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00816-x.

Neutatz:2024:AHC

[1134] Felix Neutatz, Marius Lindauer, and Ziawasch Abedjan. AutoML in heavily constrained applications. VLDB Journal: Very Large Data Bases, 33 (4):957–979, July 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-023-00820-1.

Meduri:2024:AAL

[1135] Venkata Vamsikrishna Meduri, Abdul Quamar, Chuan Lei, Xiao Qin, and Berthold Reinwald. Alfa: active learning for graph neural network-based semantic schema alignment. VLDB Journal: Very Large Data Bases, 33 (4):981–1011, July 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-023-00822-z.

Olteanu:2024:GRD

[1136] Dan Olteanu, Nils Vortmeier, and – Dor– de Živanović. Givens rotations for QR decomposition, SVD and PCA over database joins. VLDB Journal: Very Large Data Bases, 33 (4):1013–1037, July 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-023-00818-9.

Paganelli:2024:MFA

[1137] Matteo Paganelli, Donato Tiano, and Francesco Guerra. A multifacet analysis of BERT-based entity matching models. VLDB Journal: Very Large Data Bases, 33 (4):1039–1064, July 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00824-x.

Luo:2024:MPM

[1138] Yongping Luo, Peiquan Jin, Zhaole Chu, Xiaoliang Wang, Yigui Yuan, Zhou Zhang, Yun Luo, Xufei Wu, and Peng Zou. Morphtree: a polymorphic main-memory learned index for dynamic workloads. VLDB Journal: Very Large Data Bases, 33 (4):1065–1084, July 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00823-y.

Trummer:2024:DBM

[1139] Immanuel Trummer. DB-BERT: making database tuning tools “read” the manual. VLDB Journal: Very Large Data Bases, 33(4):1085–1104, July 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-023-00831-
y.

Huynh:2024:TFR

[1140] Andy Huynh, Harshal A. Chaudhari, Evimaria Terzi, and Manos Athanassoulis. Towards flexibility and robustness of LSM trees. VLDB Journal: Very Large Data Bases, 33 (4):1105–1128, July 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-023-00826-9.

Redyuk:2024:ADD

[1141] Sergey Redyuk, Zoi Kaoudi, Sebastian Schelter, and Volker Markl. Assisted design of data science pipelines. VLDB Journal: Very Large Data Bases, 33 (4):1129–1153, July 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00835-2.

Vu:2024:LBF

[1142] Tin Vu, Alberto Belussi, Sara Migliorini, and Ahmed Eldawy. A learningbased framework for spatial join processing: estimation, optimization and tuning. VLDB Journal: Very Large Data Bases, 33 (4):1155–1177, July 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00836-1.

Song:2024:SST

[1143] Yuanfeng Song, Raymond Chi-Wing Wong, and Xuefang Zhao. Speechto-SQL: toward speech-driven SQL query generation from natural language question. VLDB Journal: Very Large Data Bases, 33 (4):1179–1201, July 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00837-0.

Shahbazi:2024:REI

[1144] Nima Shahbazi and Abolfazl Asudeh. Reliability evaluation of individual predictions: a data-centric approach. VLDB Journal: Very Large Data Bases, 33(4):1203–1230, July 2024.

CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-024-00857-
w.

Xu:2024:SGD

[1145] Lijie Xu, Shuang Qiu, Binhang Yuan, Jiawei Jiang, Cedric Renggli, Shaoduo Gan, Kaan Kara, Guoliang Li, Ji Liu, Wentao Wu, Jieping Ye, and Ce Zhang. Stochastic gradient descent without full data shuffle: with applications to in-database machine learning and deep learning systems. VLDB Journal: Very Large Data Bases, 33(5): 1231–1255, September 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00845-0.

Karegar:2024:DAI

[1146] Reza Karegar, Melicaalsadat Mirsafian, Parke Godfrey, Lukasz Golab, Mehdi Kargar, Divesh Srivastava, and Jaroslaw Szlichta. Discovering approximate implicit domain orders through order dependencies. VLDB Journal: Very Large Data Bases, 33(5): 1257–1282, September 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00847-y.

Chang:2024:DDT

[1147] Jiwon Chang, Bohan Cui, Fatemeh Nargesian, Abolfazl Asudeh, and H. V. Jagadish. Data distribution tailoring revisited: cost-efficient integration of representative data. VLDB Journal: Very Large Data Bases, 33(5): 1283–1306, September 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00849-w.

Chen:2024:LAL

[1148] Xingguang Chen, Rong Zhu, Bolin Ding, Sibo Wang, and Jingren Zhou. Lero: applying learning-to-rank in query optimizer. VLDB Journal: Very Large Data Bases, 33(5):1307– 1331, September 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00850-3.

Preti:2024:HDO

[1149] Giulia Preti, Gianmarco De Francisci Morales, and Francesco Bonchi. Hyperdistance oracles in hypergraphs. VLDB Journal: Very Large Data Bases, 33(5): 1333–1356, September 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00851-2.

Shi:2024:ECE

[1150] Gongyu Shi, Geng Wang, Shi-Feng Sun, and Dawu Gu. Efficient cryptanalysis of an encrypted database supporting data interoperability. VLDB Journal: Very Large Data Bases, 33(5): 1357–1375, September 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00852-1.

Zhang:2024:SDT

[1151] Chen Jason Zhang, Yunrui Liu, Pengcheng Zeng, Ting Wu, Lei Chen, Pan Hui, and Fei Hao. Similarity-driven and task-driven models for diversity of opinion in crowdsourcing markets. VLDB Journal: Very Large Data Bases, 33(5):1377– 1398, September 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00853-0.

Wang:2024:EAR

[1152] Kai Wang, Minghao Cai, Xiaoshuang Chen, Xuemin Lin, Wenjie Zhang, Lu Qin, and Ying Zhang. Efficient algorithms for reachability and path queries on temporal bipartite graphs. VLDB Journal: Very Large Data Bases, 33(5): 1399–1426, September 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00854-z.

Xu:2024:EEA

[1153] Yichen Xu, Chenhao Ma, Yixiang Fang, and Zhifeng Bao. Efficient and effective algorithms for densest subgraph discovery and maintenance. VLDB Journal: Very Large Data Bases, 33(5): 1427–1452, September 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00855-y.

Wang:2024:PBC

[1154] Zhibin Wang, Longbin Lai, Yixue Liu, Bing Shui, Chen Tian, and Sheng Zhong. Parallelization of butterfly counting on hierarchical memory. VLDB Journal: Very Large Data Bases, 33(5):1453–1484, September 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00778-024-00856-
                                        x.

Song:2024:SHT

[1155] Haoze Song, Wenchao Zhou, Heming Cui, Xiang Peng, and Feifei Li. A survey on hybrid transactional and analytical processing. VLDB Journal: Very Large Data Bases, 33(5): 1485–1515, September 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00858-9.

Peng:2024:MMC

[1156] Peng Peng, Shengyi Ji, M. Tamer Özsu, and Lei Zou. Minimum motifcut: a workload-aware RDF graph partitioning strategy. VLDB Journal: Very Large Data Bases, 33(5): 1517–1542, September 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00860-1.

Xia:2024:GBB

[1157] Yifei Xia, Feng Zhang, Qingyu Xu, Mingde Zhang, Zhiming Yao, Lv Lu, Xiaoyong Du, Dong Deng, Bingsheng He, and Siqi Ma. GPUbased butterfly counting. VLDB Journal: Very Large Data Bases, 33(5): 1543–1567, September 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00861-0.

Kitsios:2024:FGL

[1158] Xenophon Kitsios, Panagiotis Liakos, Katia Papakonstantinopoulou, and Yannis Kotidis. Flexible grouping of linear segments for highly accurate lossy compression of time series data. VLDB Journal: Very Large Data Bases, 33(5):1569– 1589, September 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00862-z.

Pan:2024:SVD

[1159] James Jie Pan, Jianguo Wang, and Guoliang Li. Survey of vector database management systems. VLDB Journal: Very Large Data Bases, 33(5): 1591–1615, September 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00864-x.

Liang:2024:FSF

[1160] Zhiyu Liang and Hongzhi Wang.

FedST: secure federated shapelet transformation for time series classification. VLDB Journal: Very Large Data Bases, 33(5):1617–1641, September 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-024-00865-
w.

Yang:2024:FRC

[1161] Yifei Yang, Xiangyao Yu, Marco Serafini, Ashraf Aboulnaga, and Michael Stonebraker. FlexpushdownDB: rethinking computation pushdown for cloud OLAP DBMSs. VLDB Journal: Very Large Data Bases, 33(5): 1643–1670, September 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00867-8.

Neuhof:2024:OBF

[1162] Franziska Neuhof, Marco Fisichella, George Papadakis, Konstantinos Nikoletos, Nikolaus Augsten, Wolfgang Nejdl, and Manolis Koubarakis. Open benchmark for filtering techniques in entity resolution. VLDB Journal: Very Large Data Bases, 33(5): 1671–1696, September 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00868-7.

Liu:2024:WUG

[1163] Zirui Liu, Fenghao Dong, Chengwu Liu, Xiangwei Deng, Tong Yang, Yikai Zhao, Jizhou Li, Bin Cui, and Gong Zhang. WavingSketch: an unbiased and generic sketch for finding top-k items in data streams. VLDB Journal: Very Large Data Bases, 33(5): 1697–1722, September 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00869-6.

Zhang:2024:FES

[1164] Xingyi Zhang, Jinchao Huang, Fangyuan Zhang, and Sibo Wang. FICOM: an effective and scalable active learning framework for GNNs on semisupervised node classification. VLDB Journal: Very Large Data Bases, 33(5): 1723–1742, September 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00870-z.

Wu:2024:AZS

[1165] Xinle Wu, Xingjian Wu, Bin Yang, Lekui Zhou, Chenjuan Guo, Xiangfei Qiu, Jilin Hu, Zhenli Sheng, and Christian S. Jensen. AutoCTS++: zero-shot joint neural architecture and hyperparameter search for correlated time series forecasting. VLDB Journal: Very Large Data Bases, 33(5): 1743–1770, September 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00872-x.

Lin:2024:CRR

[1166] Hong Lin, Ke Chen, Dawei Jiang, Lidan Shou, and Gang Chen. Correction to: “Refiner: a reliable and efficient incentive-driven federated learning system powered by blockchain”. VLDB Journal: Very Large Data Bases, 33 (5):1771, September 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00866-9.  See

[1127].

May:2024:SIM

[1167] Norman May, Spyros Blanas, and Danica Porobic. Special issue: modern hardware. VLDB Journal: Very Large Data Bases, 33(6):1773– 1774, November 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00841-4.

Nicholson:2024:HME

[1168] Hamish Nicholson, Periklis Chrysogelos, and Anastasia Ailamaki. HPCache: memory-efficient OLAP through proportional caching revisited. VLDB Journal: Very Large Data Bases, 33(6): 1775–1791, November 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-023-00828-7.

Li:2024:PAL

[1169] Tianyu Li, Badrish Chandramouli, and Samuel Madden. Performant almostlatch-free data structures using epoch protection in more depth. VLDB Journal: Very Large Data Bases, 33(6): 1793–1812, November 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00859-8.

Wang:2024:OLB

[1170] Ruihong Wang, Chuqing Gao, Jianguo Wang, Prishita Kadam, M. TamerÖzsu, and Walid G. Aref. Optimizing LSM-based indexes for disaggregated memory. VLDB Journal: Very Large Data Bases, 33(6):1813– 1836, November 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00863-y.

Fan:2024:EST

[1171] Zhuochen Fan, Bowen Ye, Ziwei Wang, Zheng Zhong, Jiarui Guo, Yuhan Wu, Haoyu Li, Tong Yang, Yaofeng Tu, Zirui Liu, and Bin Cui. Enabling space-time efficient range queries with REncoder. VLDB Journal: Very Large Data Bases, 33(6): 1837–1859, November 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00873-w.

He:2024:DCV

[1172] Yizhang He, Kai Wang, Wenjie Zhang, Xuemin Lin, and Ying Zhang. Discovering critical vertices for reinforcement of large-scale bipartite networks. VLDB Journal: Very Large Data Bases, 33(6):1861– 1886, November 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00871-y.

Wang:2024:DDA

[1173] Zeyu Wang, Qitong Wang, Peng Wang, Themis Palpanas, and Wei Wang. DumpyOS: a data-adaptive multi-ary index for scalable data series similarity search. VLDB Journal: Very Large Data Bases, 33(6): 1887–1911, November 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00874-9.

Li:2024:VFA

[1174] Yiran Li, Gongyao Guo, Jieming Shi, Renchi Yang, Shiqi Shen, Qing Li, and Jun Luo. A versatile framework for attributed network clustering via Knearest neighbor augmentation. VLDB Journal: Very Large Data Bases, 33(6): 1913–1943, November 2024. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00875-8.

Pan:2025:AMB

[1175] Dong Pan, Xu Zhou, Wensheng Luo, Zhibang Yang, Qing Liu, Yunjun Gao, and Kenli Li. Accelerating maximum biplex search over large bipartite graphs. VLDB Journal: Very Large Data Bases, 34 (1):??, January 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00882-9.

Arroyuelo:2025:ERP

[1176] Diego Arroyuelo, Adrián Gómez-Brandón, and Gonzalo Navarro. Evaluating regular path queries on compressed adjacency matrices. VLDB Journal: Very Large Data Bases, 34(1):??, January 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00885-6.

Hu:2025:PRF

[1177] Zheng Hu, Cong Xu, and Weiguo Zheng. A powerful reducing framework for accelerating set intersections over graphs. VLDB Journal: Very Large Data Bases, 34 (1):??, January 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00881-w.

Liu:2025:EOR

[1178] Hao Liu and Raymond Chi-Wing Wong. On efficient 3D object retrieval. VLDB Journal: Very Large Data Bases, 34(1):??, January 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-024-00884-
7.

Zeakis:2025:DAP

[1179] Alexandros Zeakis, George Papadakis, Dimitrios Skoutas, and Manolis Koubarakis. An in-depth analysis of pre-trained embeddings for entity resolution. VLDB Journal: Very Large Data Bases, 34 (1):??, January 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00879-4.

Hu:2025:GPM

[1180] Lin Hu, Yinnian Lin, Lei Zou, and M. Tamer Özsu. A graph pattern mining framework for large graphs on GPU. VLDB Journal: Very Large Data Bases, 34 (1):??, January 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00883-8.

Shin:2025:UIL

[1181] Jaewoo Shin, Libin Zhou, Jianguo Wang, and Walid G. Aref. An update-intensive LSM-based R-tree index. VLDB Journal: Very Large Data Bases, 34(1):??, January 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-024-00876-
7.

Georgiadis:2025:RIO

[1182] Thanasis Georgiadis, Eleni Tzirita Zacharatou, and Nikos Mamoulis. Raster interval object approximations for spatial intersection joins. VLDB Journal: Very Large Data Bases, 34(1):??, January 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00887-4.

Mondal:2025:SME

[1183] Manuel Mondal, Mourad Khayati, Hông-Ân Sandlin, and Philippe Cudré-Mauroux. A survey of multimodal event detection based on data fusion. VLDB Journal: Very Large Data Bases, 34 (1):??, January 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00878-5.

Messaoud:2025:TBK

[1184] Aghiles Ait Messaoud, Sonia Ben Mokhtar, and Anthony Simonet-Boulogne. Tee-based key-value stores: a survey. VLDB Journal: Very Large Data Bases, 34 (1):??, January 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00877-6.

Ahmed:2025:RTA

[1185] Waqas Ahmed, Leticia Gómez, Alejandro Vaisman, and Esteban Zimányi. Reconciling tuple and attribute timestamping for temporal data warehouses. VLDB Journal: Very Large Data Bases, 34(1):??, January 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-024-00889-
2.

Guo:2025:DQO

[1186] Yunyan Guo, Guoliang Li, Ruilin Hu, and Yong Wang. In-database query optimization on SQL with ML predicates. VLDB Journal: Very Large Data Bases, 34(1):??, January 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-024-00888-
3.

Mann:2025:STS

[1187] Willi Mann, Nikolaus Augsten, Christian S. Jensen, and Mateusz Pawlik. SWOOP: top-k similarity joins over set streams. VLDB Journal: Very Large Data Bases, 34 (1):??, January 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00880-x.

Song:2025:ETS

[1188] Yitong Song, Bin Yao, Zhida Chen, Xin Yang, Jiong Xie, Feifei Li, and Mengshi Chen. Efficient top-k spatialrange-constrained approximate nearest neighbor search on geo-tagged high-dimensional vectors. VLDB Journal: Very Large Data Bases, 34(1):??, January 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00894-5.

Kim:2025:ESC

[1189] Hyunju Kim, Heechan Moon, Fanchen Bu, Jihoon Ko, and Kijung Shin. Estimating simplet counts via sampling. VLDB Journal: Very Large Data Bases, 34(2):??, March 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00890-9.

Mulder:2025:ONG

[1190] Thomas Mulder, George Fletcher, and Nikolay Yakovets. Optimizing navigational graph queries. VLDB Journal: Very Large Data Bases, 34(2):??, March 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00892-7.

Liu:2025:HGM

[1191] Qiyu Liu, Maocheng Li, Yuxiang Zeng, Yanyan Shen, and Lei Chen. How good are multi-dimensional learned indexes? An experimental survey. VLDB Journal: Very Large Data Bases, 34(2):??, March 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00893-6.

Jiang:2025:EHL

[1192] Changkun Jiang, Heze Lao, Chaorui Zhang, Ji Cheng, Chen Jason Zhang, and Jianqiang Li. HeteroStamp: leveraging heterogeneous social interactions for mobility prediction-enhanced cost-aware spatiotemporal crowdsensing. VLDB Journal: Very Large Data Bases, 34(2):??, March 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00891-8.

Tong:2025:HFE

[1193] Yongxin Tong, Yuxiang Zeng, Yang Song, Xuchen Pan, Zeheng Fan, Chunbo Xue, Zimu Zhou, Xiaofei Zhang, Lei Chen, Yi Xu, Ke Xu, and Weifeng Lv. Hu-Fu: efficient and secure spatial queries over data federation. VLDB Journal: Very Large Data Bases, 34(2):??, March 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00896-3.

Tao:2025:DPE

[1194] Yuchao Tao, Amir Gilad, Ashwin Machanavajjhala, and Sudeepa Roy. Differentially private explanations for aggregate query answers. VLDB Journal: Very Large Data Bases, 34(2):??, March 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00895-4.

Trummer:2025:GHC

[1195] Immanuel Trummer. Generating highly customizable Python code for data processing with large language models. VLDB Journal: Very Large Data Bases, 34(2):??, March 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-025-00900-4.

Peng:2025:AAS

[1196] Jingshu Peng, Qiyu Liu, Zhao Chen, Yingxia Shao, Yanyan Shen, Lei Chen, and Jiannong Cao. From Sancus to Sancus^{q}: staleness and quantizationaware full-graph decentralized training in graph neural networks. VLDB Journal: Very Large Data Bases, 34(2):??, March 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00897-2.

Skavantzos:2025:TBC

[1197] Philipp Skavantzos and Sebastian Link. Third and Boyce–Codd normal form for property graphs. VLDB Journal: Very Large Data Bases, 34(2):??, March 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-025-00902-2.

Helt:2025:CCC

[1198] Jeffrey Helt, Abhinav Sharma, Daniel J. Abadi, Wyatt Lloyd, and Jose M. Faleiro. C5: cloned concurrency control that always keeps up. VLDB Journal: Very Large Data Bases, 34(2):??, March 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-025-00901-3.

Burckhardt:2025:NEE

[1199] Sebastian Burckhardt, Badrish Chandramouli, Chris Gillum, David Justo, Konstantinos Kallas, Connor McMahon, Christopher S. Meiklejohn, and Xiangfeng Zhu. Netherite: efficient execution of serverless workflows. VLDB Journal: Very Large Data Bases, 34(2):??, March 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00898-1.

Yu:2025:QHC

[1200] Yuanhang Yu, Dong Wen, Michael Yu, Lu Qin, Ying Zhang, Wenjie Zhang, and Xuemin Lin. Querying historical K-cores in large temporal graphs. VLDB Journal: Very Large Data Bases, 34(2):??, March 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-025-00903-1.

Miao:2025:ESH

[1201] Xupeng Miao, Hailin Zhang, Yining Shi, Xiaonan Nie, Zhi Yang, Yangyu Tao, Jie Jiang, and Bin Cui. Efficient and scalable huge embedding model training via distributed cache management. VLDB Journal: Very Large Data Bases, 34(3):??, May 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-025-00908-w.

Peng:2025:CAP

[1202] Huanhuan Peng, Xiaoye Miao, Jinshan Zhang, Yunjun Gao, Shuiguang Deng, and Jianwei Yin. Costaware prediction service pricing with incomplete information. VLDB Journal: Very Large Data Bases, 34(3):??, May 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-025-00909-9.

Zhang:2025:MCO

[1203] Zhuoxing Zhang and Sebastian Link. Mixed covers: optimizing updates and queries using minimal keys and functional dependencies. VLDB Journal: Very Large Data Bases, 34(3):??, May 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-025-00910-2.

Liu:2025:DFA

[1204] Chunwei Liu, Anna Pavlenko, Matteo Interlandi, and Brandon Haynes. Data formats in analytical DBMSs: performance trade-offs and future directions. VLDB Journal: Very Large Data Bases, 34(3):??, May 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-025-00911-1.

Li:2025:JOR

[1205] Rui Li, Zongyan He, and Jeffrey Xu Yu. Join optimization revisited: a novel DP algorithm for join&sort order selection. VLDB Journal: Very Large Data Bases, 34(3):??, May 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-025-00906-y.

Boniol:2025:VEE

[1206] Paul Boniol, Ashwin K. Krishna, Marine Bruel, Qinghua Liu, Mingyi Huang, Themis Palpanas, Ruey S.

Tsay, Aaron Elmore, Michael J. Franklin, and John Paparrizos. VUS: effective and efficient accuracy measures for time-series anomaly detection. VLDB Journal: Very Large Data Bases, 34(3):??, May 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-025-00907-x.

Yin:2025:LLI

[1207] Ziqi Yin, Shanshan Feng, Shang Liu, Gao Cong, Yew Soon Ong, and Bin Cui. LIST: learning to index spatio-textual data for embedding based spatial keyword queries. VLDB Journal: Very Large Data Bases, 34(3):??, May 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-024-00886-5.

Luo:2025:EIS

[1208] Qi Luo, Wenjie Zhang, and Liping Wang. Efficient indexing and searching of constrained core in hypergraphs. VLDB Journal: Very Large Data Bases, 34(3):??, May 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-025-00915-x.

Li:2025:GEG

[1209] Pengfei Li, Yong Zhang, and Hua Lu. GRELA: Exploiting graph representation learning in effective approximate query processing. VLDB Journal: Very Large Data Bases, 34(3):??, May 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-025-00914-y.

Ji:2025:TID

[1210] Daomin Ji, Hui Luo, and J. Shane Culpepper. Table integration in data lakes unleashed: pairwise integrability judgment, integrable set discovery, and multi-tuple conflict resolution. VLDB Journal: Very Large Data Bases, 34(3):??, May 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-025-00917-9.

Li:2025:DTY

[1211] Qian Li, Peter Kraft, and Michael Stonebraker. DBOS: three years later. VLDB Journal: Very Large Data Bases, 34(3):??, May 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-024-00899-0.

Li:2025:PPL

[1212] Shuaimin Li, Xuanang Chen, and Lei Chen. prompt4vis: prompting large language models with example mining for tabular data visualization. VLDB Journal: Very Large Data Bases, 34(4):??, July 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-025-00912-0.

Budiu:2025:DAI

[1213] Mihai Budiu, Leonid Ryzhyk, and Val Tannen. DBSP: automatic incremental view maintenance for rich query languages. VLDB Journal: Very Large Data Bases, 34(4):??, July 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-025-00922-y.

Tang:2025:TPC

[1214] Dixin Tang, Alan Fekete, and Aditya G. Parameswaran. Transactional panorama: a conceptual framework for user perception in analytical visual interfaces (extended version). VLDB Journal: Very Large Data Bases, 34(4):??, July 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-025-00923-x.

Nikookar:2025:MRR

[1215] Sepideh Nikookar, Sohrab Namazi Nia, and Behrooz Omidvar-Tehrani. Model reusability in reinforcement learning. VLDB Journal: Very Large Data Bases, 34(4):??, July 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-025-00920-0.

Haubenschild:2025:OCA

[1216] Michael Haubenschild and Viktor Leis. OLTP in the cloud: architectures, tradeoffs, and cost. VLDB Journal: Very Large Data Bases, 34(4):??, July 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-025-00913-z.

Zhang:2025:HHK

[1217] Jun Zhang, Jue Wang, and Xuejian Gong. HMI: hierarchical knowledge management for efficient multi-tenant inference in pretrained language models. VLDB Journal: Very Large Data Bases, 34(4):??, July 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-025-00919-7.

Zhang:2025:ECF

[1218] Zhihao Zhang, Jianpeng Qi, and Yanwei Yu. Efficiently counting four-node motifs in large-scale temporal graphs. VLDB Journal: Very Large Data Bases, 34(4):??, July 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-025-00926-8.

Zhou:2025:TIO

[1219] Xinjing Zhou, Xiangpeng Hao, and Michael Stonebraker. Tiered-indexing: Optimizing access methods for skew. VLDB Journal: Very Large Data Bases, 34(4):??, July 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-025-00928-6.

Gao:2025:EAU

[1220] Xiangyu Gao, Xingxing Xiao, and Jianzhong Li. Efficient algorithms for uncertain restricted skyline query processing. VLDB Journal: Very Large Data Bases, 34(4):??, July 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-025-00925-
9.

Li:2025:ATS

[1221] Peng Li, Yeye He, and Surajit Chaudhuri. Auto-tables: synthesizing multistep transformations to relationalize tables without using examples. VLDB Journal: Very Large Data Bases, 34(4):??, July 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-025-00921-z.

Meilicke:2025:CAB

[1222] Christian Meilicke, Melisachew Wudage Chekol, and Heiner Stuckenschmidt. Correction: Anytime bottom-up rule learning for large-scale knowledge graph completion. VLDB Journal: Very Large Data Bases, 34(4):??, July 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-025-00918-8.  See

[1099].

Bonifati:2025:TQT

[1223] Angela Bonifati, Stefania Dumbrava, and Dominik Tomaszuk. Threshold queries in theory and in the wild. VLDB Journal: Very Large Data Bases, 34(4):??, July 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-025-00916-w.

Bonte:2025:LSR

[1224] Pieter Bonte, Christophe Callé, and Riccardo Tommasini. Languages and systems for RDF stream processing, a survey. VLDB Journal: Very Large Data Bases, 34(4):??, July 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-025-00927-
7.

Zhao:2025:ETR

[1225] Hongyao Zhao, Wei Lu, and Xiaoyong Du. An efficient two-round distributed transaction processing approach over heterogeneous networks. VLDB Journal: Very Large Data Bases, 34(4):??, July 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-025-00929-5.

Freire:2025:ESI

[1226] Juliana Freire, Fatma Özcan, and Xuemin Lin. Editorial for special issue: VLDB 2022. VLDB Journal: Very Large Data Bases, 34(4):??, July 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-025-00930-y.

Hou:2025:ESG

[1227] Jiamin Hou, Zhanhao Zhao, and Xiaoyong Du. An efficient and scalable graph database with built-in temporal support. VLDB Journal: Very Large Data Bases, 34(4):??, July 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-025-00932-w.

Cambria:2025:MGR

[1228] Francesco Cambria, Francesco Invernici, and Stefano Ceri. Mine Graph Rule: a new GQL operator for mining association rules in property graph databases. VLDB Journal: Very Large Data Bases, 34(4):??, July 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-025-00934-8.

Graur:2025:SQN

[1229] Dan Graur, Ingo Müller, and Gustavo Alonso. The status-quo in nested data processing for high-energy physics. VLDB Journal: Very Large Data Bases, 34(4):??, July 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-025-00924-w.

Parciak:2025:MAF

[1230] Marcel Parciak, Sebastiaan Weytjens, and Stijn Vansummeren. Measuring approximate functional dependencies: a comparative study. VLDB Journal: Very Large Data Bases, 34(4):??, July 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-025-00931-x.

Saha:2025:BIV

[1231] Arkaprava Saha, Xiangyu Ke, and Laks V. S. Lakshmanan. Beyond influence: voting theory for opinion maximization. VLDB Journal: Very Large Data Bases, 34 (5):??, September 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-025-00933-9.

Ayad:2025:TIL

[1232] Lorraine A. K. Ayad, Grigorios Loukides, and Solon P. Pissis. Text indexing for long patterns using locally consistent anchors. VLDB Journal: Very Large Data Bases, 34 (5):??, September 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-025-00935-7.

Tian:2025:DTD

[1233] Anxin Tian, Alexander Zhou, and Chen Zhang. Distributed truss decomposition over large directed graphs. VLDB Journal: Very Large Data Bases, 34(5):??, September 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00778-025-00938-
                                        4.

Sun:2025:EMS

[1234] Renjie Sun, Chen Chen, and Xuemin Lin. Efficient maximum signed biclique and biplex identification in signed bipartite graphs. VLDB Journal: Very Large Data Bases, 34 (5):??, September 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-025-00939-3.

Li:2025:ROC

[1235] Jiajia Li, Qiulin An, and Xiaofang Zhou. Route optimization with collective spatial keywords: a skyline-based approach. VLDB Journal: Very Large Data Bases, 34 (5):??, September 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-025-00940-w.

Zhang:2025:SAQ

[1236] Yunjia Zhang, Yannis Chronis, and Theodoros Rekatsinas. Simple adaptive query processing vs. learned query optimizers: Observations and analysis. VLDB Journal: Very Large Data Bases, 34(5):??, September 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

URL https://link.springer.com/
article/10.1007/s00778-025-00936-
6.

Zhang:2025:DDL

[1237] Qi Zhang, Rong-Hua Li, and Guoren Wang. Density decomposition on large static and dynamic graphs: algorithms and applications. VLDB Journal: Very Large Data Bases, 34 (6):??, November 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-025-00942-8.

Peng:2025:GSS

[1238] Jinfeng Peng, Hanghai Cui, and Ge Yu.

GARF^{+}: self-supervised and interpretable data cleaning with sequence generative adversarial networks. VLDB Journal: Very Large Data Bases, 34 (6):??, November 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-025-00941-9.

Oulefki:2025:BMB

[1239] Samira Oulefki, Lamia Berkani, and Aicha Mokhtari. BioGITOM: Matching biomedical ontologies with graph isomorphism transformer. VLDB Journal: Very Large Data Bases, 34 (6):??, November 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-025-00943-7.

Kang:2025:PTS

[1240] Hongbo Kang, Yiwei Zhao, and Phillip B. Gibbons. PIM-tree: a skew-resistant index for processingin-memory. VLDB Journal: Very Large Data Bases, 34(6):??, November 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00778-025-00937-
                                        5.

Arroyuelo:2025:CST

[1241] Diego Arroyuelo, Daniela Campos, and Domagoj Vrgoc. CompactLTJ: Space & time efficient leapfrog triejoin on graph databases. VLDB Journal: Very Large Data Bases, 34 (6):??, November 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00778-025-00945-5.

Zhong:2025:FLD

[1242] Kai Zhong, Luming Sun, and Hong Chen. FOSS: a learned doctor for query optimization. VLDB Journal: Very Large Data Bases, 34 (6):??, November 2025. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL https:

//link.springer.com/article/10.
1007/s00778-025-00947-3.